A Bibliography of Publications in Computer Networks and ISDN Systems

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

29 August 2024 Version 1.14

Title word cross-reference

(m, k) [1965]. + [713]. ++ [589]. 3 [783, 1494]. = [713]. ^{+} [762]. ^{3} [593]. _{i} [588, 972]. n [1868].

-ary [1868]. -Firm [1965]. -LMP [557]. -persistent [588, 972]. -tree [762].

/WWW [574].

100-Mbps [2026]. 1000 [2026]. 1000/100 [2026]. 1000/100-Mbps [2026]. 10021 [207]. 100Mbps [1839]. 100VG [1412]. 100VG-AnyLAN [1412]. 10th [2033]. 11 [157–159]. 11th [28, 2034]. 12 [1612, 1613]. 12083 [1784]. 12th [2035]. 13th [2036]. 14th [2037]. 15th [2038]. 16th [2039]. 17th [2040]. 18th [2041]. 1994 [1587]. 1995 [795, 917]. 1997 [2051]. 19th [2043]. 1De [1928]. 1De-Bruijn [1928]. 1Gb [1956]. 1Gb/s [1956]. 1st [2045].

2 [437, 1361, 1974, 2015]. 2-D [1182]. 2/3 [796]. 20th [2046]. 21st [122, 2049]. 22nd [2052]. 23rd [2054]. 2D [1058]. 2HTML [659]. 2nd [2042].

3 [1409]. 3-dimensional [1307]. 3-Tier [1409]. 34CZ [1277]. 3D [1052, 1134, 1155, 1166, 1290, 1309]. 3D-modeling [1166]. 3L [446]. 3M [1348]. 3rd [376].

4.4 [445]. 4th [182].

5th [2044].

6 [710].

7 [2051]. 7th [2053].

8 [1608]. 802 [40, 69, 113, 279, 1368]. 802.6 [408]. ’86 [43, 44]. 8th [78].

9 [1474]. ’95 [442, 795, 917, 973]. ’97 [2052].

A-ISDN [239]. Ability [1587]. ABR [965, 1087, 1088, 1284– 1286, 1288, 1289, 1334, 1894, 1954, 1999, 2000]. Abstract [93]. Academic [195, 249, 448, 449, 451, 463, 470, 680, 690, 1277, 1663, 1665]. accept [832]. Acceptable [336]. acceptance [1167]. Access [8, 23, 27, 61, 65, 68, 71, 87, 109, 166, 169, 177, 210, 215, 240, 247, 274, 299, 309, 324, 392, 411, 441, 524, 560, 580, 582, 583, 598, 600, 603, 608, 633, 637, 722, 756, 757, 763, 779, 787, 790, 797, 800, 836, 846, 852, 865, 884, 889, 930, 937, 948, 968, 1000, 1002, 1024, 1031, 1035, 1065, 1081, 1129, 1139, 1153, 1206, 1226, 1228, 1244, 1261, 1271, 1408–1410, 1561, 1571, 1580, 1596, 1598, 1601, 1644, 1648, 1664, 1668, 1669, 1721, 1725, 1737, 1755, 1792, 1836, 1867, 1887, 1913, 1946, 1966, 1984, 1985]. Accessibility [1212, 1213, 1655]. accessible [1028]. Accessing [1001, 1739]. Accessories [1453]. accident [1297]. accounting [1146]. accumulative [1248]. Acknowledgment [1248, 1287]. Acquisition [1555]. Across [14, 441, 621, 1123, 1739]. ACT [361, 380]. action [492]. action-based [492]. Actions [161, 1021]. Active [855, 882, 998, 1048, 1118, 1657, 2008]. Activities [255, 359, 550]. activity [1129]. activity-centred [1129]. actors [1150]. Ad [1996]. Adaptability [1560]. adaptation [374, 1039]. adapters [689]. Adaptive [21, 319, 389, 486, 514, 590, 803, 804, 964, 1122, 1321, 1323, 1853, 1907, 1947, 1975, 1988, 1999, 2004, 2016]. Added [30]. Adding [768, 1184, 1185, 1484, 1659]. Address [56, 242, 262, 627, 1921, 2003]. Addresses [94, 97]. Addressing [94, 188, 244, 371, 682, 1461]. Adjustment [1361]. Administration [1517, 1629, 1748, 1749]. Administration/HTTPD [1748]. Administrative [14, 315]. Administrators [1641]. Admission [290, 604, 752, 911, 939, 962– 964, 1413, 1863, 1902]. Adopting [1663]. adoption [1268]. ADPCM [267]. Advance [38, 1273, 1980]. Advanced [29, 103, 156, 715, 931, 1177, 1742, 1744, 1757, 1951]. Advances [155, 811, 910]. advantages [1213]. Adventures [1711]. advertisement [894]. advertising [894]. Affecting [17]. after [462, 795, 917, 1150]. Age [1269, 1624]. Agent [24, 813, 895, 1018, 1172, 1265, 1321, 1980]. Agent-based [895, 1980]. Agents [780, 1449, 1450, 1727, 1796, 1797]. aggregate [613]. Agricultural [1518]. AHA [1031]. AI [848]. Aid [1489]. Aided [33]. aids [1040]. air [868]. algebraic [380]. algebras [1223]. Algorithm [3, 39, 111, 223, 230, 388, 513, 544, 586, 663, 944, 1049, 1063, 1085, 1125, 1326, 1385, 1421, 1571, 1823, 1868, 1893, 1943, 1947, 1953, 1986, 2002, 2004]. Algorithm-Based [1986]. Algorithms [74, 211, 228, 240, 276, 319, 321, 615, 628, 951, 1056, 1060, 1230, 1290, 1335, 1384, 1853, 1863, 1895, 2007]. alive [1025]. ALIWEB [655]. All-Optical [608, 1232, 1830]. Allocation [42, 214, 218, 224, 263, 392, 425, 510, 511, 533, 611, 614, 615, 1093, 1225, 1363]. ALOHA [85, 317, 1373]. alternate [429]. Alternating [1966]. Alternative [512, 1364, 1840]. Alternatives [294, 543, 1831]. Amiga [1589]. Amigo [156]. among [487, 998, 1130, 1270]. Amphi [1628]. AMS [1542]. Analysis [4, 7, 27, 37, 59, 60, 63, 82, 83, 127, 130, 160, 213, 221, 228, 229, 240, 406, 408, 411, 422, 424, 430, 433, 528, 535, 541, 562, 628, 631, 640, 646, 696, 725, 742, 778, 797, 803, 805, 829, 835, 915, 942, 945, 971, 977, 1011, 1083, 1090, 1092, 1219, 1220, 1229, 1271, 1295, 1297, 1316, 1337, 1343, 1346, 1357, 1375, 1390, 1410, 1424, 1431, 1439, 1440, 1481, 1818, 1830, 1832, 1833, 1844, 1848, 1860, 1878, 1882, 1884, 1899, 1903, 1906, 1907, 1920, 1935, 1936, 1982]. analytic [522, 970]. Analytical [219, 697, 799, 1076, 1355]. analyze [697]. analyzing [1101]. anatomy [1105]. ancient [1055]. Ancillary [1495]. Anglo [1455]. Anglo-Saxon [1455]. animals [1299]. animated [1171]. Animation [663, 1053, 1791]. annotations [720, 857]. annoyance [1174]. Annual [2044, 2052, 2054]. ANSA [431]. answers [506]. Anti [1695]. Anti-net [1695]. Anticipating [13]. ANU [1124]. AnyLAN [1412]. anyone [1035]. anytime [1035]. Anywhere [1035]. Apache [862, 1127]. API [862, 1551, 1809]. Apollo [105]. Apple [1758]. AppleTalk [1847]. applets [883]. Application [139, 171, 278, 284, 297, 316, 636, 830, 845, 886, 1004, 1042, 1060, 1075, 1086, 1125, 1162, 1187, 1265, 1382, 1383, 1443, 1488, 1640, 1665, 1693, 1772, 1802, 1805, 1835, 1851, 1920, 1963, 1987, 2019]. Application-Driven [1382]. Application-independent [1162]. Application-Level [845, 1086, 1920, 1987]. Application-Oriented [1265, 1805]. Application-Specific [1802]. Application-Tailored [1383]. Application-to-Application [1851]. Applications [89, 229, 292, 399, 400, 405, 457, 525, 605, 738, 846–848, 856, 877, 950, 991, 1021, 1023, 1041, 1166, 1194, 1234, 1241, 1263, 1265, 1273, 1378, 1398, 1426, 1429, 1431, 1433, 1447, 1488, 1544, 1645, 1651, 1676, 1678, 1798, 1804, 1815, 1855, 1865, 1893, 1909, 1940, 2007]. applied [1297]. Applying [383, 958, 1069]. Approach [47, 107, 147, 159, 270, 273, 474, 493, 588, 620, 694, 743, 753, 810, 856, 874, 968, 1037, 1046, 1075, 1118, 1216, 1251, 1378, 1382, 1395, 1402, 1415, 1443, 1808, 1827, 1910, 1977, 2005, 2025]. Approaches [155, 233, 671, 1752, 1905].

Approaching [1528]. Approximate [229, 522, 799]. Approximation [300, 1355, 1858]. April [2047, 2051, 2053]. Aquarelle [1238]. Arbitrary [1362, 1728]. arbitrated [1081]. Arbitration [169, 1916]. Archie [655]. Archie-like [655]. architectural [958, 966, 1051, 1252]. Architecture [31, 140, 267, 277, 313, 355, 512, 534, 557, 594, 624, 685, 792, 806, 840, 847, 873, 885, 937, 954, 955, 957, 959, 974, 1042, 1099, 1114, 1267, 1314, 1320, 1323, 1349, 1366, 1381, 1388, 1427, 1430, 1442, 1686, 1709, 1799, 1820, 1842, 1866, 1909, 1912, 1927, 1955, 1980, 1985, 2009– 2011, 2019, 2030]. Architectures [236, 396, 444, 614, 765, 823, 1215, 1917, 1918]. Archives [1671]. Area [3, 7, 34, 37, 40, 47, 56, 69, 86, 110, 147, 226, 229, 272, 276, 292, 299, 321, 396, 407, 425, 441, 532, 542, 576, 580, 596, 617, 627, 647, 810, 839, 978, 1081, 1094, 1178, 1330, 1352, 1366, 1367, 1373, 1391, 1417, 1438, 1507, 1511, 1584, 1597, 1634, 1776, 1826, 1840, 1846, 1958, 2004]. Arequipa [1336]. Ariadne [1107]. ARQ [281, 370, 698, 1248]. Array [787, 1380, 1598]. Arrival [516, 829, 947, 1902]. arrivals [389, 578, 696]. Art [887, 1302, 1456, 1464, 1466]. articles [1036]. Artificial [288, 1299]. Artist [1463]. arts [1302]. ary [1868]. Asia [1662]. ASN.1 [140, 381, 523]. aspect [609]. Aspects [138, 253, 429, 679, 1513, 1870]. Assembly [162]. Assess [2007]. Assessing [1202]. Assessment [132, 543, 742, 758, 771, 909, 1364, 1568]. assessments [500]. Assets [1240]. Assignment [56, 60, 167, 237, 603]. assist [1148, 1172]. Assistance [1644]. assistant [884, 1176]. Assisted [1123, 1183, 1720]. associated [1101]. Associative [1718]. assurance [876]. Assure [45]. Assuring [259]. Astronomy [784, 1468, 1469]. Astrophysics [1470, 1473]. AstroWeb [1471]. Asynchronous [402, 403, 406, 544, 592, 999, 1363, 1762, 1776, 1934]. AT&T [1546]. Atelier [360]. Atlantic [1607]. ATM [290, 295, 389, 404, 405, 430, 438, 510, 530, 531, 538, 547, 566, 568, 578, 607, 614–616, 645– 648, 650, 689, 695, 700, 756, 765, 799, 800, 809, 840, 910–912, 924, 927, 938, 941, 949, 967, 1080, 1087, 1091, 1094, 1226, 1234, 1250, 1289, 1334, 1336, 1337, 1345, 1350, 1384, 1385, 1387, 1431, 1434, 1435, 1817–1819, 1824, 1828, 1830, 1833, 1848, 1849, 1859, 1864, 1877, 1883, 1885, 1895, 1900, 1902, 1903, 1923, 1934, 1936, 1938, 1955, 1962, 1964, 1966, 1970, 1974, 1985, 1986, 1989, 1990, 2000, 2002, 2011, 2012, 2031, 2032]. ATM [293, 374, 402, 440, 522, 533, 536, 537, 625, 629, 762, 803, 826, 827, 837, 923, 940, 944, 947, 951, 953, 962, 964, 978, 1074, 1076, 1083, 1084, 1086, 1089, 1092, 1093, 1220, 1233, 1284, 1287, 1288, 1346, 1353, 1386, 1413, 1416, 1418, 1429, 1433, 1820–1822, 1826, 1829, 1831, 1832, 1852, 1862, 1863, 1865, 1880, 1882, 1884, 1891, 1899, 1901, 1906, 1913, 1919, 1921, 1922, 1924, 1925, 1935, 1941, 1944, 1945, 1948, 1950, 1961, 1977, 1978, 1988, 1999, 2005, 2013, 2028]. ATM-Based [440, 629, 927, 1435, 1880, 1899, 1913, 1964]. ATM-FDDI [1387]. ATM-to-DQDB [1353]. ATM-UBR [1966]. ATM/STM [568]. Atmospheric [309, 1504]. ATMR [592]. atomic [1021]. Attacking [39]. Attacks [758, 1960]. Attention [1652]. Attention-Holding [1652]. audiences [1274]. Audio [747, 1031, 1134, 1136, 1187, 1200, 1304, 1564, 1785, 1870, 1911]. audio-video [1304]. audio/video [1136, 1187]. auditing [1146]. augmentation [1246]. augmented [1127, 1254]. AURORA [482]. Australia [679, 1662, 2053]. Australian [1149]. Austria [470]. Authentication [23, 457, 732, 1261, 1960, 1995]. Authenticity [1466]. Author [853, 1800]. Author-oriented [853]. Author-Publisher-Reader [1800].

Authoring

[360, 710, 711, 714, 860, 984, 1300, 1808, 1811]. Authority [1459]. Authorization [733, 1735]. Authors [1810]. Auto [781, 1452]. Auto-FAQ [781, 1452]. Automated [1102, 1304, 1535, 1701, 1720, 2023]. Automatic [26, 824, 1101, 1173, 1191, 1308]. Automatically [879, 1737]. Automating [1188, 1479]. Autonet [483]. Autonomous [2004]. Availability [203]. Available [826, 1334]. Avenue [2042]. Average [1629]. Avoidance [227, 228, 234, 235, 397, 422, 533, 599, 601, 1423]. Award [368]. Aware [1421]. awe [1791].

B [419, 509, 611, 644, 762, 798, 912, 927, 1906, 1913, 1942]. B-ISDN [419, 509, 611, 644, 912, 1411, 1913, 1942]. B-ISDNs [798]. B-Tree [1906]. B-WiN [927]. Babel [674]. Back [842, 1900]. Back-Pressure [1900]. Backbone [148, 149, 163, 344, 438, 439, 515, 819, 937, 1833]. Background [2000]. backpressure [614]. Balancing [408, 529, 938, 1280, 1378, 1391, 1846]. Bandwidth [55, 71, 214, 323, 392, 408, 425, 430, 510, 533, 534, 593, 611, 615, 627, 801, 803, 841, 890, 1081, 1282, 1363, 1370, 1385, 1391, 1435, 1637, 1838, 1911, 1969, 1974, 1998, 2024]. Bandwidth-limited [1637]. Bank [1385]. Banking [788, 921, 1522]. Banyan [539, 1882, 1903]. Banyan-Based [1882, 1903]. Base [896, 1275, 1671]. Based [93, 107, 221, 223, 230, 271, 303, 392, 492, 510, 527, 531, 555, 607, 626, 656, 703, 733, 747, 755, 788, 818, 834, 871, 877, 895, 917, 919, 927, 938, 963, 988, 998, 1007, 1037, 1094, 1107, 1122, 1154, 1166, 1169, 1226, 1250, 1254, 1293, 1295, 1298, 1378, 1413, 1430, 1446, 1514, 1516, 1522, 1523, 1540, 1571, 1591, 1603, 1635, 1673, 1678, 1682, 1687, 1751, 1760, 1779, 1789, 1796, 1798, 1808, 1819, 1836, 1850, 1867, 1874, 1880, 1882, 1888, 1892, 1897, 1899, 1902– 1904, 1910, 1911, 1913, 1928, 1929, 1938, 1954, 1963, 1964, 1979, 1986, 2001, 2002, 2009]. based [263, 276, 380, 414, 440, 484, 494, 559, 595, 629, 658, 708, 724, 726, 775, 794, 795, 918, 920, 925, 953, 965, 1011, 1039, 1047, 1051, 1053, 1063, 1074, 1090, 1099, 1114, 1183, 1195, 1264, 1268, 1435, 1606, 1630, 1754, 1980, 2020]. Basic [267, 404, 749, 956]. Basics [901]. basis [703, 1291]. batch [578, 696]. BATS [1447]. be [317, 1336]. bearer [939]. beasts [673]. become [1211]. Becoming [672]. been [426]. Behavior [27, 82, 281, 394, 892, 939, 1204, 1418, 1868]. behavioral [429]. behaviors [1036]. Behaviour [322, 956, 1184, 1318]. behavioural [492]. Being [1463]. Belgium [150]. Bell [212, 1607]. Benchmark [1324]. Benchmarks [1578]. Beneficial [1445]. Beowulf [1455]. Bernoulli [1355]. Best [1083, 1988]. Better [1102, 1629, 1695]. Between [69, 212, 464, 838, 1201, 1218, 1223, 1321, 1722, 1939, 1999]. Beyond [190, 250, 789, 900, 1030, 1095, 1451, 1482, 1503, 1514, 1815]. Bézier [1046]. BGI [1581]. bibliographic [451, 1161]. BIDS [451]. Big [1730]. Billboards [1696]. Binary [169, 1301, 1581, 1916]. Biology [454, 1492, 1493, 1632]. Biomedical [1682]. Birds [1495]. BISDN [645]. Bit [284, 297, 588, 805, 826, 1081, 1085, 1334, 1908, 1934]. Bit-stream [284]. Black [2021]. BLAM [1916]. blazing [716]. Blocked [240]. blocking [633, 840, 1217]. BMSec [812, 922]. Board [255, 1509, 1510, 1839]. bogs [673]. Bonestell [1464]. bookmarking [1005]. bookmarks [879]. Books [1800]. Boomerang [1812]. Borealis [863]. Boston [1635]. Botanic [1458]. both [963]. bottleneck [512]. Bottlenecks [385, 1850]. Bound [1893]. Boundaries [14]. bounded [320]. box [1295]. box-counting [1295]. Branch [166]. Branch-Switching [166]. Bridge [113, 239, 321, 386, 459, 1606, 1939]. bridges [525, 1182]. Bridging [69].

Bringing [1187, 1645, 1668, 1765, 1782, 1791]. Brisbane [2053]. Britannica [1664]. Broadband [202, 266, 294, 511, 518, 533, 534, 563, 581, 643, 649, 752, 754, 796, 839, 937, 966, 1065, 1222, 1247, 1249, 1278, 1347, 1447, 1828, 1907]. Broadcast [72, 211, 422, 763, 1227, 1379, 1949]. broadcasting [606, 815]. Broken [874]. Broken-Link’ [874]. broker [1163, 1242]. brokering [1259]. Browser [737, 857, 877, 894, 1030, 1495, 1681, 1813]. browser-based [877]. Browsers [772, 1453, 1557]. Browsing [708, 734, 861, 1064, 1133, 1188, 1451, 1762, 1774, 1795]. Bruijn [1928]. brush [1047]. BSCW [1039, 1763]. BSD [445]. BSS [616]. BUBL [1667]. Bucket [430, 535, 808, 977, 1287, 1385]. Buffer [162, 224, 281, 299, 510, 536, 613, 614, 646, 695, 696, 829, 977, 1430, 1859, 1890, 2028]. buffer-management [646]. Buffered [84, 539, 915, 1233, 1900]. Buffering [614, 762, 941, 1434, 1876, 1900, 1911]. bufferless [1337]. Buffers [322, 578, 640, 945, 1228, 1361]. Builder [735]. Building [347, 743, 871, 1266, 1488, 1593, 1606, 1668, 1671]. Bulletin [1509, 1510]. Bundle [1908]. Buoys [787, 1598]. Burst [88, 799, 1888]. Burstiness [625, 829]. Bursting [1956]. Bursts [742, 2021]. Bursty [289, 389, 430, 742, 756]. Bus [4, 6, 8–10, 166, 389, 418, 521, 580, 943, 1081, 1390, 1391, 1836]. Business [98, 146, 893, 1066, 1212, 1213, 1519, 1534, 1536, 1883]. Buy [892, 1525]. BWB [740]. byl [159].

C [226, 1622]. C-net [226]. Cable [66, 164, 832]. CAC [1978]. Cache [727, 844, 886–888, 992, 994, 1112, 1157, 1186, 1204, 1280, 1316, 1318, 1321, 1322, 1325, 1329, 1331, 1332, 1953]. caches [1113, 1326]. Caching [262, 654, 771, 843, 993, 995, 996, 1114, 1115, 1315, 1317, 1319, 1321, 1323, 1327, 1328, 1330, 1568, 1571, 1766, 1767]. CAL [1183]. calculate [1076]. Calculations [1897, 1905]. calculus [1059]. Call [366, 632, 963, 1217, 1822, 1863, 1902]. Cambridge [18, 515]. Caml [883]. Camp [1369, 1401]. Camp-On [1369, 1401]. Campus [333, 1448, 1513–1516, 1958]. Campus-Wide [1448, 1513–1515]. can [1102, 1336]. Canada [1739]. Canadian [1517]. Candleweb [1791]. CANSIM [1739]. Capabilities [1412, 1594]. Capability [222, 733]. capability-based [733]. Capacity [38, 41, 167, 386, 593, 1439, 1444]. Capture [1424, 1861, 1868]. Care [1676]. Carrier [65]. Cascaded [1029]. Case [474, 500, 668, 788, 962, 995, 1013, 1231, 1305, 1353, 1365, 1446, 1522, 1602, 1640, 1646, 1648, 1671]. Cash [1806]. cast [1821]. Catalog [1475]. catalogue [1239]. catalogues [1181]. cataloguing [727]. Catastrophe [82]. categories [1009]. CATV [110]. CATV-Type [110]. Causal [2030]. CAVEview [1754]. CBC [531]. CCI [1570, 1789]. CCI-Based [1789]. CCITT [16, 89, 220, 232, 1340]. CCNet [1670]. CD [4, 9, 11, 53, 54, 82, 114, 565, 664, 1407, 1422, 1956, 1967, 2026]. CD-based [221, 263]. CD-ROMs [664]. CDMA [608]. CEC [158, 190, 338, 624]. Cell [406, 581, 590, 607, 625, 700, 800, 914, 963, 1083, 1293, 1339, 1430, 1849, 1900, 1905, 1925, 1945, 1948, 1962, 1986, 1988, 1999, 2031]. cell-based [1293]. Cell-Level [1962]. cell-switching [607]. Cellular [563, 1842, 1930]. Censorship [1461]. Center [1594, 1708]. centered [1098, 1103]. Centralized [1971]. centre [550, 1240, 1302]. centred [1129]. centric [985]. Century [122]. CERN [769, 1555, 1748]. CERT [362]. certificate [833]. certification [922]. CERTs [926]. CFTP [1327]. CGI [1022]. chains [1223]. Challenge [81, 252, 261, 332]. challenges [334, 1212]. Champaign [1670]. changes [880]. Changing [1601, 1861]. Channel [8, 42, 68, 86, 87, 166, 263, 593, 757, 761, 1368– 1371, 1402, 1647, 1836, 1871, 1872, 1984]. Channels [65, 393, 760, 1320]. Chapter [710]. Characteristics [36, 86, 262, 312, 1397, 1433]. Characterization [613, 625, 771, 1395, 1420, 1568, 1829, 1875]. characterizations [1147]. Characterizing [734]. Charging [297, 827, 990]. charter [1577]. Charts [904, 905]. check [391, 833]. check/generator [391]. Chemistry [669, 1491, 1493]. CHEOPS [357]. Chesley [1464]. Chest [1886]. CHIC [1267]. CHIC-Pilot [1267]. Chicago [2042]. China [192]. chip [914]. choices [581]. Christmas [484]. Christmas-tree [484]. Chronicle [1770]. CIESIN [1601]. Circuit [37, 59, 218, 409, 804, 1428]. Circuit-Switched [59]. Circuit/Packet [37]. Circuits [67, 487, 2025]. Cities [1669]. City [1275, 1669, 1740]. CLAN [478]. Clara [2051]. Claris [1477]. Class [62, 141, 911, 1296, 1603, 1867, 1886]. Class-4 [62]. Class-Chest [1886]. Classes [841, 1397]. Classical [1944]. classification [1173, 1291]. Classifying [1816]. Classroom [718, 785, 1199, 1609]. Clean [1203]. Clearing [98]. Clearinghouse [1529, 1688]. click [990, 1030]. Client [528, 656, 846, 861, 1341, 1550, 1678, 1732, 1783, 1809]. client-based [656]. Client-Side [1550, 1783]. Client/Server [846, 1341]. Clients [732, 773, 1552]. Climate [1594, 1599, 1600]. climates [1192]. clip [233]. clipping [1056]. CLNS [340]. clock [766]. clocks [527]. Clockwork [1698]. Closer [1429]. Cluster [997, 1404, 1979, 2018]. Clustering [176, 1010, 1015]. CNM [1347, 1348]. co [1113]. co-operating [1113]. coax [1219]. coaxial [149]. Code [177, 608, 769, 1556]. code-division [608]. Coded [2013].

Coding [53, 1853]. coherence [844]. coherency [1112]. Collaboration [774, 998, 1244, 1501, 1505, 1511, 1660, 1760, 1761]. Collaborative [775, 929, 1048, 1064, 1115, 1165, 1169, 1205, 1279, 1613, 1751, 1763]. Collaboratory [1504]. Collector [1532]. College [1469, 1513, 1518]. Collision [9, 84, 169, 215, 227, 234, 235, 422, 521, 545, 599, 601, 802, 1421, 1474, 1967]. collision-avoidance [422, 601]. Collision-free [169, 215, 802]. color [1309]. colorimetric [1133]. colour [1312]. Combined [736, 968]. Combining [1620, 1718]. Come [225]. command [1041]. command-line [1041]. CommBridge [2019]. Commerce [1163, 1260, 1807, 2020]. Commercial [500, 801, 1521, 1524, 1645, 1706, 1707, 1800]. Commission [190]. commodity [579]. Common [107, 440, 769, 1191, 1345, 1430, 1556]. Communicating [133, 751, 866, 1632, 1797]. Communication [6, 12, 20, 52, 58, 78, 79, 119, 121, 124, 128– 131, 136, 142, 146, 161, 171, 217, 318, 335, 342, 347, 358, 376, 401, 415, 428, 501, 516, 526, 556, 633, 759, 811, 932, 981, 988, 1209, 1304, 1359, 1376, 1405, 1416, 1427, 1438, 1508, 1531, 1697, 1739, 1857, 1930, 1932, 1937, 1943, 1946, 1968, 2011, 2029]. Communications [81, 98, 100, 127, 143, 163, 253, 314, 337, 354, 370, 382, 474, 563, 619, 628, 822, 823, 955, 1083, 1248, 1388, 1423, 1734, 1872, 1881, 1919, 1934, 2010]. Communities [702, 1152, 1275]. Community [249, 254, 335, 359, 451, 465, 469, 548, 821, 927, 1141, 1257, 1517, 1536, 1610, 1619, 1654, 1670]. community-shared [1141]. company [127, 212]. Comparative [605, 635, 689, 808, 1250, 1905, 1933, 1991]. Comparing [559]. Comparison [32, 37, 262, 321, 682, 755, 965, 1218, 1321, 1335, 1340, 1368, 1386, 1403, 1411, 1773, 1831, 1934, 1999, 2015, 2026]. Comparisons [88, 271]. compatibility [948]. Compatible [25, 950]. compilation [1101]. Complementary [39]. complete [1067]. Complex [39, 1109, 1561]. compliant [1022]. components [341, 1239]. Compositing [1311]. Comprehensive [264, 455, 1251]. Compression [1585]. Computational [750, 1493, 1575]. computed [1340]. Computer [5, 20, 23, 33, 49, 51, 53, 67, 78, 79, 81, 83, 111, 120, 158, 179, 189, 228, 237– 239, 262, 266, 273, 316, 348, 362, 409, 438, 448, 455, 499, 692, 758, 1058, 1060, 1123, 1183, 1290, 1298, 1302, 1311, 1313, 1461, 1508, 1545, 1574, 1615, 1673, 1776, 1832, 1850, 1899, 2033, 2038]. Computer-Aided [33]. computer-based [1298]. Computer-Mediated [1302, 1508]. Computers [1377]. Computing [300, 337, 431, 1381, 1511, 1582, 1617, 1947, 1959, 2018]. concentration [484]. Concentrator [1358]. concentrators [341, 832]. Concept [108, 807, 830, 953, 1270, 1718, 1771]. concepts [348, 956]. Conceptual [725, 831, 1164]. Concerns [1496, 1715]. concise [353, 458]. concurrent [492, 610]. condition [317]. Conditions [1428, 1973]. Conference [5, 43, 51, 78, 119, 376, 550, 795, 917, 1255, 1379, 2033–2054]. Conferencing [120, 499, 550, 621, 1437, 1528, 1620, 1760, 1920, 1931]. Configuration [528, 1097, 1374, 1383, 1887]. Configurations [1840]. Configuring [1801, 1847]. confirmed [1733]. Confman [1255]. Conformance [117, 236, 259, 264, 303, 523, 700, 932, 934, 1246]. Congested [1977]. Congestion [21, 50, 228, 397, 487, 531, 533– 535, 577, 839, 910, 943, 949, 1087, 1286, 1351, 1385, 1821, 1899, 1913, 1954]. Congestion-Control [50]. Congress [28, 2042]. Connect [1077, 1405]. Connected [49, 339, 889]. Connecting [1670]. Connection [226, 444, 527, 585, 968, 1128, 1201, 1230, 1236, 1922]. Connectionless [447, 629, 978]. Connections [1336, 1821, 1945, 1971, 1981, 1983]. Connectivity [248, 552, 1019, 1139, 1345]. CONS [340, 445]. CONS/CLNS [340]. conscious [995]. Consequences [1156, 1520]. Considerations [410, 862, 1289, 1417, 1476]. Considering [1893]. consistency [727]. consistent [1126]. Consortia [1529]. Constellation [1798]. constraint [523]. Constraints [911, 1362]. Constructing [230, 270, 717, 877, 1021, 1543]. Construction [526, 1017, 1479, 1908]. consulting [810]. contemporary [479]. Content [724, 985, 1143, 1155, 1320, 1699, 1745]. Content-centric [985]. Contention [87, 614, 1360]. contents [879]. context [523, 955, 1171]. Contextual [1651]. continuation [395, 588]. continuation-bit [588]. continuation-of-message [395]. Continuing [103]. Continuous [284, 560, 623, 746, 834, 1061, 1914]. Contract [1545]. contribution [445, 472]. contributions [795, 917]. Control [23, 36, 50, 53, 59, 71, 218, 287, 290, 293, 319, 328, 405, 411, 427, 472, 487, 509, 511, 531, 532, 534, 535, 569, 577, 587, 597, 603, 604, 633, 647, 648, 685, 752, 756, 796, 824, 834, 839, 842, 870, 910, 911, 943, 949, 951, 962, 964, 965, 968, 1038, 1087, 1090, 1094, 1226, 1228, 1236, 1255, 1284, 1286, 1351, 1352, 1360, 1389, 1396, 1400, 1405, 1409, 1434, 1436, 1502, 1536, 1792, 1821, 1822, 1855, 1859, 1863, 1875, 1887, 1892, 1894, 1897, 1899, 1901, 1902, 1911, 1913– 1915, 1922, 1954, 1974, 1985, 1999, 2023, 2025]. Controlled [667, 759, 939, 1089]. controller [949]. controlling [800]. Controls [21, 637]. Controversies [153]. Convection [1649]. Convection-Diffusion [1649]. Conventional [1718]. convergence [690]. conversational [1030]. conversion [1702]. Converters [189, 916]. Converting [1488]. convey [1055]. COOL [1180]. Coollists [782, 1719]. Cooperate [1766]. cooperation [346, 1039]. Cooperative [348, 712, 1254, 1330, 1511, 1582, 1811, 1940]. coordinate [847]. Coordinates [1054, 1059]. Coordinating [1533]. COPEARMS [1269]. Coping [311]. Copyrighted [1715]. CORBA [983, 1549, 1939]. CorbaWeb [875]. CoReview [713]. Cornerstone [1755]. Corporate [865, 867, 1512, 1531, 1532, 1540, 1543, 1697]. correct [1016]. correctness [940]. Correlated [1091, 1875]. correlation [614, 979]. Correlations [1899]. COSINE [181, 187, 196– 199, 256, 352, 353, 447, 456, 462, 471, 548]. COSSC [1472]. COST [32, 157–159, 624, 828, 1062]. COST-11 [157–159, 624]. Cost-quality [828]. costs [1246]. Counter [531, 1394]. Countering [1652]. Counting [1295, 1933]. County [1670]. Couplers [1983]. Course [871, 1305, 1622]. courses [871, 1208]. Courseware [717, 1122, 1605, 1617, 1623]. court [1148]. courts [1148]. coverage [935]. covered [1577]. CoWeb [885]. CoWeb-architecture [885]. crawlers [1106]. crawling [1110]. CRC [391]. created [1330]. Creating [113, 718, 778, 1106, 1144, 1481, 1699]. Creation [1536]. Credit [388, 1859, 2020]. Credits [982]. Critical [1459, 1626]. CRMA [595, 969, 1079, 1957]. CRMA-II [595, 969]. Cross [777, 979, 1077, 1243, 1722]. cross-correlation [979]. cross-language [1243]. cross-references [777]. crowd [1308]. CRS4 [478]. cruise [453]. CRWR [1220]. Crypto [1071]. cryptographic [543]. CrystalWeb [714]. CS [1616]. CSCW [348]. Csma [9, 22, 53, 54, 82, 114, 221, 263, 565, 1407, 1422, 1956, 1967, 2026]. CSMA-CD [11]. CSMA/CD [4, 9, 53, 54, 82, 114, 221, 263, 565, 1407, 1422, 1956, 1967, 2026]. CSMA/CD-based [221, 263]. cube [1063]. cube-based [1063]. Cultural [1157, 1457, 1461, 1463]. Current [158, 181, 301, 1612]. Curriculum [1610]. curve [913]. Cusp [82]. Custom [1678]. Customer [424, 567, 629, 966, 1347, 1348]. customer-premise [424]. Customers [1549]. Customizable [1015]. customization [507]. Customize [1668]. cut [539, 541]. cut-through [539, 541]. cutting [1056]. Cyber [1148, 1293]. Cybermouse [1680]. CyberProf [1776]. Cyberspace [781, 1452, 1457, 1465]. Cyberview [1753]. cyborgs [1156]. Cycle [590, 1397, 1834]. Cycles [1933]. Cyclic [1410]. Cypress [149]. Czech [449, 2044].

D [783, 1182, 1494]. Daemon [1580]. daemons [856]. Daisy [1232]. Darmstadt [2047]. dash [1184]. Data [11, 25, 34, 60, 76, 83, 99, 119, 123, 142, 163, 164, 183, 214, 232, 263, 282, 309, 311, 319, 335, 361, 373, 418, 428, 433, 451, 472, 474, 478, 486, 532, 585, 629, 699, 708, 723, 739, 759, 767, 787, 790, 896, 918, 977, 978, 1026, 1303, 1342, 1347, 1390, 1450, 1469, 1474, 1475, 1487, 1489, 1526, 1532, 1555, 1561, 1565, 1596, 1598, 1600, 1601, 1641, 1646, 1736, 1737, 1814, 1818, 1850, 1872, 1874, 1915, 1988, 2028]. Database [100, 823, 850, 865, 1022, 1149, 1465, 1499, 1639, 1640, 1648, 1739]. Databases [176, 666, 728, 864, 1638, 1718, 1774]. Datagram [1434]. DB [1774]. DBP [1965]. DBP-M [1965]. DBR [1974]. DCE [704, 1735, 1939]. DCP [587]. DDT [687]. Deadline [1362]. Deadlines [1851]. deadlock [916]. dealing [136]. Decade [1420]. Deceit [660]. Decentering [1459]. Decentralized [87, 1230]. Decimal [1173]. DeckScape [737]. declarative [1291, 1294]. decoder [140]. Decoration [1781]. Decrease [228]. default [571]. Definition [306, 413, 678, 723, 909, 1829, 1869]. Deflection [396, 606, 755, 1876]. Deformable [1057]. deformation [1061]. Delaware [1458]. Delay [152, 320, 410, 524, 541, 614, 633, 836, 971, 995, 1229, 1355, 1893, 1948, 1998, 2024]. delay-conscious [995]. Delay/Loss [2024]. Delays [85, 486, 522, 627, 993, 1967]. delegation [1073]. Deliver [1474]. Delivering [1123, 1257]. Delivery [72, 83, 164, 507, 551, 674, 920, 1098, 1136, 1342, 1439, 1541, 1676, 1712, 1962, 2030]. DELPHI [1555]. Demand [60, 313, 818, 1081, 1136, 1256, 1439, 1857, 1880, 1912]. Demand-Priority [1857]. Demise [1698]. demonstration [1188]. Department [1305, 1469]. Dependent [646, 947, 1082, 2000]. Deploying [856, 1818]. deployment [864, 1314]. Derivation [322, 522]. Deriving [496, 916, 1886]. describing [726, 1096]. Description [126, 142, 220, 275, 474, 490, 946, 1221, 1252]. Descriptive [369]. Design [46, 61, 74, 118, 171, 179, 227, 234, 273, 285, 398, 427, 505, 543, 645, 688, 692, 693, 748, 752, 786, 796, 808, 809, 862, 886, 914, 938, 942, 948, 953, 956, 965, 1011, 1051, 1080, 1103, 1127, 1192, 1213, 1240, 1256, 1285, 1289, 1302, 1317, 1340, 1346, 1364, 1371, 1374, 1379, 1399, 1476, 1566, 1603, 1637, 1650, 1679, 1789, 1798, 1804, 1824, 1830, 1831, 1840, 1878, 1889, 1910, 1918, 1939, 1950, 1955, 1963, 1979, 1985, 2012, 2022, 2027]. Designing [292, 731, 1040, 1608, 1652, 1887, 1924]. DESIRE [1111, 1258, 1266]. Desktop [619, 1275, 1437, 1659]. desperate [1197]. Desperately [571]. Destination [262, 384, 1289, 1911]. Detail [480]. detailed [1216]. DeTeBerkom [812, 922]. detecting [1253]. Detection [54, 1986]. Determine [2016]. Deutsche [868]. develop [958]. developers [1213]. Developing [264, 415, 507, 1208, 1265, 1513, 1544, 1798]. Development [62, 77, 145, 171, 304, 416, 470, 490, 707, 999, 1067, 1122, 1168, 1240, 1243, 1263, 1371, 1402, 1544, 1623, 1626, 1670, 1706]. Developments [16, 301, 1573]. Device [1002]. Device-independent [1002]. Devices [1186, 1831]. Dewey [1173]. DFN [927]. Diagnosis [26]. diagnostic [683].

Diagnostics [1594]. Dial [622, 1992]. Dial-in [1992]. dial-up [622]. DIALECT [919]. Different [39, 275, 838, 1356, 1428]. difficult [1228]. Diffusion [1649]. Digestor [1002]. digests [1322]. Digital [17, 66, 76, 146, 174, 285, 500, 563, 575, 668, 919, 1269, 1597, 1616, 1725]. Dikon [342]. dimension [1659]. dimensional [618, 1307]. Dimensioning [510, 536, 825, 886]. Dimensions [1642]. DINGO [496]. dinner [330]. Directory [172, 187, 208, 249, 298, 361, 477, 546, 555, 556, 626, 630, 693, 699, 816, 817, 833, 1262]. Disabilities [1206, 1637, 1655]. disabled [1039]. Discard [1925, 1966]. Discarding [829, 1083, 1964]. discipline [1089]. disclosure [1243]. Discover [724]. Discovering [1651, 1724]. Discovery [724, 779, 1124, 1189, 1238, 1530, 1543, 1606, 1625, 1689, 1721, 1769]. Discrete [535, 976]. Discretionary [23, 637]. Discretized [319]. discussing [1577]. Discussion [620, 1507]. Disjointly [41]. Disk [1380]. Dispatcher [1128]. Dispenser [1532]. Dispersity [485]. Display [621, 1677, 1701]. displets [1006]. Dissection [783, 1494]. Disseminating [1646]. Distance [870, 1350, 1619, 1620, 1628, 1829, 1919]. Distillation [890, 1450, 1689]. distortion [979]. Distributed [2, 39, 48, 73, 74, 118, 120, 148, 151, 157, 165, 170, 223, 237, 276, 280, 291, 298, 312, 326, 367, 376, 420, 473, 496, 526, 544, 587, 593, 623, 657, 676, 684, 691, 693, 707, 714, 742, 744– 749, 847, 848, 850, 855, 865, 959, 983, 992, 1053, 1075, 1099, 1111, 1114, 1134, 1165, 1229, 1257, 1267, 1273, 1320, 1329, 1331, 1349, 1366, 1381, 1408, 1489, 1502, 1690, 1726, 1735, 1764, 1810, 1814, 1834, 1835, 1851, 1854, 1940, 1943, 1976, 2001, 2007]. distributed-control [587]. Distribution [87, 484, 543, 684, 706, 791, 833, 1377, 1448, 1469, 1475, 1595, 1647, 1898, 1951]. distributions [517]. diverse [1574]. Division [177, 210, 274, 484, 604, 608].

DMTS [1854]. DNS [684]. Do [251, 1102]. Document [203, 301, 551, 573, 651, 674, 813, 1132, 1158, 1196, 1321, 1572]. Documentation [308, 703, 1591, 1726]. Documented [1638]. Documents [658, 730, 771, 782, 843, 849, 858, 995, 1118, 1126, 1141, 1144, 1151, 1243, 1483, 1485, 1558, 1560, 1568, 1657, 1719, 1722, 1784]. does [144]. Domain [315, 994, 1013, 1237, 1259]. DQDB [394, 395, 408, 513, 514, 520, 525, 585, 586, 636, 697, 740, 1229, 1353–1356, 1409, 1411, 1838]. DQDB/B [1411]. DQDB/B-ISDN [1411]. DQFS [2006]. Dragons [1662]. Drawings [1576]. Driven [983, 1041, 1382, 2023]. driver [741]. Drop [793, 1710, 1988]. Drop-in [793, 1710]. DSAs [817]. DTM [392]. Dual [148, 384, 580, 762, 943, 1077, 1081, 1246, 1357, 1381]. dual-bus [580]. Dual-hop [1077]. Dual-state [1246]. Dumb [250]. Duplex [299, 1425, 2026]. during [1310]. Duty [1665]. Dyna [1803]. Dynamic [56, 74, 82, 167, 221, 263, 392, 393, 425, 472, 486, 511, 532, 564, 837, 852, 951, 972, 1013, 1020, 1094, 1181, 1296, 1351, 1384, 1409, 1560, 1567, 1571, 1572, 1677, 1825, 1981, 1988]. Dynamically [1844]. Dynamics [561, 980, 1575]. DynaWeb [1701].

E-DCP [587]. e-mail [477]. Each [1926]. Early [1455, 1966]. EARN [464, 468, 558]. EARN/INTERNET [464]. Earth [790, 1065, 1596, 1610]. ease [1167]. Easing [143]. Easy [1477]. Economic [81, 931, 1247, 1520]. economics [893]. ecosystems [1299]. EDGAR [1646]. EDI [699, 1268]. EDIL [674]. edit [712]. editing [1141]. Editor [306, 481, 877, 1042, 1476, 1478]. Editorial [307, 325, 329, 377, 412, 432, 489, 670, 961]. editors [503]. EDT [416]. Education [103, 175, 191, 337, 433, 461, 664, 718, 878, 920, 1122, 1152, 1275, 1283, 1290, 1313, 1606, 1614, 1616, 1619, 1620, 1674, 1675, 1777].

Educational [715, 717, 719, 821, 1189, 1193, 1775]. edutainment [878]. EEPG [343]. Effect [85, 539, 614, 1233, 1422, 1424, 1994]. Effective [510, 1100, 1124, 1450]. Effectiveness [1142, 1344, 1962]. Effects [838, 1140, 1445, 1857, 1874, 1967]. Efficiency [1142, 1960]. Efficient [111, 370, 884, 915, 944, 1056, 1110, 1137, 1236, 1303, 1372, 1823, 1876, 1948, 1949, 1970]. Effort [1083, 1988]. Efforts [279]. EFR [1993]. EFTs [150]. Election [42]. electro [512]. electro-optic [512]. Electronic [1, 122, 138, 436, 464, 466, 679, 699, 726, 838, 867, 897, 1163, 1242, 1245, 1260, 1302, 1455, 1469, 1486, 1509, 1604, 1681, 1713, 1714, 1784, 1800, 1807, 1991, 2019, 2020]. Elements [799, 915, 1830]. ELFI [1259]. Embedded [1001, 1413]. Embedding [1372, 1553, 1827]. EMBnet [454]. Emergency [362]. Emerging [1276, 1425]. Empirical [1448]. Employing [177]. Empowering [1624]. emulator [520]. enabled [983]. Enabling [1033]. Encapsulation [1915]. encoder [140]. encoder-decoder [140]. Encoding [951, 1786]. End [522, 553, 675, 967, 1270, 1281, 1433, 1463, 1633, 1645, 1664, 1965, 1969, 2024]. End-Stations [1433]. End-to-End [522, 967, 1281, 1965, 1969, 2024]. enforcement [430]. Engine [729, 1012, 1015, 1104, 1105, 1116, 1120, 1149, 1493, 1773]. Engineering [296, 343, 428, 497, 886, 1034, 1244, 1275, 1298, 1502, 1513, 1536, 1576, 1588, 1603]. engineers [436]. engines [1016, 1131]. Enhanced [480, 707, 730, 760, 861, 921, 1901, 1968]. Enhancements [1731]. Enhancing [704, 1206]. Enterprise [131, 1365, 1446, 1815, 2019]. Enticing [892]. entity [304]. Entrez [1492]. entry [573]. Environment [64, 65, 91, 113, 179, 213, 247, 264, 294, 339, 350, 371, 372, 415, 417, 473, 570, 637, 701, 711, 714, 719, 778, 865, 871, 895, 920, 929, 983, 1039, 1053, 1062, 1067, 1076, 1092, 1192, 1207, 1254, 1367, 1405, 1418, 1480, 1481, 1510, 1527, 1535, 1561, 1601, 1693, 1775, 1777, 1778, 1813, 1845, 1910, 2009, 2029]. Environmental [1724]. Environments [310, 365, 642, 1058, 1064, 1068, 1099, 1153, 1195, 1279, 1381, 1400, 1752]. EOSDIS [790, 1596]. equal [1330]. Equalizer [1392]. EQuB [2021]. equilibrium [317, 517]. equipment [424, 966]. Erasure [586, 1356]. Ergodic [319]. ERICA [1999]. ERMS [1269]. Error [834, 842, 1352, 1389, 1897, 1911, 1973, 1986]. Error-Control [1897]. errors [842]. Esprit [442, 469]. Establishing [67]. Estelle [151, 305, 379, 416, 417, 491, 496]. Esterel [304]. ESTIM [417]. estimates [993]. Estimation [1948]. ET-LOTOS [946]. Ethernet’ [25, 631, 950, 1424, 2021, 2026]. Ethernets [764, 1861, 1920, 1976]. Ethersim [1086]. Ethical [780, 1449]. Ethnography [1658]. ETINU [1067]. EUnet [194]. Euro [460]. Euro-networking [460]. EUROMATH [100, 206]. EuropaNET [479]. Europe [43, 98, 338, 463, 926, 1071]. European [95, 100, 103, 182, 183, 190, 200, 343, 350, 433, 443, 445, 454, 469, 475, 502, 550, 819, 2044]. EUV [1473]. Evaluating [520, 542, 1333, 1919]. Evaluation [68, 90, 154, 230, 272, 494, 576, 752, 799, 1039, 1080, 1082, 1088, 1281, 1285, 1388, 1401, 1413, 1826, 1871, 1872, 1916, 1918, 1955, 1985]. event [976, 998, 1099]. event-based [998, 1099]. Evolution [231, 232, 583, 674, 1817, 1879]. Evolutionary [174]. Exact [322, 517]. Example [630, 678, 693]. Examples [378, 1302, 1565]. exams [1191]. Excavating [1503]. Exchange [99, 552, 605, 661, 921]. executable [385]. Execution [93]. Exercise [204]. exhaustive [229]. Exhibit [1684]. Exhibition [1686]. Exhibits [1456]. Existing [1066, 1321, 1488]. Exorcising [856]. Exp1 [1321]. Expanding [339]. expansion [617, 1164]. expectations [508]. Experience [39, 483, 694, 1349, 1689, 1835]. Experiences [305, 345, 508, 626, 741, 812, 922, 923, 1140, 1266, 1402, 1694, 1717, 1969]. Experiment [305, 781, 785, 994, 1452, 1528, 1555, 1609]. Experimental [179, 478, 737, 1826, 1880, 1989]. Experimentation [1358, 1817]. Experiments [1341]. Expert [1651, 1718, 1843]. Explicit [121, 1954]. exploitation [727]. Exploiting [1405]. Exploration [19, 1018]. exploratory [1192]. Explore [1599]. Explorer [1618]. exponentially [742]. Exposed [1434]. EXPRESS [215]. Extended [113, 1949]. Extending [523, 697, 773, 774, 860, 1006, 1505, 1552, 1588, 1809]. Extensibility [1560, 1589]. Extensible [1116]. Extension [92, 587, 1550, 1566, 1969]. Extensions [431, 1224]. External [816, 1553]. Extraction [1600]. EzMath [1185].

Fabric [1295, 1337, 1878, 1882]. Facilitate [1601]. Facilitating [1109]. Facilities [164]. facsimile [122]. Factors [159]. Faculty [1668]. Fail [276, 1366]. Fail-Safe [276, 1366]. Fair [19, 590, 1229, 1715]. Fairness [395, 537, 562, 969, 1375, 2024]. FAITH [641]. fall [1705]. familiar [207]. family [271]. FAPL [116]. FAQ [781, 1452]. FAQs [794, 1523]. FARA [11]. FARA/CSMA [11]. FARA/CSMA-CD [11]. Fasnet [271]. Fasnet-based [271]. Fast [88, 484, 1049, 1082, 1139, 1661, 1878, 1957]. FastLane [1538]. Fault [26, 170, 645, 935, 1415, 1820, 1843, 1906, 1949]. Fault-Tolerance [1415]. Fault-Tolerant [645, 1820, 1906, 1949]. FCP [1871]. FDDI [55, 341, 584, 945, 1357, 1359– 1364, 1387, 1397, 1404– 1407, 1833, 1858, 1860, 1870, 1874, 1934]. FDDI-2 [1361]. FDDI-Based [1874].

FDDI-II [1363, 1364, 1833]. FDQ [1229]. feasible [647, 656]. Feather [1495]. feature [1253, 1289]. Features [1009, 1252, 1405, 1742, 1757]. FEC [1962, 1964]. Federal [77, 449, 1644]. Feedback [121, 800, 1164, 2014]. FERRET [1600]. Ferry [155, 233]. few [506]. fewer [1338]. Fiber [418, 580, 588, 608, 1219, 1369, 1373, 1391, 1983, 2025]. fiber-coax [1219]. fiber-optic [580, 608]. Fibre [109, 210, 274, 1368, 1370, 1371, 1402, 1871, 1872]. FieldBuses [1405, 1990]. FIFO [767, 1399]. Fifth [2050]. Fighter [1544]. figures [740]. File [13, 64, 219, 247, 499, 576, 1380, 1445, 1461, 1584]. Files [1576, 1977]. Filling [885, 1307]. Filtered [1792]. filtering [1205, 1326]. Final [624]. Financial [1268, 1535, 1646]. Finding [517, 757, 1009, 1525, 1717]. Finite [303, 386, 578, 613, 829, 915, 935, 1235, 1445]. Firm [1965]. First [662, 708, 1094]. First-order [1094]. Fish [1017, 1728]. Fish-Search [1728]. fittest [434]. Five [1469]. Fixed [1926]. Fixing [874]. fledged [523]. Flexible [286, 420, 685, 695, 732, 808, 996, 1383, 1968]. Flooding [47]. Flow [319, 427, 532, 597, 633, 739, 842, 1087, 1090, 1094, 1287, 1400, 1434, 1859, 1894, 1997]. Fluid [1575]. FMMRA [1087]. focus [337]. focused [824]. focusing [1186]. Folded [1836]. Fonts [1586]. Footprint [818]. Footsteps [716]. force [135]. Form [1061, 1567]. Formal [142, 275, 303, 413, 474, 490, 493, 495, 545, 750, 795, 917, 932, 933, 976, 1221, 1222, 1231, 1284, 1869]. Formalization [905, 952]. format [116, 651, 1096]. Formation [1469]. Formatting [1487]. forming [217]. Forms [141, 324, 882, 885, 921, 1044, 1644, 1718]. Forte [973]. Forum [268, 1649]. forwarding [1339]. Foundation [1538, 1756]. Foundations [1593]. Four [559, 1662]. Fourth [2048]. fractal [1295]. Frame [375, 423, 439, 440, 639, 1396, 1956]. FrameMaker [658]. Frames [1915]. Framework [320, 387, 414, 492, 534, 647, 815, 864, 866, 909, 966, 988, 998, 999, 1003, 1106, 1165, 1260, 1263, 1567, 1588, 1798, 1841, 1914]. France [2050]. Free [169, 215, 299, 802, 916, 943, 1061, 1661]. Freedom [311]. freeWAIS [730]. freeWAIS-sf [730]. freshness [1100]. Frog [783, 1494]. Front [1633]. FTAM [441, 465]. FTP [1327]. fulfilling [1258]. Full [299, 523, 593, 1108, 1425, 1737, 2026]. Full-Duplex [299, 1425, 2026]. full-fledged [523]. Full-Text [1108, 1737]. fully [426]. fun [1174]. Function [404, 1885]. Functional [219]. functionality [730, 885]. Functions [484, 838, 1400]. Fund [1647]. Funding [678, 1259]. Further [538, 837]. Future [143, 283, 301, 1169, 1216, 1466, 1647, 1686, 1694]. Fuzzy [214, 1824, 1877, 1942].

G [306, 610, 721, 821]. G-LOTOS [306, 610]. G-MING [821]. Gain [2031]. Gala [330]. Gala-dinner [330]. Galleries [1456]. Gallery [1464, 1606]. games [878]. gap [690]. Gardens [1458]. Gateway [50, 464, 684, 1189, 1387, 1411, 1499, 1581, 1590, 1835, 1879]. Gateways [189, 270, 830, 989, 1022, 1258, 1555, 1780]. Gb [540, 595, 695]. Gb/s [540, 595, 695]. Gbit/s [512]. GEM [1189]. General [20, 541, 641, 1217, 1567, 1914]. generalization [219]. Generalized [757, 1928]. Generating [1177, 1737]. Generation [173, 407, 502, 694, 859, 934, 936, 944, 1035, 1063, 1102, 1231, 1308, 1572, 1946]. generative [743]. Generator [391, 1772]. Generic [681, 864, 875, 1085, 1163, 1971]. Geneva [346, 2045]. geographic [891, 1192]. geographical [682, 1157]. geometric [1298]. geometry [1054]. Geomview [1753]. geospatial [1025]. GeoWeb [1489]. German [1159, 1274]. Germany [77, 1661, 2047]. gestation [549]. get [339].

GFR [2032]. Gigabit

[356, 392, 482, 1236, 1375, 1380, 1874, 2026]. Gigabit/s [1375]. GIX [552]. Global [99, 133, 248, 339, 355, 476, 552, 570, 627, 833, 1065, 1244, 1323, 1528, 1582, 1685, 1779, 1799]. go [842]. go-back- [842]. good [678]. Gopher [1531]. Government [1159, 1213, 1274, 1545, 1671, 1688, 1689, 2019]. Gracious [1925]. Grade [73, 160, 1069]. Grade-Of-Service [160]. Graffiti [1463]. Graph [861, 1096, 1218, 1928]. Graphic [91, 1795]. GRaphical [306, 441, 610, 776, 1144, 1723]. Graphically [1591]. Graphically-Based [1591]. Graphics [783, 1045, 1051, 1060, 1290, 1298, 1305, 1311, 1313, 1494, 1781]. Grass [1512]. Grassroots [866]. Greater [821]. Green [190]. Grif [860]. GRiNS [1144]. Ground [1465]. Group [159, 314, 343, 347, 501, 720, 988, 1376, 1378, 1423, 1469, 1762]. Group-based [1378]. Groups [156]. groupware [987]. Growing [679, 1661]. Grows [1645]. Growth [1319, 1662]. GSM [1995]. GSN [1927]. GSQL [1499]. Guarantee [593, 1965, 1978]. Guaranteed [1999]. Guaranteeing [1362, 1851]. Guarantees [1089, 2032]. Guessing [1960]. Guest [307, 325, 329, 412, 432, 481, 489, 503, 670]. GUI [1794]. guided [872, 1107]. Guidelines [77, 379, 1365].

hacker [1197]. handling [1251]. Handoff [1945]. Happening [1758]. Hard [86, 2029]. hardware [698, 972]. Harvest [779, 888, 1721]. having [710]. HDF [790, 1596]. HDLC [304]. Head [861, 1793]. headed [1521]. healing [1074]. Health [1270, 1676]. Helix [914]. helminthiasis [327]. help [671, 1210]. helps [1018]. Heritage [1457]. Heterogeneous [280, 317, 466, 691, 829, 1161, 1281, 1384, 1902, 1997]. Heuristic [1905]. Hewlett [346]. Hidden [931, 1994]. Hierarchical [15, 167, 176, 429, 511, 555, 743, 807, 872, 881, 1115, 1303, 1346, 1366, 1912, 2030]. hierarchy [575, 1319]. High [55, 200, 210, 271, 274, 277, 292, 309, 310, 337, 338, 354, 355, 358, 397, 401, 427, 438, 442, 472, 479, 485, 486, 488, 506, 521, 532, 542, 549, 568, 580, 582, 588, 590, 596, 598, 616, 633, 642, 698, 759, 762, 797, 808, 819, 821, 836, 927, 943, 968, 974, 979, 989, 1078, 1094, 1133, 1201, 1217, 1277, 1282, 1296, 1351, 1352, 1373, 1380, 1381, 1390, 1396, 1398–1400, 1403, 1414, 1418, 1427, 1447, 1672, 1697, 1819, 1829, 1830, 1834, 1844, 1853, 1856, 1867, 1873, 1875, 1876, 1888, 1919, 1933, 1957, 1967, 2001, 2010, 2027]. High-level [277]. High-Performance [292, 762, 974, 1381, 1447]. High-Speed [200, 210, 271, 274, 309, 310, 358, 397, 442, 472, 485, 488, 506, 532, 568, 580, 582, 588, 598, 633, 642, 759, 797, 819, 927, 943, 968, 979, 1078, 1094, 1217, 1277, 1351, 1352, 1373, 1398– 1400, 1418, 1819, 1834, 1844, 1856, 1873, 1875, 1876, 1888, 1919, 2001, 2027]. Higher [201, 1054]. Highlights [1683]. Highly [596, 1032, 1875]. Highway [1599]. HIPPI [1401, 1919]. HIPPI-ATM [1919]. histograms [1309]. historical [1024]. histories [1321]. History [1604, 1654, 1795]. History-of-Use [1654]. Hits [1027, 1330]. Hoc [659, 1996]. Holding [1652]. Home [788, 1124, 1522, 1910, 1951]. homepage [1013]. Homing [1357]. homogeneous [1059]. Hong [1306]. Hop [223, 1077, 1926, 1944, 1983]. Hopwise [706]. hospital [929]. Host [165, 559, 686, 1388, 1732, 1859, 1923]. hosting [997]. hosts [560]. hot [835]. HotCon [810]. Hotel [2033–2036, 2038, 2042]. Hotline [810]. HotMetal [1706]. house [1544]. HP [1203]. HPCC [337]. HTML [651, 768, 782, 858, 859, 882, 885, 1006, 1031, 1116, 1117, 1203, 1476–1479, 1482, 1484, 1487, 1488, 1496, 1547, 1550, 1553, 1657, 1701– 1704, 1719, 1743, 1744, 1808].

HTML/SGML [1704]. HTTP

[653, 770, 848, 1138, 1191, 1509, 1562, 1578, 1579, 1731, 1796, 1802]. HTTP-Based [1796]. HTTPD [1544, 1748]. hub [1250, 1426]. Human [159, 892, 1058, 1156, 1601, 1637, 1655, 1756, 1776]. Human-Computer [1058, 1776]. human-machine [1156]. Hungarian [254]. Hungary [450]. Hurst [2016]. Hybrid [4, 10, 11, 60, 218, 518, 580, 836, 1219, 1842, 1917]. Hybrinet [10]. Hyper [721, 742, 1016, 1182, 1528]. hyper-exponentially [742]. Hyper-G [721]. Hyper-Media [1528]. Hyperactive [1491]. Hypercard [1679]. HyperCode [1548]. Hypercube [1372]. hyperdocument [1178]. hypergraph [1218]. HyperKids [1631]. hyperlink [850, 1101]. hyperlinked [1118]. hyperlinks [1007]. Hypermedia [703, 775, 898, 984, 1121, 1153, 1628, 1631, 1690, 1746, 1751, 1805]. hypermedia-based [703]. Hypertext [657, 659, 665, 776, 852, 1194, 1211, 1455, 1480, 1492, 1628, 1677, 1679, 1723, 1726, 1800]. hypertextual [1105].

I/O [1871]. IAFA [1768]. IBM

[464, 540, 1757]. ICCC [44]. Identification [165, 1720]. Identity [1466]. IDLE [722]. IEEE [40, 69, 113, 166, 279, 408, 525, 1967, 1994]. IFIP [79]. II [595, 969, 1042, 1363, 1364, 1443, 1833]. illustrated [955]. Image [164, 863, 1060, 1133, 1295, 1308, 1309, 1312, 1512, 1550, 1783, 1977]. images [929]. Imaging [1466, 1486]. IMHP [686]. IMJ [1136]. immersive [1058]. Impact [253, 771, 809, 835, 1213, 1441, 1568, 1672, 1849, 1862]. Impacts [132, 419, 1247, 1899]. Impairments [17]. implement [495]. Implementation [13, 56, 62, 74, 137, 138, 147, 196– 199, 256, 303, 305, 414, 455, 500, 630, 637, 698, 753, 934, 948, 1212, 1239, 1256, 1317, 1367, 1379, 1470, 1509, 1566, 1573, 1657, 1663, 1790, 1839, 1887, 1889, 1939, 1952, 2027]. Implementations [45, 496, 1272]. Implemented [1681]. Implementing [414, 548, 1432, 1534, 1656]. Implications [108, 823, 1439]. implicit [121]. Implosion [1423]. Implosion-Avoidance [1423]. Imprecise [1977]. Improve [1325, 1369, 1900]. Improved [418, 631, 642, 1359]. improvement [513]. Improvements [710, 1651]. Improving [770, 1022, 1328, 1559, 1616, 1734, 1937, 2014]. Impulse [1652]. In-house [1544]. In-place [1141]. INCA [2010]. including [2016]. incomputer [121]. incomputer-mediated [121]. Incorporating [395, 1564]. Increase [228]. Increase/Decrease [228]. Incremental [304, 544, 1018]. independence [842, 1304]. independent [685, 696, 842, 1002, 1162]. Index [782, 817, 1007, 1574, 1719]. Index-based [1007]. indexed [817]. indexing [655, 1108, 1111, 1267, 1574]. Individualised [1622]. Individualizing [1454]. Industrial [109, 131, 132, 460, 690, 1240]. Industry [1529, 1691]. Inexpensive [1807]. infancy [331]. Influence [581, 740, 1157]. Influences [1355]. INFN [241]. INFNet [241, 260, 684]. InfoHarness [1716]. informal [811]. Information [1, 29, 72, 118, 136, 223, 353, 395, 457, 573, 656, 661, 665, 672, 703, 721, 722, 725, 755, 779, 818, 854, 857, 866, 867, 884, 898, 913, 1011, 1012, 1014, 1016, 1025, 1055, 1100, 1129, 1139, 1153, 1164, 1165, 1176, 1189, 1209, 1240–1242, 1257, 1259, 1266, 1274, 1396, 1467, 1470, 1472, 1514– 1516, 1529, 1530, 1534, 1540, 1541, 1549, 1597, 1599, 1601, 1609, 1614, 1624, 1629, 1632, 1635, 1643, 1644, 1653, 1660, 1667, 1678, 1682, 1688, 1689, 1693, 1700, 1707, 1716, 1718, 1721, 1724, 1728, 1734, 1755, 1763, 1797, 1799, 1801, 1915]. information [785, 814, 818, 1098]. Information/Control [1396]. InfoSleuth [1727]. infostructures [657]. Infrastructure [507, 821, 913, 922, 987, 1320, 1543, 1796, 1801, 1961]. infrastructures [1066]. Ingrid [1801]. inhabiting [1064]. Initiated [2004]. Initiative [216, 337, 387, 1786]. initiatives [348]. Inline [1576]. Innovation [1663]. Innovative [1623]. Input [614, 633, 640, 762, 1120, 1233, 1335]. input-output-buffering [614]. input-queued [1335]. Inquirus [1104]. Inquiry [1651]. INS [231]. Insertion [299, 406, 407, 590, 1391]. inspection [1295]. inspire [1067]. Instantaneous [71]. Institutional [1697]. instruction [1179]. Instructional [1068, 1679]. Instructions [1553]. Integrated [37, 100, 110, 172, 174, 213, 214, 263, 287, 365, 376, 393, 411, 415, 417, 424, 427, 455, 500, 511, 512, 550, 568, 587, 759, 763, 810, 839, 966, 979, 1092, 1385, 1394, 1852, 1866, 1941, 2005, 2006, 2010, 2023]. Integrating [666, 707, 738, 831, 867, 873, 1164, 1526, 1561, 1594, 1605, 1803]. Integration [11, 25, 90, 350, 373, 418, 574, 1044, 1353, 1531, 1716, 2019]. Integrity [83, 238, 1214]. intellectual [731, 1154, 1269]. Intelligence [288, 768, 1484]. Intelligent [285, 383, 577, 609, 701, 831, 1174, 1192, 1244, 1252, 1682, 1726, 1727, 1776]. IntelliTeaching [1672]. intensity [1307]. intensive [419]. Inter [14, 226, 315, 459, 691, 979, 1237]. Inter-Administrative [315]. inter-bridge-security [459]. inter-domain [1237]. inter-media [691]. Inter-network [226]. Inter-Organizational [14]. inter-source [979]. Interaction [125, 413, 487, 747, 1042, 1058, 1251, 1253, 1756]. Interactions [1252, 1601]. Interactive [309, 664, 665, 718, 722, 783, 791, 878, 900, 918, 919, 985, 1038, 1044, 1136, 1179, 1209, 1283, 1459, 1461, 1494, 1509, 1554, 1575, 1595, 1606, 1628, 1630, 1649, 1673, 1684, 1685, 1739, 1770]. Interactively [858]. Interactivity [1567].

Interchange [699, 813, 1463]. Interconnected [3, 321, 342, 528, 1437]. Interconnecting [172, 1990]. Interconnection [40, 107, 179, 397, 472, 488, 566, 915, 1401, 1406, 1426, 1833, 1884, 1918, 1935, 1968]. Interdisciplinary [159]. Interest [216]. Interface [178, 267, 439, 729, 786, 1030, 1039, 1144, 1178, 1182, 1399, 1558, 1581, 1638, 1650, 1655, 1718, 1776, 1780, 1859, 1963]. Interface-Parasite [1780]. Interfaces [6, 212, 816, 900, 1109, 1142, 1177, 1637, 1653]. Interfacing [1380]. Interfering [65]. Interim [156]. interleaved [689]. interleaving [484]. Intermediaries [1143]. intermediation [869]. Internal [464, 786, 840, 1337, 1544, 1650]. International [17, 28, 78, 80, 286, 335, 376, 382, 644, 1611, 1989, 2042, 2047, 2048, 2050, 2051, 2053]. internationalisation [681]. Internet [255, 327, 334, 339, 349, 364, 453, 464, 552, 571, 677, 678, 682, 684, 686, 828, 920, 921, 1000, 1017, 1039, 1065, 1066, 1069, 1072, 1128, 1148, 1188, 1206, 1234, 1255, 1268, 1271, 1276, 1320, 1528, 1543, 1645, 1646, 1662, 1666, 1668, 1687, 1693, 1695, 1699, 1768, 1816, 1837, 1953, 2009, 2044]. Internet-based [1039, 1268]. Internet2 [1320]. Internetting [978]. Internetwork [50, 328, 342]. Internetworked [1941]. Internetworking [185, 260, 923, 1419, 1828, 1866, 1938]. internetworks [841, 1695, 1964]. Interoperability [259, 387, 435, 465]. interoperable [268]. interoperation [363]. interpretation [638]. Interruptions [322]. Intertest [387]. Interval [1892]. Interval-Based [1892]. Interworking [45, 232, 340, 372, 1921]. Intra [1881]. Intra-Media [1881]. intranet [1140, 1183]. Intranets [2023]. Introduce [1345]. Introducing [508, 845, 1664, 1791]. Introduction [115, 151, 286, 302, 378, 380, 546, 612, 745, 902, 903, 946, 961, 1741].

Introductory [1615]. Intrusive [1734]. invalidation [1112]. Investigation [109, 129, 849]. investigations [806]. Investment [1530]. IP [261, 342, 364, 559, 560, 683, 806, 923, 924, 997, 1087, 1417, 1418, 1845, 1847, 1921, 1944, 1975]. IP-OSI [365]. IP-X.25 [363]. IP/IPng [923]. IPng [923]. IPv6 [1272]. IPX [1847]. IR [1392, 1737]. IR-Wireless [1392]. Irregular [1372]. ISDN [16, 30– 32, 35, 36, 43, 108, 144, 153, 178, 219, 239, 266– 268, 383, 419, 443, 500, 509, 581, 611, 641, 643, 644, 649, 796, 832, 912, 1411, 1887, 1913, 1942]. ISDN-Concept [108]. ISDN-PABX [35]. ISDNs [160, 798]. ISI [451, 1537]. ISO [13, 45, 80, 115, 207, 350, 445, 744, 745, 1063, 1784]. iso-surface [1063]. ISO/OSI [45]. Isochronous [393, 760, 1081, 1339, 1343, 1363]. ISODE [258]. issue [44, 969]. Issues [74, 94, 184, 294, 295, 530, 569, 680, 817, 822, 824, 931, 1071, 1157, 1249, 1254, 1285, 1374, 1415, 1466, 1496, 1498, 1534, 1577, 1626, 1655, 1715]. Iterative [1335, 1827]. IX [302]. IXI [257].

Jamming [986]. JANET [195, 364]. Japan [20, 376, 994, 1660]. jargon [354]. Jasper [1797]. JAVA [865, 983, 988, 992, 1048, 1107, 1120, 1150, 1154, 1173, 1290, 1788]. Java-based [988, 1107, 1154]. JCAT [1048]. JESSICA [1121]. Joint [195, 2044]. Jordan [1056]. Journal [1508, 1681, 1713, 1714, 2048]. Jukebox [1136]. June [2044]. Just [1781]. JW [1636].

Kerberos [277]. kernel [623]. Kernels [165]. Key [543, 833, 922, 1071, 1961]. kill [1316, 1569]. Kiosk [1687]. Kit [1017, 1563]. knowledge [129, 1258, 1275]. knowledge-base [1275]. Known [1984]. Kong [1306]. Krakatoa [1770]. KView [1443]. Kyoto [376].

labelled [934]. Laboratory [446, 540, 1535, 1690]. LAN [113, 179, 185, 210, 227, 234, 274, 346, 372, 540, 562, 1345, 1365, 1380, 1392, 1393, 1403, 1425, 1444, 1446, 1642, 1817, 1834, 1835, 1844, 1847, 1857, 1867, 1921, 1938, 1941, 1949, 2029]. LAN-ATM [1938]. LAN-Based [1446]. LAN/MAN [540, 1834]. LAN/WAN [185, 372, 562, 1817, 1835]. Lands [1696]. landscape [1276]. Langley [1692]. Language [89, 115, 116, 126, 151, 220, 380, 523, 610, 723, 1023, 1119, 1243, 1383, 1462, 1622, 1630, 1718, 1725, 1805]. Languages [90, 1170]. LANs [54, 109, 235, 265, 342, 391, 445, 512, 521, 528, 582, 583, 588, 595, 598, 759, 797, 802, 1077, 1343, 1406, 1422, 1426, 1428, 1434, 1435, 1437, 1440, 1860, 1866, 1941, 1993, 2025]. LANs/MANs [588, 759]. laparoscopic [1057]. Large [35, 67, 118, 148, 204, 313, 688, 898, 1042, 1105, 1137, 1266, 1280, 1281, 1376, 1426, 1591, 1639, 1665, 1803, 1847, 1967, 1998]. Large-Bandwidth [1998]. Large-Scale [898, 1105, 1266, 1426, 1639, 1665]. laryngoscopy [1310]. Lasers [1929]. Late [245]. Latency [770, 890, 1361, 1559, 1587]. lattice [1307]. launch [1705]. Launches [1642]. Laws [1210]. Layer [40, 57, 109, 244, 374, 638, 683, 704, 767, 1390, 1885, 1940, 1969, 1992, 2003, 2011]. Layered [1415]. layers [446, 557]. LB [1226]. LB-SELA [1226]. LCN [2052]. LCN’98 [2054]. leaky [430, 535, 808, 977]. leaky-bucket [430, 535]. Learned [820, 1635, 1656]. Learners [1199, 1202]. Learning [319, 378, 570, 870, 878, 1036, 1123, 1153, 1160, 1169, 1183, 1195, 1350, 1385, 1462, 1602, 1672, 1778, 1829, 1863, 1910]. lectures [919]. Legacy [723, 846, 1526]. length [524, 578, 588]. less [69]. Lessons [460, 820, 1574, 1635, 1656, 1786]. Level [277, 278, 694, 732, 845, 968, 1086, 1296, 1325, 1366, 1788, 1920, 1962, 1987]. Leveraging [781, 1452, 1536]. Levy [1474]. Librarian [1665]. Libraries [1592, 1663, 1725]. Library [769, 1292, 1296, 1556, 1616, 1664– 1667, 1712, 1737, 1835]. life [462, 675, 1321]. lifecycle [1034]. Light [1978]. Lighthearted [1463]. Lightwave [605, 607, 1218, 1227, 1926, 1928]. Lightweight [728, 856]. Like [72, 655, 900, 1812, 1909]. limit [1153]. Limitations [1767]. Limited [1445, 1637, 1929]. limits [1095]. Linda [1779]. Line [702, 832, 1041, 1059, 1062, 1126, 1171, 1208, 1479]. Linear [300, 1052, 2002]. Linear-Based [2002]. Linear-Time [300]. Lines [98, 524]. Link [167, 214, 230, 853, 874, 944, 952, 1102, 1162, 1211, 1214, 1370, 1390, 1636, 1810]. Link-by-Link [214]. link-state [944]. linkage [1139]. Linking [852, 1318, 1799]. links [1180, 1277]. lip [1304]. lip-sync [1304]. Lisp [1191]. LispWeb [848]. list [1577, 1635]. LITBASE [1638]. literature [729]. Live [1562, 1600]. LMP [557]. Load [53, 529, 567, 759, 938, 993, 1280, 1378, 1846]. load-controlled [759]. loading [632]. Local [3, 5, 7, 34, 37, 40, 47, 49, 51– 53, 56, 66, 69, 86, 147, 226, 229, 299, 321, 407, 419, 576, 580, 627, 1241, 1346, 1352, 1367, 1373, 1391, 1400, 1414, 1418, 1438, 1541, 1792, 1832, 1840, 1846, 1899, 2004, 2033– 2041, 2043, 2046, 2049, 2052, 2054]. Local-Area [86, 407]. Local/Metropolitan [1414]. Locality [262, 1874]. localized [1018]. Location [1489, 1947]. LOEWE [497]. Log [736]. Logarithmic [1916]. LOGGIE [306]. Logic [399, 1917, 2025]. Logical [492, 1989, 2012]. Long [225, 363, 1899, 1919, 1998, 2000]. Long-Delay [1998]. Long-Range [2000]. Long-Term [1899]. look [538]. looking [1636]. Loop [49, 170]. Loopback [1358]. loosely [889]. Lose [1334]. Lose-it [1334]. Loss [160, 581, 800, 805, 1536, 1849, 1900, 1948, 2024, 2031]. lotos [115, 305, 306, 378, 414, 494, 495, 497, 610, 739, 946, 1223, 1224]. Lots [1467]. Low [524, 836, 1062, 1065, 1217, 1788, 1911]. Low-Bandwidth [1911]. low-earth-orbit [1065]. low-speed [524]. Lowell [2054]. Lower [244, 446, 557, 704]. lower-layer [704]. Loyola [1674]. LSAM [1329].

M [1965]. MAC [40, 109, 271, 587, 589, 590, 595, 802, 969, 1364, 1368, 1994, 2027]. machine [303, 1156]. machines [997, 1235]. Macintosh [1486, 1880]. Made [504, 1477]. Magellan [1475]. MAID [1240]. Mail [138, 252, 464, 466, 477, 480, 838, 921, 1377, 1991]. mailing [1577]. Main [843]. Maintaining [657]. Maintenance [727, 1479]. major [199]. make [1210]. Making [445, 526, 656, 1766]. mall [1293]. MAN [513, 514, 525, 540, 820, 1353, 1834]. MAN/WAN [1353]. manageable [1032]. Managed [1441]. Management [96, 124, 157, 209, 219, 243, 287, 288, 320, 344– 346, 375, 393, 473, 499, 527, 534, 550, 557, 567, 638, 643, 646, 753, 760, 807, 809, 812, 815, 827, 831, 853, 897, 910, 922, 928, 991, 1073, 1075, 1088, 1140, 1165, 1175, 1191, 1227, 1255, 1265, 1293, 1347, 1348, 1385, 1386, 1435, 1441– 1444, 1535, 1642, 1677, 1727, 1745, 1823, 1824, 1855, 1859, 1890, 1947, 2005, 2023, 2028]. Manager [388, 1843, 1976]. managers [1033]. Managing [134, 777, 1586, 1663, 1722]. Manchester [821]. Manhattan [596, 605]. Manifesto [768, 1484]. Manipulating [789, 1143]. manipulation [918]. MANs [271, 384, 391, 582, 583, 588, 595, 598, 606, 759, 1409, 1828]. Manufacturing [1852]. Map [1597, 1771, 1783]. mapped [1062]. Mapping [97, 556, 684, 791, 916, 1020, 1125, 1595, 1771]. Maps [1025, 1550]. March [1587]. market [1066]. Marketing [1734]. Marketplace [1]. Markov [636, 829, 1223]. Markup [859, 1190, 1784]. Mass [1699]. Massachusetts [2054]. Master [1619]. match [526]. match-making [526]. matching [1137]. Material [1068, 1715]. Materials [671, 1189, 1305]. Math [1185, 1604]. Mathematical [1590, 1784]. Mathematicians [100]. Mathematics [1778]. Maturation [1512]. maturity [331]. Max [768, 1484]. Maximum [1307, 1893]. May [1587, 2045, 2050]. Mbit [391]. Mbit/s [391]. Mbps [437, 1908, 2026]. MC02 [1959]. MD [593]. MDMA [1580]. means [690]. Measured [160, 265]. Measurement [297, 421]. Measurements [437, 1249, 1404, 1587, 1848]. Measures [284]. Measuring [851, 1131, 1324]. Mechanism [161, 408, 514, 637, 740, 1038, 1293, 1877, 1881, 1894, 1970, 1978]. Mechanisms [55, 158, 621, 1273, 1285, 1286, 1360, 1411, 1837]. Medi [1183]. Medi-CAL [1183]. Media [582, 583, 623, 691, 706, 746, 834, 1183, 1239, 1519, 1528, 1746, 1865, 1881, 1887, 1914]. Media-access [582]. Mediated [121, 1302, 1508]. mediation [1008]. Medical [929, 1674, 1675, 1678]. Medium [122, 363, 392, 580, 1043, 1699]. MEDLINE [1718]. Meeting [465, 1507, 1760, 1965]. Meets [1196, 1779]. Megabit [356]. Meganet [342]. Memory [312, 843, 1114, 1430, 1543, 1940, 1979]. Memory-based [1114]. MEMS [1533]. MendelWeb [1604]. Mercuri [1826]. Mercury [1540]. Merging [894]. merit [453]. mesh [1316]. meshes [1049, 1280, 1332]. Message [278, 395, 766, 904, 905, 916, 1043]. Messages [588, 1362]. Messaging [156]. Meta [306, 725, 857, 1104, 1843]. meta-information [725, 857]. Meta-Model [1843]. meta-tool [306]. Metabroker [1163]. MetaCenter [1683]. MetaCrawler [1773]. Metadata [1008, 1095–1097, 1189, 1193, 1768]. Metamorphoses [1460]. Metamorphosizing [1460]. MetaNet [597, 1374]. Metaring [594, 1375]. meteor [1042]. Method [166, 171, 300, 542, 752, 1103, 1115, 1120, 1295, 1916, 1936]. Methodologies [533]. Methodology [272, 908, 1863, 2007]. Methods [303, 795, 907, 917, 932, 976, 1222, 1271, 1281, 1561, 1673, 1737, 1860, 2016]. Metropolitan [110, 276, 292, 396, 425, 542, 596, 617, 810, 1081, 1366, 1414]. Mh [24]. MHS [180, 458, 473]. MICE [550]. Michigan [1736, 2042]. Micro [623, 999]. Micro-kernel [623]. micro-servers [999]. Microcomputer [1593]. Microsystem [786, 1650]. microwave [1277]. Middle [1465]. Middleware [1963]. Migrating [1305]. Migration [16, 143, 186, 258, 467, 1407, 1445]. military [1069]. Millicent [1807]. Mind [1608]. MING [821]. Mini [240]. Mini-slots [240]. minimizing [1246]. Minimum [223, 544, 1858, 1999]. Minimum-Hop [223]. Mining [1014]. MiniPay [990]. Minn [2037]. Minneapolis [2033–2041, 2043, 2046, 2049, 2052]. Minnesota [1350, 2033–2036, 2038– 2041, 2043, 2046, 2049, 2052]. miss [1027]. miss-es [1027]. Missing [1211, 1214]. Mixed [34]. Mobile [502, 559, 560, 686, 772, 823, 884, 1033, 1086, 1231, 1265, 1557, 1794, 1796, 1868, 1931, 1932, 1947, 1959]. Mobility [1930]. Mode [402, 403, 406, 444, 447, 592, 1888]. Model [219, 388, 525, 563, 697, 703, 733, 744, 745, 747– 750, 831, 1355, 1427, 1516, 1641, 1800, 1843, 1872, 1909, 1942, 2013]. modeler [1294]. Modeling [613, 761, 1046, 1047, 1086, 1151, 1166, 1219, 1247, 1292, 1301, 1862, 1917, 1923, 1930]. Modelling [283, 742, 791, 1091, 1291, 1296, 1298, 1310, 1595]. Models [31, 35, 38, 316, 385, 518, 519, 576, 743, 861, 953, 960, 967, 975, 1057, 1062, 1224, 1249]. modes [1294]. modified [764]. Modular [856, 1346, 1777]. Modulated [829]. module [1127, 1702]. Modules [566, 1679]. Molecular [454, 1492, 1493]. Molecules [1491]. MOMspider [657]. Monitoring [558, 953, 1958]. Montréal [795, 917]. monuments [1055]. MOO [775, 1751]. MOO-based [775, 1751]. Moored [787, 1598]. morphing [1052]. Mosaic [713, 717, 785, 787, 790, 1458, 1459, 1465, 1502, 1503, 1509, 1527, 1531, 1532, 1534, 1544, 1545, 1551, 1561, 1569, 1570, 1575, 1588, 1589, 1591, 1592, 1594, 1596, 1598, 1600, 1609, 1611, 1617, 1620, 1621, 1623, 1628, 1642, 1647, 1651, 1657, 1664, 1665, 1676, 1685, 1687, 1718, 1728, 1729, 1740–1742, 1752, 1754, 1793, 2042]. Mosaic-Based [1687, 1754]. Mosaic/WWW [1534, 1561]. MosaicForms [1648]. Move [256]. Moving [1429, 1817]. MOWS [992]. MPEG [924, 1183, 1862, 1891, 1908, 1974, 1975]. MPEG-2 [1183, 1974]. MPEG-I [1891]. MRAP [1093]. MS [1488]. MS-Windows [1488]. Multi [49, 169, 365, 392, 512, 618, 634, 694, 821, 861, 911, 1015, 1082, 1225, 1239, 1366, 1773, 1793, 1821, 1838, 1885, 1983]. Multi-access [169]. Multi-cast [1821]. multi-class [911]. multi-client [861]. Multi-Connected [49]. multi-dimensional [618]. Multi-Engine [1015, 1773]. multi-Gbit [512]. multi-Gbit/s [512]. Multi-Gigabit [392]. Multi-Head [861, 1793]. Multi-Layer [1885]. Multi-Level [694, 1366]. multi-media [1239]. multi-protocol [365]. Multi-Segment [1838]. multi-service [821, 1082, 1225]. Multi-Star [1983]. Multi-Tail [861, 1793]. multi-token [634]. Multiaccess [590, 1883]. Multicast [314, 650, 706, 806, 924, 981, 1004, 1080, 1187, 1220, 1328, 1329, 1378, 1864, 1889, 1893– 1895, 1897, 1898, 1972, 1973, 2015]. Multicasting [1896, 2008]. multichannel [425, 606, 645, 754, 941, 942]. Multifarious [24]. multifiber [1232]. Multihoming [273]. Multihop [61, 68, 579, 604, 607, 835, 1227, 1827, 1927, 1928]. Multilingual [661, 1161, 1566, 1586]. Multimedia [252, 266, 358, 360, 420, 431, 499, 550, 574, 618, 620– 622, 688, 691, 950, 959, 975, 1012, 1076, 1135, 1136, 1151, 1155, 1158, 1174, 1190, 1222, 1240– 1243, 1264, 1273, 1283, 1313, 1341, 1342, 1344, 1349, 1351, 1378, 1379, 1382, 1398, 1426, 1427, 1429, 1438– 1440, 1458, 1510, 1534, 1562, 1580, 1620, 1628, 1829, 1852, 1854, 1855, 1878, 1879, 1881, 1889, 1909, 1920, 1963, 1970, 1976, 2008, 2012]. multiparty [747]. Multipath [1935]. Multipeer [2011]. Multiple [6, 177, 210, 274, 486, 608, 757, 763, 797, 816, 1092, 1397, 1410, 1440, 1926, 1946, 1970]. multiple-access [608, 797]. Multiplexer [293, 424, 578, 805, 1076, 1091, 1396, 1925, 2031]. Multiplexers [524, 613, 949, 1936]. Multiplexing [538, 550, 742, 947, 1081, 1225, 1250, 2031]. Multipoint [1248, 1288, 1436, 1922, 1971]. Multipoint-to-multipoint [1288]. multiprotocol [439, 937]. Multirate [53]. Multiresolution [1301]. Multiserver [229]. multiservices [1093]. Multistage [522, 915, 1884, 1935]. Multitasking [488]. Multithreaded [1580, 1835]. multiwavelength [1227]. Museum [1685, 1687]. Museums [1696]. Music [1782]. Mutual [1647]. mWeb [1003]. my [1569]. Myriad [729].

N [557, 842]. NABLE [230]. NAFUR [1216]. Name [111, 237, 369, 630, 705]. Naming [94, 188, 351, 371, 845]. NASA [1474, 1692]. NASK [448]. Nation [1694]. Nation-wide [1694]. National [150, 191, 198, 337, 451, 461, 549, 887, 1507, 1535, 1538]. Nations [308]. Native [1234, 1955]. Natural [136, 1718, 1725]. nav [881]. Naval [1690]. Navigating [453, 570, 880, 1171, 1757]. Navigation [573, 776, 870, 1498, 1654, 1656, 1677, 1718, 1723, 1801]. navigational [735, 1040]. navigator [875, 881, 883]. NBS [62]. NCSA [653, 717, 771, 1568, 1570, 1606, 1611, 1729, 1741, 1742]. Near [1474]. Near-Real [1474]. NECI [1104]. Need [144, 251, 671, 1142]. needs [428, 465, 1208]. Negotiating [1270]. negotiation [1216, 1974]. Net [226, 469, 1695, 1859]. NetBiochem [1490]. NetCache [1314]. NetCoach [1682]. NetDyn [980]. nets [517]. Network [10, 12, 34, 37, 56, 61, 64, 76, 94, 96, 107, 108, 110, 148, 162, 178, 191, 214, 219, 230, 238, 254, 287, 288, 300, 341, 448, 454, 469, 511, 534, 541, 549, 596, 604, 620, 632, 635, 647, 831, 1067, 1073, 1128, 1170, 1277, 1279, 1347– 1349, 1362, 1367, 1372, 1375, 1381, 1391, 1399, 1407, 1412, 1418, 1427, 1433, 1437, 1441, 1443, 1444, 1448, 1708, 1798, 1818, 1823, 1825, 1832, 1840, 1842, 1870, 1883, 1885, 1902, 1910, 1912, 1917, 1918, 1928, 1929, 1937, 1938, 1958, 1976, 1979, 1984, 1985, 2003, 2009, 2010, 2014, 2016, 2018, 2022, 2027]. network [195, 226, 232, 285, 294, 299, 310, 337, 345, 383, 422, 427, 430, 437, 444, 447, 453, 461, 464, 474, 478, 507, 515, 532, 560, 561, 563, 565, 567, 572, 575, 576, 596, 601, 605, 629, 634, 637, 671, 681, 685, 692, 697, 712, 754, 807, 810, 825, 919, 927, 945, 948, 949, 953, 978, 980, 999, 1033, 1093, 1230, 1232, 1250, 1252, 1253, 1265, 1278, 1325, 1330, 1333, 1569, 1670, 1674, 2015]. Network-Based [919, 1910]. network-level [1325]. network-wide [697]. Networked [333, 1189, 1313, 1377, 1382, 1534]. Networking [47, 193, 200, 338, 392, 436, 438, 442, 443, 449, 450, 460, 463, 470, 475, 476, 479, 558, 676, 679, 680, 931, 1244, 1427, 1909, 2044]. networkload [743]. Networks [4–6, 8, 14, 21, 23, 30, 33, 40, 41, 52, 53, 59, 60, 67, 69, 73, 83, 111, 152, 174, 175, 177, 183, 189, 215, 217, 228, 229, 251, 262, 290, 292, 295, 376, 404, 910, 1334, 1342, 1351, 1352, 1359, 1370, 1373, 1379, 1385, 1395, 1414, 1417, 1438, 1441, 1819, 1830, 1833, 1846, 1849, 1853, 1856, 1864, 1868, 1875, 1877, 1887, 1891, 1895, 1899, 1902, 1904, 1923, 1927, 1930, 1932, 1933, 1943, 1945, 1962, 1965, 1982, 1989, 1995, 1996, 2001, 2002, 2006, 2008, 2012, 2033– 2041, 2043, 2046, 2049, 2052, 2054]. Networks [3, 7, 15, 17, 48, 49, 68, 70, 76, 77, 106, 124, 145, 154, 158, 163, 167, 170, 172, 212, 272, 276, 282, 309, 318, 319, 332, 395, 396, 401, 406, 407, 409, 411, 441, 510, 530, 531, 538, 564, 566, 590, 599, 602, 607, 615, 650, 700, 756, 830, 834, 1081, 1087, 1094, 1218, 1226, 1249, 1281, 1356, 1357, 1366, 1386, 1392, 1397, 1398, 1416, 1820, 1821, 1850, 1852, 1863, 1865, 1873, 1876, 1884, 1888, 1890, 1901, 1907, 1919, 1926, 1935, 1950, 1957, 1961, 1967, 1975, 1977, 1978, 1983, 1992, 1997– 1999, 2004, 2005, 2013]. networks [51, 123, 147, 226, 231, 237, 266, 273, 286, 316, 317, 320, 321, 326, 373, 375, 383, 386, 418, 421, 425, 429, 433, 434, 455, 472, 485, 486, 500, 506, 516, 518, 522, 529, 533, 537, 539, 542, 556, 577, 579, 580, 586, 600, 608, 609, 617, 621, 627, 633, 646, 648, 752, 758, 760, 763, 800, 808, 827, 836, 839, 915, 939, 943, 944, 951, 953, 964, 968, 978, 979, 1074, 1078, 1079, 1082, 1083, 1086, 1089, 1092, 1152, 1217, 1220, 1225, 1227, 1229, 1236, 1282, 1289, 1338, 2000]. Networkshop [95, 182]. NETworth [1647]. Neural [949, 1392, 1885, 1902]. Neural-Network-Based [1902]. neutral [1257]. Newspaper [1036, 1037, 1770]. newspaper-based [1037]. newspapers [1241]. Next [407, 475, 677, 1035, 1198, 1944]. next-generation [1035]. NFS [1933]. nif [881]. nif-T-nav [881]. Ninth [5]. NIU [268]. NIU-forum [268]. NLANR [1319]. NLM [1718]. No [1665]. NOAA [1694]. Node [46, 697, 1355, 1356, 1861]. Nodes [586, 1970, 1984]. noisy [541]. Non [172, 229, 429, 554, 637, 689, 840, 1108, 1359, 1734]. non-blocking [840]. non-discretionary [637]. Non-exhaustive [229]. non-hierarchical [429]. non-interleaved [689]. Non-Intrusive [1734]. Non-OSI [172]. Non-Real-Time [1359]. non-technical [554]. non-textual [1108]. Nonblocking [1337, 1948]. NORDUnet [345]. Normal [1857]. North [1666]. Notation [242, 382]. Notations [1553].

Notification [987]. Novel [1395, 1881]. November [2052]. ns [2015]. ns-2 [2015]. NSAP [948]. NSF [1538]. NSFNET [344]. NSSDC [708]. NSTP [987]. NTT’s [231]. Nuclear [1691]. Number [635, 1090, 1929, 1983]. numbering [232]. Numerical [1860, 1905, 1936].

O [1871]. Oak [1535]. Object

[92, 693, 753, 875, 888, 1034, 1075, 1121, 1149, 1151, 1765, 1774, 1843, 2017]. Object-Oriented [92, 693, 753, 1034, 1075, 1121, 1151, 1765, 1774, 2017]. object-relational [1149]. objectives [1078]. Objects [855, 1150, 1764, 1816, 1835]. Obstacles [476]. Occupancy [1355]. Oceanographic [1814]. Oct [2037]. October [795, 917, 2033– 2036, 2038, 2039, 2042, 2043, 2046, 2049, 2054]. ODA [359]. ODP [366, 624, 750, 958]. Off [732, 1388]. Off-Host [1388]. Office [126, 133, 137, 1446, 1910, 1989, 1991]. Officer [197]. Olias [1704]. Omninet [226]. OmniPort [1526]. OMNIS [729]. OMNIS/Myriad [729]. Omniware [1787]. OMNIWeb [708]. on-demand [1136, 1256]. On-Line [702, 1126, 1171, 1208, 1479]. One [380, 997, 1243, 1694, 1695, 1837, 1861]. ONE-IP [997]. One-Time [1837]. Ongoing [174, 1706]. Online [892, 1036, 1067, 1159, 1274, 1537, 1555, 1664, 1669, 1713, 1715]. Only [1861]. onto [132]. OODBMS [1633]. OODBs [873]. Open [13, 301, 326, 367, 383, 744– 749, 854, 959, 996, 1165, 1249, 1442, 1760]. Operating [212, 1113, 1952]. Operation [298, 344, 1412, 1921]. operational [463]. Operations [1527]. operator [1253]. Opportunities [1534]. opportunity [426]. Optic [418, 512, 580, 588, 608, 1391, 2025]. Optical [60, 109, 210, 274, 600–604, 608, 754, 802, 1232, 1370, 1373, 1830, 1856, 1927, 1982, 2025]. Optimal [42, 61, 409, 526, 529, 578, 1084, 1228, 1827, 1864, 1950, 1953, 1957, 2022]. Optimization [73, 579, 1227, 1370, 1903, 1942]. Optimizing [757]. options [819]. ORACLE [1772]. ORACLE/WWW [1772]. orbit [1065]. Orchestration [1453]. order [921, 1094, 1129, 1224]. ordered [970]. ordering [1110]. Oregon [1671, 1817]. Organisational [335]. Organization [80]. Organizational [14, 129, 838, 1268, 1534]. Organizational-Wide [1534]. Organizing [785, 866, 879, 1005, 1393, 1609]. ORGUE [360]. Oriented [92, 284, 585, 693, 739, 753, 853, 968, 1034, 1066, 1075, 1121, 1151, 1265, 1765, 1774, 1805, 2017]. origin [1132]. origins [348]. ORNL [1535]. Orwell [591]. OSDL [92]. OSI [45, 57, 75, 94, 117, 139, 172, 192, 194, 209, 243– 247, 269, 331, 347, 363, 365, 371, 372, 379, 387, 417, 426, 442, 444, 446, 457, 467, 469, 501, 523, 543, 545, 546, 557, 693, 753, 812, 948]. OSI-compatibility [948]. OSI-management [812]. OSIRIDE [216, 387]. OSIRIDE-Interest [216]. OSIRIDE-Intertest [387]. OTCA [2027]. Other [341, 857, 1718, 1735]. our [1199]. Outlining [1657]. Output [114, 322, 484, 614, 640, 696, 915, 941, 1091, 1395, 1487, 1900]. output-buffered [915]. outsiders [930]. overflow [911]. Overhead [2006]. overheads [581]. overlap [1131]. Overload [54, 218, 824, 1228, 1356]. Overview [17, 89, 157, 181, 337, 381, 515, 594, 597, 1430]. Ovid [1460]. Own [1665].

p [82, 588, 972]. P-Persistent [82]. P3P [1210]. P8 [456]. PABX [35]. Pace [103]. Pacific [571]. Packard [346]. Packet [21, 33, 37, 46, 61, 68, 70, 73, 88, 149, 152, 154, 162, 183, 212, 215, 282, 284, 318, 320, 401, 409, 427, 484, 524, 542, 579, 602, 615, 621, 628, 632, 640, 696, 752, 829, 834, 842, 945, 971, 1232, 1361, 1422, 1430, 1448, 1825, 1856, 1867, 1878, 1882, 1884, 1890, 1904, 1924, 1935, 1962, 1966].

Packet-Switched [33, 73, 621, 1856, 1890, 1904]. Packet-Switching [46, 70, 632, 834, 971]. Packet/Cell [1430]. Packetized [296, 1352]. Page [993, 1009, 1124, 1486, 1812]. pages [626, 817, 881, 925, 1130, 1141, 1184, 1185, 1203, 1479]. PageSpace [847]. painting [1047]. Pair [1823]. Palaeobotanic [1648]. Pan [200, 469, 857]. Pan-Browser [857]. Pan-European [200, 469]. Panel [1577, 1707, 1712]. panelists [795, 917]. Paper [190, 658, 795, 917, 1485, 1812]. Papers [398, 2044, 2045, 2051]. Paradigm [1000, 1076, 1136, 1924]. Paradigms [136]. PARADISE [352]. Parallel [148, 169, 967, 1617]. Parallel-serial [169]. Parameter [742, 2016]. Parameters [61, 912, 1340]. parametric [1046]. Parasite [1014, 1780]. Paris [2050]. part [635]. Partial [1224]. partially [970]. particular [671]. Partition [15]. Partitioned [1982]. Partitioning [565, 1301, 1988]. Partitions [63]. Partnerships [1668]. Passcode [1837]. Passing [166]. Passive [600, 1218, 1373, 1982]. Password [457, 1960]. past [460]. Path [410, 502, 803, 953, 1138, 1950]. Pathology [1679]. Paths [12, 486, 1074]. Pattern [1984]. Patterns [581, 852, 1569, 1571]. PayMe [1806]. Payment [869, 1806, 2020]. PCBA [1430]. PDAs [772, 1557]. PDQRAP [1408]. peak [700]. Peer [897]. Penn [1518]. People [866, 1206, 1637, 1655, 1717]. perceived [671, 1167]. Performance [7, 18, 37, 68, 70, 88, 154, 227, 234, 265, 271, 284, 291– 293, 321, 337, 373, 376, 386, 408, 411, 421, 422, 424, 430, 518, 528, 628, 632, 640, 646, 691, 696, 761, 803, 806, 840, 886, 977, 1083, 1090, 1115, 1286, 1324, 1337, 1343, 1355, 1356, 1369, 1374, 1380, 1381, 1388, 1396, 1398, 1401, 1413, 1414, 1417, 1428, 1433, 1437, 1440, 1447, 1507, 1825, 1831, 1832, 1844, 1845, 1853, 1860, 1871, 1882, 1884, 1891, 1894, 1895, 1906, 1907, 1916, 1918– 1920, 1929, 1933–1936, 1941, 1955, 1958, 1984, 1994, 1999, 2000, 2010, 2015]. performance [320, 388, 423, 494, 513, 515, 520, 539, 549, 567, 576, 596, 598, 642, 683, 689, 697, 740, 762, 797, 799, 821, 822, 912, 938, 965, 974, 1022, 1076, 1080, 1086, 1218, 1250, 1254, 1271, 1285, 1318, 1579]. performance-analysis [1271]. Period [1629]. permanencevia [1127]. Permutation [1338]. Persistent [82, 114, 588, 972]. person [1577]. Personal [239, 609, 822, 823, 1009, 1036, 1106, 1182, 1377, 1946]. personalization [1100]. Personalized [622, 1164, 1770]. perspective [130, 232, 436, 571, 1169, 1253]. perspectives [433, 584]. Pesce [1497]. PET [496]. Petri [517]. PGP [1789]. Phase [181, 196–199, 256]. Philosophy [1696]. Phoenicia [1516]. Photonic [227, 234]. Physical [2012]. Physically [1047]. Physically-based [1047]. physics [708]. Pico [1842]. Pico-Cellular [1842]. Picture [1730, 1877]. Pictures [441, 1472, 1725]. Piece [1052]. Piece-wise [1052]. piecewise [1046]. Piggyback [631, 1112]. Pilot [204, 249, 257, 349, 352, 359, 447, 929, 1267]. Piloting [457, 550]. ping [683]. Pipelined [1430]. pipelining [689]. place [1141]. Placement [543]. Places [1143, 1756]. Plane [1059, 1985]. Planetary [1475]. Planning [32, 33, 108, 343, 1444, 1619]. Plans [194, 232]. plant [703]. Plants [1458]. Platform [440, 710, 983, 1154, 1210, 1303, 1443, 1716, 1880, 1959]. playing [1144]. Plaza [2033–2036, 2038]. Point [948, 1030, 1248, 1922, 1965, 2027]. Point-to-Multipoint [1248, 1922]. Point-to-Point [1965, 2027]. pointer [575]. points [413, 633, 1271]. Poisson [578]. POL34 [1278]. Poland [448]. Policies [460, 542, 614, 812, 931, 962, 1233, 1334, 1413]. Policing [404, 514, 1831, 1877, 1885]. Policy [50, 315, 336, 340, 571, 578, 636, 679, 839, 1093, 1445, 1507, 1517, 1988, 2023]. Policy-Driven [2023]. Polish [1278]. Polling [72, 229, 316, 389, 587, 757]. polygon [1056]. Polymap [1783]. polynomial [1046]. Pooh [710]. POP [1991]. poppy [1130]. populating [1064]. Portable [1835]. Possible [159, 174, 356, 502]. post [659, 1991]. post-hoc [659]. Poster [368]. Potentials [1767]. Power [1791]. PR [215]. PR-EXPRESS [215]. Practical [125, 1118, 1156, 1969, 1972]. pragmatic [493]. Prague [2044]. PRC [680]. predict [1138, 1167]. Predicting [1279]. predictive [487, 839]. preemption [1230]. Preface [390]. preferences [1036]. Prefix [564]. preliminary [515]. premise [424]. premises [629, 966]. Preprocessing [1553]. Present [192, 193, 1467]. Presentation [126, 769, 984, 1003, 1457, 1458, 1472, 1483, 1556, 1643]. presentations [1174]. Preserving [1945]. President [1507]. Pressure [1900]. Pricing [794, 1225, 1523, 1904]. Primary [12, 178]. Principle [214, 405, 1867, 1915]. principled [1006]. Principles [297, 400, 965]. Print [1487]. printing [1703]. Priorities [70, 510, 805, 1091, 1344, 1395, 1440]. prioritized [646, 1408]. Priority [536, 829, 971, 1384, 1431, 1438, 1819, 1857, 1882, 1938]. Priority-Based [1819, 1938]. Privacy [480, 677, 921, 1210, 1262]. Private [30, 1992]. PRO [1706]. Probabilistic [1917]. probabilities [522]. Probability [160]. Probing [2001]. Problem [15, 545, 874, 1228, 1863]. Problems [159, 334, 512, 649, 1253, 1528, 1579, 1649]. Procedure [173, 366, 1981]. Proceedings [2034–2036, 2038–2043, 2046–2048, 2052– 2054]. Process [897, 1091, 1223, 1395, 1405, 1849]. Processes [114, 751, 829, 947, 1148]. Processing [48, 148, 170, 301, 367, 744– 749, 959, 1060, 1162, 1308, 1567, 1718, 1779]. Processor [1121, 1824]. Procurement [1537]. producing [1143]. Product [184, 324, 1700]. production [1140].

Products [203, 387, 1475]. profile [1137, 1176]. profiles [1138]. programmable [391]. Programme [103, 104]. Programming [884, 1119, 1188, 1622, 1787, 1805, 1963]. programming-by-demonstration [1188]. Programs [1565, 1673]. Progress [19]. progressive [1312]. Project [105, 141, 157, 158, 180, 197, 249, 257, 349, 353, 442, 454, 456, 471, 624, 812, 922, 994, 1242, 1266, 1320, 1460, 1461, 1489, 1538, 1606, 1646, 1666, 1727, 1735]. projection [1307]. Projects [929, 1612, 1613]. promise [1215]. Promondia [988]. Promotion [1697]. Proof [842, 940]. Propagation [85, 486, 1967]. Properties [492, 1070, 1862]. Property [731, 1154, 1269, 1900]. Proposal [239, 797, 831, 1894]. Proposed [359, 637, 1402]. proposes [710]. Prospects [288]. Protection [54, 361, 1154, 1262]. Protocol [4, 11, 19, 25, 31, 42, 55, 56, 60, 63, 69, 71, 109, 112, 113, 116, 171, 173, 215, 221, 222, 233, 263, 269, 275, 281, 299, 303, 305, 365, 384, 398, 400, 407, 418, 459, 490, 491, 504, 523, 527, 562, 587– 590, 592, 593, 595, 631, 634, 641, 642, 686, 690, 720, 757, 766, 797, 802, 916, 921, 950, 952, 969, 972, 974, 987, 1008, 1081, 1236, 1364, 1376, 1383, 1390, 1394, 1402, 1409, 1410, 1414, 1806, 1807, 1834, 1836, 1838, 1856, 1887– 1889, 1897, 1952, 1955, 1968, 1971, 1994, 1995, 2003, 2018, 2027]. Protocol [57, 62, 1389, 1394, 1408, 1896, 1991]. Protocol-less [69]. Protocols [45, 52, 58, 213, 224, 235, 271, 358, 364, 373, 374, 379, 401, 415, 417, 505, 506, 521, 559, 580, 598, 600, 698, 704, 796, 836, 932, 939, 975, 1070, 1356, 1373, 1400, 1423, 1867, 1886, 1960, 1972, 1991, 2015, 2029]. prototype [540, 653, 807, 1172]. prototypes [1073]. Prototyping [93]. Provide [320, 1761, 2005]. provider [672, 682]. provider-rooted [682]. Providers [435, 1707]. Providing [560, 626, 790, 816, 866, 981, 1024, 1037, 1065, 1078, 1089, 1549, 1565, 1596, 1629, 1735, 1737, 1931, 1932, 2032]. proving [548]. provision [335, 383]. Provisioning [1442]. ProWeb [1001]. Proxies [652, 1733, 1767]. Proxy [993, 1112– 1115, 1157, 1317, 1324, 1329, 1330, 1802]. Pseudo [1000, 1339]. Pseudo-isochronous [1339]. Pseudo-serving [1000]. PSF [498]. PSTN [508]. PSTV [302]. PSTV-IX [302]. Public [20, 30, 33, 77, 145, 183, 232, 282, 478, 678, 720, 922, 978, 1131, 1262, 1270, 1653, 1669, 1695, 1696, 1961, 2009]. Publication [1469]. Publisher [1257, 1800]. publisher-neutral [1257]. Publishers [1242, 1810]. Publishing [572, 701, 792, 793, 1121, 1160, 1170, 1244, 1245, 1257, 1496, 1639, 1699, 1708– 1711, 1755, 1800]. Purdue [1305]. purpose [641]. push [1098, 1100]. push-type [1100]. Putting [723, 913, 1485, 1746].

Q [593]. QoS

[912, 1078, 1281, 1315, 1963, 1996, 2023].

QoS-based [1963]. Quadro [602]. Quality [282, 283, 471, 472, 569, 828, 838, 876, 951, 963, 1037, 1216, 1336, 1437, 1781, 1849, 1855, 1981, 2001, 2021]. Quality-of-Service [2001]. Quantitative [758, 1224, 1335]. Quasi [842, 1846]. quasi-independence [842]. Quasi-Static [1846]. queries [1109]. query [1023, 1109, 1164, 1177]. querying [880, 1149]. quest [1016]. question [128]. questions [128, 506]. Queue [524, 578, 829, 1094, 1229]. queue-length-threshold [578]. queued [1335]. Queueing [525, 593, 636, 840, 1408, 1431]. queues [800]. queuing [484].

R&D [254]. Radio

[61, 68, 215, 579, 830, 952]. radiosity [1303]. Radisson [2036, 2038]. RAE [180]. Raggett [1706]. Ramada [2042]. Ramada-Congress [2042]. Random [27, 61, 87, 322, 324, 741, 836, 1408]. Range [2000]. Rapid [507, 1402]. rapidly [679]. RARE [95, 338, 447, 548]. RARE/CEC [338]. RARE/COSINE [447, 548]. Rate [178, 267, 297, 633, 700, 805, 826, 965, 1081, 1085, 1089, 1090, 1094, 1226, 1284, 1334, 1905, 1934, 1954, 1974, 1999, 2032]. Rate-Based [965, 1090, 1094, 1226, 1954]. rate-controlled [1089]. rates [695]. Ratio [406, 2031]. Rationale [92]. rationalisation [659]. RAVE [899]. RBF [1392]. RDF [1215]. Re [162, 593, 1275, 1536, 1974]. Re-Assembly [162]. Re-engineering [1275, 1536]. Re-negotiation [1974]. re-use [593]. Reaching [553]. Reactive [1354, 1913]. Reader [125, 1800]. Readers [1810]. Ready [256]. Real [25, 86, 284, 288, 632, 746, 787, 789, 811, 841, 890, 891, 899, 908, 924, 950, 988, 1085, 1140, 1263, 1311, 1342, 1359, 1412– 1414, 1416, 1474, 1496, 1503, 1564, 1785, 1826, 1848, 1851, 1857, 1889– 1893, 1920, 1974, 1977, 2029, 2030]. Real-Time [25, 86, 284, 632, 746, 787, 811, 841, 890, 891, 899, 924, 988, 1085, 1140, 1263, 1342, 1359, 1412–1414, 1416, 1564, 1785, 1826, 1848, 1851, 1857, 1889, 1891–1893, 1920, 1974, 1977, 2029]. Real-World [1496]. realised [426]. Reality [1055, 1215, 1254, 1297, 1754]. realizing [268]. Really [144, 251, 357]. Realtime [1598, 1620]. Reassembly [152, 284]. Receiver [281]. reception [916]. recognition [627]. Recommendation [242]. recommendations [340]. Reconfigurable [1844]. Reconfiguration [26, 71, 1074, 1227, 1858]. reconstruction [628]. Recovery [170, 842]. Reduce [1629, 2006]. Reduced [87, 972]. Reducing [323, 627, 890, 1967]. Reduction [1987]. Redundancy [1911]. Redundancy-Based [1911]. Reengineering [1640]. REFEREE [991]. Reference [671, 744, 745, 747, 749, 1013, 1127]. References [777, 1722]. refinement [1049]. Refining [1998]. refraction [1062]. regional [898, 1111]. regionalization [468]. Register [407, 1391]. Register-insertion [407]. Regular [68]. Regulating [1287, 1997]. Regulation [841, 1071, 1210, 1923, 1969, 2024]. Related [238, 784, 1468]. Relating [957]. relation [1316]. relational [1149]. relations [414, 934]. relationships [838]. relative [1131]. relaxation [1109]. relay [375, 439, 440, 639, 654]. relaying [371]. relays [562]. Release [796, 1867]. relevance [1164]. Relevant [1601]. Reliability [55, 300, 705, 1357, 1823, 1917, 1982]. Reliable [165, 596, 706, 767, 921, 939, 1021, 1026, 1423, 1896, 1898, 1949, 1972, 1973, 2015]. Relieve [1629]. remarks [125]. Remedial [1409]. Remote [366, 1179, 1502, 1813]. Rendering [1052, 1062, 1116, 1134]. RENDEZVOUS [1583]. renegotiable [1336]. Repeat [281, 642]. Repeater [2026]. repetition [1102]. Replacement [1953]. replicated [980]. Replication [845, 2017]. Report [43, 156, 338, 343, 376, 1260, 1574, 1634, 1689]. Reporting [791, 1595]. Repositories [880, 1803]. Repository [723, 1590, 1683]. Representation [413, 1008]. Representational [1641]. Representing [1658]. Republic [77, 449, 2044]. Request [1332, 1836]. requests [1138]. Requirement [404]. Requirements [93, 162, 282, 315, 746, 824, 890, 909, 912, 1190, 1815, 1818, 1965, 1981]. Research [106, 125, 191, 335, 337, 338, 348, 359, 433, 436, 443, 448, 449, 461, 463, 465, 540, 549, 837, 1073, 1257, 1259, 1504, 1533, 1588, 1594, 1606, 1653, 1690]. Research-based [1606]. researchers [550]. Reservation [38, 405, 761, 801, 803, 836, 1081, 1236, 1273, 1410, 1838, 1951, 2002]. Reservations [1216, 1929, 1980]. Reset [1834]. Residential [1247, 1883]. Resistant [1960]. Resolution [9, 84, 111, 369, 614, 1117, 1133, 1421, 1921, 1944, 2003]. resolving [1253]. Resource [2, 511, 724, 725, 757, 796, 827, 1093, 1101, 1124, 1225, 1238, 1604, 1626, 1769, 1907, 1966, 2005, 2023]. Resources [726, 825, 1108, 1173, 1245, 1554, 1574, 1612, 1614, 1615, 1668, 1806]. respect [700]. Response [362]. responsible [1000]. Responsive [1042]. Rest [1420]. Restaurant [1635]. restricted [800]. restructuring [858]. Results [158, 624, 662, 709, 1759]. retransmission [581, 834, 940]. retransmission-based [834]. Retransmit [1993]. Retrieval [308, 656, 708, 729, 730, 1012, 1164, 1176, 1189, 1243, 1480, 1492, 1530, 1682, 1718, 1729, 1865]. Return [1530]. Reuse [395, 586, 627, 1079, 1592, 1957]. Reusing [1239]. Reversible [217]. Review [159, 180, 184, 196, 209, 651, 897, 1507, 1635]. Review-Based [1635]. revisited [980]. RFC [684]. RFC822 [97]. RIB [1391]. ride [1180]. Ridge [1535]. rights [1269]. Ring [7, 10, 18, 26, 55, 112, 154, 299, 323, 373, 384, 386, 407, 421, 592, 595, 634, 754, 760, 1232, 1341, 1344, 1394, 1395, 1414, 1440, 1938, 2027]. ring-based [595]. rings [393]. Risk [758, 1665]. RISP [2003]. RLP1 [952]. RMC [1804]. Road [476, 1697]. Robin [1901]. Robot [1451]. Robotic [1684]. Robust [112, 285, 1571, 1929, 1984]. Role [190, 268, 461, 956, 1480, 1670]. Romp [1463]. ROMs [664]. Room [1461]. rooted [682]. Roots [1512]. Rotation [1394]. Round [1897, 1901]. Round-Trip [1897]. Route [163, 1237]. routed [1338]. router [801, 1128]. Routes [217, 1066]. Routing [3, 59, 74, 223, 276, 286, 290, 315, 319, 365, 371, 410, 429, 444, 466, 485, 542, 564, 597, 604, 606, 649, 724, 767, 799, 804, 835, 968, 1077, 1217, 1237, 1325, 1332, 1338, 1827, 1864, 1876, 1895, 1942, 1943, 1981, 1996, 2001, 2002, 2004]. RSVP [923]. RTD [457]. RTMW [1187]. Rules [684, 1809]. runtime [1702].

s [391, 512, 540, 583, 589, 595, 695, 1375, 1956]. Safe [276, 1210, 1366]. safeguarding [1269].

SAFENET [1415]. said [710]. salient [1009]. Sampler [1687]. Sampling [1281, 1613, 2016]. Santa [2051]. SARA [2004]. Sardinia [898]. Satellite [357, 370, 818, 1248, 1264, 1845, 2000]. satellite-based [818]. satellites [1065]. Saxon [1455]. Scalable [653, 720, 981, 1072, 1128, 1232, 1806]. Scale [148, 204, 313, 688, 898, 947, 1105, 1137, 1266, 1280, 1426, 1639, 1665]. scales [1088, 1092]. Scaling [817, 1038, 1515, 1956]. scan [1062]. scan-line [1062]. Scenario [835, 1855]. scenarios [356, 1282]. Schedulers [1986]. Scheduling [533, 578, 588, 759, 962, 963, 1250, 1335, 1384, 1819, 1890, 1907, 1970, 2006]. Schema [1808]. Schema-Based [1808]. Scheme [8, 27, 69, 169, 370, 393, 759, 760, 763, 803, 938, 940, 942, 943, 964, 977, 1079, 1220, 1248, 1284, 1287, 1312, 1352, 1409, 1897, 1900, 1901, 1913, 1925, 1954, 1957, 1993, 2006, 2020, 2024]. Schemes [68, 262, 324, 389, 564, 646, 965, 1351, 1386, 1438, 1874, 1999]. Scholarly [792, 1709]. School [1160, 1534, 1607, 1608, 1626]. Science [918, 1475, 1538, 1574, 1604, 1610, 1615, 1683, 1740]. Science/Math/History [1604]. Sciences [136, 784, 790, 1468, 1518, 1596]. Scientific [310, 1278, 1653]. Scrapbook [1188]. screen [135]. SCSI [1871]. SDL [89, 91, 92, 220, 305, 612, 901, 902, 906, 907, 909]. SDL-92 [612, 902]. SDN [474]. SDRS [1626]. Seamless [1044, 1932]. Search [1015, 1016, 1026, 1104, 1105, 1111, 1125, 1131, 1149, 1158, 1178, 1450, 1727, 1728, 1773, 1937]. Searching [656, 730, 1019, 1064, 1161, 1728]. Seas [1667]. second [709]. secondary [1160]. sector [433]. Secure [328, 639, 813, 921, 1070, 1145, 1146, 1260, 1806]. secured [921]. Securing [1790]. Security [150, 158, 251, 277, 326, 455–459, 548, 624, 812, 989, 1069, 1270, 1367, 1500, 1730– 1732, 1747, 1788, 1789, 1873, 1960, 1961]. seed [1307]. seeking [571]. seen [710]. Segment [1838]. Segmentation [221]. Seismic [99]. SELA [1226]. Selected [2044, 2045]. Selection [12, 163, 739, 1178, 1825]. Selective [122, 281, 642, 1049, 1643, 1964, 1966, 1988, 2001]. Selective-Repeat [281]. Self [671, 1074, 1210, 1393, 1673, 1697, 1801, 1987]. Self-Configuring [1801]. self-healing [1074]. self-help [671]. Self-Organizing [1393]. Self-Promotion [1697]. self-regulation [1210]. Self-Similarity [1987]. Self-Study [1673]. Semantic [859, 1727, 1729]. semantics [750, 933, 958]. semi [767]. semi-FIFO [767]. SEMPER [1260]. Sense [65]. sensitive [946]. Separate [1915]. September [376, 2038, 2040, 2041]. Sequence [904, 905, 936, 1945]. sequenced [766]. sequences [236, 1235, 1311]. sequential [751]. serial [169]. Server [64, 528, 555, 576, 653, 668, 688, 771, 846, 848, 862, 863, 992, 999, 1112, 1126, 1139, 1157, 1191, 1327, 1341, 1380, 1509, 1518, 1529, 1568, 1572, 1587, 1597, 1608, 1626, 1634, 1656, 1678, 1680, 1701, 1704, 1705, 1732]. Server-Side [1572]. server/client [528]. Servers [705, 722, 732, 989, 998, 999, 1115, 1237, 1593, 1616, 1745, 1748, 1749, 1766, 1802]. Service [32, 73, 141, 154, 160, 172, 249, 265, 282, 283, 313, 318, 352, 353, 361, 424, 444, 447, 451, 471, 472, 477, 507, 512, 585, 626, 629, 636, 639, 761, 817, 821, 826, 833, 925, 948, 987, 997, 1005, 1037, 1051, 1074, 1082, 1089, 1111, 1216, 1225, 1251, 1270, 1288, 1289, 1334, 1336, 1453, 1667, 1805, 1810, 1841, 1849, 1855, 1865, 1869, 1909, 1941, 1974, 1981, 1989, 1999, 2001, 2006, 2009, 2021, 2032]. serviceability [1333]. Services [30, 38, 58, 75, 77, 96, 174, 187, 201, 202, 208, 280, 284, 285, 296, 335, 379, 419, 435, 440, 442, 456, 457, 474, 500, 508, 511, 546, 548, 558, 568, 587, 622, 630, 635, 665, 681, 699, 704, 763, 788, 796, 810, 825, 854, 899, 963, 966, 970, 981, 1032, 1128, 1247, 1262, 1266, 1274, 1343, 1347, 1361, 1368, 1394, 1521, 1522, 1628, 1635, 1663, 1669, 1671, 1712, 1735, 1828, 1866, 1879, 1912, 1954, 1961, 2005, 2006, 2009, 2023]. Servicing [289, 953]. Serving [247, 721, 1000, 1554, 1597]. Session [57, 732]. Sessioneer [732]. Sessions [1700]. Set [214, 860, 1806]. Sets [916, 1489]. Setting [1621]. SFgate [730]. SGML [768, 860, 1482– 1484, 1547, 1701, 1704, 1705, 1803]. shaper [808]. Shaping [1432, 1920, 1987]. Shared [41, 536, 695, 840, 1141, 1293, 1763, 1887, 1940, 1979]. shared-buffer [695]. Shared-Media [1887]. Sharing [2, 866, 1005, 1050, 1209, 1534, 1763]. shark [1125]. shark-search [1125]. sheets [1029, 1190]. shelf [732]. SHEN [1731]. Shoemaker [1474]. Shoemaker-Levy [1474]. shoppers [892, 1172]. Shorthand [242]. shower [1042]. shuffle [605, 1338]. Shuffle-Exchange [605]. ShuffleNets [835]. shuffleout [1080]. Siberia [1276]. siblings [1316]. Side [1550, 1572, 1783]. Sides [39]. sifting [1013]. signal [614]. signaling [294, 419, 798]. signaling-intensive [419]. Signalling [374, 635]. Signals [36, 423, 1870]. significance [230]. Silent [135]. SILS [419]. Similarity [1987]. Simple [64, 504, 606, 1056, 1321, 1336, 1385, 1478, 1571, 1855, 1991]. Simplifications [519]. simplified [275]. Simulated [227, 234]. Simulation [48, 68, 213, 278, 318, 417, 419, 575, 676, 967, 1057, 1092, 1250, 1270, 1310, 1371, 1387, 1407, 1591, 1680, 1842, 1904, 2014]. Simulations [1605, 1903]. Simulator [1086, 2015]. simultaneously [885]. Singapore [1201]. SingAREN [1201]. Single [410, 579, 696, 697, 914, 945, 1926, 1983]. Single-Hop [1926, 1983]. single-node [697]. single-packet [945]. single-stage [696]. Sink [47]. Site [199, 569, 1020, 1096, 1106, 1124, 1125, 1641, 1652]. site-specific [1106]. SiteHelper [1018]. sites [1103]. Situation [181, 192, 193]. six [500]. Sixth [2051]. size [1131]. sizes [409]. Skipping [8]. slice [837]. sliding [842]. Slot [395, 586, 1079, 1355, 1924, 1957]. slots [240]. Slotted [114, 588, 754, 1373, 1938]. Slotted-Ring [1938]. Slovak [449]. Slow [1998]. Slow-Start [1998]. Small [46, 1883]. Smart [1261, 1790]. SMDS [440, 650, 1348]. SMIL [1144]. Smooth [1998]. smoothing [1085]. smoothness [1089]. SMTP [1991]. SNMPv2 [807]. SOAPs [702]. Sociable [1506]. Social [679, 1740]. societal [931]. Society [29, 2044]. Society/5th [2044]. Socio [138]. Socio-technical [138]. Sociological [1513]. Sociology [136]. Soft [1317, 1890, 1978]. Soft-Real-Time [1890]. Software [2, 62, 93, 171, 213, 264, 446, 707, 1590– 1592, 1822, 1898]. solid [1292]. Solution [512, 602, 1084, 1419, 1422, 1424]. Solutions [159, 334]. Solving [1861]. Some [125, 174, 253, 506, 519, 649, 740, 794, 838, 1071, 1340, 1523]. SONET [423]. sorting [1056]. sound [1300]. Source [979, 1289, 1720, 1831, 2004, 2013]. Source-Initiated [2004]. source/virtual [1289]. Sourcerer [1720]. Sources [297, 742, 1024, 1090, 1161, 1824, 2031]. South [1667, 2042]. Space [311, 474, 484, 708, 784, 1196, 1301, 1468]. space-division [484]. Spanning [113, 544]. SPARC [1850]. SPARC-Based [1850]. Sparrow [1141]. spatial [627, 1300]. Speaker [1304]. Special [44]. specialists [142]. specialized [848]. Species [791, 1132, 1595]. Specific [1106, 1176, 1802]. Specification [90, 115, 151, 181, 220, 269, 398, 493, 747, 954, 955, 957–959, 1231, 1284, 1869]. Specifications [93, 275, 380, 414, 495, 545, 935, 1547]. specified [417]. specify [379]. Specifying [382, 960, 1097]. Spectrum [177, 726]. speech [330, 1029]. Speed [200, 201, 210, 271, 274, 309, 310, 338, 354, 355, 358, 397, 401, 427, 438, 442, 472, 479, 485, 486, 488, 506, 521, 524, 532, 542, 568, 580, 582, 588, 598, 616, 633, 642, 698, 759, 797, 808, 819, 927, 943, 968, 979, 1078, 1094, 1201, 1217, 1277, 1351, 1352, 1373, 1390, 1398– 1400, 1403, 1418, 1427, 1577, 1819, 1830, 1834, 1844, 1853, 1856, 1867, 1873, 1875, 1876, 1888, 1919, 1933, 1957, 1967, 2001, 2027]. Speeds [1874]. speedup [1337]. SPHINX [1106]. Spice [1184]. Spinning [1627]. spot [835]. Spread [177]. SQL [1737]. SR [1248]. SSCMP [766]. Stability [21, 22]. stable [317]. Stack [1868]. Stage [696, 1426]. Standalone [1941]. Standard [501, 545, 626, 674, 1402, 1475]. Standardisation [101]. Standardization [80, 644, 932, 1249]. Standardized [264, 1487]. Standards [13, 17, 139, 192, 203, 209, 279, 301, 643, 1073, 1097, 1267, 1365]. Star [104, 422, 600, 601, 802, 836, 1218, 1373, 1469, 1983]. Stars [1982]. Start [1998]. Starter [1563]. Starting [569]. Starvation [299, 1422]. starvation-free [299]. State [19, 27, 301, 303, 646, 887, 935, 944, 1235, 1246, 1518, 1644, 1671]. state-dependent [646]. States [317, 461, 1612]. Static [1846]. Station [235, 311, 636, 972, 1926, 1929]. Stations [945, 1433]. Statistical [293, 424, 524, 538, 613, 1078, 1225, 1862]. Statistics [34, 1739, 2016]. Status [216, 584, 795, 917, 1260]. Steady [27]. Step [32, 58, 677]. Step-Wise [58]. stepping [366]. steps [475, 727]. Still [1877]. STM [568]. Stochastic [517, 1223, 1999]. stone [366]. Stops [1463]. Storage [308, 499, 688, 1320, 1445, 1912]. Strategies [37, 318, 351, 614, 734, 755, 835, 893, 1250, 1855, 1904, 1907]. Strategy [411, 430, 658, 1401, 1908]. stratification [830]. Stream [284, 939, 1802]. Streaming [1975]. Streamlines [1545]. Streams [278, 625, 947, 1908, 2008]. Street [596, 605]. Strike [1544]. Stringing [14]. stripping [384]. strong [1261]. Structural [960, 1014]. Structure [68, 183, 707, 762, 798, 1096, 1101, 1632, 1737, 1822]. structure-enhanced [707]. Structured [63, 666, 711, 730, 864, 971, 1096, 1811, 1952]. structured-priority [971]. structures [132, 335, 526, 1303]. structuring [866, 956]. Students [1156, 1198, 1611, 1612, 1624]. Studies [1040, 1667, 1891]. Study [6, 178, 278, 318, 373, 419, 500, 513, 561, 578, 605, 635, 636, 668, 788, 808, 829, 892, 962, 970, 980, 1013, 1115, 1216, 1250, 1270, 1298, 1305, 1365, 1396, 1407, 1414, 1446, 1448, 1522, 1602, 1629, 1640, 1646, 1648, 1671, 1673, 1853, 1895, 1904, 1905, 1922, 1933, 1989, 1991, 2024]. Studying [1569]. style [1029, 1184, 1190]. Sub [456, 1434]. Sub-Datagram [1434]. Sub-Project [456]. Subject [1258]. subjectivity [1156]. Subnets [212]. Subnetwork [15, 408, 740, 1838]. Subscriber [1839]. Subscribers [832, 1883]. Substantiating [878]. Substantive [1697]. Substrate [1787]. Subsystem [26]. Subtle [74]. Suburban [1666]. successes [1212]. Successful [137]. sufficient [317]. suites [264, 933]. Summary [343, 1147, 1587]. Sun [786, 1650]. SunWeb [786, 1650]. superchannel [763]. Supercomputers [253]. supercomputing [1207]. Supercube [605]. Superhighway [814]. SuperJANET [549]. SUPERLAN [512]. Supernodes [177]. Support [2, 64, 91, 258, 314, 444, 495, 620, 623, 650, 699, 833, 857, 895, 920, 950, 966, 1034, 1051, 1067, 1160, 1208, 1275, 1325, 1383, 1440, 1515, 1544, 1550, 1576, 1588, 1706, 1826, 1828]. Supported [348]. Supporting [154, 161, 490, 681, 746, 850, 872, 945, 1032, 1183, 1395, 1763, 1878]. Supports [1502]. supra [542]. supra-high-speed [542]. surface [1063]. Surfing [1652]. surgery [1057]. Surrounding [153]. Survey [49, 458, 662, 709, 765, 811, 910, 1028, 1054, 1168, 1759]. survival [434]. Sustainable [1243, 1905]. SVC [1818]. swapping [1171]. Swim [1665]. Swiss [191]. Switch [149, 191, 227, 234, 235, 422, 484, 536, 568, 614, 645, 695, 765, 809, 837, 914, 938, 941, 1080, 1233, 1285, 1337, 1346, 1369, 1819, 1900, 1903, 1906, 1924]. Switched [33, 59, 73, 183, 285, 320, 427, 478, 621, 1425, 1428, 1825, 1856, 1861, 1871, 1872, 1879, 1890, 1904, 2026, 2029]. Switches [640, 643, 696, 840, 1084, 1335, 1884, 1935, 1948]. Switching [36, 37, 46, 70, 88, 152, 166, 218, 282, 419, 522, 539, 541, 566, 602, 607, 616, 632, 762, 799, 822, 834, 915, 971, 1220, 1250, 1426, 1430, 1830, 1878, 1882, 1884, 1922, 1935, 1970]. SWITCHmail [188]. Switzerland [788, 1522, 2045]. Swoop [1772]. Symmetric [1984]. Symposia [860, 1809]. symposium [338]. sync [1304]. Synchronizable [1235]. Synchronization [423, 691, 1583, 1841, 1881, 2007]. Synchronized [527, 766, 1135]. Synchronous [52, 575, 774, 870, 987, 1362, 1363, 1501, 1505]. Syntactic [1010]. Synthesis [63, 976]. System [36, 66, 76, 93, 99, 100, 118, 124, 137, 150, 164, 218, 277, 312, 313, 350, 455, 635, 641, 666, 684, 708, 724, 736, 775, 779, 822, 850, 866, 869, 898, 971, 984, 1020, 1034, 1067, 1088, 1107, 1160, 1209, 1231, 1238, 1255, 1256, 1293, 1295, 1308, 1349, 1367, 1377, 1379, 1393, 1401, 1431, 1437, 1447, 1470, 1502, 1515, 1555, 1583, 1584, 1610, 1618, 1625, 1626, 1629, 1666, 1682, 1721, 1736, 1748, 1749, 1751, 1760, 1763, 1769, 1777, 1814, 1854, 1871, 1880, 1920, 1922, 1952, 2017]. Systematic [270, 1910]. Systems [13, 85, 86, 114, 129, 151, 157, 159, 166, 219, 229, 280, 289, 291, 326, 376, 382, 397, 419, 420, 488, 492, 499, 502, 566, 567, 610, 623, 624, 691, 693, 748, 906–908, 918, 934, 946, 959, 960, 976, 1001, 1011, 1038, 1041, 1075, 1100, 1219, 1222, 1228, 1264, 1343, 1382, 1413, 1430, 1442, 1488, 1514, 1516, 1678, 1718, 1772, 1779, 1822, 1831, 1851, 1862, 1946, 1953, 1976, 2028].

T [881]. T.V. [164]. Tabu [1937]. tag [860]. tags [860]. Tail [522, 861, 1793]. tailorable [1020, 1142]. Tailored [1125, 1383]. Tailoring [1472]. Taking [1275, 1697, 1815]. tall [1130]. tangled [660]. TAO [787, 1598]. Targeting [1198]. Tariff [183]. Task [1195]. Task-based [1195]. tasks [126, 1188]. Tax [1644]. TAXI [1839]. Taxing [1644]. Tcl [1197]. TCP [75, 261, 342, 363– 365, 560, 1087, 1286, 1287, 1417, 1418, 1845, 1847, 1966, 1993, 1997, 1998, 2000, 2032]. TCP/IP [261, 342, 363– 365, 560, 1087, 1417, 1418, 1845, 1847]. TCP/IP-OSI [365]. TCP/IP-X.25/OSI [363]. TDM [392]. Tdma [60]. teacher [1208]. Teachers [1199, 1612]. Teaching [1160, 1534, 1628, 1630, 1776]. Team [362, 1603]. Team-Based [1603]. Technical [138, 554, 819, 1212, 1305, 1348, 1513, 1574, 1634]. TEchnique [392, 795, 799, 917, 1131, 1651, 1915, 1965]. Techniques [61, 142, 169, 275, 351, 490, 582, 583, 606, 997, 1058, 1060, 1083, 1221, 1308, 1321, 1572, 1718, 1876, 1997]. Technologies [354, 1403, 1543, 1693]. Technology [143, 146, 438, 679, 687, 869, 913, 1179, 1264, 1530, 1597, 1692, 1765]. Tele [1528, 1628]. Tele-Amphi [1628]. Tele-Conferencing [1528]. Telecommunication [38, 108, 145]. Telecommunications [17, 29, 93, 134, 190, 609, 821]. Teleconferencing [60]. Telematic [159]. telematics [132]. telepathology [859]. Telephone [66, 128, 960]. Telephony [1231, 1941]. telepresence [1254]. Telerobotic [1813]. Telescopes [1684]. Teleservice [16, 818, 928]. Teleteaching [1068, 1282]. Teletext [72]. Teletext-Like [72]. Teletraffic [28, 295, 519, 530, 823]. Teletrust [102]. Television [66]. TeleWeb [889]. teleworkers [1039]. Teleworking [1854]. Template [1117]. Templates [1768]. Temporal [399]. temporary [1182]. Ten [331, 1277]. TEN-34CZ [1277]. Tenet [620]. Tensorial [1059]. Tenth [51]. Ter [157–159, 624]. terabit [583]. terabit/s [583]. Term [363, 1102, 1899]. Terminal [64, 101, 141, 222, 250, 280, 689, 1823].

Terminal-Pair [1823]. Terminals [350, 1994]. terrain [1049]. Terrestrial [2000]. Test [173, 236, 264, 641, 694, 739, 741, 933, 934, 936, 1235, 1246, 1270, 1502]. Testbed [482, 1610, 1826]. Testing [45, 76, 117, 155, 233, 236, 303, 523, 690, 700, 907, 932, 934, 936, 1272]. Tests [382, 387, 935]. tether [675]. Text [122, 659, 1101, 1108, 1455, 1459, 1480, 1737, 1786]. textbooks [1048]. Textual [1108, 1459]. Texture [1539]. textures [1052]. Their [123, 253, 566, 830, 1252, 1790, 1862]. theorem [516]. theory [303, 1268]. there [462]. Thesaurus [1720]. Thesaurus-Assisted [1720]. things [1129]. Thinking [1459]. Thinternet [675]. Third [95, 502, 1759, 1946, 2047]. Third-Generation [1946]. THORN [204]. those [207]. threats [238]. Three [398, 929, 1616]. Threshold [536, 578, 938, 977, 1094, 1431]. threshold-based [938]. Thresholds [1970]. Throughput [22, 240, 385, 437, 579, 590, 835, 836, 939, 1355, 1356, 1868]. Throughput-Delay [1355]. ticker [1200]. TIDY [1203, 1706]. Tier [1409]. TIFF [1977]. Time [22, 25, 86, 284, 288, 300, 535, 632, 746, 787, 811, 837, 841, 890, 891, 899, 908, 924, 946, 947, 950, 988, 1082, 1085, 1088, 1092, 1140, 1263, 1342, 1359, 1397, 1412–1414, 1416, 1474, 1564, 1785, 1826, 1837, 1848, 1851, 1857, 1858, 1889– 1893, 1897, 1920, 1924, 1974, 1977, 2029, 2030]. time-dependent [1082]. time-scale-dependent [947]. time-sensitive [946]. Timed [936, 1867, 1901]. Timed-Round-Robin [1901]. Timeline [1415]. timely [1098]. Timeout [87]. Timer [527, 638, 1888]. Timer-Based [527, 1888]. Times [1644]. TINA [1909, 2009]. TINA-based [2009]. TINA-Like [1909]. Title [1636, 1738]. Title/Link [1636]. TkWWW [1451]. today [845]. TOGA [787, 1598].

TOGA-TAO [787, 1598]. together [341]. Token [4, 7, 8, 10, 55, 166, 323, 386, 421, 634, 1341, 1344, 1394, 1395, 1397, 1440, 1929]. Token-Based [1929]. Token-CSMA [4]. Token-CSMA/CD [4]. Token-Cycle [1397]. Token-Passing [166]. Token-Ring [1394]. Token-Skipping [8]. Tokens [1790]. Tolerance [1415]. Tolerant [645, 1263, 1820, 1906, 1949]. tongue [1310]. Too [245, 1199]. Tool [306, 494, 495, 681, 709, 726, 776, 777, 833, 871, 1015, 1097, 1122, 1126, 1242, 1454, 1492, 1501, 1528, 1591, 1630, 1722, 1723, 1804, 1810, 2023]. toolkit [498, 704, 1166]. Tools [39, 142, 360, 490, 496, 710, 909, 1055, 1168, 1266, 1291, 1300, 1471, 1495, 1628, 1702]. toolset [416, 1001]. Top [75]. topics [1182]. Topological [617, 692]. Topologies [49, 598, 603, 1218, 1937]. Topology [109, 113, 176, 211, 273, 1217, 1232, 1319, 1373, 1421, 1836, 2022]. Topology-Aware [1421]. topology-design [273]. TOPS [1479]. Toroidal [276]. Toroidal-based [276]. Total [1442]. tour [1107]. Touring [570, 1306]. tours [872]. Town [1507]. tradeoffs [828]. trader [958]. trading [420]. Traditional [1673]. Traffic [32, 34– 36, 38, 59, 124, 178, 288, 296, 297, 312, 328, 375, 405, 509, 529, 609, 613, 614, 643, 647, 648, 756, 771, 803, 808, 809, 829, 834, 835, 841, 891, 910, 911, 947, 950, 965, 968, 979, 1081, 1085, 1087– 1089, 1219, 1227, 1249, 1279– 1281, 1286, 1297, 1344, 1359, 1363, 1384, 1386, 1397, 1428, 1432, 1438, 1568, 1569, 1824, 1829, 1848, 1857, 1862, 1875, 1878, 1890, 1899, 1901, 1920, 1923, 1970, 1987, 2000, 2012, 2013, 2016]. traffics [430]. Trail [716]. Trail-blazing [716]. Trails [702, 1800]. Training [671, 1666, 1776, 1777]. Transaction [1070, 1779]. Transactions [1524]. Transceivers [1926]. Transducers [1802]. Transfer [13, 278, 402, 403, 406, 592, 987, 1389, 1692, 1850, 1991]. Transformation [1483, 1701, 1702]. Transforming [1041].

Transient [322]. transit [328]. Transition [194, 356, 502, 934, 1679]. Transitioning [664, 1537]. Translating [1784]. Translation [912]. Translator [1477]. Transmission [17, 112, 267, 533, 602, 767, 942, 971, 1072, 1312, 1870, 1892, 1914, 1977]. Transmissions [1911]. Transmitting [1891]. Transport [62, 75, 107, 224, 265, 275, 294, 442, 621, 970, 975, 1400, 1447, 1814, 1865, 1886, 1896, 1969]. TRASEX [150]. Tree [113, 217, 227, 234, 235, 240, 484, 529, 599, 603, 762, 1358, 1378, 1893, 1906]. Tree-forming [217]. trees [544, 1301]. Trends [145, 174, 184, 216, 795, 917, 1071, 1221, 1302, 1425]. Trial [1348, 1350, 1829]. Trip [1897]. Troubleshooting [385]. TRP [895]. Truncation [1967]. Trunk [2002]. Trust [991, 1175]. TSE [895]. TTCM [382]. TTCN [933]. Tunable [2006]. Tunneling [687, 1145, 1992]. Tutorial [92, 207, 243, 354, 398, 399, 403, 504, 505, 904]. TVNet [164]. TWDM [1827, 1926]. Twenty [1243]. Twenty-One [1243]. Two [65, 66, 373, 521, 529, 606, 764, 968, 976, 1056, 1426, 1695, 1855, 1861, 2011]. Two-Layer [2011]. two-level [968]. Two-Node [1861]. Two-Stage [1426]. Two-Way [66, 529]. Type [110, 1100]. Types [1616].

U.S. [106, 337]. UARC [1504]. Ubiquitous [894]. UBR [1966]. Udine [573]. UIUC [1513]. UK [249, 451, 887, 1028]. UL [1978]. UL-CAC [1978]. Ultra [1978]. Ultrafast [608]. unapproachable [673]. uncertainty [129]. undefined [1182]. Understanding [1291, 1530]. Unfairness [1409]. UNI [1913]. Unidirectional [166]. Unified [27, 722, 1574]. Uniform [769, 866, 1556]. UniGuide [1149]. Unit [391, 463, 1921]. United [308, 461, 1612]. Uniting [1611]. Units [1436]. Universal [573, 822, 1068, 1213, 1787, 1924]. Universe [1698]. universities [1149]. University [175, 179, 346, 461, 929, 1156, 1195, 1305, 1350, 1458, 1621, 1674, 1736, 1740, 1817]. Unix [445, 710]. unpredictable [394]. unspecified [916]. unsupervised [1063]. Untangling [1480]. UPC [1411]. Update [206, 571, 1126]. updates [944]. Upgrading [832]. Upper [1504]. URI [1573]. URIs [815, 1127]. URL [1110]. URLs [1005]. US-Japan [1660]. USA [193]. Usability [1039, 1040, 1651]. Usage [477, 794, 1026, 1273, 1523, 1904]. Usage-Based [794, 1523, 1904]. Use [288, 336, 441, 490, 593, 798, 838, 1068, 1167, 1179, 1193, 1231, 1334, 1474, 1569, 1591, 1599, 1601, 1612, 1654, 1690, 1715, 1768, 1863]. Use-it [1334]. Useful [1614, 1627]. usefulness [1167]. User [24, 32, 98, 101, 178, 436, 553, 662, 709, 786, 813, 816, 852, 912, 1000, 1036, 1098, 1103, 1109, 1142, 1176, 1196, 1275, 1318, 1419, 1558, 1589, 1650, 1651, 1671, 1694, 1734, 1759, 1915, 1931]. User-centered [1098]. User-Network [178]. user-responsible [1000]. Users [84, 123, 186, 205, 434, 435, 554, 1067, 1270, 1637, 1645, 1664]. Uses [715, 1621, 1740]. Using [61, 71, 72, 82, 164, 166, 172, 214, 275, 278, 306, 357, 379, 385, 487, 517, 551, 560, 574, 630, 665, 694, 709, 717, 719, 782, 791, 805, 860, 865, 915, 981, 1021, 1048, 1065, 1066, 1077, 1138, 1140, 1148, 1167, 1173, 1220, 1265, 1290, 1301, 1303, 1308, 1313, 1321, 1325, 1336, 1338, 1345, 1369, 1373, 1385, 1392, 1401, 1444, 1489, 1502, 1525, 1534, 1544, 1549, 1575, 1584, 1595, 1603, 1617, 1629, 1685, 1719, 1752, 1761, 1773, 1789, 1795, 1800, 1843, 1852, 1860, 1871, 1885, 1911, 1926, 1937, 1944, 1951, 1967, 1970, 1992, 2011, 2015, 2021]. Utility [1113, 1706]. Utilizing [1458, 1527, 2009].

Validation [19, 39, 57, 58, 398, 400, 505, 906, 933, 952]. validity [833]. Value [30, 1205, 1585]. Value-Added [30]. Variability [1599]. Variable [297, 319, 588, 805, 1081, 1085, 1090, 1934]. variable-length [588]. variance [320]. Variations [1204]. various [1088]. vBNS [1201]. VBR [297, 1890, 1908, 2000, 2013, 2031]. VCR [900]. VCR-like [900]. VD [1289]. Vector [1576]. Vendor [685]. Verification [417, 493, 504, 527, 1371]. verifying [492]. Versatile [687, 1783]. Version [1127, 1706]. Versioning [1761]. Versus [1673]. Vertically [1952]. VESAR [493]. Via [309, 684, 706, 783, 1001, 1225, 1306, 1426, 1494, 1503, 1533, 1632, 1718, 1727]. Vice [1507]. Video [60, 313, 499, 615, 747, 811, 900, 985, 1072, 1083, 1085, 1136, 1183, 1187, 1256, 1304, 1311, 1343, 1377, 1437, 1785, 1853, 1862, 1865, 1870, 1879, 1880, 1891, 1912, 1931, 1951, 1975, 1988, 2013]. Video-on-Demand [1880, 1912]. Videoconferencing [1436]. VideoTalk [983]. videotex [125, 219]. View [80, 98, 101, 106, 197–199, 618, 735, 1307, 1972]. Viewed [1433]. Viewing [1025]. views [816]. Vince [685]. Viper [1166]. Virginia [1607]. Virtual [22, 67, 101, 141, 222, 350, 409, 487, 676, 718, 783, 803, 804, 898, 913, 953, 1055, 1058, 1064, 1074, 1099, 1195, 1199, 1279, 1289, 1293, 1296, 1297, 1299, 1329, 1494, 1517, 1590, 1619, 1752, 1754, 1756, 1887, 1950, 1992]. Virus [758, 1632]. viruses [238]. visual [1055, 1109]. Visualising [1064]. Visualization [310, 491, 778, 891, 1045, 1155, 1166, 1309, 1319, 1481, 1575, 1600, 1750, 1753]. visualize [1290]. Visualizing [735, 1019, 1100, 1171]. VLNET [1050]. VLSI [1824]. Voice [11, 25, 53, 60, 154, 162, 214, 263, 318, 373, 418, 628, 632, 756, 759, 951, 1343, 1352, 1448, 2031]. voice-call [632]. Voice-Data [11, 418, 759]. Voice/Data [25, 214, 263, 373]. Volatile [1467]. VRML [1099, 1290, 1303]. vs [1289, 1678]. VS/VD [1289]. Vulnerabilities [1837]. VuNet [1349].

W3 [722, 773, 1552, 1605, 1678]. W3-access [722]. W3C [1210]. W3Gate [930]. W3Kit [1753]. W3Objects [874, 1765]. WAIS [1489, 1531, 1724]. WAN [185, 355, 372, 562, 1353, 1420, 1817, 1835]. Want [205, 1525, 1717]. waste [943]. waste-free [943]. watching [1027]. Waterloo [175]. WATERS [1634]. wave [1180]. Wavelength [210, 274, 603, 604, 607, 1077, 1338, 1825]. Wavelength-based [607]. wavelength-division [604]. wavelength-routed [1338]. wavelengths [1338]. WAXweb [775]. Way [66, 225, 529, 1006, 1111, 1265]. WCP [1126]. WDM [601, 836, 1218, 1372, 1373, 1825, 1929, 1984]. Web [452, 569, 570, 655, 656, 707, 717, 721, 727, 731, 734, 737, 771, 791, 793, 844, 845, 847, 851, 861, 867, 876, 877, 886, 930, 981, 991, 996, 998, 1002, 1010, 1019, 1023, 1028, 1037, 1040, 1042, 1064, 1103, 1105, 1122, 1131, 1139, 1141– 1143, 1145, 1154, 1155, 1158, 1164, 1166, 1210– 1212, 1214, 1261, 1283, 1292, 1315, 1473, 1506, 1541, 1549, 1554, 1555, 1568, 1590, 1593– 1595, 1612, 1661, 1671, 1684, 1686, 1694, 1696, 1707, 1710, 1724, 1728, 1735, 1755, 1783, 1785, 1786, 1791, 1794, 1795, 1811, 2042, 2045, 2048, 2050, 2051, 2053]. Web [652, 654, 657, 666, 668, 672, 704, 715, 718, 733, 778, 780, 782, 789, 814, 864, 871, 872, 880, 891, 899, 985, 986, 988– 990, 1004, 1016, 1021, 1027, 1032, 1034, 1044, 1053, 1095, 1096, 1107, 1109, 1113, 1119, 1124, 1125, 1149, 1167–1169, 1171, 1173– 1175, 1179, 1326, 1449, 1450, 1462, 1463, 1467, 1468, 1480, 1481, 1493, 1503, 1516, 1520, 1521, 1533, 1542, 1546, 1556, 1613, 1617, 1640, 1641, 1643, 1656, 1670, 1675, 1682, 1697, 1701, 1704, 1705, 1708, 1719, 1735, 1737, 1739, 1760, 1765, 1766, 1770, 1771, 1781, 1787, 1790, 1805, 2047]. Web [661, 662, 664, 665, 667, 709, 711, 714, 716, 723, 726, 735, 768, 769, 783, 786, 843, 850, 856, 857, 860, 883, 887, 889, 893, 983, 995, 1011, 1012, 1014, 1018, 1038, 1041, 1045, 1050, 1051, 1055, 1106, 1130, 1133, 1146, 1184–1188, 1194, 1203, 1205, 1207, 1215, 1245, 1305, 1312, 1323, 1454, 1456, 1470, 1474, 1475, 1484, 1485, 1489, 1491, 1494, 1512, 1515, 1524, 1525, 1537, 1564, 1565, 1584, 1597, 1615, 1632, 1650, 1652, 1658, 1663, 1669, 1681, 1691, 1692, 1698, 1720, 1747, 1750, 1753, 1757, 1764, 1774, 1782, 1798, 1803, 1813]. Web [660, 710, 738, 792, 849, 870, 992, 1009, 1066, 1190, 1500, 1536, 1679, 1709, 1712, 1762, 1769, 1789]. Web-based [726, 1011, 1122, 1166, 1169, 1682, 1760, 1798]. web2mush [1554]. WebAssist [1176]. WebCanal [1004]. Webcasting [1137]. WebComposition [1034]. WebCrawler [1717]. WebCutter [1020]. WebExplorer [1757]. WebGALAXY [1030]. WebGUIDE [880]. WebL [1119]. Webland [1463]. Weblet [1627]. WebLinker [777, 1722]. Weblint [876]. WebMake [707]. WebMaker [658]. WebMap [776, 1723]. Webmaster [1563]. WebOFDAV [1171]. WebOOGL [1753]. WebQuery [1019]. WebQuest [878]. WebRule [998]. Webs [1513]. Webspace [1064]. WebWriter [877, 1042]. Weight [1978]. weighted [526]. welcome [657]. Welle [868]. West [920, 1348, 1607]. Where [1521]. which [710]. White [626, 817, 1812]. Who [144]. Wide [272, 396, 647, 654, 662, 672, 769, 791– 793, 849, 857, 864, 872, 876, 886, 889, 891, 985, 1002, 1018, 1028, 1040, 1045, 1051, 1107, 1149, 1154, 1155, 1167, 1175, 1178, 1190, 1214, 1292, 1305, 1312, 1417, 1448, 1456, 1462, 1467, 1473, 1474, 1489, 1507, 1511, 1513– 1515, 1520, 1534, 1542, 1546, 1556, 1584, 1594, 1595, 1597, 1617, 1632, 1658, 1663, 1671, 1675, 1681, 1684, 1691, 1692, 1696, 1698, 1709, 1710, 1720, 1728, 1735, 1739, 1750, 1753, 1762, 1766, 1769, 1776, 1785, 1786, 1790, 1791, 1795, 1811, 1813, 1826, 2042, 2045, 2048, 2050, 2051, 2053]. wide [441, 452, 532, 542, 652, 656, 661, 663, 665, 668, 697, 711, 714, 715, 717, 733– 735, 738, 768, 814, 839, 871, 896, 978, 1012, 1024, 1038, 1053, 1094, 1330, 1470, 1484, 1485, 1516, 1555, 1634, 1694, 2047]. Wide-Area [532, 647, 1094, 1330, 1507, 1511, 1584, 1597]. widgets [1300]. Wild [673]. Will [1210, 1211, 1569]. WILT [1630]. WiN [927]. Window [319, 350, 532, 842, 1660]. Windowed [683]. Windowing [222, 1233]. windows [486, 1488]. winner [368]. Wired [1497]. Wireless [563, 797, 1086, 1093, 1392, 1393, 1932, 1945, 1958, 1959, 1993, 1996, 2004]. Wires [14]. WISCONSIN [1324, 1713]. Wise [58, 1052]. Within [113, 131, 457, 543, 1089, 1584, 1814, 2023]. without [1174]. Word [1725]. Words [1472]. Work [130, 341, 348, 445, 712, 929, 1466, 1628, 1693, 1940]. workbench [497]. Workflow [895, 928]. workgroup [811]. Working [159, 1294]. workload [1157]. Workshop [669, 1547]. Workspace [1763]. Workstation [334, 350, 1400, 1404, 1979]. Workstations [250, 332]. World [428, 452, 466, 652, 654, 656, 662, 663, 665, 668, 672, 711, 714, 715, 717, 733– 735, 769, 789, 791, 793, 814, 864, 871, 872, 876, 886, 891, 896, 985, 1002, 1012, 1018, 1024, 1040, 1053, 1107, 1149, 1154, 1155, 1167, 1175, 1214, 1292, 1293, 1296, 1305, 1311, 1456, 1462, 1467, 1473, 1485, 1489, 1496, 1503, 1516, 1520, 1542, 1546, 1555, 1556, 1594, 1595, 1597, 1601, 1607, 1617, 1618, 1671, 1675, 1684, 1692, 1696, 1698, 1710, 1728, 1735, 1739, 1750, 1753, 1766, 1785, 1786, 1790, 1791, 1795, 1811, 1813, 2042, 2045, 2047, 2048, 2050, 2051, 2053]. World [661, 738, 768, 792, 849, 857, 889, 1038, 1045, 1051, 1190, 1312, 1470, 1474, 1484, 1515, 1584, 1632, 1658, 1663, 1681, 1691, 1709, 1720, 1762, 1769]. World-Wide [452, 652, 656, 661– 663, 665, 668, 711, 714, 715, 717, 733– 735, 738, 768, 769, 814, 896, 1012, 1038, 1053, 1470, 1484, 1485, 1516, 1555, 2045, 2047]. World-Wide-Web [1051, 1542]. worlds [1050]. worldwide [285, 1697]. Worth [1725]. WPS [574]. wrapped [1338]. wrapped-around [1338]. write [125].

Writing [658]. WSDM [1103].

Wudangshan [1658]. WWW [658, 704– 706, 708, 709, 712, 719, 729, 772, 774, 775, 777, 784, 788, 846, 871, 881, 884, 890, 894, 900, 918, 1001, 1025, 1026, 1111, 1114, 1123, 1126, 1135, 1147, 1181, 1202, 1209, 1306, 1328, 1331, 1453, 1483, 1492, 1501, 1505, 1510, 1514, 1518, 1522, 1526, 1527, 1534, 1535, 1540, 1544, 1557, 1559– 1561, 1566, 1576, 1577, 1583, 1586, 1602–1604, 1611, 1616, 1620, 1629, 1630, 1633, 1638, 1644, 1660, 1667, 1680, 1688, 1693, 1718, 1722, 1730, 1734, 1745, 1749, 1751, 1759, 1763, 1772, 1777– 1779, 1792, 1797, 1803, 1804, 1806, 1809, 2048]. WWW [574, 873, 1035, 1192, 1519, 1587, 1601, 1623, 1761]. WWW-Based [708, 788, 871, 918, 1522, 1540, 1779]. WWW-SPEED [1577]. WWW/Mosaic [1620]. WWW94 [669]. WWWinda [1453].

X [350, 1657]. X-Mosaic [1657]. X.25 [225, 437, 635, 638]. X.25/OSI [363]. X.400 [97, 207, 242, 349, 477, 551, 555, 1991]. X.500 [208, 298, 352, 361, 477, 554, 574, 626, 630, 699, 817, 925]. XML [1096, 1116, 1117, 1197, 1215]. XML/HTML [1117]. Xpress [1389, 1896]. XTND [1477]. XTP [547, 1841, 1845, 1869]. XTV [713].

Y-NET [469]. Yanomamo [1658]. year [1027, 1198]. years [331, 795, 917]. Yellow [925]. York [1740]. Yuri [1547].

Z.105 [903]. Zealand [888]. Zurich [540].

References

Turoff:1985:EIM

[1] Murray Turoff and Sanjit Chinai. Electronic information marketplace. Computer Networks and ISDN Systems, 9 (2):79–90, February 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Mamrak:1985:SSD

[2] Sandra A. Mamrak and Dennis W. Leinbaugh. Software support for distributed resource sharing. Computer Networks and ISDN Systems, 9(2): 91–107, February 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Sloman:1985:RAI

[3] M. S. Sloman and X. Andriopoulos. Routing algorithm for interconnected local area networks. Computer Networks and ISDN Systems, 9 (2):109–130, February 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gopal:1985:AHT

[4] P. M. Gopal and J. W. Wong. Analysis of a hybrid token-CSMA/CD protocol for bus networks. Computer Networks and ISDN Systems, 9(2): 131–141, February 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Anonymous:1985:NCL

[5] Anonymous. Ninth conference on local computer networks. Computer Networks and ISDN Systems, 9(3):175– 240, March 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chlamtac:1985:SCI

[6] I. Chlamtac and O. Ganz. Study of communication interfaces for multiple bus networks. Computer Networks and ISDN Systems, 9(3):177–189, March 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Sethi:1985:PAT

[7] Adarshpal S. Sethi and Tuncay Saydam. Performance analysis of token ring local area networks. Computer Networks and ISDN Systems, 9 (3):191–200, March 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Jayasumana:1985:TCA

[8] Anura P. Jayasumana and P. David Fisher. Token-skipping channel access scheme for bus networks. Computer Networks and ISDN Systems, 9 (3):201–208, March 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Molloy:1985:CRC

[9] Michael K. Molloy. Collision resolution on the CSMA/CD bus. Computer Networks and ISDN Systems, 9 (3):209–214, March 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Ibe:1985:HHB

[10] Oliver C. Ibe and Pin-Yee Chen. Hybrinet: a hybrid bus and token ring network. Computer Networks and ISDN Systems, 9(3):215–221, March 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Goel:1985:HFP

[11] R. K. Goel and A. K. Elhakeem. Hybrid FARA/CSMA-CD protocol for voice-data integration. Computer Networks and ISDN Systems, 9(3):223– 240, March 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Tcha:1985:SPP

[12] Dong wan Tcha and Kiyoshi Maruyama. On the selection of primary paths for a communication network. Computer Networks and ISDN Systems, 9(4):257–265, April 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Petersen:1985:AIF

[13] Niels K. Petersen and Tom Skovgaard. Anticipating the ISO file transfer standards in an Open Systems Implementation. Computer Networks and ISDN Systems, 9(4):267–280, April 1985. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Estrin:1985:INS

[14] Deborah L. Estrin. Inter-organizational networks: Stringing wires across administrative boundaries. Computer Networks and ISDN Systems, 9(4):281– 295, April 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Perlman:1985:HNS

[15] Radia Perlman. Hierarchical networks and the subnetwork partition problem. Computer Networks and ISDN Systems, 9(4):297–303, April 1985. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bence:1985:TDC

[16] Antony T. Bence. Teleservice developments in CCITT — migration to the ISDN. Computer Networks and ISDN Systems, 9(5):329–337, May 1985. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

McLintock:1985:OIS

[17] Robert W. McLintock. Overview of international standards for transmission impairments affecting digital telecommunications networks. Computer Networks and ISDN Systems, 9(5):339–344, May 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Sorensen:1985:CRP

[18] Soren-Aksel Sorensen. Cambridge Ring performance. Computer Networks and ISDN Systems, 9(5):345–352, May 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gouda:1985:PVF

[19] Mohamed G. Gouda and Ji-Yun Han. Protocol validation by fair progress state exploration. Computer Networks and ISDN Systems, 9(5):353–361, May 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Hirota:1985:CCJ

[20] Ken’ichiro Hirota. Computer communication and Japan’s general public. Computer Networks and ISDN Systems, 10(1):1–5, August 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Fernow:1985:SAC

[21] James P. Fernow and Mohamed El-Sayed. Stability of adaptive congestion controls in packet networks. Computer Networks and ISDN Systems, 10(1):7–18, August 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Meditch:1985:STV

[22] J. S. Meditch and C. T. A. Lea. Stability and throughput in virtual time csma. Computer Networks and ISDN Systems, 10(1):19–26, August 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Karger:1985:ADA

[23] Paul A. Karger. Authentication and discretionary access control in computer networks. Computer Networks and ISDN Systems, 10(1):27–37, August 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Rose:1985:MMU

[24] Marshall T. Rose, Einar A. Stefferud, and Jerry N. Sweet. Mh: a multifarious user agent. Computer Networks and ISDN Systems, 10(2):65–80, September 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chlamtac:1985:ECP

[25] Imrich Chlamtac. Ethernet compatible protocol for real-time voice/data integration. Computer Networks and ISDN Systems, 10(2):81–96, September 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Ebihara:1985:FDA

[26] Yoshihiko Ebihara, Katsuo Ikeda, Tomoo Nakamura, Shigeo Nakatuska, and Michihiro Ishizaka. Fault diagnosis and automatic reconfiguration for a ring subsystem. Computer Networks and ISDN Systems, 10(2):97–109, September 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Onozato:1985:UAS

[27] Yoshikuni Onozato. Unified analysis of steady state behavior in random access schemes’. Computer Networks and ISDN Systems, 10(2):111–122, September 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Anonymous:1985:ITC

[28] Anonymous. 11th International Teletraffic Congress. Computer Networks and ISDN Systems, 10(3–4):135–229, October–November 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kitahara:1985:AIS

[29] Yasusada Kitahara. Advanced information society and telecommunications. Computer Networks and ISDN Systems, 10(3–4):141–145, October– November 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gimpelson:1985:IVA

[30] Lester A. Gimpelson. ISDN and value-added services in public and private networks. Computer Networks and ISDN Systems, 10(3–4):147–156, October–November 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Potter:1985:IPA

[31] R. M. Potter. ISDN protocol and architecture models. Computer Networks and ISDN Systems, 10(3–4):157– 166, October/November 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Soto:1985:SIP

[32] Oscar Gonzalez Soto and Teodoro Borja. Step to ISDN planning: from user traffic to service cost comparison. Computer Networks and ISDN Systems, 10(3–4):167–174, October– November 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Burattin:1985:CPP

[33] A. Burattin, U. Mazzei, and C. Mazzetti. Computer-aided planning of public packet-switched networks. Computer Networks and ISDN Systems, 10(3– 4):175–183, October–November 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Marshall:1985:SMD

[34] W. T. Marshall and S. P. Morgan. Statistics of mixed data traffic on a local area network. Computer Networks and ISDN Systems, 10(3–4):185– 194, October–November 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Also in Eleventh International Teletraffic Conference Proceedings, M. Akiyama, ed., Elsevier Science Publishers B. V. (North-Holland), IAC, September 1985, pp. 4.1A.1.1-7.

Hofstetter:1985:TML

[35] Helga Hofstetter and Dietmar Weber. Traffic models for large ISDN-PABX’s. Computer Networks and ISDN Systems, 10(3–4):195–202, October–November 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Murata:1985:TCC

[36] Hatsuho Murata. Traffic characteristics of control signals in an ISDN switching system. Computer Networks and ISDN Systems, 10(3–4):203–210, October–November 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Goldner:1985:ICP

[37] Ernst-Heinrich H. Göldner. Integrated circuit/packet switching local area network — performance analysis and comparison of strategies. Computer Networks and ISDN Systems, 10(3–4):211– 219, October–November 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Roberts:1985:TMT

[38] James Roberts and Keqiang Liao. Traffic models for telecommunication services with advance capacity reservation. Computer Networks and ISDN Systems, 10(3–4):221–229, October– November 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Groz:1985:ACD

[39] Roland Groz, Claude Jard, and Claire Lassudrie. Attacking a complex distributed algorithm from different sides: an experience with complementary validation tools. Computer Networks and ISDN Systems, 10(5):245–257, December 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Berntsen:1985:MLI

[40] Janet A. Berntsen, James R. Davin, Daniel A. Pitt, and Neil G. Sullivan. MAC layer interconnection of IEEE 802 local area networks. Computer Networks and ISDN Systems, 10 (5):259–273, December 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Lagarias:1985:CDS

[41] J. C. Lagarias, A. M. Odlyzko, and D. B. Zagier. On the capacity of disjointly shared networks. Computer Networks and ISDN Systems, 10(5): 275–285, December 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chlamtac:1985:OEP

[42] I. Chlamtac and W. R. Franta. Optimal election protocol for channel allocation. Computer Networks and ISDN Systems, 10(5):287–305, December 1985. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Anonymous:1986:IEC

[43] Anonymous. ISDN Europe 86 conference report. Computer Networks and ISDN Systems, ??(??):183–189, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Enslow:1986:SII

[44] P. H. Enslow, Jr. Special issue: ICCC ’86. Computer Networks and ISDN Systems, ??(??):207–275, November 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Linn:1986:TAI

[45] R. J. Linn. Testing to assure interworking of implementations of ISO/OSI protocols. Computer Networks and ISDN Systems, 11(??):277–286, 1986. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Patel:1986:DSP

[46] A. Patel and M. Purser. The design of a small packet-switching node. Computer Networks and ISDN Systems, 11(??):287–295, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Hutchinson:1986:FSN

[47] Norm Hutchinson, Terry Patten, and Brian Unger. Flooding sink — a new approach to local area networking. Computer Networks and ISDN Systems, 11(1):1–14, January 1986. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Tsuchiya:1986:SDP

[48] M. Tsuchiya and Douglas C. Shannon. Simulation of distributed processing networks. Computer Networks and ISDN Systems, 11(1):15–27, January 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Raghavendra:1986:SML

[49] C. S. Raghavendra and J. A. Silvester. Survey of multi-connected loop topologies for local computer networks. Computer Networks and ISDN Systems, 11 (1):29–42, January 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Varakulsiripunth:1986:CPI

[50] Ruttikorn Varakulsiripunth, Norio Shiratori, and Shoichi Noguchi. Congestion-control policy on the internetwork gateway. Computer Networks and ISDN Systems, 11(1):43–58, January 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Anonymous:1986:TCL

[51] Anonymous. Tenth Conference on Local Computer Networks. Computer Networks and ISDN Systems, 11 (2):87–146, February 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Ibe:1986:SCP

[52] Oliver C. Ibe. Synchronous communication protocols for local networks. Computer Networks and ISDN Systems, 11(2):89–97, February 1986. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Frost:1986:MVC

[53] Victor S. Frost, Edward M. Friedman, and Gary J. Minden. Multirate voice coding for load control on CSMA/ CD local computer networks. Computer Networks and ISDN Systems, 11 (2):99–110, February 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Wang:1986:CCL

[54] Paul T. R. Wang and Michael F. McGurrin. CSMA/CD LANs with overload detection and protection. Computer Networks and ISDN Systems, 11(2):111–120, February 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Johnson:1986:RMF

[55] Marjory J. Johnson. Reliability mechanisms of the FDDI high bandwidth token ring protocol. Computer Networks and ISDN Systems, 11(2):121– 131, February 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Loucks:1986:IDA

[56] W. M. Loucks, W. I. Kwak, and Z. G. Vranesic. Implementation of a dynamic address assignment protocol in a local area network. Computer Networks and ISDN Systems, 11 (2):133–146, February 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

West:1986:VOS

[57] Colin H. West. Validation of the OSI Session Layer Protocol. Computer Networks and ISDN Systems, 11 (3):173–182, March 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Ural:1986:SVC

[58] Hasan Ural and Robert L. Probert. Step-wise validation of communication protocols and services. Computer Networks and ISDN Systems, 11 (3):183–202, March 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chemouil:1986:ACT

[59] Prosper Chemouil, Janusz Filipiak, and Paul Gauthier. Analysis and control of traffic routing in circuit-switched networks. Computer Networks and ISDN Systems, 11(3):203–217, March 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Elhakeem:1986:AHD

[60] A. K. Elhakeem, B. Grela-M’Poko, Mustafa Ali, and J. F. Hayes. Analysis of a hybrid (demand assignment tdma) protocol for video teleconferencing — voice data optical networks. Computer Networks and ISDN Systems, 11 (3):219–241, March 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chang:1986:ODP

[61] Chung-Ju Chang and Jin-Fu Chang. Optimal design parameters in a multihop packet radio network using random access techniques. Computer Networks and ISDN Systems, 11(5):337– 351, May 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chong:1986:SDI

[62] H. Y. Chong. Software development and implementation of NBS Class-4 Transport Protocol. Computer Networks and ISDN Systems, 11(5):353– 365, May 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Choi:1986:PAS

[63] Tat Y. Choi and Raymond E. Miller. Protocol analysis and synthesis by structured partitions. Computer Networks and ISDN Systems, 11(5):367– 381, May 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Srinivasan:1986:SFS

[64] B. Srinivasan and A. L. Ananda. Simple file server for a terminal support network environment. Computer Networks and ISDN Systems, 11(5):383–389, May 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Cidon:1986:CSA

[65] Israel Cidon and Raphael Rom. Carrier sense access in an environment of two interfering channels. Computer Networks and ISDN Systems, 12(1):1–10, August 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Saito:1986:LDT

[66] Tadao Saito, Kazutaka Murakami, and Hiroshi Inose. Local digital telephone system on two-way cable television. Computer Networks and ISDN Systems, 12(1):11–25, August 1986. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Baratz:1986:EVC

[67] Alan E. Baratz and Jeffrey M. Jaffe. Establishing virtual circuits in large computer networks. Computer Networks and ISDN Systems, 12(1):27–37, August 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Tobagi:1986:PEC

[68] Fouad A. Tobagi and David H. Shur. Performance evaluation of channel access schemes in multihop packet radio networks with regular structure by simulation. Computer Networks and ISDN Systems, 12(1):39–60, August 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kummer:1986:PLS

[69] P. Kummer, R. Tasker, N. Linge, and E. Ball. A protocol-less scheme for bridging between IEEE 802 local area networks. Computer Networks and ISDN Systems, 12(2):81–87, 1986. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Tropper:1986:PPP

[70] Carl Tropper. Priorities and performance in packet-switching networks. Computer Networks and ISDN Systems, 12(2):89–98, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Todd:1986:ACP

[71] Terence D. Todd. An access control protocol using instantaneous bandwidth reconfiguration. Computer Networks and ISDN Systems, 12(2):99– 106, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Ammar:1986:TID

[72] Mostafa H. Ammar. Teletext-like information delivery using broadcast polling. Computer Networks and ISDN Systems, 12(2):107–115, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chardaire:1986:GSO

[73] Pierre Chardaire and Marc Lesk. Grade of service and optimization of distributed packet-switched networks. Computer Networks and ISDN Systems, 12(3):139–146, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Jaffe:1986:SDI

[74] Jeffrey M. Jaffe, Alan E. Baratz, and A. Segall. Subtle design issues in the implementation of distributed, dynamic routing algorithms. Computer Networks and ISDN Systems, 12(3): 147–158, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Rose:1986:OTS

[75] Marshall T. Rose and Dwight E. Cass. OSI transport services on top of the TCP. Computer Networks and ISDN Systems, 12(3):159–173, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Also available as NRTC Technical Paper #700.

Togawa:1986:NTS

[76] Takashi Togawa, Mineo Nishiwaki, and Ken-Ichiro Yoshida. Network testing system for digital data networks. Computer Networks and ISDN Systems, 12(3):175–182, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Schwarz-Schilling:1986:GDP

[77] Christian Schwarz-Schilling. Guidelines for the development of public networks and services in the Federal Republic of Germany. Computer Networks and ISDN Systems, 12(4):209– 215, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Anonymous:1986:ICC

[78] Anonymous. 8th International Conference on Computer Communication. Computer Networks and ISDN Systems, 12(4):209–238, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Uhlig:1986:ICC

[79] Ronald P. Uhlig. IFIP and computer communication. Computer Networks and ISDN Systems, 12(4):225– 229, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Rankine:1986:VIO

[80] L. J. Rankine. View from the International Organization for Standardization (ISO). Computer Networks and ISDN Systems, 12(4):231–233, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gassmann:1986:CCE

[81] Hans-Peter Gassmann. Computer communications: the economic challenge. Computer Networks and ISDN Systems, 12(4):235–238, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Yokohira:1986:ADB

[82] Tokumi Yokohira, Takeshi Nishida, and Hideo Miyahara. Analysis of dynamic behavior in P-persistent CSMA/ CD using cusp catastrophe. Computer Networks and ISDN Systems, 12(5): 277–289, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Lai:1986:AID

[83] Wai Sum Lai. Analysis of the integrity of data delivery in computer networks. Computer Networks and ISDN Systems, 12(5):291–310, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chang:1986:CRB

[84] Jin-Fu Chang and Jiun Wang. Collision resolution for buffered users. Computer Networks and ISDN Systems, 12(5):311–327, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Onozato:1986:EPD

[85] Yoshikuni Onozato, Jin Liu, Shigeru Shimamoto, and Shoichi Noguchi. Effect of propagation delays on Aloha systems. Computer Networks and ISDN Systems, 12(5):329–337, 1986. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Ramamritham:1987:CCL

[86] Krithi Ramamritham. Channel characteristics in local-area hard real-time systems. Computer Networks and ISDN Systems, 13(1):3–13, 1987. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gerr:1987:TDR

[87] Neil L. Gerr. Timeout distribution for reduced contention in a decentralized random access channel. Computer Networks and ISDN Systems, 13(1):15–19, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

OReilly:1987:BFP

[88] Peter O’Reilly. Burst and fast packet switching: Performance comparisons. Computer Networks and ISDN Systems, 13(1):21–32, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Also published in Proc. INFOCOM’86, pp. 653–666.

Saracco:1987:CSO

[89] R. Saracco and P. A. J. Tilanus. CCITT SDL: Overview of the language and its applications. Computer Networks and ISDN Systems, 13(2):65–74, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bruijning:1987:EIS

[90] J. Bruijning and SPECS Consortium. Evaluation and integration of specification languages. Computer Networks and ISDN Systems, 13(2):75–89, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kossmann:1987:GSS

[91] Heinz Kossmann. A graphic SDL support environment. Computer Networks and ISDN Systems, 13(2):91–96, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Moller-Pedersen:1987:RTO

[92] Birger Møller-Pedersen, Dag Belsnes, and Hans-Petter Dahle. Rationale and tutorial on OSDL: An object-oriented extension of SDL. Computer Networks and ISDN Systems, 13(2):97–117, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Wakahara:1987:PST

[93] Yasushi Wakahara and Atsushi Ito. Prototyping system for telecommunications software based on abstract execution of requirements specifications. Computer Networks and ISDN Systems, 13(2):119–128, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Tucker:1987:NAI

[94] Jeremy Tucker. Naming and addressing issues — OSI network addresses. Computer Networks and ISDN Systems, 13(3):149–153, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Anonymous:1987:RTE

[95] Anonymous. Rare’s Third European Networkshop. Computer Networks and ISDN Systems, 13(3):149–249, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bovenga:1987:MNS

[96] Piet Bovenga. Management of network services. Computer Networks and ISDN Systems, 13(3):155–160, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Henken:1987:MXR

[97] Gerrit Henken. Mapping of X.400 and RFC822 addresses. Computer Networks and ISDN Systems, 13(3):161– 164, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Anonymous:1987:CLU

[98] Anonymous. Clearing the lines — a user’s view on business communications in Europe. Computer Networks and ISDN Systems, 13(3):165– 167, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Pihl:1987:GSS

[99] Jorgen Pihl. Global system for seismic data exchange. Computer Networks and ISDN Systems, 13(3):217– ??, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Topsoe:1987:EID

[100] Flemming Topsoe. Euromath — the integrated database and communications system for European mathematicians. Computer Networks and ISDN Systems, 13(3):219–??, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gilmore:1987:UVV

[101] Brian Gilmore. User view of virtual terminal standardisation. Computer Networks and ISDN Systems, 13(3):229– 233, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Rihaczek:1987:T

[102] Karl Rihaczek. Teletrust. Computer Networks and ISDN Systems, 13(3): 235–239, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bieber:1987:PEP

[103] Jacques Bieber. Pace — a European Programme of Advanced Continuing Education. Computer Networks and ISDN Systems, 13(3):241–243, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Pascal:1987:SP

[104] Olivier Pascal. Star programme. Computer Networks and ISDN Systems, 13(3):245–??, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Pascal:1987:AP

[105] Olivier Pascal. Apollo Project. Computer Networks and ISDN Systems, 13 (3):247–??, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Merola:1987:VUR

[106] Sandy Merola. View of U.S. research networks. Computer Networks and ISDN Systems, 13(3):249–??, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Narayanan:1987:CTB

[107] N. A. Narayanan. Common transport based approach to network interconnection. Computer Networks and ISDN Systems, 13(4–5):257–267, 1987. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Hackbarth:1987:IIC

[108] K. D. Hackbarth. Implications of the ISDN-concept on telecommunication network planning. Computer Networks and ISDN Systems, 13(4–5):269–274, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Senior:1987:TML

[109] J. M. Senior, W. M. Walker, and A. Ryley. Topology and MAC layer access protocol investigation for industrial optical fibre LANs. Computer Networks and ISDN Systems, 13(4–5):275–289, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Ibe:1987:ICM

[110] Oliver C. Ibe. Integrated CATV-type metropolitan area network. Computer Networks and ISDN Systems, 13(4–5): 291–299, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Ahamad:1987:EAN

[111] Mustaque Ahamad. An efficient algorithm for name resolution in computer networks. Computer Networks and ISDN Systems, 13(4–5):301–311, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Pachl:1987:RRT

[112] J. Pachl and L. Casey. Robust ring transmission protocol. Computer Networks and ISDN Systems, 13(4–5):313– 321, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Linge:1987:BPC

[113] N. Linge, E. Ball, R. Tasker, and P. Kummer. Bridge protocol for creating a spanning tree topology within an IEEE 802 extended LAN environment. Computer Networks and ISDN Systems, 13(4–5):323–332, 1987. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Takagi:1987:OPS

[114] Hideaki Takagi and Masayuki Murata. Output processes of slotted persistent CSMA and CSMA/CD systems. Computer Networks and ISDN Systems, 13(4–5):333–346, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bolognesi:1987:IIS

[115] Tommaso Bolognesi and Ed Brinksma. Introduction to the ISO specification language LOTOS. Computer Networks and ISDN Systems, 14(1):25–59, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Nash:1987:FPL

[116] Simon C. Nash. Format and Protocol Language (FAPL). Computer Networks and ISDN Systems, 14(1):61–77, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Rayner:1987:OCT

[117] D. Rayner. OSI conformance testing. Computer Networks and ISDN Systems, 14(1):79–98, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Stevens:1987:DLD

[118] R. J. Stevens, R. T. Greenwood, G. Alvisi, and M. Deschammps. Design of a large distributed information system. Computer Networks and ISDN Systems, 14(2–5):123–135, 1987. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Anonymous:1987:CCD

[119] Anonymous. Conference on communication and data communication. Computer Networks and ISDN Systems, 14(2–5):123–395, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Palme:1987:DCC

[120] Jacob Palme. Distributed computer conferencing. Computer Networks and ISDN Systems, 14(2–5):137–145, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

SeverinsonEklundh:1987:EIF

[121] Kerstin Severinson Eklundh. Explicit and implicit feedback incomputermediated communication. Computer Networks and ISDN Systems, 14(2–5): 147–153, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Maeght:1987:SFE

[122] Phillppe E. Maeght and Anton W. Jolkovski. Selective facsimile: the electronic text medium of the 21st Century? Computer Networks and ISDN Systems, 14(2–5):155–158, 1987. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Saint-Remi:1987:DNT

[123] J. Saint-Remi. Data networks and their users. Computer Networks and ISDN Systems, 14(2–5):159–165, 1987. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Forestier:1987:CNT

[124] Marc Forestier, Catherine Berenbliit, and Corinne Mercier. Communication networks in a traffic management system. Computer Networks and ISDN Systems, 14(2–5):167–177, 1987. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Basevi:1987:SRV

[125] Enrica Basevi. Some remarks on videotex interaction. how to write for a new reader: a practical research.

Computer Networks and ISDN Systems, 14(2–5):179–185, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Maritorena:1987:PDL

[126] Christian Maritorena. Presentation of a description language for office tasks. Computer Networks and ISDN Systems, 14(2–5):187–197, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kaivers:1987:CAC

[127] Reiner Kaivers. Communications analysis in the company. Computer Networks and ISDN Systems, 14(2–5):199– 205, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Claisse:1987:TQQ

[128] Gerard Claisse and Frantz Rowe. Telephone in question: questions on communication. Computer Networks and ISDN Systems, 14(2–5):207–219, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Cecez-Kecmanovic:1987:OKS

[129] Dubravka Cecez-Kecmanovic. Organizational knowledge systems: an investigation on knowledge communication uncertainty. Computer Networks and ISDN Systems, 14(2–5):221–230, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Lacoste:1987:CPW

[130] Michele Lacoste. Communication in the perspective of work analysis. Computer Networks and ISDN Systems, 14 (2–5):231–235, 1987. CODEN CNISE9.

ISSN 0169-7552 (print), 1879-2324 (electronic).

Pape:1987:CWI

[131] Detlef Pape. Communication within an industrial enterprise. Computer Networks and ISDN Systems, 14(2–5):237– 244, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Moens:1987:AIT

[132] Luc Moens and Gerard Valenduc. Assessment of impacts of telematics onto industrial structures. Computer Networks and ISDN Systems, 14(2–5):245– 249, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Pocock:1987:CGO

[133] Rosalie A. Pocock. Communicating in the global office. Computer Networks and ISDN Systems, 14(2–5):251–258, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Sutherland:1987:MT

[134] Ewan Sutherland and Yves V. H. Morieux. Managing telecommunications. Computer Networks and ISDN Systems, 14(2–5):259–266, 1987. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Baert:1987:SFS

[135] Patrick Baert, Christel Cuypers, and Jan DeSchampheleire. Silent force of the screen. Computer Networks and ISDN Systems, 14(2–5):267–270, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Philippart:1987:PNS

[136] Eric Philippart. Paradigms of natural sciences dealing with information and communication in sociology. Computer Networks and ISDN Systems, 14(2–5):271–278, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Tamine:1987:SIO

[137] J. Tamine. Successful implementation of an office system. Computer Networks and ISDN Systems, 14(2–5):279–282, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Latrille:1987:SAE

[138] James Latrille. Socio-technical aspects of electronic mail implementation. Computer Networks and ISDN Systems, 14(2–5):283–290, 1987. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Woodcock:1987:TAO

[139] A. Woodcock. Towards the application of OSI standards. Computer Networks and ISDN Systems, 14(2–5):291–295, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Caneschi:1987:AAE

[140] F. Caneschi and E. Merelli. Architecture for an ASN.1 encoder-decoder. Computer Networks and ISDN Systems, 14(2–5):297–303, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bilgen:1987:FCV

[141] Semih Bilgen, Mustafa Aykut, and Orkun Ekinci. Forms class virtual terminal service project. Computer Networks and ISDN Systems, 14(2–5):305– 309, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Thooft:1987:FDT

[142] G. Thooft. Formal description techniques: communication tools for data communication specialists. Computer Networks and ISDN Systems, 14(2–5): 311–321, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Cochrane:1987:EMF

[143] D. Cochrane. Easing the migration into the future of communications technology. Computer Networks and ISDN Systems, 14(2–5):323–329, 1987. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Franck:1987:WRD

[144] Reinhold Franck. Who really does need ISDN? Computer Networks and ISDN Systems, 14(2–5):331–337, 1987. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Frantzen:1987:TDP

[145] Viktor Frantzen. Trends in the development of public telecommunication networks. Computer Networks and ISDN Systems, 14(2–5):339–358, 1987. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Jordaan:1987:DTB

[146] M. J. Jordaan and H. Van Kampen. Digital technology for business communication. Computer Networks and ISDN Systems, 14(2–5):359–364, 1987.

CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Janssens:1987:IAL

[147] G. K. Janssens, J. Raes, and D. Staellens. Implementation approach for local area networks. Computer Networks and ISDN Systems, 14(2–5):365–372, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Endrizzi:1987:DBN

[148] Adriano Endrizzi. DUAL backbone network: distributed and parallel processing on a large scale. Computer Networks and ISDN Systems, 14(2–5):373– 381, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Comer:1987:CCB

[149] Douglas Comer, Thomas Narten, and Rajendra Yavatkar. Cypress coaxial backbone packet switch. Computer Networks and ISDN Systems, 14(2–5): 383–388, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

VanHeurck:1987:TNS

[150] Philippe Van Heurck. TRASEX: national security system for EFTs in Belgium. Computer Networks and ISDN Systems, 14(2–5):389–395, 1987. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Budkowski:1987:IES

[151] S. Budkowski and P. Dembinski. An introduction to Estelle: A specification language for distributed systems. Computer Networks and ISDN Systems, 14(3):3–23, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Evequoz:1988:RDP

[152] Claude Evequoz and Carl Tropper. On reassembly delay in packet switching networks. Computer Networks and ISDN Systems, 15(1):1–25, 1988. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bolger:1988:CSI

[153] Thomas E. Bolger. Controversies surrounding ISDN. Computer Networks and ISDN Systems, 15(1):27–30, 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Pattavina:1988:PER

[154] Achille Pattavina and Giovanni Pacifici. Performance evaluation of ring networks supporting a packet voice service. Computer Networks and ISDN Systems, 15(1):31–45, 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Zeng:1988:NAF

[155] H. X. Zeng, Q. Li, X. F. Du, and C. S. He. New advances in ferry testing approaches. Computer Networks and ISDN Systems, 15(1):47–54, 1988. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Santo:1988:AAM

[156] Horst Santo. Amigo: Advanced messaging in groups — interim report. Computer Networks and ISDN Systems, 15(1):55–60, 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Langsford:1988:DSM

[157] A. Langsford. Distributed systems management — overview of a COST-11 ter project. Computer Networks and ISDN Systems, 15(1):61–65, 1988. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Muftic:1988:SMC

[158] Sead Muftic. Security mechanisms for computer networks. current results of the CEC COST-11 ter project. Computer Networks and ISDN Systems, 15(1):67–71, 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

VanDerVeer:1988:IAH

[159] Gerrit C. Van Der Veer, Steve Guest, Pim Haselager, Peter Innocent, Eddy McDaid, Lars Oestericher, Michael J. Tauber, Ulfert Vos, and Yvonne Waern. Interdisciplinary approach to human factors in telematic systems. A review of the problems and possible solutions byl a COST-11 ter working group. Computer Networks and ISDN Systems, 15(1):73–80, 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Saito:1988:AML

[160] Hiroshi Saito. Analysis of measured loss probability and grade-of-service in ISDNs. Computer Networks and ISDN Systems, 15(2):87–96, 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Rothermel:1988:CMS

[161] Kurt Rothermel. A communication mechanism supporting actions. Computer Networks and ISDN Systems, 15 (2):97–108, 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

MehmetAli:1988:RAB

[162] M. K. Mehmet Ali, C. M. Woodside, and J. F. Hayes. Re-assembly buffer requirements in a packet voice network. Computer Networks and ISDN Systems, 15(2):109–119 (or 109–120??), 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Narasimhan:1988:RSB

[163] Sridhar Narasimhan, Hasan Pirkul, and Prabuddha De. Route selection in backbone data communications networks. Computer Networks and ISDN Systems, 15(2):121–133 (or 121–135??), 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Karshmer:1988:TID

[164] Arthur I. Karshmer, James Phelan, and James Thomas. TVNet: An image and data delivery system using cable T.V. facilities. Computer Networks and ISDN Systems, 15(2):135– 151 (or 135–152??), 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chanson:1988:HIR

[165] S. T. Chanson and K. Ravindran. Host identification in reliable distributed kernels. Computer Networks and ISDN Systems, 15(3):159–176, 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Todd:1988:BSU

[166] T. D. Todd and L. Li. Branch-switching in unidirectional channel systems using the IEEE 802.4 Token-Passing Bus Access Method. Computer Networks and ISDN Systems, 15(3):177–188, 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Nance:1988:LCA

[167] R. E. Nance and R. L. Moose, Jr. Link capacity assignment in dynamic hierarchical networks. Computer Networks and ISDN Systems, 15(3):189– 202, 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Paepcke:1988:XXX

[168] A. Paepcke, R. D. Crawford, R. Jamp, C. A. Freeman, and F. J. Lee. ?? Computer Networks and ISDN Systems, 15(3):203–216, 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Panda:1988:PSB

[169] D. K. Panda and T. Viswanathan. A parallel-serial binary arbitration scheme for collision-free multi-access techniques. Computer Networks and ISDN Systems, 15(3):217–224, 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Yanney:1988:FRD

[170] Raif M. Yanney and John P. Hayes. Fault recovery in distributed processing loop networks. Computer Networks and ISDN Systems, 15(4):229–243 (or 229–244??), September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Shiratori:1988:SDM

[171] Norio Shiratori, Kaoru Takahashi, and Shoichi Noguchi. A software design method and its application to protocol and communication software development. Computer Networks and ISDN Systems, 15(4):245–267 (or 245– 268??), September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Zatti:1988:ION

[172] Stefano Zatti and Philippe Janson. Interconnecting OSI and non-OSI networks using an integrated directory service. Computer Networks and ISDN Systems, 15(4):269–283 (or 269–284??), September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Sabnani:1988:PTG

[173] Krishan Sabnani and Anton Dahbura. A protocol test generation procedure. Computer Networks and ISDN Systems, 15(4):285–297 (or 285– 298??), September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Leakey:1988:ISD

[174] D. M. Leakey. Integrated services digital networks: Some possible ongoing evolutionary trends. Computer Networks and ISDN Systems, 15(5): 303–312, October 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Cowan:1988:NEU

[175] D. D. Cowan, S. Fenton, J. W. Graham, and T. M. Stepien. Networks for education at the University of Waterloo. Computer Networks and ISDN Systems, 15(5):313–327, October 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Jaffe:1988:HCT

[176] Jeffrey M. Jaffe. Hierarchical clustering with topology databases. Computer Networks and ISDN Systems, 15 (5):329–339, October 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Birk:1988:NES

[177] Yitzhak Birk and Fouad A. Tobagi. ‘Supernodes’ in networks employing spread spectrum with Code Division Multiple Access. Computer Networks and ISDN Systems, 15(5):341–357, October 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Miyake:1988:TSP

[178] Kou Miyake. Traffic study of primary rate ISDN user-network interface. Computer Networks and ISDN Systems, 15(5):359–367, October 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Yan:1988:DEC

[179] Xiong Yan. Design of an experimental computer LAN interconnection in university environment. Computer Networks and ISDN Systems, 15 (5):369–373, October 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Hansen:1988:RMP

[180] A. Hansen. RAE MHS project review. Computer Networks and ISDN Systems, 16(1–2):8–12, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Hutton:1988:OCS

[181] James Hutton. Overview of the current situation (COSINE specification phase). Computer Networks and ISDN Systems, 16(1–2):21–23, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Anonymous:1988:EN

[182] Anonymous. 4th European networkshop. Computer Networks and ISDN Systems, 16(1–2):21–174, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Cornillie-Braun:1988:TSE

[183] Anne Cornillie-Braun and Wulfdieter Bauerfeld. Tariff structure of European public packet switched data networks. Computer Networks and ISDN Systems, 16(1–2):24–30, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Dand:1988:PRT

[184] Andrew Dand. Product review: Trends and issues. Computer Networks and ISDN Systems, 16(1–2):31–33, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Clyne:1988:LWI

[185] Les Clyne. LAN/WAN internetworking. Computer Networks and ISDN Systems, 16(1–2):34–39, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Wilhelm:1988:MU

[186] Martin Wilhelm. Migration for users. Computer Networks and ISDN Systems, 16(1–2):40–43, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Mendoza:1988:DSC

[187] Eduardo Mendoza. Directory services and COSINE. Computer Networks and ISDN Systems, 16(1–2):44–47, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Lubich:1988:NAS

[188] Hannes Lubich and Bernhard Plattner. Naming and addressing in SWITCHmail. Computer Networks and ISDN Systems, 16(1–2):48–54, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Fluckiger:1988:GCC

[189] François Fluckiger. Gateways and converters in computer networks. Computer Networks and ISDN Systems, 16 (1–2):55–59 (or 55–60??), September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Richter:1988:REC

[190] Jorgen A. Richter and Nicholas K. Newman. The role of the European Commission in Telecommunications: The CEC Green Paper and beyond. Computer Networks and ISDN Systems, 16(1–2):64–74, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Plattner:1988:SNN

[191] Bernhard Plattner. The Swiss National Network for Research and Education (SWITCH). Computer Networks and ISDN Systems, 16(1–2): 75–82, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Zhao:1988:PSO

[192] Xiaofan Zhao. Present situation of OSI standards in China. Computer Networks and ISDN Systems, 16(1– 2):83–88, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Wolff:1988:PNS

[193] Stephen Wolff. The present networking situation in the USA. Computer Networks and ISDN Systems, 16(1–2):89– 91 (or 89–93??), September 1988. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Karrenberg:1988:EOT

[194] Daniel Karrenberg. EUnet and OSI transition plans. Computer Networks and ISDN Systems, 16(1–2):94–100, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Smith:1988:JAN

[195] I. L. Smith. Joint Academic Network (JANET). Computer Networks and ISDN Systems, 16(1–2):101–105 (or 101–107??), September 1988. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Linington:1988:CIP

[196] Peter F. Linington. The COSINE implementation phase review. Computer Networks and ISDN Systems, 16(1–2): 108–115, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Beale:1988:CIP

[197] J. S. Beale. COSINE implementation phase: COSINE project officer’s view. Computer Networks and ISDN Systems, 16(1–2):116–118, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Malagardis:1988:CIP

[198] Nicholas Malagardis. COSINE implementation phase: National view. Computer Networks and ISDN Systems, 16 (1–2):119–120, September 1988. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Carpenter:1988:CIP

[199] Brian E. Carpenter. COSINE implementation phase: The view from a major site. Computer Networks and ISDN Systems, 16(1–2):121–122 (or 121–123??), September 1988. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

VanBinst:1988:PEH

[200] Paul Van Binst. Pan-European highspeed networking. Computer Networks and ISDN Systems, 16(1–2): 124–128, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Koudelka:1988:HSS

[201] Otto Koudelka and Mervyn Hine. Higher speed services. Computer Networks and ISDN Systems, 16(1–2): 129–134, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Burkle:1988:BS

[202] Martin Bürkle. Broadband services. Computer Networks and ISDN Systems, 16(1–2):135–136 (or 135–137??), September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Scheller:1988:DSA

[203] Angela Scheller. Document standards: Availability and products. Computer Networks and ISDN Systems, 16(1–2): 138–142, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Kille:1988:TLS

[204] Steve Kille. The THORN large scale pilot exercise. Computer Networks and ISDN Systems, 16(1–2): 143–145, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Mount:1988:WUW

[205] Richard P. Mount. What users want. Computer Networks and ISDN Systems, 16(1–2):146–149, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

McKenna:1988:UE

[206] Peter McKenna. Update on EURO-MATH. Computer Networks and ISDN Systems, 16(1–2):150–152, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Craigie:1988:IXT

[207] Jim Craigie. ISO 10021 — X.400(88): A tutorial for those familiar with X.400(84). Computer Networks and ISDN Systems, 16(1–2):153–160, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Huitema:1988:XDS

[208] C. Huitema. The X.500 directory services. Computer Networks and ISDN Systems, 16(1–2):161–166, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Jeffree:1988:ROM

[209] Tony Jeffree. A review of OSI management standards. Computer Networks and ISDN Systems, 16(1–2): 167–174, September 1988. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Cusworth:1989:WDM

[210] S. D. Cusworth, J. M. Senior, and A. Ryley. Wavelength division multiple access on a high-speed optical fibre LAN. Computer Networks and ISDN Systems, ??(??):323–333, 1989/ 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Humbelt:1989:TBA

[211] P. A. Humbelt and S. R. Soloway. Topology broadcast algorithms. Computer Networks and ISDN Systems, 16 (3):179–186, January 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Turman:1989:BOC

[212] B. Turman and R. Rubin. Bell Operating Company packet interfaces between networks and subnets. Computer Networks and ISDN Systems, 16 (3):187–196, January 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Aggarwal:1989:SAP

[213] S. Aggarwal and Z. Har’el. Simulation analysis of protocols in an integrated software environment. Computer Networks and ISDN Systems, 16 (3):197–216, January 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Tcha:1989:LLB

[214] D.-W. Tcha and C.-Y. Jin. Linkby-link bandwidth allocation in an integrated voice/data network using the fuzzy set principle. Computer Networks and ISDN Systems, 16(3): 217–228, January 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Fratta:1989:PEC

[215] L. Fratta and J. Wozniak. PR-EXPRESS: Collision-free access protocol for packet radio networks. Computer Networks and ISDN Systems, 16 (3):229–242, January 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Lenzini:1989:OII

[216] L. Lenzini. The OSIRIDE-interest initiative: Status and trends. Computer Networks and ISDN Systems, 16 (3):243–256, January 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Cockburn:1989:TFR

[217] Alistair A. R. Cockburn. Tree-forming reversible routes in communication networks. Computer Networks and ISDN Systems, 16(4):267–279, March 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Zukerman:1989:CAO

[218] Moshe Zukerman. Circuit allocation and overload control in a hybrid switching system. Computer Networks and ISDN Systems, 16(4):281–298, March 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Cheung:1989:FNM

[219] To-Yat Cheung and Mike Sablatash. Functional network model for analytical file management in ISDN systems from generalization of videotex systems. Computer Networks and ISDN Systems, 16(4):299–310, March 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Belina:1989:CSD

[220] Ferenc Belina and Dieter Hogrefe. The CCITT Specification and Description Language SDL. Computer Networks and ISDN Systems, 16(4):311– 341, March 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Ko:1989:ACC

[221] Chi Chung Ko, Kim Mun Lye, Kee Chaing Chua, and Fa Toh Yap. Analysis of a CSMA/CD-based protocol with dynamic segmentation. Computer Networks and ISDN Systems, 16 (5):347–355 (or 347–356??), May 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Han:1989:VTP

[222] Sunyoung Han, Kilnam Chon, and Dongman Lee. A virtual terminal protocol with windowing capability. Computer Networks and ISDN Systems, 16 (5):357–365 (or 357–366??), May 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Garcia-Luna-Aceves:1989:MHR

[223] Jose Joaquin Garcia-Luna-Aceves. A minimum-hop routing algorithm based on distributed information. Computer Networks and ISDN Systems, 16(5):367–382, May 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Zissopoulos:1989:BAT

[224] Athanassios Zissopoulos and Carl Tropper. On buffer allocation in transport protocols. Computer Networks and ISDN Systems, 16(5):383– 393 (or 383–394??), May 1989. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Burg:1989:XCL

[225] Fred M. Burg and Patrick Puges. X.25: It’s come a long way. Computer Networks and ISDN Systems, 16(5):395–404, May 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Yan:1989:ICL

[226] Xiong Yan and Kuo-Liang Chen. Internetwork connection of the local area networks C-net and Omninet. Computer Networks and ISDN Systems, 16(5):405–414, May 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Morris:1989:TLC

[227] S. Morris, T. Suda, and T. Nguyen. A tree LAN with collision avoidance: photonic switch design and simulated performance. Computer Networks and ISDN Systems, 17(??):89–100, 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chiu:1989:AID

[228] Dah-Ming Chiu and Raj Jain. Analysis of the increase/decrease algorithms for congestion avoidance in computer networks. Computer Networks and ISDN Systems, 17(1):1–14, June 1989. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kamal:1989:AAN

[229] Ahmed E. Kamal and V. Carl Hamacher. Approximate analysis of non-exhaustive multiserver polling systems with applications to local area networks. Computer Networks and ISDN Systems, 17(1):15–27, June 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Ikeuchi:1989:NCA

[230] Noriyuki Ikeuchi and Makoto Yoshida. Network constructing algorithm based on link significance evaluation: NABLE. Computer Networks and ISDN Systems, 17(1):29–36, June 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Niwa:1989:ENN

[231] Akio Niwa and Hiroshi Tokunaga. Evolution of NTT’s networks towards INS. Computer Networks and ISDN Systems, 17(1):37–45, June 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Matsubara:1989:ECN

[232] Mark M. Matsubara. Evolution of CCITT numbering plans and network interworking: public data network perspective. Computer Networks and ISDN Systems, 17(1):47–57, June 1989.

CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Zeng:1989:FCA

[233] Hua-Xin Zeng, Samuel T. Chanson, and Brian R. Smith. On ferry clip approaches in protocol testing. Computer Networks and ISDN Systems, 17(2):77–88, July 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Suda:1989:TLCa

[234] Tatsuya Suda, Steve Morris, and Thieu Nguyen. A tree LAN with collision avoidance: Photonic switch design and simulated performance. Computer Networks and ISDN Systems, 17(2):89–100, July 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Suda:1989:TLCb

[235] Tatsuya Suda and Steve Morris. Tree LANs with collision avoidance: station and switch protocols. Computer Networks and ISDN Systems, 17(2):101–110, July 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Sarikaya:1989:CTA

[236] Behcet Sarikaya. Conformance testing: architectures and test sequences. Computer Networks and ISDN Systems, 17(2):111–126, July 1989. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Segall:1989:DNA

[237] Adrian Segall and Inder S. Gopal. Distributed name assignment in computer networks. Computer Networks and ISDN Systems, 17(2):127–139, July 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chess:1989:CVR

[238] David M. Chess. Computer viruses and related threats to computer and network integrity. Computer Networks and ISDN Systems, 17(2):141–148, July 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

VanTill:1989:APB

[239] Jaap Van Till. A-ISDN proposal to bridge ‘personal computers’ and ‘ISDN’. Computer Networks and ISDN Systems, 17(2):149–152, July 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Oie:1989:TAB

[240] Y. Oie, S. Nishio, and T. Hasegawa. Throughput analysis of blocked access tree algorithms with mini-slots. Computer Networks and ISDN Systems, 17(3):201–221, 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Castelli:1989:II

[241] E. Castelli. The INFN and the INFNet. Computer Networks and ISDN Systems, 17(4–5):252–255, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Grimm:1989:RSX

[242] R. Grimm and D. Heagert. Recommendation for a shorthand X.400 address notation. Computer Networks and ISDN Systems, 17(4–5): 263–269, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Sluman:1989:TOM

[243] C. Sluman. A tutorial on OSI management. Computer Networks and ISDN Systems, 17(4–5):270–278, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Clyne:1989:LLO

[244] L. Clyne. Lower layer OSI addressing. Computer Networks and ISDN Systems, 17(4–5):279–283, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Carpenter:1989:OTL

[245] B. E. Carpenter. Is OSI too late? Computer Networks and ISDN Systems, 17 (4–5):284–286, October 10 1989. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Linington:1989:WO

[246] P. F. Linington. Why OSI? Computer Networks and ISDN Systems, 17(4–5): 287–290, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Black:1989:FAS

[247] W. Black. File access and serving in an OSI environment. Computer Networks and ISDN Systems, 17(4–5): 294–299, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bauerfeld:1989:GC

[248] W. Bauerfeld and P. Holleczek. Global connectivity. Computer Networks and ISDN Systems, 17(4–5):300–304, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Craigie:1989:UAC

[249] J. Craigie. UK Academic Community Directory Service Pilot Project. Computer Networks and ISDN Systems, 17 (4–5):305–311, October 10 1989. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gilmore:1989:WBD

[250] B. Gilmore. Workstations: Beyond the dumb terminal. Computer Networks and ISDN Systems, 17(4–5): 312–314, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Grimm:1989:SND

[251] R. Grimm. Security on networks: Do we really need it? Computer Networks and ISDN Systems, 17(4–5): 315–323, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Huitema:1989:CMM

[252] C. Huitema. The challenge of multimedia mail. Computer Networks and ISDN Systems, 17(4–5):324–327, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Herchuelz:1989:SSA

[253] P. Herchuelz. Supercomputers: Some aspects of their impact on communications. Computer Networks and ISDN Systems, 17(4–5):328–331, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bakonyi:1989:NHR

[254] P. Bakonyi and L. Csaba. Network for the Hungarian R&D community. Computer Networks and ISDN Systems, 17 (4–5):332–336, October 10 1989. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Cerf:1989:IAB

[255] V. Cerf. The Internet Activities Board. Computer Networks and ISDN Systems, 17(4–5):337–343, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Tindemans:1989:CRM

[256] P. Tindemans. COSINE: Ready to move into implementation phase. Computer Networks and ISDN Systems, 17 (4–5):344–345, October 10 1989. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Michau:1989:IPP

[257] C. Michau. The IXI pilot project. Computer Networks and ISDN Systems, 17 (4–5):346–347, October 10 1989. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Onions:1989:ISM

[258] J. Onions. ISODE: In support of migration. Computer Networks and ISDN Systems, 17(4–5):362–366, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Lloyd:1989:CAI

[259] M. Lloyd. Conformance: Assuring interoperability. Computer Networks and ISDN Systems, 17(4–5): 367–370, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Ghiselli:1989:II

[260] A. Ghiselli. Internetworking in INFNet. Computer Networks and ISDN Systems, 17(4–5):371–375, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Lovdal:1989:CTI

[261] E. Lovdal. The challenge of TCP/ IP. Computer Networks and ISDN Systems, 17(4–5):376–380, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Jain:1989:CDA

[262] R. Jain. Characteristics of destination address locality in computer networks: A comparison of caching schemes. Computer Networks and ISDN Systems, 18(??):243–254, 1989/ 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

URL http://www.cis.ohio-state.
edu/~jain/papers/cache_ps.htm.

Sharrock:1989:CCB

[263] S. M. Sharrock, S. Ghanta, H. C. Du, and K. J. Maly. A CSMA/CD-based, integrated voice/data protocol with dynamic channel allocation. Computer Networks and ISDN Systems, 18(1): 1–18, November 24, 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Probert:1989:CSE

[264] R. L. Probert, H. Ural, and M. W. A. Hornbeek. A comprehensive software environment for developing standardized conformance test suites. Computer Networks and ISDN Systems, 18 (1):19–30, November 24, 1989. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Svobodova:1989:MPT

[265] L. Svobodova. Measured performance of transport service in LANs. Computer Networks and ISDN Systems, 18 (1):31–45 (or 31–46??), November 24, 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Popescu-Zeletin:1989:BIM

[266] R. Popescu-Zeletin. From broadband ISDN to multimedia computer networks. Computer Networks and ISDN Systems, 18(1):46–54, November 24, 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Lin:1989:IBR

[267] T.-P. Lin. ISDN basic rate interface architecture for ADPCM transmission. Computer Networks and ISDN Systems, 18(1):55–64, November 24, 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Holman:1989:RNF

[268] T. Holman, S. Wakid, and T. DeWitt. The role of the NIU-forum in realizing interoperable ISDN. Computer Networks and ISDN Systems, 18(1): 65–72, November 24, 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bochmann:1990:PSO

[269] G. v. Bochmann. Protocol specification for OSI. Computer Networks and ISDN Systems, 18(??):167–184, 1990. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Biersack:1990:SAC

[270] E. W. Biersack. A systematic approach for constructing gateways. Computer Networks and ISDN Systems, 18(2): 79–96, 28 February 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bondavalli:1990:MPH

[271] A. Bondavalli, M. Conti, E. Gregori, L. Lenzini, and L. Strigni. MAC protocols for high-speed MANs: Performance comparisons for a family of Fasnet-based protocols. Computer Networks and ISDN Systems, 18(2):97– 114, 28 February 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Adas:1990:EMW

[272] A. A. A. Adas. Evaluation methodology for wide area networks. Computer Networks and ISDN Systems, 18 (2):115–132, 28 February 1990. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Orda:1990:MCN

[273] A. Orda and R. Rom. Multihoming in computer networks: A topologydesign approach. Computer Networks and ISDN Systems, 18(2):133– 142, 28 February 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Cusworth:1990:WDM

[274] S. D. Cusworth, J. M. Senior, and A. Ryley. Wavelength division multiple access on a high-speed optical fibre LAN. Computer Networks and ISDN Systems, 18(5):323–333 (or 323–334??), June 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bochmann:1990:SST

[275] G. v. Bochmann. Specifications of a simplified transport protocol using different formal description techniques. Computer Networks and ISDN Systems, 18(5):335–378, June 1990. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Wong:1990:DFS

[276] J. S. K. Wong and Y. Kang. Distributed and fail-safe routing algorithms in toroidal-based metropolitan area networks. Computer Networks and ISDN Systems, 18(5):355–378, June 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Russel:1990:HLS

[277] D. Russel. High-level security architecture and the Kerberos system. Computer Networks and ISDN Systems, 19(??):201–214, 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Brachman:1990:SSA

[278] B. J. Brachman and S. T. Chanson. A simulation study of application level message transfer using message streams. Computer Networks and ISDN Systems, 19(2):79–94, 1990. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gibson:1990:ISE

[279] R. W. Gibson. IEEE 802 standards efforts. Computer Networks and ISDN Systems, 19(2):95–104, 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Nessett:1990:TSH

[280] D. M. Nessett and G. M. Lee. Terminal services in heterogeneous distributed systems. Computer Networks and ISDN Systems, 19(2):105–128, 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bruneel:1990:RBB

[281] H. Bruneel, J. De Vriendt, and C. Ysebaert. Receiver buffer behavior for the selective-repeat ARQ protocol. Computer Networks and ISDN Systems, 19 (2):129–142, 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Park:1990:QSR

[282] J. L. Park. Quality of service requirements in packet switching public data networks. Computer Networks and ISDN Systems, 20(1–5):3–6, 1990. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). ITC Specialist Seminar, 25–29 September 1989, Adelaide, Australia.

Ackerley:1990:FQS

[283] R. G. Ackerley. The future of quality of service modelling. Computer Networks and ISDN Systems, 20(1–5):7–14, 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). ITC Specialist Seminar, 25–29 September

Bradlow:1990:PMR

[284] H. S. Bradlow. Performance measures for real-time continuous bit-stream oriented services: Application to packet reassembly. Computer Networks and ISDN Systems, 20(1–5):15–26, 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). ITC Specialist Seminar, 25–29 September 1989.

Kashper:1990:RDS

[285] A. N. Kashper, S. S. Katz, and G. R. Ash. Robust design for switched digital services in a worldwide intelligent network. Computer Networks and ISDN Systems, 20(1–5):81–88, 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kashper:1990:IFR

[286] A. Kashper. Introduction of flexible routing for international networks. Computer Networks and ISDN Systems, 20(1–5):137–141, 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chemouil:1990:INM

[287] P. Chemouil and J. Filipiak. Integrated network management and control. Computer Networks and ISDN Systems, 20(1–5):143–153, 1990. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Warfield:1990:PUA

[288] B. Warfield and P. Sember. Prospects for the use of artificial intelligence in real time network traffic management. Computer Networks and ISDN Systems, 20(1–5):163–169, 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Forys:1990:SBS

[289] L. J. Forys and D. E. Smith. Servicing of bursty systems. Computer Networks and ISDN Systems, 20(1–5):171–177, 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). ITC Specialist Seminar, 25–29 September 1989, Adelaide, Australia.

Dziong:1990:ACR

[290] Z. Dziong, J. Choquette, K. Q. Liao, and L. Mason. Admission control and routing in ATM networks. Computer Networks and ISDN Systems, 20(1–5): 189–196, 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). ITC Specialist Seminar, 25–29 September 1989, Adelaide, Australia.

Kleinrock:1990:DSP

[291] Leonard Kleinrock. On distributed systems performance. Computer Networks and ISDN Systems, 20(1–5): 209–215, December 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bannister:1990:DMA

[292] Joseph A. Bannister, Luigi Fratta, and Mario Gerla. Designing metropolitan area networks for high-performance applications. Computer Networks and ISDN Systems, 20(1–5):223–230, December 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Woodruff:1990:CAS

[293] G. N. Woodruff, R. Kositpaiboon, G. Fitzpatrick, and P. Richards. Control of ATM statistical multiplexer performance. Computer Networks and ISDN Systems, 20(1–5):351–360, 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). ITC Specialist Seminar, 25–29 September 1989, Adelaide, Australia.

Skoog:1990:AIN

[294] R. A. Skoog and A. R. Modarressi. Alternatives and issues for network signaling transport in a broadband environment. Computer Networks and ISDN Systems, 20(1–5):361–368, 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kawashima:1990:TIA

[295] K. Kawashima and H. Saito. Teletraffic issues in ATM networks. Computer Networks and ISDN Systems, 20(1–5): 369–375, 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). ITC Specialist Seminar, 25–29 September 1989, Adelaide, Australia.

Gall:1990:PTE

[296] P. Le Gall. Packetized traffic engineering for new services. Computer Networks and ISDN Systems, 20(1–5):425– 433, 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). ITC Specialist Seminar, 25–29 September 1989, Adelaide, Australia.

Veiro:1990:TMV

[297] B. Veiro. Traffic measurement on variable bit rate (VBR) sources with application to charging principles. Computer Networks and ISDN Systems, 20(1–5): 435–445, 1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). ITC Specialist Seminar, 25–29 September 1989, Adelaide, Australia.

Smetaniuk:1991:DOX

[298] Bohdan Smetaniuk. Distributed operation of the X.500 directory. Computer Networks and ISDN Systems, 21(1):17–40, March 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Simha:1991:SFA

[299] Rahul Simha and Yoram Ofek. A starvation-free access protocol for a full-duplex buffer insertion ring local area network. Computer Networks and ISDN Systems, 21(2):109–120, April 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Belovich:1991:LTA

[300] Steve G. Belovich and Vijaya K. Konangi. A linear-time approximation method for computing the reliability of a network. Computer Networks and ISDN Systems, 21(2):121–127, April 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bormann:1991:SOD

[301] U. Bormann and C. Bormann. Standards for open document processing: current state and future developments. Computer Networks and ISDN Systems, 21(3):149–163, May 1991. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Brinksma:1991:IP

[302] E. Brinksma, G. Scollo, and C. A. Vissers. Introduction to the PSTV-IX. Computer Networks and ISDN Systems, 22(1):1–??, August 1991. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bosik:1991:FSM

[303] B. S. Bosik and M. U. Uyar. Finite state machine based formal methods in protocol conformance testing: from theory to implementation. Computer Networks and ISDN Systems, 22(1):7–??, August 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Berry:1991:IDH

[304] G. Berry and G. Gonthier. Incremental development of an HDLC entity in Esterel. Computer Networks and ISDN Systems, 22(1):35–??, August 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Murphy:1991:EEL

[305] S. C. Murphy, P. Gunningberg, and J. P. J. Kelly. Experiences with Estelle, LOTOS and SDL: a protocol implementation experiment. Computer Networks and ISDN Systems, 22(1):51–??, August 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bolognesi:1991:DGG

[306] T. Bolognesi, O. Hagsand, D. Latella, and B. Pehrson. The definition of a graphical G-LOTOS editor using the meta-tool LOGGIE. Computer Networks and ISDN Systems, 22(1):61–??, August 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Heinzmann:1991:GE

[307] P. Heinzmann and H. Rudin. Guest editorial. Computer Networks and ISDN Systems, 22(2):83–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Piguet:1991:SRD

[308] P. Piguet. Storage and retrieval of documentation at the United Nations. Computer Networks and ISDN Systems, 22(2):87–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Hibbard:1991:IAD

[309] W. Hibbard, D. Santek, and G. Tripoli. Interactive atmospheric data access via high-speed networks. Computer Networks and ISDN Systems, 22(2):103– ??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Vaziri:1991:SVH

[310] A. Vaziri. Scientific visualization in high-speed network environments. Computer Networks and ISDN Systems, 22(2):111–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Johnson:1991:CDS

[311] M. J. Johnson. Coping with data from Space Station Freedom. Computer Networks and ISDN Systems, 22(2):131– ??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Smith:1991:TCD

[312] J. M. Smith and D. J. Farber. Traffic characteristics of a distributed memory system. Computer Networks and ISDN Systems, 22(2):143–154, September 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Sincoskie:1991:SAL

[313] W. D. Sincoskie. System architecture for a large scale video on demand service. Computer Networks and ISDN Systems, 22(2):155–162, September 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Ngoh:1991:MSG

[314] L. H. Ngoh. Multicast support for group communications. Computer Networks and ISDN Systems, 22(3):165– ??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Estrin:1991:PRI

[315] D. Estrin. Policy requirements for Inter-Administrative Domain Routing. Computer Networks and ISDN Systems, 22(3):179–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Takagi:1991:APM

[316] H. Takagi. Application of polling models to computer networks. Computer Networks and ISDN Systems, 22(3): 193–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Wang:1991:HAN

[317] X. Wang, J. Kaniyil, Y. Onozato, and S. Noguchi. Heterogeneous ALOHA networks: a sufficient condition for all equilibrium states to be stable.

Computer Networks and ISDN Systems, 22(3):213–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chang:1991:SSS

[318] C. J. Chang, J. W. Wong, and J. H. Chiu. A simulation study on the service strategies for packet voice communication networks. Computer Networks and ISDN Systems, 22(3):225– ??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Vasilakos:1991:VWF

[319] A. V. Vasilakos, C. A. Moschonas, and C. T. Paximadis. Variable window flow control and ergodic discretized learning algorithms for adaptive routing in data networks. Computer Networks and ISDN Systems, 22(4):235–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Jin:1991:PMF

[320] S. Jin, D. R. Vaman, and D. Sinha. A performance management framework to provide bounded packet delay and variance in packet switched networks. Computer Networks and ISDN Systems, 22(4):249–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Lee:1991:PCB

[321] D. Y. Lee and J. Y. Lee. Performance comparison of bridge algorithms in interconnected local area networks. Computer Networks and ISDN Systems, 22(4):265–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bruneel:1991:EDT

[322] H. Bruneel. Exact derivation of transient behaviour for buffers with random output interruptions. Computer Networks and ISDN Systems, 22(4):277– 285, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chung:1991:BRT

[323] P. Chung and A. K. Elhakeem. A bandwidth reducing token ring. Computer Networks and ISDN Systems, 22(4): 287–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

vanDijk:1991:PFR

[324] N. M. van Dijk. Product forms for random access schemes. Computer Networks and ISDN Systems, 22(4):303– 317, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Rudin:1991:GE

[325] H. Rudin. Guest editorial. Computer Networks and ISDN Systems, 22(5): 321–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Janson:1991:SON

[326] Phil Janson and Refik Molva. Security in open networks and distributed systems. Computer Networks and ISDN Systems, 22(5):323–346, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Reynolds:1991:HI

[327] J. K. Reynolds. The helminthiasis of the Internet. Computer Networks and ISDN Systems, 22(5):347–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Estrin:1991:SCT

[328] D. Estrin and G. Tsudik. Secure control of transit internetwork traffic. Computer Networks and ISDN Systems, 22(5):363–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

VanBinst:1991:GE

[329] P. Van Binst. Guest editorial. Computer Networks and ISDN Systems, 23 (1–3):1–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Hunke:1991:GS

[330] H. Hunke. Gala-dinner speech. Computer Networks and ISDN Systems, 23 (1–3):3–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Pouzin:1991:TYO

[331] L. Pouzin. Ten years of OSI — maturity or infancy? Computer Networks and ISDN Systems, 23(1– 3):11–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chesson:1991:CWN

[332] G. Chesson. The challenge of workstations to networks. Computer Networks and ISDN Systems, 23(1–3): 15–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Levrat:1991:NC

[333] B. Levrat. The networked campus. Computer Networks and ISDN Systems, 23(1–3):19–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Caldwell:1991:WIP

[334] L. G. Caldwell. Workstation to internet; problems, solutions, and challenges. Computer Networks and ISDN Systems, 23(1–3):25–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Ullmann:1991:OSP

[335] K. Ullmann. Organisational structures for the provision of international data communication services in the research community. Computer Networks and ISDN Systems, 23(1– 3):29–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Hutton:1991:AUP

[336] J. S. Hutton and A. Jeffree. Acceptable use policy. Computer Networks and ISDN Systems, 23(1–3):33–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bostwick:1991:HOU

[337] W. E. Bostwick. HPCC. an overview of the U.S. High Performance Computing and Communications initiative with focus on the National Research and Education Network. Computer Networks and ISDN Systems, 23(1– 3):37–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

VanBinst:1991:RRC

[338] P. Van Binst. Report on the RARE/ CEC symposium on high speed networking for research in Europe. Computer Networks and ISDN Systems, 23 (1–3):41–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gerich:1991:EIG

[339] E. Gerich. Expanding the Internet to a global environment but … how to get connected? Computer Networks and ISDN Systems, 23(1– 3):43–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Clyne:1991:CCI

[340] L. Clyne. CONS/CLNS interworking — policy and recommendations. Computer Networks and ISDN Systems, 23 (1–3):47–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Franzen:1991:FCH

[341] M. Franzen. FDDI concentrators and how they work together with other FDDI network components. Computer Networks and ISDN Systems, 23(1– 3):51–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kongsli:1991:TII

[342] T. A. Kongsli. TCP/IP internetwork communication through LANs interconnected by Dikon Meganet. Computer Networks and ISDN Systems, 23 (1–3):57–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Neggers:1991:EEP

[343] K. Neggers. European Engineering Planning Group (EEPG) — summary report. Computer Networks and ISDN Systems, 23(1–3):63–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gerich:1991:MON

[344] E. Gerich. Management and operation of the NSFNET backbone. Computer Networks and ISDN Systems, 23 (1–3):69–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Stockman:1991:NEN

[345] B. Stockman. NORDUnet experiences in network management. Computer Networks and ISDN Systems, 23(1– 3):73–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Solomon:1991:LMC

[346] C. Solomon, J. F. L’Haire, and J. F. Paccini. LAN management by cooperation: Hewlett–Packard and the University of Geneva. Computer Networks and ISDN Systems, 23(1–3): 79–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Benford:1991:BGC

[347] S. Benford. Building group communication on OSI. Computer Networks and ISDN Systems, 23(1–3):87–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Wilson:1991:CSC

[348] P. Wilson. Computer Supported Cooperative Work (CSCW): origins, concepts and research initiatives. Computer Networks and ISDN Systems, 23 (1–3):91–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Cargille:1991:IXP

[349] C. A. Cargille, R. A. Hagens, A. Hansen, and L. H. Landweber. The Internet X.400 pilot project. Computer Networks and ISDN Systems, 23 (1–3):97–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Dyer:1991:IXW

[350] J. Dyer. The integration of the X Window System and ISO Virtual Terminals for a European workstation environment. Computer Networks and ISDN Systems, 23(1–3):101–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Huitema:1991:NST

[351] C. Huitema. Naming: strategies and techniques. Computer Networks and ISDN Systems, 23(1–3):107–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Goodman:1991:PCX

[352] D. Goodman. PARADISE: the CO-SINE X.500 pilot service. Computer Networks and ISDN Systems, 23(1– 3):111–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Knight:1991:CCI

[353] G. Knight. The COSINE CONCISE information service project. Computer Networks and ISDN Systems, 23(1– 3):115–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Burren:1991:HSC

[354] J. W. Burren. High speed communications — a tutorial on the jargon and technologies. Computer Networks and ISDN Systems, 23(1–3): 119–124, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Despres:1991:GHS

[355] R. Despres. Global high speed WAN architecture for the 90’s. Computer Networks and ISDN Systems, 23(1–3): 125–128, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Fluckiger:1991:MGP

[356] F. Fluckiger. From megabit to gigabit: possible transition scenarios. Computer Networks and ISDN Systems, 23 (1–3):129–138, November 1991. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Altaber:1991:CRU

[357] J. Altaber, B. Carpenter, S. Cannon, D. Davids, M. Hine, C. Isnard, J. M. Jouanigot, S. Olofsson, B. Segal, N. Stheneur, et al. CHEOPS: really using a satellite. Computer Networks and ISDN Systems, 23(1– 3):139–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bauerfeld:1991:MCH

[358] W. Bauerfeld and H. Westbrock. Multimedia communication with highspeed protocols. Computer Networks and ISDN Systems, 23(1–3):143–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kirstein:1991:POP

[359] P. T. Kirstein and G. Montaser-Kohsari. Proposed ODA pilot activities for the research community. Computer Networks and ISDN Systems, 23 (1–3):153–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bessiere:1991:MAT

[360] C. Bessiere, J. L. Leonhardt, and R. Zeiliger. Multimedia authoring tools: Atelier ORGUE. Computer Networks and ISDN Systems, 23(1– 3):157–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Hill:1991:XDS

[361] J. M. Hill. The X.500 Directory Service and the Data Protection Act. Computer Networks and ISDN Systems, 23 (1–3):163–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Harvey:1991:CCE

[362] C. C. Harvey. CERT — Computer Emergency Response Team. Computer Networks and ISDN Systems, 23(1– 3):167–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Sales:1991:TSI

[363] B. Sales. TCP/IP-X.25/OSI interoperation: from the medium term to the long term. Computer Networks and ISDN Systems, 23(1–3): 171–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Crowcroft:1991:TSI

[364] J. Crowcroft and R. Bradshaw. TCP/ IP Internet protocols and JANET. Computer Networks and ISDN Systems, 23(1–3):177–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Callon:1991:IRM

[365] R. W. Callon. Integrated routing for multi-protocol TCP/IP-OSI environments. Computer Networks and ISDN Systems, 23(1–3):185–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Robinson:1991:RPC

[366] D. Robinson. Remote Procedure Call: a stepping stone towards ODP. Computer Networks and ISDN Systems, 23 (1–3):191–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bowen:1991:ODP

[367] D. Bowen. Open distributed processing. Computer Networks and ISDN Systems, 23(1–3):195–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Wacker:1991:PAW

[368] C. Wacker. Poster: Award winner. Computer Networks and ISDN Systems, 23(1–3):203–??, November 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Neufeld:1991:DNR

[369] G. Neufeld. Descriptive name resolution. Computer Networks and ISDN Systems, 23(4):211–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Leung:1991:NEA

[370] C. H. C. Leung, Y. Kikumoto, S. A. Sorensen, and S. O. Falaki. A new efficient ARQ scheme for satellite communications. Computer Networks and ISDN Systems, 23(4):229–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Schepers:1991:NAR

[371] H. J. J. H. Schepers, O. B. P. Rikkert de Koe, G. M. J. Havermans, and D. K. Hammer. Naming, addressing, routing and relaying in the OSI environment. Computer Networks and ISDN Systems, 23(4):241–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Schepers:1991:LSW

[372] H. J. J. H. Schepers, O. B. P. Rikkert de Koe, G. M. J. Havermans, and D. K. Hammer. LAN/WAN interworking in the OSI environment. Computer Networks and ISDN Systems, 23(4):253– ??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Yang:1991:PST

[373] Q. Yang, D. Ghosal, and S. K. Tripathi. Performance study of two protocols for voice/data integration on ring networks. Computer Networks and ISDN Systems, 23(4):267–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Stassinopoulos:1991:AAL

[374] G. I. Stassinopoulos and I. S. Venieris. ATM adaptation layer protocols for signalling. Computer Networks and ISDN Systems, 23(4):287–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Platt:1991:TMF

[375] A. Platt and M. J. Morse. Traffic management in frame relay networks. Computer Networks and ISDN Systems, 23(4):305–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Akyildiz:1991:RIC

[376] I. F. Akyildiz. Report on the 3rd International Conference on The Performance of Distributed Systems and Integrated Communication Networks, 10–12 September 1991, Kyoto, Japan. Computer Networks and ISDN Systems, 23(4):317–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Vuong:1991:E

[377] S. T. Vuong. Editorial. Computer Networks and ISDN Systems, 23(5):323– ??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Logrippo:1991:ILL

[378] L. Logrippo, M. Faci, and M. Haj-Hussein. An introduction to LOTOS: learning by examples. Computer Networks and ISDN Systems, 23(5):325– ??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Sijelmassi:1991:GUE

[379] R. Sijelmassi and R. J. Linn. Guidelines for using Estelle to specify OSI services and protocols. Computer Networks and ISDN Systems, 23(5):343– ??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

deMeer:1991:IAS

[380] J. de Meer, R. Roth, and S. Vuong. Introduction to algebraic specifications based on the language ACT ONE. Computer Networks and ISDN Systems, 23(5):363–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Neufeld:1991:OA

[381] G. Neufeld and S. Vuong. An overview of ASN.1. Computer Networks and ISDN Systems, 23(5):393–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Probert:1991:TIN

[382] R. L. Probert and O. Monkewich. TTCM: the international notation for specifying tests of communications systems. Computer Networks and ISDN Systems, 23(5):417–??, 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Popescu-Zeletin:1992:AON

[383] R. Popescu-Zeletin and T. Magedanz. Applying open network provision to ISDN and intelligent networks. Computer Networks and ISDN Systems, 24(1):1–??, March 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bondavalli:1992:DSD

[384] A. Bondavalli, L. Strigini, and M. Sereno. Destination stripping dual ring: a new protocol for MANs. Computer Networks and ISDN Systems, 24(1):15–??, March 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Zinky:1992:TTB

[385] J. A. Zinky and J. Etkin. Troubleshooting throughput bottlenecks using executable models. Computer Networks and ISDN Systems, 24(1):33–??, March 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Murata:1992:PTR

[386] M. Murata and H. Takagi. Performance of token ring networks with a finite capacity bridge. Computer Networks and ISDN Systems, 24(1):45–??, March 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Lenzini:1992:ITO

[387] L. Lenzini and F. Zoccolini. Interoperability tests on OSI products in the framework of the OSIRIDE-Intertest initiative. Computer Networks and ISDN Systems, 24(1):65–??, March 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Heyman:1992:PMC

[388] D. P. Heyman. A performance model of the credit manager algorithm. Computer Networks and ISDN Systems, 24(1):81–??, March 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Karlsson:1992:APS

[389] J. M. Karlsson, H. G. Perros, and I. Viniotis. Adaptive polling schemes for an ATM bus with bursty arrivals. Computer Networks and ISDN Systems, 24(1):93–103, March 1992. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Pujolle:1992:P

[390] G. Pujolle and O. Spaniol. Preface. Computer Networks and ISDN Systems, 24(2):107–??, April 1992. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Birch:1992:PMC

[391] J. Birch, L. G. Christensen, and M. Skov. A programmable 800 mbit/s CRC check/generator unit for LANs and MANs. Computer Networks and ISDN Systems, 24(2):109–??, April 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gauffin:1992:MGN

[392] L. Gauffin, L. Håkansson, and B. Pehrson. Multi-gigabit networking based on DTM — A TDM medium access technique with dynamic bandwidth allocation. Computer Networks and ISDN Systems, 24(2):119–130, April 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Cohen:1992:NSD

[393] R. Cohen and A. Segall. A new scheme for dynamic management of isochronous channels in integrated rings. Computer Networks and ISDN Systems, 24(2):131–??, April 1992. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Phung:1992:UBD

[394] V. P. T. Phung and R. Breault. On the unpredictable behavior of DQDB. Computer Networks and ISDN Systems, 24(2):145–??, April 1992. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Banerjee:1992:ICM

[395] S. Banerjee and B. Mukherjee. Incorporating continuation-of-message information, slot reuse, and fairness in DQDB networks. Computer Networks and ISDN Systems, 24(2):153– 169, April 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Borgonovo:1992:DNA

[396] F. Borgonovo and L. Fratta. Deflection networks: architectures for metropolitan and wide area networks. Computer Networks and ISDN Systems, 24(2):171–??, April 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Towsley:1992:CAH

[397] D. Towsley, S. Fdida, and H. Santoso. Congestion avoidance in highspeed interconnection systems. Computer Networks and ISDN Systems, 24(2):185–??, April 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Logrippo:1992:TTP

[398] L. Logrippo, R. L. Probert, and H. Ural. Three tutorial papers on protocol specification, validation, and design. Computer Networks and ISDN Systems, 24(3):199–??, May 01, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gotzhein:1992:TLA

[399] R. Gotzhein. Temporal logic and applications– a tutorial. Computer Networks and ISDN Systems, 24(3): 203–??, May 01, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

West:1992:PVP

[400] C. H. West. Protocol validation — principles and applications. Computer Networks and ISDN Systems, 24 (3):219–??, May 01, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Doshi:1992:CPH

[401] B. T. Doshi and P. K. Johri. Communication protocols for high speed packet networks. Computer Networks and ISDN Systems, 24(3):243–??, May 01, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Rudin:1992:ATM

[402] H. Rudin. The ATM — Asynchronous Transfer Mode. Computer Networks and ISDN Systems, 24(4):277–??, May 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

LeBoudec:1992:ATM

[403] J. Y. Le Boudec. The asynchronous transfer mode: A tutorial. Computer Networks and ISDN Systems, 24(4):279–309, May 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Guillemin:1992:BRP

[404] F. Guillemin and A. Dupuis. A basic requirement for the policing function in ATM networks. Computer Networks and ISDN Systems, 24(4):311– 320, May 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Boyer:1992:RPA

[405] P. E. Boyer and D. P. Tranchier. A reservation principle with applications to the ATM traffic control. Computer Networks and ISDN Systems, 24(4):321–334, May 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Anagnostou:1992:CIR

[406] M. E. Anagnostou, M. E. Theologou, and E. N. Protonotarios. Cell insertion ratio analysis in asynchronous transfer mode networks. Computer Networks and ISDN Systems, 24(4):335–??, May 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bahaa-El-Din:1992:RPN

[407] W. H. Bahaa-El-Din and M. T. Liu. Register-insertion: A protocol for the next generation of ring local-area networks. Computer Networks and ISDN Systems, 24(5):349–??, June 1992. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bisdikian:1992:PAI

[408] C. Bisdikian. A performance analysis of the IEEE 802.6 (DQDB) subnetwork with the bandwidth balancing mechanism. Computer Networks and ISDN Systems, 24(5):367–385, June 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Evequoz:1992:OPS

[409] C. Evequoz and C. Tropper. Optimal packet sizes in virtual circuit computer networks. Computer Networks and ISDN Systems, 24(5):387–??, June 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Simha:1992:SPR

[410] R. Simha and B. Narahari. Single path routing with delay considerations. Computer Networks and ISDN Systems, 24(5):405–??, June 1992. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Atmaca:1992:PAA

[411] T. Atmaca, G. Pujolle, and F. Sha. Performance analysis of an access control strategy in integrated networks. Computer Networks and ISDN Systems, 24(5):421–??, June 1992. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chanson:1992:GE

[412] S. Chanson, H. Rudin, K. Turner, and S. Vuong. Guest editorial. Computer Networks and ISDN Systems, 25 (1):1–??, August 03, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gotzhein:1992:FDR

[413] R. Gotzhein. Formal definition and representation of interaction points. Computer Networks and ISDN Systems, 25 (1):3–??, August 03, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Leduc:1992:FBI

[414] G. Leduc. A framework based on implementation relations for implementing LOTOS specifications. Computer Networks and ISDN Systems, 25(1): 23–??, August 03, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Schneider:1992:IED

[415] J. M. Schneider, L. F. Mackert, G. Zorntlein, R. J. Velthuys, and U. Bar. An integrated environment for developing communication protocols. Computer Networks and ISDN Systems, 25(1):43–??, August 03, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Budkowski:1992:EDT

[416] S. Budkowski. Estelle development toolset (EDT). Computer Networks and ISDN Systems, 25(1):63–??, August 03, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Courtiat:1992:EIE

[417] J.-P. Courtiat and P. De Saqui-Sannes. ESTIM: an integrated environment for the simulation and verification of OSI protocols specified in Estelle. Computer Networks and ISDN Systems, 25 (1):83–??, August 03, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Mukherjee:1992:IVI

[418] B. Mukherjee and S.-K. Kao. An improved voice-data integration protocol for fiber optic bus networks. Computer Networks and ISDN Systems, 25 (2):103–??, August 17, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Liu:1992:ISL

[419] S. S. Liu. Impacts of signalingintensive local services (SILS) on B-ISDN switching systems: a simulation study. Computer Networks and ISDN Systems, 25(2):121–??, August 17, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Macartney:1992:FTD

[420] A. J. Macartney and G. S. Blair. Flexible trading in distributed multimedia systems. Computer Networks and ISDN Systems, 25(2):145–??, August 17, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Tusch:1992:PMT

[421] J. Tusch. Performance measurement in token ring networks. Computer Networks and ISDN Systems, 25(2):

159–??, August 17, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Nam:1992:PAB

[422] S. H. Nam and C. K. Un. Performance analysis of broadcast star network with collision-avoidance switch. Computer Networks and ISDN Systems, 25(2): 169–??, August 17, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Luan:1992:FSP

[423] Z. Luan, J. F. Hayes, and M. K. Mehmet Ali. Frame synchronization performance of SONET signals. Computer Networks and ISDN Systems, 25 (2):183–??, August 17, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chang:1992:PAS

[424] C.-J. Chang and S.-Y Wang. Performance analysis of a statistical multiplexer for integrated service in the customer-premise equipment. Computer Networks and ISDN Systems, 25 (2):191–??, August 17, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Maly:1992:DAB

[425] K. J. Maly, E. C. Foudriat, R. Mukkamala, C. M. Overstreet, and D. Game. Dynamic allocation of bandwidth in multichannel metropolitan area networks. Computer Networks and ISDN Systems, 25(2):203–??, August 17, 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kalin:1992:OOB

[426] T. Kalin and D. Barber. Has the OSI opportunity been fully realised? Computer Networks and ISDN Systems, 25 (3):227–??, September 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Thilakam:1992:DFC

[427] R. Krishna Thilakam and A. Jhunjhunwala. The design and flow control of a high speed, integrated, packet switched network. Computer Networks and ISDN Systems, 25(3):241–??, September 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Frieder:1992:CND

[428] O. Frieder and R. L. Shuey. Communication needs in a data engineering world. Computer Networks and ISDN Systems, 25(3):259–??, September 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kaniyil:1992:BAA

[429] J. Kaniyil, Y. Onozato, and S. Noguchi. On the behavioral aspects of alternate routing in non-hierarchical networks. Computer Networks and ISDN Systems, 25(3):275–??, September 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kim:1992:PAL

[430] Y. H. Kim, B. C. Shin, and C. K. Un. Performance analysis of leaky-bucket bandwidth enforcement strategy for bursty traffics in an ATM network.

Computer Networks and ISDN Systems, 25(3):295–??, September 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Coulson:1992:EAM

[431] G. Coulson, G. S. Blair, N. Davies, and N. Williams. Extensions to ANSA for multimedia computing. Computer Networks and ISDN Systems, 25(3): 305–??, September 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

VanBinst:1992:GE

[432] P. Van Binst. Guest editorial. Computer Networks and ISDN Systems, 25 (4–5):327–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Ullmann:1992:DNE

[433] K. Ullmann. Data networks for the European research and education sector — analysis and perspectives. Computer Networks and ISDN Systems, 25 (4–5):329–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Stals:1992:UNF

[434] B. Stals. Users and networks — survival of the fittest? Computer Networks and ISDN Systems, 25(4– 5):335–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Rysavy:1992:USP

[435] F. R. Rysavy. Users and services providers: interoperability. Computer Networks and ISDN Systems, 25(4– 5):339–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bishop:1992:ENE

[436] A. P. Bishop. Electronic networking for engineers: research from a user perspective. Computer Networks and ISDN Systems, 25(4–5): 344–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Holleczek:1992:TMM

[437] P. Holleczek and T. Baumgarten. Throughput measurements in a 2 Mbps X.25 network. Computer Networks and ISDN Systems, 25(4–5):351–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

DePrycker:1992:ATB

[438] M. De Prycker. ATM technology: a backbone for high speed computer networking. Computer Networks and ISDN Systems, 25(4–5): 357–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Heinanen:1992:FRM

[439] J. Heinanen. Frame relay as a multiprotocol backbone interface. Computer Networks and ISDN Systems, 25 (4–5):363–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Martinez:1992:FRS

[440] D. Martinez. Frame relay and SMDS services on a common ATM-based platform. Computer Networks and ISDN Systems, 25(4–5):370–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Cullen:1992:UFA

[441] J. Cullen, B. Day, and M. Planes. The use of FTAM to access graphical pictures across wide area networks. Computer Networks and ISDN Systems, 25 (4–5):377–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Danthine:1992:EPO

[442] A. Danthine. Esprit Project OSI 95. new transport services for highspeed networking. Computer Networks and ISDN Systems, 25(4–5): 384–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Sterba:1992:IER

[443] M. Sterba. ISDN in European research networking. Computer Networks and ISDN Systems, 25(4–5): 400–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kia:1992:RAS

[444] R. Najmabadi Kia and B. Sales. Routing architectures for the support of the OSI connection mode network service. Computer Networks and ISDN Systems, 25(4–5):405–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Husemann:1992:ICL

[445] D. Husemann. ISO CONS in LANs — making it all work. A European contribution to 4.4 BSD Unix. Computer Networks and ISDN Systems, 25 (4–5):411–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Mannie:1992:SLO

[446] E. Mannie and B. Sales. 3L, a software laboratory for the OSI lower layers. Computer Networks and ISDN Systems, 25(4–5):420–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Reijs:1992:RCC

[447] V. Reijs. RARE/COSINE connectionless mode network service pilot. Computer Networks and ISDN Systems, 25 (4–5):426–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bem:1992:NAC

[448] D. J. Bem. NASK — research and academic computer network in Poland. Computer Networks and ISDN Systems, 25(4–5):431–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gruntorad:1992:RAN

[449] J. Gruntorad. Research and academic, networking in the Czech and Slovak Federal Republic. Computer Networks and ISDN Systems, 25(4– 5):438–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Turchany:1992:NH

[450] G. Turchany. Networking in Hungary. Computer Networks and ISDN Systems, 25(4–5):444–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Morrow:1992:BIN

[451] T. Morrow. BIDS ISI: a new national bibliographic data service to the UK academic community. Computer Networks and ISDN Systems, 25(4– 5):448–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Berners-Lee:1992:WWW

[452] T. J. Berners-Lee, R. Cailliau, and J.- F. Groff. The World-Wide Web. Computer Networks and ISDN Systems, 25 (4–5):454–459, November 1992. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kelleher:1992:NIM

[453] L. A. Kelleher. Navigating the Internet — a merit network cruise. Computer Networks and ISDN Systems, 25 (4–5):460–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Doelz:1992:EPM

[454] R. Doelz. The EMBnet Project — European molecular biology network. Computer Networks and ISDN Systems, 25(4–5):464–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Muftic:1992:ICI

[455] S. Muftic. Implementation of the comprehensive Integrated Security System for computer networks. Computer Networks and ISDN Systems, 25 (4–5):469–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Purser:1992:CSP

[456] M. Purser. COSINE Sub-Project P8: security services. Computer Networks and ISDN Systems, 25(4–5): 476–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kirstein:1992:PAS

[457] P. T. Kirstein and P. Williams. Piloting authentication and security services within OSI applications for RTD information (PASSWORD). Computer Networks and ISDN Systems, 25(4– 5):483–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Wu:1992:MSC

[458] S. Wu. MHS security — a concise survey. Computer Networks and ISDN Systems, 25(4–5):490–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Lipp:1992:IP

[459] P. Lipp and R. Posch. An interbridge-security protocol. Computer Networks and ISDN Systems, 25(4– 5):496–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Fluckiger:1992:EIP

[460] F. Fluckiger. Euro-networking and industrial policies: lessons from the past. Computer Networks and ISDN Systems, 25(4–5):501–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Roberts:1992:URU

[461] M. M. Roberts. The university role in the United States National Research and Education Network. Computer Networks and ISDN Systems, 25(4– 5):512–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Davies:1992:TLA

[462] D. Davies. Is there life after COSINE. Computer Networks and ISDN Systems, 25(4–5):518–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Greisen:1992:OUR

[463] F. Greisen. The operational unit for research and academic networking in Europe. Computer Networks and ISDN Systems, 25(4–5):521–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Pinse:1992:EMG

[464] D. Pinse and P. Sylvester. An electronic mail gateway between EARN/ INTERNET and the IBM internal network. Computer Networks and ISDN Systems, 25(4–5):526–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Ashby:1992:FIM

[465] P. Ashby. FTAM interoperability: meeting the needs of the research community. Computer Networks and ISDN Systems, 25(4–5):531–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Alvestrand:1992:EMR

[466] H. T. Alvestrand. Electronic mail routing in a heterogeneous world. Computer Networks and ISDN Systems, 25 (4–5):535–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Wallace:1992:OM

[467] B. Wallace. OSI migration. Computer Networks and ISDN Systems, 25 (4–5):540–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bovio:1992:RE

[468] D. Bovio. The regionalization of EARN. Computer Networks and ISDN Systems, 25(4–5):546–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Berrino:1992:YEP

[469] C. Berrino and Manuello D. Y-NET — the Esprit pan-European community OSI network. Computer Networks and ISDN Systems, 25(4–5): 554–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kunft:1992:DAN

[470] W. Kunft. Development in academic networking in Austria. Computer Networks and ISDN Systems, 25(4– 5):561–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Wijgerde:1992:CQS

[471] J. Wijgerde. The COSINE quality of service project. Computer Networks and ISDN Systems, 25(4–5): 566–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Simoni:1992:IHD

[472] N. Simoni and S. Znaty. Interconnection of high-speed data networks: contribution of dynamic control and quality of service. Computer Networks and ISDN Systems, 25(4–5): 570–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Auge:1992:DMM

[473] E. Auge and P. Brun. Distributed management of an MHS environment. Computer Networks and ISDN Systems, 25(4–5):577–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Geruts:1992:FAD

[474] P. Geruts and G. Theis. A formal approach to the description of communications services: the case of the Space Data Network (SDN). Computer Networks and ISDN Systems, 25 (4–5):583–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Neggers:1992:NSE

[475] K. Neggers. Next steps for European networking. Computer Networks and ISDN Systems, 25(4–5): 592–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Dixon:1993:ORG

[476] T. Dixon. Obstacles on the road to global networking. Computer Networks and ISDN Systems, 25(Supplement 1): S9–??, September 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Computer Networks for Research in Europe.

Eppenberger:1993:XDS

[477] U. Eppenberger. X.500 directory service usage for X.400 e-mail. Computer Networks and ISDN Systems, 25(Supplement 1):S3–??, September 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Computer Networks for Research in Europe.

Malara:1993:CAC

[478] P. Malara, R. Tirler, and A. Trannoy. CLAN at CRS4: an experimental Public Switched Data Network. Computer Networks and ISDN Systems, 25(Supplement 1):S17–??, September 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Computer Networks for Research in Europe.

Nielsen:1993:EHS

[479] Svend Møller Nielsen. EuropaNET – contemporary high speed networking. Computer Networks and ISDN Systems, 25(Supplement 1):S25–??, September 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Computer Networks for Research in Europe.

Zegwaart:1993:PEM

[480] Erik Zegwaart. Privacy Enhanced Mail in more detail. Computer Networks and ISDN Systems, 25(Supplement 2):S63– S71, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL zegwaart@surfnet.
nl;   gopher://erasmus.rare.nl:
70/00cnre/cnre2/S063-S071.txt.

Sidhu:1993:AGE

[481] D. P. Sidhu. About the guest editor. Computer Networks and ISDN Systems, 25(6):598–??, January 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Clark:1993:AGT

[482] David D. Clark, Bruce S. Davie, David J. Farber, Inder S. Gopal, Bharath K. Kadaba, W. David Sincoskie, Jonathan M. Smith, and David L. Tennenhouse. The AU-RORA gigabit testbed. Computer Networks and ISDN Systems, 25(6):599– 621, January 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Rodeheffer:1993:EA

[483] T. L. Rodeheffer. Experience with Autonet. Computer Networks and ISDN Systems, 25(6):623–??, January 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Wang:1993:CSO

[484] W. Wang and F. A. Tobagi. The Christmas-tree switch: an output queuing space-division fast packet switch based on interleaving distribution and concentration functions. Computer Networks and ISDN Systems, 25 (6):631–??, January 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Maxemchuk:1993:DRH

[485] N. F. Maxemchuk. Dispersity routing in high-speed networks. Computer Networks and ISDN Systems, 25 (6):645–??, January 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Mitra:1993:DAW

[486] D. Mitra and J. B. Seery. Dynamic adaptive windows for high speed data networks with multiple paths and propagation delays. Computer Networks and ISDN Systems, 25(6):663–??, January 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Ko:1993:IAV

[487] K.-T. Ko, P. P. Mishra, and S. K. Tripathi. Interaction among virtual circuits using predictive congestion control. Computer Networks and ISDN Systems, 25(6):681–??, January 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chlamtac:1993:MHI

[488] I. Chlamtac and M. G. Kienzle. Multitasking in high-speed interconnection systems. Computer Networks and ISDN Systems, 25(6):701–??, January 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Quemada:1993:GE

[489] J. Quemada. Guest editorial. Computer Networks and ISDN Systems, 25 (7):719–??, February 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Chanson:1993:TSU

[490] S. T. Chanson, A. A. F. Loureiro, and S. T. Vuong. On tools supporting the use of formal description techniques in protocol development. Computer Networks and ISDN Systems, 25 (7):723–??, February 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Amer:1993:PVE

[491] P. D. Amer and D. New. Protocol visualization in Estelle. Computer Networks and ISDN Systems, 25 (7):741–??, February 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

DeNicola:1993:ABF

[492] R. De Nicola, A. Fantechi, S. Gnesi, and G. Ristori. An action-based framework for verifying logical and behavioural properties of concurrent systems. Computer Networks and ISDN Systems, 25(7):761–778, February 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Algayres:1993:VPA

[493] B. Algayres, V. Coelho, L. Doldi, H. Garavel, and Y. Lejeune. VESAR: a pragmatic approach to formal specification and verification. Computer Networks and ISDN Systems, 25(7): 779–??, February 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Miguel:1993:LBP

[494] C. Miguel, A. Fernandez, J. M. Ortuno, and L. Vidaller. A LOTOS based performance evaluation tool. Computer Networks and ISDN Systems, 25 (7):791–??, February 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Manas:1993:TSI

[495] J. A. Manas, T. De Miguel, J. Salvachua, and A. Azcorra. Tool support to implement LOTOS formal specifications. Computer Networks and ISDN Systems, 25(7):815–??, February 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Sijelmassi:1993:PDT

[496] R. Sijelmassi and B. Strausser. The PET and DINGO tools for deriving distributed implementations from Estelle. Computer Networks and ISDN Systems, 25(7):841–??, February 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Karjoth:1993:LLE

[497] G. Karjoth, C. Binding, and J. Gustafsson. LOEWE: A LOTOS engineering workbench. Computer Networks and ISDN Systems, 25(7):853–??, February 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Veltink:1993:PT

[498] G. Veltink. The PSF toolkit. Computer Networks and ISDN Systems, 25 (7):875–??, February 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Rangan:1993:VCF

[499] P. V. Rangan. Video conferencing, file storage, and management in multimedia computer systems. Computer Networks and ISDN Systems, 25(8):901–??, March 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Thachenkary:1993:ISD

[500] C. S. Thachenkary. Integrated Services Digital Networks (ISDN): six case study assessments of a commercial implementation. Computer Networks and ISDN Systems, 25(8):921–??, March 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Benford:1993:SOG

[501] S. Benford and J. Palme. A standard for OSI group communication. Computer Networks and ISDN Systems, 25(8):933–??, March 1993. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Grillo:1993:TTG

[502] D. Grillo, R. J. G. MacNamee, and B. Rashidzadeh. Towards third generation mobile systems: a European possible transition path. Computer Networks and ISDN Systems, 25(8):947– ??, March 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Jonsson:1993:AGE

[503] B. Jonsson, J. Parrow, and B. Pehrson. About the guest editors. Computer Networks and ISDN Systems, 25(9): 968–??, April 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gouda:1993:PVM

[504] M. G. Gouda. Protocol verification made simple: a tutorial. Computer Networks and ISDN Systems, 25(9):969–??, April 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Holzmann:1993:DVP

[505] G. J. Holzmann. Design and validation of protocols: a tutorial. Computer Networks and ISDN Systems, 25 (9):981–1017, April 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). also in: Proc. 11th PSTV91, INWG/IFIP, Stockholm, Sweden.

Partridge:1993:PHS

[506] Craig Partridge. Protocols for highspeed networks: some questions and a few answers. Computer Networks and ISDN Systems, 25(9):1019–1028, April 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Richards:1993:RSD

[507] P. S. Richards. Rapid service delivery and customization in a developing network infrastructure. Computer Networks and ISDN Systems, 25(10):1031–??, May 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Mizusawa:1993:EEI

[508] J.-I. Mizusawa and M. Akiyama. Experiences and expectations of introducing PSTN new services. Computer Networks and ISDN Systems, 25(10):1041–??, May 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Roberts:1993:TCB

[509] J. W. Roberts. Traffic control in the B-ISDN. Computer Networks and ISDN Systems, 25(10):1055–??, May 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Dziong:1993:EBA

[510] Z. Dziong, K.-Q. Liao, and L. Mason. Effective bandwidth allocation and buffer dimensioning in ATM based networks with priorities. Computer Networks and ISDN Systems, 25(10): 1065–??, May 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bolla:1993:DHC

[511] R. Bolla and F. Davoli. Dynamic hierarchical control of resource allocation in an integrated services broadband network. Computer Networks and ISDN Systems, 25(10):1079–??, May 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Popescu:1993:ASE

[512] A. Popescu and R. P. Singh. An alternative solution to the electrooptic and service bottleneck problems in integrated multi-gbit/s LANs: the SUPERLAN architecture. Computer Networks and ISDN Systems, 25 (10):1089–1105, May 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Yokotani:1993:SPI

[513] T. Yokotani, H. Sato, and S. Nakatsuka. A study on a performance improvement algorithm in DQDB MAN.

Computer Networks and ISDN Systems, 25(10):1107–1117, May 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Lombardo:1993:APM

[514] A. Lombardo, S. Palazzo, D. Panno, R. Pignatelli, and L. Susanna. An adaptive policing mechanism for a DQDB MAN. Computer Networks and ISDN Systems, 25(10):1119–1126, May 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Greaves:1993:CBN

[515] David J. Greaves and Krzysztof Zieliński. The Cambridge Backbone Network. An overview and preliminary performance. Computer Networks and ISDN Systems, 25(10):1127–1133, May 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

VanDijk:1993:ATC

[516] N. M. Van Dijk. On the arrival theorem for communication networks. Computer Networks and ISDN Systems, 25(10):1135–??, May 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Henderson:1993:FUE

[517] W. Henderson. Finding and using exact equilibrium distributions for stochastic Petri nets. Computer Networks and ISDN Systems, 25(10):1143–??, May 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Rosenberg:1993:PMH

[518] C. Rosenberg and A. Le Bon. Performance models for hybrid broadband networks. Computer Networks and ISDN Systems, 25(10):1155–??, May 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Sokolov:1993:STM

[519] N. Sokolov, I. Ekhriel, R. Rerle, and S. Brusilovsky. On some teletraffic models simplifications. Computer Networks and ISDN Systems, 25 (10):1165–1173, May 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Yap:1993:EED

[520] M. T. Yap and D. Hutchison. An emulator for evaluating DQDB performance. Computer Networks and ISDN Systems, 25(11):1177–1204, June 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Dobosiewicz:1993:TCP

[521] W. Dobosiewicz, P. Giburzynski, and P. Rudnicki. On two collision protocols for high speed bus LANs. Computer Networks and ISDN Systems, 25 (11):1205–1225, June 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Steyaert:1993:EDM

[522] B. Steyaert, H. Bruneel, G. H. Petit, and E. Desmet. End-to-end delays in multistage ATM switching networks: approximate analytic derivation of tail probabilities. Computer Networks and ISDN Systems, 25(11):1227–??, June 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Varvitsiotis:1993:EAF

[523] A. P. Varvitsiotis and G. I. Stassinopoulos. Extending ASN.1 into a fullfledged constraint language in the context of OSI protocol conformance testing. Computer Networks and ISDN Systems, 25(11):1243–??, June 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bruneel:1993:PDQ

[524] H. Bruneel. Packet delay and queue length for statistical multiplexers with low-speed access lines. Computer Networks and ISDN Systems, 25(12):1267– ??, July 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Bisdikian:1993:QMA

[525] C. Bisdikian. A queueing model with applications to bridges and the DQDB (IEEE 802.6) MAN. Computer Networks and ISDN Systems, 25 (12):1279–1289, July 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Nakajima:1993:COC

[526] A. Nakajima. Construction of optimal communication structures for weighted distributed match-making. Computer Networks and ISDN Systems, 25(12): 1291–??, July 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Biersack:1993:TCM

[527] E. W. Biersack and D. C. Feldmeier. A timer-based connection management protocol with synchronized clocks and its verification. Computer Networks and ISDN Systems, 25(12):1303–??, July 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Du:1993:PAI

[528] J. Du, R. H. Deng, and C. C. Ko. Performance analysis of interconnected LANs with server/client configuration. Computer Networks and ISDN Systems, 25(12):1321–??, July 1993. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Li:1993:OLB

[529] J. Li and H. Kameda. Optimal load balancing in tree networks with twoway traffic. Computer Networks and ISDN Systems, 25(12):1335–??, July 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Akyildiz:1993:TIA

[530] Ian F. Akyildiz, Izhak Rubin, and Kazem Sohraby. Teletraffic issues in ATM networks. Computer Networks and ISDN Systems, 26(1):1– 4, September 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1993&volume=26&
                                        issue=1&aid=1187.

Chlamtac:1993:CBC

[531] Imrich Chlamtac and Tao Zhang. A counter based congestion control (CBC) for ATM networks. Computer Networks and ISDN Systems, 26(1): 5–27, September 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1993&volume=26&
                                        issue=1&aid=1188.

Hahne:1993:DWF

[532] E. L. Hahne, C. R. Kalmanek, and S. P. Morgan. Dynamic window flow control on a high-speed widearea data network. Computer Networks and ISDN Systems, 26(1):29– 41, September 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=1&aid=1189. Summarized in

13th International Teletraffic Conference Proceedings, Copenhagen, Denmark, June 1991, pp. 867–872.

Sriram:1993:MBA

[533] K. Sriram. Methodologies for bandwidth allocation, transmission scheduling, and congestion avoidance in broadband ATM networks. Computer Networks and ISDN Systems, 26(1):43– 59, September 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=1&aid=1190.

Gun:1993:BMC

[534] Levent Gün and Roch Guérin. Bandwidth management and congestion control framework of the broadband network architecture. Computer Networks and ISDN Systems, 26(1):61– 78, September 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/

                                        browse.cgi?year=1993&volume=26&
                                        issue=1&aid=1191.

Wu:1993:DTA

[535] Guo-Liang Wu and Jon W. Mark. Discrete time analysis of leaky-bucket congestion control. Computer Networks and ISDN Systems, 26(1):79– 94, September 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1993&volume=26&
                                        issue=1&aid=1192.

Meyer:1993:DAS

[536] J. F. Meyer, S. Montagna, and R. Paglino. Dimensioning of an ATM switch with shared buffer and threshold priority. Computer Networks and ISDN Systems, 26(1):95– 108, September 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1993&volume=26&
                                        issue=1&aid=1193.

Zukerman:1993:FAN

[537] Moshe Zukerman and Sammy Chan. Fairness in ATM networks. Computer Networks and ISDN Systems, 26(1): 109–117, September 1, 1993. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1993&volume=26&
                                        issue=1&aid=1194.

Bonomi:1993:FLS

[538] Flavio Bonomi, Sergio Montagna, and Roberto Paglino. A further look at statistical multiplexing in ATM networks. Computer Networks and ISDN Systems, 26(1):119– 138, September 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=1&aid=1195.

Widjaja:1993:ECT

[539] I. Widjaja, A. Leon-Garcia, and H. T. Mouftah. The effect of cutthrough switching on the performance of buffered banyan networks. Computer Networks and ISDN Systems, 26 (1):139–159, September 1, 1993. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=1&aid=1196.

Zurfluh:1993:IZR

[540] E. A. Zurfluh, R. D. Cideciyan, P. Dill, R. Heller, W. Lemppenau, P. Mueller, H. R. Schindler, and P. Zafiropulo. The IBM Zurich Research Laboratory’s 1.13 Gb/s LAN/MAN prototype. Computer Networks and ISDN Systems, 26 (2):163–183, October 1, 1993. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=2&aid=1181.

Chang:1993:DAN

[541] Chung-Ju Chang, Jung-Hui Chiu, and Shyh-Jong Lin. The delay analysis in a noisy general cut-through switching network. Computer Networks and ISDN Systems, 26(2):185– 194, October 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1993&volume=26&
                                        issue=2&aid=1182.

Khasnabish:1993:NME

[542] Bhumip Khasnabish. A new method for evaluating packet routing policies in supra-high-speed metropolitan (or wide) area networks. Computer Networks and ISDN Systems, 26(2): 195–216, October 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1993&volume=26&
                                        issue=2&aid=1183.

Fumy:1993:PCK

[543] Walter Fumy and Matthias Leclerc. Placement of cryptographic key distribution within OSI: design alternatives and assessment. Computer Networks and ISDN Systems, 26(2): 217–225, October 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1993&volume=26&
                                        issue=2&aid=1184.

Tsin:1993:IDA

[544] Yung H. Tsin. Incremental distributed asynchronous algorithm for minimum spanning trees. Computer Networks and ISDN Systems, 26(2): 227–232, October 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=2&aid=1185.

Massart:1993:CPO

[545] T. Massart. A collision problem in OSI standard formal specifications. Computer Networks and ISDN Systems, 26 (2):233–238, October 1, 1993. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=2&aid=1186.

Bumbulis:1993:IOD

[546] P. J. Bumbulis, D. D. Cowan, C. M. Durance, and T. M. Stepien. An introduction to the OSI Directory Services. Computer Networks and ISDN Systems, 26(2):239– 249, October 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=2&aid=1198.

Kure:1993:XA

[547] Øivind Kure and Ingvild Sorteberg. XTP over ATM. Computer Networks and ISDN Systems, 26(3):253–262 (or 253–267??), November 1, 1993. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=3&aid=1269.

Barry:1993:IPS

[548] Susan Barry, Patricia McQuillan, Michael Purser, and Jonathan Moffett. Implementing and proving security services for the RARE/ COSINE community. Computer Networks and ISDN Systems, 26(3):263– 267, November 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=3&aid=1267.

Cooper:1993:SGH

[549] Robert Cooper. SuperJANET. the gestation of a high performance national research network. Computer Networks and ISDN Systems, 26(3): 269–274, November 1, 1993. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=3&aid=1272.

Handley:1993:MIC

[550] Mark J. Handley, Peter T. Kirstein, and M. Angela Sasse. Multimedia integrated conferencing for European researchers (MICE): piloting activities and the conference management and multiplexing centre. Computer Networks and ISDN Systems, 26(3):

275–290, November 1, 1993. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=3&aid=1259.

Tuck:1993:UXD

[551] Bill Tuck and Ruth Moulton. Using X.400 for document delivery. Computer Networks and ISDN Systems, 26 (3):291–296, November 1, 1993. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=3&aid=1270.

Stockman:1993:GCG

[552] Bernhard Stockman. Global connectivity — the Global Internet eXchange (GIX). Computer Networks and ISDN Systems, 26(3):297– 303, November 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=3&aid=1258.

Huizer:1993:REU

[553] Erik Huizer, Janneke Abbema, Maria Heijne, and Ton Verschuren. Reaching out to the end user. Computer Networks and ISDN Systems, 26(3): 305–310, November 1, 1993. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/

                                        browse.cgi?year=1993&volume=26&
                                        issue=3&aid=1257.

Plagemann:1993:XNT

[554] Thomas Plagemann and Bernhard Plattner. X.500 for non-technical users. Computer Networks and ISDN Systems, 26(3):311–315, November 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/comnet/
                                        cas_sub/browse/browse.cgi?year=
                                        1993&volume=26&issue=3&aid=1266.

Afifi:1993:HDB

[555] Hossam Afifi. A hierarchical directory based X.400 server. Computer Networks and ISDN Systems, 26(3): 316–325, November 1, 1993. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1993&volume=26&
                                        issue=3&aid=1265.

Mansfield:1993:MCN

[556] Glenn Mansfield, K. Jayanthi, K. Higuchi, Thomas Johannsen, Y. Nemoto, and S. Noguchi. Mapping communication networks in the directory. Computer Networks and ISDN Systems, 26(3): 327–335, November 1, 1993. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1993&volume=26&
                                        issue=3&aid=1264.

Colin:1993:LAO

[557] Michel Colin and Bernard Sales. (N)- LMP: an architecture for the OSI lower layers management. Computer Networks and ISDN Systems, 26(3): 337–342, November 1, 1993. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=3&aid=1260.

Bovio:1993:MEN

[558] Daniele Bovio and Greg Lloyd. Monitoring EARN networking services. Computer Networks and ISDN Systems, 26(3):343–348, November 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/comnet/
cas_sub/browse/browse.cgi?year=
1993&volume=26&issue=3&aid=1263.

Myles:1993:CFI

[559] Andrew Myles and David Skellern. Comparing four IP based mobile host protocols. Computer Networks and ISDN Systems, 26(3):349– 355, November 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=3&aid=1268.

Perkins:1993:PCN

[560] Charles Perkins. Providing continuous network access to mobile hosts using TCP/IP. Computer Networks and ISDN Systems, 26(3):357– 369, November 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/

                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1993&volume=26&
                                        issue=3&aid=1262.

Sanghi:1993:SND

[561] Dheeraj Sanghi, Olafur Gudmundsson, and Ashok K. Agrawala. Study of network dynamics. Computer Networks and ISDN Systems, 26(3):371– 378, November 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1993&volume=26&
                                        issue=3&aid=1261.

Monteiro:1993:FAL

[562] E. Monteiro, F. Boavida, and V. Freitas. A fairness analysis of LAN/ WAN protocol relays. Computer Networks and ISDN Systems, 26(3):379– 388, November 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1993&volume=26&
                                        issue=3&aid=1271.

Younger:1993:MBC

[563] E. J. Younger, K. H. Bennett, and R. Hartley-Davies. A model for a broadband cellular wireless network for digital communications. Computer Networks and ISDN Systems, 26(4): 391–402, December 1, 1993. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1993&volume=26&
                                        issue=4&aid=1197.

Bakker:1993:PRS

[564] Erwin M. Bakker, Jan van Leeuwen, and Richard B. Tan. Prefix routing schemes in dynamic networks. Computer Networks and ISDN Systems, 26 (4):403–421, December 1, 1993. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=4&aid=1199.

Pitts:1993:CCN

[565] Robert A. Pitts, Ralph Martinez, and Larry C. Schooley. CSMA/CD with network partitioning. Computer Networks and ISDN Systems, 26(4):423– 432, December 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=4&aid=1200.

Gupta:1993:SMA

[566] Anil K. Gupta, Luis Orozco Barbosa, and N. D. Georganas. Switching modules for ATM switching systems and their interconnection networks. Computer Networks and ISDN Systems, 26 (4):433–445, December 1, 1993. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1993&volume=26&
issue=4&aid=1201.

Moharram:1993:CNM

[567] Omayma E. Moharram. Customer network management systems: load and performance. Computer Networks and ISDN Systems, 26(4):447– 457, December 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1993&volume=26&
                                        issue=4&aid=1202.

Jiang:1993:HSI

[568] Xi Jiang and James S. Meditch. A high-speed integrated services ATM/STM switch. Computer Networks and ISDN Systems, 26(4):459– 477, December 1, 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1993&volume=26&
                                        issue=4&aid=1203.

Baker:1994:SWS

[569] T. Baker, I. Bruning, L. Klein, and M. Lenz. Starting a Web site: issues of quality control. Computer Networks and ISDN Systems, 26(Supplement 2– 3):S55–S61, ???? 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Crowcroft:1994:TNG

[570] Jon Crowcroft and Gordon Joly. Touring and navigating a global learning environment — the Web. Computer Networks and ISDN Systems, 26 (Supplement 2–3):S71–S74, ???? 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Huston:1994:DSD

[571] G. Huston. Desperately seeking default: Internet policy update: A perspective from the Pacific. Computer Networks and ISDN Systems, 26 (Supplement 1):S3–??, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Computer Networks for Research in Europe.

Morrow:1994:PNP

[572] T. M. Morrow and A. M. Mumford. Publishing on the network. Computer Networks and ISDN Systems, 26 (Supplement 1):S15–??, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Computer Networks for Research in Europe.

Muller:1994:UUD

[573] P. Muller and H. P. Grossmann. Udine: A universal document information and navigation entry. Computer Networks and ISDN Systems, 26(Supplement 1): S29–??, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Computer Networks for Research in Europe.

Touvet:1994:MWU

[574] J.-C. Touvet and P.-A Pays. A multimedia WPS using X.500 /WWW integration. Computer Networks and ISDN Systems, 26(Supplement 1):S39– ??, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Computer Networks for Research in Europe.

Owen:1994:SDH

[575] Henry L. Owen and Thomas M. Klett. Synchronous digital hierarchy network pointer simulation. Computer Networks and ISDN Systems, 26(5): 481–491, January 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=5&aid=1204.

Rao:1994:MPE

[576] B. Madhu Rao and Sub Ramakrishnan. Models for performance evaluation of a local area network with a file server. Computer Networks and ISDN Systems, 26(5):493– 509, January 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=5&aid=1205.

Pham:1994:CCI

[577] Xuan Huy Pham and Richard Betts. Congestion control for intelligent networks. Computer Networks and ISDN Systems, 26(5):511–524, January 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/comnet/
cas_sub/browse/browse.cgi?year=
1994&volume=26&issue=5&aid=1206.

Chang:1994:SOQ

[578] Chung-Ju Chang, Po-Chou Lin, and Jia-Ming Chen. Study on optimal queue-length-threshold scheduling policy for an ATM multiplexer with finite buffers and batch Poisson arrivals. Computer Networks and ISDN Systems, 26(5):525– 540, January 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/

tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=5&aid=1207.

Wang:1994:TOS

[579] Jonathan L. Wang and John A. Silvester. Throughput optimization in single commodity multihop packet radio networks. Computer Networks and ISDN Systems, 26(5):541– 562, January 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=5&aid=1209.

Kamal:1994:HMA

[580] A. E. Kamal and B. W. Abeysundara. Hybrid medium access protocols for high-speed, dual-bus fiberoptic local area networks. Computer Networks and ISDN Systems, 26(5): 563–583, January 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=5&aid=1210.

Bondi:1994:ICL

[581] Andre B. Bondi and Wai-Sum Lai. The influence of cell loss patterns and overheads on retransmission choices in broadband ISDN. Computer Networks and ISDN Systems, 26(5):585– 598, January 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/

                                        browse.cgi?year=1994&volume=26&
                                        issue=5&aid=1212.

vanAs:1994:MATa

[582] Harmen R. van As and Harry Rudin. Media-access techniques for high-speed LANs and MANs. Computer Networks and ISDN Systems, 26(6–8): 601–602, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=26&
                                        issue=6-8&aid=1309.

vanAs:1994:MATb

[583] Harmen R. van As. Media access techniques: The evolution towards terabit/s LANs and MANs. Computer Networks and ISDN Systems, 26(6– 8):603–656, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=26&
                                        issue=6-8&aid=1318.

Davids:1994:FSP

[584] Peter Davids, Thomas Meuser, and Otto Spaniol. FDDI: status and perspectives. Computer Networks and ISDN Systems, 26(6–8):657– 677, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=26&
                                        issue=6-8&aid=1219.

Martini:1994:COD

[585] Peter Martini. Connection oriented data service in DQDB. Computer Networks and ISDN Systems, 26(6– 8):679–694, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1224.

Kamal:1994:ASR

[586] Ahmed E. Kamal, Johnny W. Wong, and Hossam S. Hassanein. An algorithm for slot reuse in DQDB networks with erasure nodes. Computer Networks and ISDN Systems, 26(6– 8):695–710, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1238.

Conti:1994:DED

[587] M. Conti, E. Gregori, and L. Lenzini. E-DCP, an extension of the distributedcontrol polling MAC protocol (DCP) for integrated services. Computer Networks and ISDN Systems, 26(6–8): 711–719, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1221.

Mukherjee:1994:CBA

[588] Biswanath Mukherjee and Ahmed E. Kamal. The continuation-bit approach and the p_{i}-persistent protocol for scheduling variable-length messages on slotted, high-speed, fiber optic LANs/MANs. Computer Networks and ISDN Systems, 26(6–8): 721–744, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1234.

Watson:1994:MP

[589] Greg Watson. The S ++ MAC protocol. Computer Networks and ISDN Systems, 26(6–8):745–755, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/cgi-bin/
cas/tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1218.

Baiocchi:1994:ACC

[590] A. Baiocchi, L. Gratta, M. Listanti, G. Pacifici, A. Roveri, and R. Winkler. The adaptive cycle cell insertion MAC protocol for high throughput and fair multiaccess networks. Computer Networks and ISDN Systems, 26 (6–8):757–770, March 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1222.

Adams:1994:O

[591] J. L. Adams. Orwell. Computer Networks and ISDN Systems, 26(6– 8):771–784, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1220.

Imai:1994:AAT

[592] Kazuo Imai, Tadashi Ito, Hideki Kasahara, and Naotaka Morita. ATMR: Asynchronous transfer mode ring protocol. Computer Networks and ISDN Systems, 26(6–8):785– 798, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1223.

Slosiar:1994:MDQ

[593] Rasti Slosiar, Martin Potts, and Reto Beeler. MD^{3}Q: A distributed queueing protocol with full channel capacity reuse and guarantee of bandwidth. Computer Networks and ISDN Systems, 26 (6–8):799–815, March 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1217.

Ofek:1994:OMA

[594] Yoram Ofek. Overview of the MetaRing architecture. Computer Networks and ISDN Systems, 26(6–8): 817–829, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/

browse.cgi?year=1994&volume=26&
issue=6-8&aid=1216.

vanAs:1994:CIM

[595] H. R. Van As, W. W. Lemppenau, H. R. Schindler, and P. Zafiropulo. CRMA-II: A MAC protocol for ringbased Gb/s LANs and MANs. Computer Networks and ISDN Systems, 26 (6–8):831–840, March 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1319.

Brassil:1994:MSN

[596] Jack Brassil, Abhijit K. Choudhury, and Nicholas F. Maxemchuk. The Manhattan Street Network: a high performance, highly reliable metropolitan area network. Computer Networks and ISDN Systems, 26(6–8): 841–858, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1233.

Ofek:1994:RFC

[597] Yoram Ofek and Moti Yung. Routing and flow control on the MetaNet: an overview. Computer Networks and ISDN Systems, 26(6–8):859– 872, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1239.

Marsan:1994:PTA

[598] M. Ajmone Marsan, G. Albertengo, C. Casetti, F. Neri, and G. Panizzardi. On the performance of topologies and access protocols for highspeed LANs and MANs. Computer Networks and ISDN Systems, 26(6– 8):873–894, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1229.

Huang:1994:CAT

[599] Hung Khei Huang and Tatsuya Suda. Collision avoidance tree networks. Computer Networks and ISDN Systems, 26(6–8):895– 911, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1230.

Sudhakar:1994:APP

[600] G. N. M. Sudhakar, M. Kavehrad, and N. D. Georganas. Access protocols for passive optical star networks. Computer Networks and ISDN Systems, 26 (6–8):913–930, March 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1231.

Karol:1994:CAW

[601] Mark J. Karol and Bernard Glance.

A collision-avoidance WDM optical star network. Computer Networks and ISDN Systems, 26(6–8): 931–943, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1226.

Chlamtac:1994:QSP

[602] Imrich Chlamtac and Andrea Fumagalli. Quadro: A solution to packet switching in optical transmission networks. Computer Networks and ISDN Systems, 26(6–8): 945–963, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1232.

Gerla:1994:OTT

[603] Mario Gerla, Milan Kovačević, and Joseph Bannister. Optical tree topologies: Access control and wavelength assignment. Computer Networks and ISDN Systems, 26(6–8): 965–983, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1227.

Fratta:1994:RAC

[604] Luigi Fratta, Flaminio Borgonovo, Joseph Bannister, and Mario Gerla. Routing and admission control in the multihop wavelength-division optical network. Computer Networks and ISDN Systems, 26(6–8):985– 1005, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1225.

Sen:1994:CSS

[605] A. Sen and P. Maitra. A comparative study of Shuffle-Exchange, Manhattan Street and Supercube network for lightwave applications. Computer Networks and ISDN Systems, 26(6– 8):1007–1022, March 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1228.

Decina:1994:TST

[606] M. Dècina, V. Trecordi, G. Zanolini, and D. Zucca. Two simple techniques for broadcasting in deflection routing multichannel MANs. Computer Networks and ISDN Systems, 26(6– 8):1023–1041, March 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=6-8&aid=1236.

Elby:1994:WBC

[607] Stuart D. Elby and Anthony S. Acampora. Wavelength-based cellswitching in ATM multihop lightwave networks. Computer Networks and ISDN Systems, 26(6–8):1043– 1062, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=26&
                                        issue=6-8&aid=1235.

Kwong:1994:UAO

[608] Wing C. Kwong and Paul R. Prucnal. Ultrafast all-optical codedivision multiple-access (CDMA) fiberoptic networks. Computer Networks and ISDN Systems, 26(6–8):1063– 1086, March 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=26&
                                        issue=6-8&aid=1237.

Saito:1994:TAP

[609] Hiroshi Saito and Takuya Asaka. Traffic aspect of personal telecommunications in intelligent networks. Computer Networks and ISDN Systems, 26(9):1089–1099, June 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=26&
                                        issue=9&aid=1213.

Bolognesi:1994:GLG

[610] T. Bolognesi, E. Najm, and P. A. J. Tilanus. G-Lotos: A graphical language for concurrent systems. Computer Networks and ISDN Systems, 26(9):1101–1127, June 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/

tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=9&aid=1214.

Cheng:1994:BAB

[611] Tee-Hiang Cheng. Bandwidth allocation in B-ISDN. Computer Networks and ISDN Systems, 26(9): 1129–1142, June 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=9&aid=1215.

Faergemand:1994:IS

[612] Ove Færgemand and Anders Olsen. Introduction to SDL-92. Computer Networks and ISDN Systems, 26(9): 1143–1167, June 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=9&aid=1240.

Yegenoglu:1994:CMA

[613] Ferit Yegenoglu and Bijan Jabbari. Characterization and modeling of aggregate traffic for finite buffer statistical multiplexers. Computer Networks and ISDN Systems, 26(9): 1169–1185, June 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=9&aid=1242.

Badran:1994:ASA

[614] Hosein F. Badran and H. T. Mouftah. ATM switch architectures with inputoutput-buffering: effect of input traffic correlation, contention resolution policies, buffer allocation strategies and delay in backpressure signal. Computer Networks and ISDN Systems, 26(9):1187–1213, June 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=9&aid=1243.

Chowdhury:1994:BAA

[615] Shyamal Chowdhury and Kazem Sohraby. Bandwidth allocation algorithms for packet video in ATM networks. Computer Networks and ISDN Systems, 26(9):1215– 1223, June 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=9&aid=1244.

Fayet:1994:HSS

[616] C. Fayet, A. Jacques, and G. Pujolle. High speed switching for ATM: the BSS. Computer Networks and ISDN Systems, 26(9): 1225–1234, June 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=9&aid=1245.

Leblanc:1994:TEM

[617] Larry J. Leblanc and Sridhar Narasimhan. Topological expansion of metropolitan area networks. Computer Networks and ISDN Systems, 26(9): 1235–1248, June 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=9&aid=1246.

Binding:1994:MVM

[618] C. Binding and H. Rudin. A multidimensional view of multimedia. Computer Networks and ISDN Systems, 26(10):1251–??, July 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Hopper:1994:CD

[619] Andy Hopper. Communications at the desktop. Computer Networks and ISDN Systems, 26(10): 1253–1265, July 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=10&aid=1256.

Ferrari:1994:NSM

[620] Domenico Ferrari, Anindo Banerjea, and Hui Zhang. Network support for multimedia. A discussion of the Tenet Approach. Computer Networks and ISDN Systems, 26(10): 1267–1280, July 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/

tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=10&aid=1255.

Jeffay:1994:TDM

[621] K. Jeffay, D. L. Stone, and F. D. Smith. Transport and display mechanisms for multimedia conferencing across packet-switched networks. Computer Networks and ISDN Systems, 26 (10):1281–1304, July 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=10&aid=1252.

Ramanathan:1994:TPM

[622] Srinivas Ramanathan, Harrick M. Vin, and P. Venkat Rangan. Towards personalized multimedia dialup services. Computer Networks and ISDN Systems, 26(10):1305– 1322, July 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=10&aid=1254.

Coulson:1994:MKS

[623] Geoff Coulson, Gordon S. Blair, and Philippe Robin. Micro-kernel support for continuous media in distributed systems. Computer Networks and ISDN Systems, 26(10): 1323–1341, July 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/

browse.cgi?year=1994&volume=26&
issue=10&aid=1253.

Muftic:1994:SAO

[624] Sead Muftic. Security architecture for ODP systems. final results of the CEC COST-11 Ter “Security” Project. Computer Networks and ISDN Systems, 26(11): 1343–1349, August 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=11&aid=1304.

Sole-Pareta:1994:BCA

[625] Josep Solé-Pareta and Jordi Domingo-Pascual. Burstiness characterization of ATM cell streams. Computer Networks and ISDN Systems, 26(11): 1351–1363, August 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=11&aid=1247.

Barker:1994:EPW

[626] Paul Barker. Experiences in providing a white pages directory service based on the X.500 standard. Computer Networks and ISDN Systems, 26 (11):1365–1374, August 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=11&aid=1248.

Simha:1994:RGA

[627] Rahul Simha and Yoram Ofek. Reducing global address recognition delays in local area networks with spatial bandwidth reuse. Computer Networks and ISDN Systems, 26(11): 1375–1384, August 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=11&aid=1249.

Cho:1994:PAR

[628] Young-Jong Cho and Chong-Kwan Un. Performance analysis of reconstruction algorithms for packet voice communications. Computer Networks and ISDN Systems, 26(11): 1385–1408, August 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=11&aid=1250.

LeBoudec:1994:CDS

[629] Jean-Yves Le Boudec, Andreas Meier, Rainer Oechsle, and Hong Linh Truong. Connectionless data service in an ATM-based customer premises network. Computer Networks and ISDN Systems, 26(11): 1409–1424, August 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=11&aid=1251.

Raghavan:1994:EIN

[630] S. V. Raghavan and R. Sriram. Example Implementation of Name Services using X.500 Directory. Computer Networks and ISDN Systems, 26 (11):1425–1436, August 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=11&aid=1280.

Khanna:1994:IPE

[631] Vimal K. Khanna and Samar Singh. An improved “Piggyback Ethernet’’ protocol and its analysis. Computer Networks and ISDN Systems, 26(11): 1437–1446, August 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=11&aid=1281.

Born:1994:PSN

[632] Richard G. Born. Packet-switching network performance under real-time voice-call loading. Computer Networks and ISDN Systems, 26(11): 1447–1456, August 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=11&aid=1282.

Rubin:1994:IRF

[633] Izhak Rubin and K. David Lin. Input rate flow control for high-speed communication networks: Blocking and delay at the access points. Computer Networks and ISDN Systems, 26 (11):1457–1473, August 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=11&aid=1284.

Kamal:1994:MTR

[634] Ahmed E. Kamal. The multi-token ring network protocol. Computer Networks and ISDN Systems, 26(12):1477– 1494, September 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=12&aid=1283.

Glitho:1994:SSN

[635] Roch H. Glitho. Signalling system number 7 network services part and X.25: A comparative study. Computer Networks and ISDN Systems, 26(12): 1495–1502, September 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=12&aid=1286.

Stavrakakis:1994:MSP

[636] Ioannis Stavrakakis and Sophia Tsakiridou. A Markov service policy with application to the queueing study of a DQDB station. Computer Networks and ISDN Systems, 26(12):1503– 1522, September 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=12&aid=1287.

Vaughn:1994:PMI

[637] Rayford B. Vaughn, Hossein Saiedian, and Elizabeth A. Unger. A proposed mechanism for implementation of non-discretionary access controls in a network environment. Computer Networks and ISDN Systems, 26(12): 1523–1531, September 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=12&aid=1288.

Khanna:1994:TMX

[638] Vimal K. Khanna. Timer management in X.25 layer 2 — An interpretation. Computer Networks and ISDN Systems, 26(12):1533– 1538, September 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=26&
issue=12&aid=1289.

Katsavos:1994:SFR

[639] Panos Katsavos and Vijay Varadharajan. A secure Frame Relay service. Computer Networks and ISDN Systems, 26(12):1539– 1558, September 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/

                                        browse.cgi?year=1994&volume=26&
                                        issue=12&aid=1290.

Youn:1994:PAP

[640] Chan Jung Youn and Kwan Un Chong. Performance analysis of packet switches with input and output buffers. Computer Networks and ISDN Systems, 26(12):1559– 1580, September 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=26&
                                        issue=12&aid=1291.

Heijink:1994:FGP

[641] Rudolf Jan Heijink. FAITH, a general purpose protocol test system for ISDN. Computer Networks and ISDN Systems, 26(12):1581– 1593, September 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=26&
                                        issue=12&aid=1292.

Peyravian:1994:ISR

[642] Mohammad Peyravian. An improved selective repeat protocol and its performance in high-speed environments. Computer Networks and ISDN Systems, 26(12):1595– 1605, September 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=26&
                                        issue=12&aid=1293.

Ammar:1994:BIS

[643] Mostafa H. Ammar, Victor O. K. Li, and Mehmet Ulema. Broadband ISDN: Standards, switches, and traffic management. Computer Networks and ISDN Systems, 27(1): 1–3, October 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=1&aid=1295.

Kim:1994:ISB

[644] Jaime Bae Kim, Tatsuya Suda, and Masaaki Yoshimura. International standardization of B-ISDN. Computer Networks and ISDN Systems, 27(1):5–27, October 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=1&aid=1296.

Kim:1994:DFT

[645] Hyong S. Kim. Design of a faulttolerant multichannel ATM switch for BISDN. Computer Networks and ISDN Systems, 27(1):29–43, October 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/cgi-bin/
cas/tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=1&aid=1297.

Huang:1994:PAP

[646] Tien-Yu Huang and Jean-Lien C. Wu. Performance analysis of prioritized state-dependent buffer-management schemes in ATM networks. Computer Networks and ISDN Systems, 27(1):45–66, October 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=1&aid=1298.

Chen:1994:FFT

[647] Wen-Tsuen Chen and Uan-Jiun Liu. A feasible framework of traffic control on an ATM wide-area network. Computer Networks and ISDN Systems, 27(1):67–84, October 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=1&aid=1299.

Khan:1994:TCA

[648] Irfan Khan and Victor O. K. Li. Traffic control in ATM networks. Computer Networks and ISDN Systems, 27(1):85–100, October 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=1&aid=1300.

Huang:1994:SRP

[649] Nen-Fu Huang, Chiung-Shien Wu, and Yi-Jang Wu. Some routing problems on Broadband ISDN. Computer Networks and ISDN Systems, 27(1): 101–116, October 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/

tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=1&aid=1301.

Crocetti:1994:SMS

[650] P. Crocetti, L. Fratta, M. Gerla, and M. A. Marsiglia. SMDS multicast support in ATM networks. Computer Networks and ISDN Systems, 27 (1):117–132, October 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=1&aid=1302.

Raggett:1994:RHD

[651] David Raggett. A review of the HTML+ document format. Computer Networks and ISDN Systems, 27(2): 135–145, November 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=2&aid=1382.

Luotonen:1994:WWW

[652] Ari Luotonen and Kevin Altis. World-Wide Web proxies. Computer Networks and ISDN Systems, 27(2):147– 154, November 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=2&aid=1380.

Katz:1994:SHS

[653] Eric Dean Katz, Michelle Butler, and Robert McGrath. A scalable HTTP server: The NCSA prototype. Computer Networks and ISDN Systems, 27 (2):155–164, November 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=2&aid=1384.

Glassman:1994:CRW

[654] Steven Glassman. A caching relay for the World Wide Web. Computer Networks and ISDN Systems, 27(2): 165–173, November 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=2&aid=1386.

Koster:1994:AAL

[655] Martijn Koster. ALIWEB — Archielike indexing in the WEB. Computer Networks and ISDN Systems, 27(2): 175–182, November 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=2&aid=1385.

DeBra:1994:IRW

[656] P. M. E. De Bra and R. D. J. Post. Information retrieval in the World-Wide Web: Making client-based searching feasible. Computer Networks and ISDN Systems, 27(2):183– 192, November 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/

tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=2&aid=1381.

Fielding:1994:MDH

[657] Roy T. Fielding. Maintaining distributed hypertext infostructures: welcome to MOMspider’s Web. Computer Networks and ISDN Systems, 27 (2):193–204, November 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=2&aid=1390.

Rousseau:1994:WDP

[658] Bertrand Rousseau and Mario Ruggier. Writing documents for paper and WWW. A strategy based on FrameMaker and WebMaker. Computer Networks and ISDN Systems, 27 (2):205–214, November 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=2&aid=1383.

Drakos:1994:THP

[659] Nikos Drakos. From text to hypertext: A post-hoc rationalisation of LaTEX2HTML. Computer Networks and ISDN Systems, 27(2):215– 224, November 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=2&aid=1394.

Whitcroft:1994:TWD

[660] Andy Whitcroft and Tim Wilkinson. A tangled Web of Deceit. Computer Networks and ISDN Systems, 27(2): 225–234, November 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=27&
                                        issue=2&aid=1389.

Takada:1994:MIE

[661] Toshihiro Takada. Multilingual information exchange through the World-Wide Web. Computer Networks and ISDN Systems, 27(2):235– 241, November 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=27&
                                        issue=2&aid=1387.

Pitkow:1994:RFW

[662] James Pitkow and Mimi Recker. Results from the First World-Wide Web user survey. Computer Networks and ISDN Systems, 27(2):243– 254, November 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=27&
                                        issue=2&aid=1393.

Ibrahim:1994:WWA

[663] Bertrand Ibrahim. World-wide algorithm animation. Computer Networks and ISDN Systems, 27(2):255– 265, November 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=2&aid=1392.

Mascha:1994:IET

[664] Michael Mascha and Gary Seaman. Interactive education: Transitioning CD-ROMs to the Web. Computer Networks and ISDN Systems, 27(2): 267–272, November 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=2&aid=1378.

Putz:1994:IIS

[665] Steve Putz. Interactive information services using World-Wide Web hypertext. Computer Networks and ISDN Systems, 27(2):273– 280, November 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=2&aid=1391.

Eichmann:1994:ISD

[666] David Eichmann, Terry McGregor, and Dann Danley. Integrating structured databases into the Web: The MORE system. Computer Networks and ISDN Systems, 27(2):281– 288, November 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/

                                        browse.cgi?year=1994&volume=27&
                                        issue=2&aid=1388.

Slater:1994:CW

[667] A. F. Slater. Controlled by the Web. Computer Networks and ISDN Systems, 27(2):289–295, November 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/comnet/
                                        cas_sub/browse/browse.cgi?year=
                                        1994&volume=27&issue=2&aid=1395.

Jones:1994:DWW

[668] Russ Jones. Digital’s World-Wide Web server: A case study. Computer Networks and ISDN Systems, 27 (2):297–306, November 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=27&
                                        issue=2&aid=1379.

Rzepa:1994:WCW

[669] Henry Rzepa. WWW94 Chemistry Workshop. Computer Networks and ISDN Systems, 27(2):317– 318, November 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=27&
                                        issue=2&aid=1398.

Plattner:1994:GE

[670] B. R. Plattner. Guest editorial. Computer Networks and ISDN Systems, 27 (3):343–??, December 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Isaacs:1994:ANT

[671] Margaret Isaacs. Approaches to network training with particular reference to a perceived need for selfhelp materials. Computer Networks and ISDN Systems, 27(3):345– 352, December 16, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=3&aid=1353.

Kelly:1994:BIP

[672] Brian Kelly. Becoming an information provider on the World Wide Web. Computer Networks and ISDN Systems, 27(3):353–360, December 16, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=3&aid=1354.

Weider:1994:WBU

[673] Chris Weider. Wild beasts and unapproachable bogs. Computer Networks and ISDN Systems, 27(3):361– 366, December 16, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=3&aid=1355.

Braid:1994:BEE

[674] Andrew Braid. From Babel to EDIL: the evolution of a standard for document delivery. Computer Networks and ISDN Systems, 27(3):367– 374, December 16, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=3&aid=1356.

Shrikumar:1994:TLE

[675] H. Shrikumar and Rehmi Post. Thinternet: life at the end of a tether. Computer Networks and ISDN Systems, 27 (3):375–385, December 16, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=3&aid=1357.

Pullen:1994:NDV

[676] J. Mark Pullen. Networking for Distributed Virtual Simulation. Computer Networks and ISDN Systems, 27(3): 387–394, December 16, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=3&aid=1358.

Turkheimer:1994:PIN

[677] Frank Türkheimer. Privacy and the Internet: The next step. Computer Networks and ISDN Systems, 27(3): 395–402, December 16, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=3&aid=1359.

Hallgren:1994:FIP

[678] Martyne M. Hallgren. Funding an Internet public good: definition and example. Computer Networks and ISDN Systems, 27(3):403– 409, December 16, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=3&aid=1360.

Ingvarson:1994:ENS

[679] Daniel Ingvarson, Dora Marinova, and Peter Newman. Electronic networking: Social and policy aspects of a rapidly growing technology. electronic networking: Policy aspects for Australia. Computer Networks and ISDN Systems, 27 (3):411–418, December 16, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=3&aid=1361.

Kuo:1994:IAN

[680] Franklin Kuo, James Ding, Cindy Zheng, and Farooq Hussain. Issues in academic networking in the PRC. Computer Networks and ISDN Systems, 27(3):419–427, December 16, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=3&aid=1362.

Jerman-Blazic:1994:TSI

[681] Borka Jerman-Blažič. Tool supporting the internationalisation of the generic network services. Computer Networks and ISDN Systems, 27(3): 429–435, December 16, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=27&
                                        issue=3&aid=1363.

Francis:1994:CGP

[682] Paul Francis. Comparison of geographical and provider-rooted Internet addressing. Computer Networks and ISDN Systems, 27(3):437– 448, December 16, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=27&
                                        issue=3&aid=1364.

Mathis:1994:WPI

[683] Matthew Mathis. Windowed ping: an IP layer performance diagnostic. Computer Networks and ISDN Systems, 27 (3):449–459, December 16, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=27&
                                        issue=3&aid=1365.

Bonetti:1994:DRM

[684] Pierluigi Bonetti, Claudio Allocchio, and Antonia Ghiselli. Distribution of RFC 1327 mapping rules via the Internet DNS: the INFNet distributed gateway system. Computer Networks and ISDN Systems, 27(3):461– 469, December 16, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=3&aid=1366.

Hoffman:1994:VVI

[685] Eric Hoffman, Allison Mankin, Maryann Perez, and S. J. Marsh. Vince: Vendor independent (and architecture flexible) network control. Computer Networks and ISDN Systems, 27(3):471– 478, December 16, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=3&aid=1367.

Perkins:1994:IMH

[686] Charles Perkins, Andrew Myles, and David B. Johnson. IMHP: A mobile host protocol for the Internet. Computer Networks and ISDN Systems, 27 (3):479–491, December 16, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1994&volume=27&
issue=3&aid=1368.

Demizu:1994:DVT

[687] Noritoshi Demizu and Suguru Yamaguchi. DDT — A versatile tunneling technology. Computer Networks and ISDN Systems, 27(3):493– 503, December 16, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=27&
                                        issue=3&aid=1369.

Buddhikot:1994:DLS

[688] Milind M. Buddhikot, Guru M. Parulkar, and Jerome R. Cox, Jr. Design of a large scale multimedia storage server. Computer Networks and ISDN Systems, 27(3):503– 517, December 16, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1994&volume=27&
                                        issue=3&aid=1370.

Kumar:1995:CPI

[689] Anurag Kumar and Robert G. Cole. Comparative performance of interleaved and non-interleaved pipelining in ATM terminal adapters. Computer Networks and ISDN Systems, 27 (4):521–535, January 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=4&aid=1294.

Lai:1995:IAP

[690] R. Lai and W. Leung. Industrial and academic protocol testing: the gap and the means of convergence. Computer Networks and ISDN Systems, 27 (4):537–547, January 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/

browse.cgi?year=1995&volume=27&
issue=4&aid=1303.

Rangan:1995:PIM

[691] P. Venkat Rangan, Srinivas Ramanathan, and Thomas Kaeppner. Performance of inter-media synchronization in distributed and heterogeneous multimedia systems. Computer Networks and ISDN Systems, 27 (4):549–565, January 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=4&aid=1305.

Srivatsa:1995:TDC

[692] S. K. Srivatsa and P. Seshaiah. On the topological design of a Computer Network. Computer Networks and ISDN Systems, 27(4):567– 569, January 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=4&aid=1306.

Bochmann:1995:OOD

[693] Gregor v. Bochmann, Stéphane Poirier, and Pierre Mondain-Monval. Objectoriented design for distributed systems: The OSI directory example. Computer Networks and ISDN Systems, 27 (4):571–590, January 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=4&aid=1307.

Kripanandan:1995:ETG

[694] R. S. Kripanandan and S. V. Raghavan. Experience in test generation using multi-level approach. Computer Networks and ISDN Systems, 27(4): 591–610, January 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=4&aid=1308.

Denzel:1995:FSB

[695] W. E. Denzel, A. P. J. Engbersen, and I. Iliadis. A flexible shared-buffer switch for ATM at Gb/s rates. Computer Networks and ISDN Systems, 27 (4):611–624, January 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=4&aid=1320.

Bisdikian:1995:PAS

[696] Chatschik Bisdikian. Performance analysis of single-stage, output buffer packet switches with independent batch arrivals. Computer Networks and ISDN Systems, 27(5): 627–652, March 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=5&aid=1322.

Jing:1995:ESN

[697] Wen Jing and Michael Paterakis. Extending the single-node DQDB analytical model to analyze networkwide performance. Computer Networks and ISDN Systems, 27(5): 653–675, March 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=5&aid=1323.

Gopal:1995:APH

[698] Inder Gopal and Raphael Rom. ARQ protocols for high speed hardware implementation. Computer Networks and ISDN Systems, 27(5): 677–689, March 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=5&aid=1324.

OMahony:1995:XDS

[699] Donal O’Mahony and Neil Weldon. X.500 directory services support for Electronic Data Interchange (EDI). Computer Networks and ISDN Systems, 27(5): 691–701, March 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=5&aid=1326.

Guillemin:1995:CCT

[700] Fabrice Guillemin, Charles Levert, and Catherine Rosenberg. Cell conformance testing with respect to the peak cell rate in ATM networks. Computer Networks and ISDN Systems, 27(5):703–726, March 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=5&aid=1329.

Pitkow:1995:TIP

[701] James E. Pitkow and R. K. Jones. Towards an intelligent publishing environment. Computer Networks and ISDN Systems, 27(6):729–737, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/comnet/sub/1995/
                                        27/6/1454.pdf.

Roscheisen:1995:STL

[702] Martin Röscheisen, Christian Mogensen, and Terry Winograd. SOAPs, trails, and on-line communities. Computer Networks and ISDN Systems, 27(6):739–749, April 3, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1455.pdf.

Kaarela:1995:IMB

[703] Kari Kaarela, Jaakko Oksanen, and Juha Takalo. An information model as a basis for hypermedia-based plant documentation. Computer Networks and ISDN Systems, 27(6): 751–764, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1449.pdf.

Lewontin:1995:DWT

[704] Steve Lewontin. The DCE Web toolkit: enhancing WWW protocols with lower-layer services. Computer Networks and ISDN Systems, 27(6): 765–771, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1995/27/6/1456.pdf.

Rowe:1995:RWN

[705] Kenneth E. Rowe and Charles K. Nicholas. Reliability of WWW name servers. Computer Networks and ISDN Systems, 27(6): 773–780, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1995/27/6/1469.pdf.

Donnelley:1995:WMD

[706] James E. Donnelley. WWW media distribution via Hopwise Reliable Multicast. Computer Networks and ISDN Systems, 27(6): 781–788, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1995/27/6/1472.pdf.

Baentsch:1995:WID

[707] Michael Baentsch, Georg Molter, and Peter Sturm. WebMake: Integrating distributed software development in a structure-enhanced Web. Computer Networks and ISDN Systems, 27(6):789–800, April 3, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1443.pdf.

Mathews:1995:NOF

[708] G. Jason Mathews and Syed S. Towheed. NSSDC OMNIWeb: The first space physics WWW-based data browsing and retrieval system. Computer Networks and ISDN Systems, 27(6):801–808, April 3, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1457.pdf.

Pitkow:1995:UWS

[709] James E. Pitkow and Margaret M. Recker. Using the Web as a survey tool: results from the second WWW user survey. Computer Networks and ISDN Systems, 27(6): 809–822, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1442.pdf.

Weber:1995:CWP

[710] Kate Weber. Chapter 6, in which Pooh proposes improvements to Web authoring tools, having seen said tools for the Unix platform. Computer Networks and ISDN Systems, 27(6): 823–829, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1465.pdf.

Quint:1995:SAE

[711] Vincent Quint, Cécile Roisin, and Irène Vatton. A structured authoring environment for the World-Wide Web. Computer Networks and ISDN Systems, 27(6):831–840, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1995/
27/6/1462.pdf.

Paoli:1995:CWN

[712] Jean Paoli. Cooperative work on the network: edit the WWW! Computer Networks and ISDN Systems, 27(6):841–847, April 3, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1995/27/6/1458.pdf.

Maly:1995:MXC

[713] K. J. Maly, H. Abdel-Wahab, R. Mukkamala, A. Gupta, A. Prabhu, H. Syed, and C. S. Vemuru. Mosaic + XTV = CoReview. Computer Networks and ISDN Systems, 27(6): 849–860, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1995/27/6/1444.pdf.

Peters:1995:CDA

[714] Ralph Peters and Christian Neuss. CrystalWeb — A distributed authoring environment for the World-Wide Web. Computer Networks and ISDN Systems, 27(6):861–870, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1995/
27/6/1448.pdf.

Ibrahim:1995:AEU

[715] Bertrand Ibrahim and Stephen D. Franklin. Advanced educational uses of the World-Wide Web. Computer Networks and ISDN Systems, 27(6): 871–877, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1438.pdf.

Nicol:1995:FTB

[716] David Nicol, Calum Smeaton, and Alan Falconer Slater. Footsteps: Trail-blazing the Web. Computer Networks and ISDN Systems, 27(6): 879–885, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1440.pdf.

Campbell:1995:CEC

[717] J. K. Campbell, S. Hurley, S. B. Jones, and N. M. Stephens. Constructing educational courseware using NCSA Mosaic and the World-Wide Web. Computer Networks and ISDN Systems, 27(6):887–896, April 3, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1451.pdf.

Dwyer:1995:CVC

[718] Dan Dwyer, Kathy Barbieri, and Helen M. Doerr. Creating a virtual classroom for interactive education on the Web. Computer Networks and ISDN Systems, 27(6): 897–904, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1995/27/6/1461.pdf.

Bilotta:1995:EEU

[719] Eleonora Bilotta, Mariano Fiorito, Dario Iovane, and Pietro Pantano. An educational environment using WWW. Computer Networks and ISDN Systems, 27(6): 905–909, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1995/27/6/1470.pdf.

LaLiberte:1995:PSG

[720] Daniel LaLiberte and Alan Braverman. A protocol for scalable group and public annotations. Computer Networks and ISDN Systems, 27(6): 911–918, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1995/27/6/1441.pdf.

Andrews:1995:SIW

[721] Keith Andrews, Frank Kappe, and Hermann Maurer. Serving information to the Web with Hyper-G. Computer Networks and ISDN Systems, 27(6):919–926, April 3, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1995/27/6/1452.pdf.

Perrochon:1995:IUW

[722] Louis Perrochon and Roman Fischer. IDLE: Unified W3-access to interactive information servers. Computer Networks and ISDN Systems, 27(6):

927–938, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1453.pdf.

Shklar:1995:PLD

[723] Leon Shklar, Kshitij Shah, and Chumki Basu. Putting legacy data on the Web: a repository definition language. Computer Networks and ISDN Systems, 27(6):939–951, April 3, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1435.pdf.

Sheldon:1995:DRD

[724] Mark A. Sheldon, Andrzej Duda, Ron Weiss, and David K. Gifford. Discover: a resource discovery system based on content routing. Computer Networks and ISDN Systems, 27(6): 953–972, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1468.pdf.

Neuss:1995:CAR

[725] Christian Neuss and Robert E. Kent. Conceptual analysis of resource meta-information. Computer Networks and ISDN Systems, 27(6): 973–984, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1466.pdf.

Vizine-Goetz:1995:SWB

[726] Diane Vizine-Goetz, Jean Godby, and Mark Bendig. Spectrum: a Web-based tool for describing electronic resources. Computer Networks and ISDN Systems, 27(6): 985–1001, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1995/27/6/1459.pdf.

Dodge:1995:WCT

[727] Chris Dodge, Beate Marx, and Hans Pfeiffenberger. Web cataloguing through cache exploitation and steps toward consistency maintenance. Computer Networks and ISDN Systems, 27 (6):1003–1008, April 3, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1995/27/6/1445.pdf.

Dobson:1995:LD

[728] S. A. Dobson and V. A. Burrill. Lightweight databases. Computer Networks and ISDN Systems, 27(6): 1009–1015, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1995/27/6/1436.pdf.

Clausnitzer:1995:WIO

[729] Alexander Clausnitzer, Pavel Vogel, and Stephan Wiesener. A WWW interface to the OMNIS/Myriad literature retrieval engine. Computer Networks and ISDN Systems, 27(6): 1017–1026, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1995/27/6/1450.pdf.

Pfeifer:1995:SSD

[730] Ulrich Pfeifer, Norbert Fuhr, and Tung Huynh. Searching structured documents with the enhanced retrieval functionality of freeWAISsf and SFgate. Computer Networks and ISDN Systems, 27(6): 1027–1036, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1446.pdf.

Norderhaug:1995:DWI

[731] Terje Norderhaug and Juliet M. Oberding. Designing a web of intellectual property. Computer Networks and ISDN Systems, 27(6): 1037–1046, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1463.pdf.

Anderson:1995:SFS

[732] Scot Anderson and Rick Garvin. Sessioneer: flexible session level authentication with off the shelf servers and clients. Computer Networks and ISDN Systems, 27(6): 1047–1053, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1464.pdf.

Kahan:1995:CBA

[733] José Kahan. A capability-based authorization model for the World-Wide Web. Computer Networks and ISDN Systems, 27(6):1055– 1064, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1995/27/6/1471.pdf.

Catledge:1995:CBS

[734] Lara D. Catledge and James E. Pitkow. Characterizing browsing strategies in the World-Wide Web. Computer Networks and ISDN Systems, 27(6): 1065–1073, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1995/27/6/1467.pdf.

Mukherjea:1995:VWW

[735] Sougata Mukherjea and James D. Foley. Visualizing the World-Wide Web with the Navigational View Builder. Computer Networks and ISDN Systems, 27(6): 1075–1087, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1995/27/6/1447.pdf.

Beckett:1995:CLS

[736] David J. Beckett. Combined log system. Computer Networks and ISDN Systems, 27(6):1089–1095 (or 1089– 1096??), April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:/

/www.cs.ukc.ac.uk/pubs/1995/155;
 http://www.elsevier.com/cas/tree/
store/comnet/sub/1995/27/6/1437.

pdf. Proceedings of the Third International World-Wide Web Conference. Slides of presentation are available at http://www.hensa.ac.uk/tools/www/logtools/www95/slides/index.html.

Brown:1995:DEW

[737] Marc H. Brown and Robert A. Shillner. DeckScape: an experimental Web browser. Computer Networks and ISDN Systems, 27(6): 1097–1104, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1460.pdf.

vanDoorn:1995:IAW

[738] Matthijs van Doorn and Anton Eliëns. Integrating applications and the World-Wide Web. Computer Networks and ISDN Systems, 27(6): 1105–1110, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1995/27/6/1439.pdf.

vanderSchoot:1995:DFO

[739] Hans van der Schoot and Hasan Ural. Data flow oriented test selection for lotos. Computer Networks and ISDN Systems, 27(7): 1111–1136, May 31, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=7&aid=1321.

Conti:1995:IBM

[740] M. Conti, E. Gregori, and L. Lenzini. Influence of the BWB mechanism on some performance figures of a DQDB subnetwork. Computer Networks and ISDN Systems, 27(7): 1137–1161, May 31, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=7&aid=1325.

West:1995:ERT

[741] Colin H. West and Angelo Tosi. Experiences with a random test driver. Computer Networks and ISDN Systems, 27 (7):1163–1174, May 31, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=7&aid=1327.

Mitrou:1995:MPA

[742] N. Mitrou, S. Vamvakos, and K. Kontovasilis. Modelling, parameter assessment and multiplexing analysis of bursty sources with hyperexponentially distributed bursts. Computer Networks and ISDN Systems, 27 (7):1175–1192, May 31, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=7&aid=1330.

Raghavan:1995:HAB

[743] S. V. Raghavan, D. Vasukiammaiyar, and Günter Haring. Hierarchical approach to building generative networkload models. Computer Networks and ISDN Systems, 27(7): 1193–1206, May 31, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/

                                        browse.cgi?year=1995&volume=27&
                                        issue=7&aid=1328.

deMeer:1995:IRM

[744] Jan de Meer. The ISO Reference Model for Open Distributed Processing. Computer Networks and ISDN Systems, 27(8):1211–1214, July 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=8&aid=1402.

Farooqui:1995:IRM

[745] Kazi Farooqui, Luigi Logrippo, and Jan de Meer. The ISO Reference Model for Open Distributed Processing: an introduction. Computer Networks and ISDN Systems, 27(8): 1215–1230, July 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=8&aid=1403.

Coulson:1995:SRT

[746] G. Coulson, G. S. Blair, J. B. Stefani, F. Horn, and L. Hazard. Supporting the real-time requirements of continuous media in open distributed processing. Computer Networks and ISDN Systems, 27(8): 1231–1246, July 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=8&aid=1350.

Gay:1995:SMA

[747] Valérie Gay, Peter Leydekkers, and Robert Huis in ’t Veld. Specification of multiparty audio and video interaction based on the Reference Model of Open Distributed Processing. Computer Networks and ISDN Systems, 27(8):1247–1262, July 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=8&aid=1349.

vanSinderen:1995:DMO

[748] Marten van Sinderen, Luís Ferreira Pires, Chris A. Vissers, and Joost-Pieter Katoen. A design model for open distributed processing systems. Computer Networks and ISDN Systems, 27(8):1263–1285, July 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=8&aid=1347.

Gotzhein:1995:TBR

[749] R. Gotzhein. Towards a Basic Reference Model of Open Distributed Processing. Computer Networks and ISDN Systems, 27(8): 1287–1304, July 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=8&aid=1346.

Najm:1995:FSO

[750] Elie Najm and Jean-Bernard Stefani. A formal semantics for the ODP computational model. Computer Networks and ISDN Systems, 27(8):1305–1329, July 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=8&aid=1348.

Pandey:1995:CSP

[751] B. D. Pandey and S. S. Pathak. On communicating sequential processes. Computer Networks and ISDN Systems, 27(9):1333–1348, August 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

                                        http://www.elsevier.com/cgi-bin/
                                        cas/tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=9&aid=1332.

Yeh:1995:DEM

[752] W. P. Yeh, J. L. Hammond, and D. Tipper. Design and evaluation of a method for admission control for broadband packet networks. Computer Networks and ISDN Systems, 27 (9):1349–1366, August 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=9&aid=1335.

Deri:1995:OOA

[753] Luca Deri and Eugenio Mattei. An object-oriented approach to the implementation of OSI management. Computer Networks and ISDN Systems, 27 (9):1367–1385, August 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=9&aid=1336.

Kang:1995:BRN

[754] Cheoul-Shin Kang, Byung-Seok Park, Jong-Dug Shin, and Je-Myung Jeong. A broadband ring network: multichannel optical slotted ring. Computer Networks and ISDN Systems, 27 (9):1387–1398, August 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=9&aid=1371.

Huang:1995:CIB

[755] Hsin-Yuan Huang, Thomas Robertazzi, and Aurel A. Lazar. A comparison of information based deflection strategies. Computer Networks and ISDN Systems, 27(9):1399– 1407, August 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=9&aid=1334.

Habib:1995:ACB

[756] Ibrahim W. Habib and Tarek N. Saadawi. Access control of bursty voice traffic in ATM networks. Computer Networks and ISDN Systems, 27(10): 1411–1427, September 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=10&aid=1333.

Bernabeu-Auban:1995:OGP

[757] José M. Bernabéu-Aubán, Mostafa H. Ammar, and Mustaque Ahamad. Optimizing a generalized polling protocol for resource finding over a multiple access channel. Computer Networks and ISDN Systems, 27(10):1429– 1445, September 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=10&aid=1351.

Soh:1995:QRA

[758] B. C. Soh, T. S. Dillon, and P. County. Quantitative risk assessment of computer virus attacks on computer networks. Computer Networks and ISDN Systems, 27(10):1447– 1456, September 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=10&aid=1400.

Kao:1995:LCS

[759] Shao kong Kao and Biswanath Mukherjee. A load-controlled scheduling scheme for integrated voicedata communication on high-speed LANs/MANs. Computer Networks and ISDN Systems, 27(10):1457– 1478, September 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=10&aid=1417.

Kamal:1995:ESM

[760] Ahmed E. Kamal. An enhanced scheme for the management of isochronous channels in ring networks. Computer Networks and ISDN Systems, 27(10): 1479–1483, September 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=10&aid=1331.

Harms:1995:PMC

[761] J. J. Harms and J. W. Wong. Performance modeling of a channel reservation service. Computer Networks and ISDN Systems, 27(11):1487– 1497, October 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=11&aid=1345.

Li:1995:BTH

[762] Jiunn-Jian Li and Cheng-Ming Weng. B^{+}-tree: a high-performance switching structure for ATM with dual input buffering. Computer Networks and ISDN Systems, 27(11):1499– 1522, October 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/

                                        browse.cgi?year=1995&volume=27&
                                        issue=11&aid=1401.

Jia:1995:SSI

[763] Feiling Jia and Biswanath Mukherjee. The superchannel scheme for integrated services on multiple access broadcast networks. Computer Networks and ISDN Systems, 27(11): 1523–1543, October 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=11&aid=1419.

Dobosiewicz:1995:TME

[764] Włodzimierz Dobosiewicz and Pawel Gburzynski. On two modified Ethernets. Computer Networks and ISDN Systems, 27(11):1545– 1564, October 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=11&aid=1420.

Awdeh:1995:SAS

[765] Ra’ed Y. Awdeh and H. T. Mouftah. Survey of ATM switch architectures. Computer Networks and ISDN Systems, 27(12):1567– 1613, November 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=27&
                                        issue=12&aid=1397.

Schmid:1995:SSS

[766] U. Schmid and A. Pusterhofer. SS-CMP: the sequenced synchronized clock message protocol. Computer Networks and ISDN Systems, 27(12): 1615–1632, November 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=12&aid=1421.

Cohen:1995:RTD

[767] Reuven Cohen and Yoram Ofek. Reliable transmission of data over a semi-FIFO routing layer. Computer Networks and ISDN Systems, 27(12):1633– 1649, November 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=27&
issue=12&aid=1423.

Sperberg-McQueen:1995:HMM

[768] C. M. Sperberg-McQueen and Robert F. Goldstein. HTML to the max: A manifesto for adding SGML intelligence to the World-Wide Web. Computer Networks and ISDN Systems, 28 (1–2):3–11, December 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=28&
issue=1-2&aid=1524.

Nielsen:1995:TUL

[769] Henrik Frystyk Nielsen and Håkon W. Lie. Towards a uniform library of common code. A presentation of the CERN World-Wide Web library. Computer Networks and ISDN Systems, 28 (1–2):13–23, December 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=28&
issue=1-2&aid=1528.

Padmanabhan:1995:IHL

[770] Venkata N. Padmanabhan and Jeffrey C. Mogul. Improving HTTP latency. Computer Networks and ISDN Systems, 28(1–2):25–35, December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/comnet/
cas_sub/browse/browse.cgi?year=
1995&volume=28&issue=1-2&aid=1530.

Braun:1995:WTC

[771] Hans-Werner Braun and Kimberly C. Claffy. Web traffic characterization: An assessment of the impact of caching documents from NCSA’s Web server. Computer Networks and ISDN Systems, 28(1–2): 37–51, December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=28&
issue=1-2&aid=1529.

Gessler:1995:PMW

[772] Stefan Gessler and Andreas Kotulla. PDAs as mobile WWW browsers. Computer Networks and ISDN Systems, 28(1–2):53–59, December 1,

1995.  CODEN CNISE9.  ISSN

0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/comnet/
cas_sub/browse/browse.cgi?year=
1995&volume=28&issue=1-2&aid=1517.

Slater:1995:EWC

[773] Alan Falconer Slater. Extending W3 clients. Computer Networks and ISDN Systems, 28(1–2): 61–68, December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=28&
issue=1-2&aid=1540.

Frivold:1995:EWS

[774] Thane J. Frivold, Ruth E. Lang, and Martin W. Fong. Extending WWW for synchronous collaboration. Computer Networks and ISDN Systems, 28 (1–2):69–75, December 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=28&
issue=1-2&aid=1516.

Meyer:1995:WMB

[775] Tom Meyer, David Blair, and Suzanne Hader. WAXweb: A MOObased collaborative hypermedia system for WWW. Computer Networks and ISDN Systems, 28(1–2): 77–84, December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=28&
issue=1-2&aid=1521.

Domel:1995:WGH

[776] Peter Dömel. WebMap: A graphical hypertext navigation tool. Computer Networks and ISDN Systems, 28 (1–2):85–97, December 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=28&
                                        issue=1-2&aid=1527.

Aimar:1995:WTM

[777] Alberto Aimar, James Casey, Nikos Drakos, Ian Hannell, Arash Khodabandeh, Paolo Palazzi, Bertrand Rousseau, and Mario Ruggier. WebLinker, a tool for managing WWW cross-references. Computer Networks and ISDN Systems, 28(1–2): 99–107, December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=28&
                                        issue=1-2&aid=1513.

Kent:1995:CWA

[778] Robert E. Kent and Christian Neuss. Creating a Web analysis and visualization environment. Computer Networks and ISDN Systems, 28(1–2): 109–117, December 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=28&
                                        issue=1-2&aid=1519.

Bowman:1995:HID

[779] C. Mic Bowman, Peter B. Danzig, Darren R. Hardy, Udi Manber, and Michael F. Schwartz. The Harvest information discovery and access system. Computer Networks and ISDN Systems, 28(1–2):119–125 (or 119– 126??), December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=28&
issue=1-2&aid=1522.

Eichmann:1995:EWA

[780] David Eichmann. Ethical Web agents. Computer Networks and ISDN Systems, 28(1–2):127–136, December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/comnet/
cas_sub/browse/browse.cgi?year=
1995&volume=28&issue=1-2&aid=1531.

Whitehead:1995:AFE

[781] Steven D. Whitehead. Auto-FAQ: An experiment in cyberspace leveraging. Computer Networks and ISDN Systems, 28(1–2):137– 146, December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=28&
issue=1-2&aid=1525.

Lim:1995:UCI

[782] Jong-Gyun Lim. Using Coollists to index HTML documents in the Web. Computer Networks and ISDN Systems, 28(1–2):147– 154, December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=28&
                                        issue=1-2&aid=1538.

Robertson:1995:VFD

[783] David Robertson, William Johnston, and Wing Nip. Virtual frog dissection: Interactive 3D graphics via the Web. Computer Networks and ISDN Systems, 28(1–2): 155–160, December 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=28&
                                        issue=1-2&aid=1541.

Egret:1995:WAR

[784] Daniel Egret and André Heck. WWW in astronomy and related space sciences. Computer Networks and ISDN Systems, 28(1–2):161– 166, December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=28&
                                        issue=1-2&aid=1514.

Wray:1995:OIM

[785] Robert E. Wray III, Ronald Chong, Joseph Phillips, Seth Rogers, William Walsh, and John Laird. Organizing information in Mosaic: A classroom experiment. Computer Networks and ISDN Systems, 28(1–2): 167–178, December 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/

tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=28&
issue=1-2&aid=1535.

Nielsen:1995:SUI

[786] Jakob Nielsen and Darrell Sano. Sun-Web: User interface design for Sun Microsystem’s internal Web. Computer Networks and ISDN Systems, 28(1– 2):179–188, December 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=28&
issue=1-2&aid=1533.

Soreide:1995:MAR

[787] N. N. Soreide, L. C. McCarty, and D. C. McClurg. Mosaic access to real-time data from the TOGA-TAO array of moored buoys. Computer Networks and ISDN Systems, 28(1– 2):189–197, December 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=28&
issue=1-2&aid=1523.

Dratva:1995:WBH

[788] Richard Dratva. WWW-based home banking services in Switzerland: A case study. Computer Networks and ISDN Systems, 28(1–2):199– 208, December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=28&
issue=1-2&aid=1539.

Goldberg:1995:BWM

[789] Ken Goldberg, Michael Mascha, Steven Gentner, Jürgen Rossman, Nick Rothenberg, Carl Sutter, and Jeff Wiegley. Beyond the Web: Manipulating the real world. Computer Networks and ISDN Systems, 28(1– 2):209–219, December 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=28&
                                        issue=1-2&aid=1518.

Meyer:1995:MHE

[790] Theodore Meyer, Ramachandran Suresh, Douglas Ilg, and Bruce Moxon. Mosaic, HDF and EOSDIS: Providing access to earth sciences data. Computer Networks and ISDN Systems, 28 (1–2):221–229, December 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=28&
                                        issue=1-2&aid=1532.

Boston:1995:ISD

[791] A. N. Boston and D. R. B. Stockwell. Interactive species distribution reporting, mapping and modelling using the World Wide Web. Computer Networks and ISDN Systems, 28(1– 2):231–238, December 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1995&volume=28&
                                        issue=1-2&aid=1537.

Weibel:1995:ASP

[792] Stuart Weibel, Eric Miller, Jean Godby, and Ralph LeVan. An architecture for scholarly publishing on the World Wide Web. Computer Networks and ISDN Systems, 28(1–2): 239–245, December 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=28&
issue=1-2&aid=1526.

Davis:1995:DPW

[793] Jim Davis and Carl Lagoze. “Dropin’’ publishing with the World Wide Web. Computer Networks and ISDN Systems, 28(1–2):247– 255, December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=28&
issue=1-2&aid=1515.

MacKie-Mason:1995:SFA

[794] Jeffrey K. MacKie-Mason and Hal R. Varian. Some FAQs about usagebased pricing. Computer Networks and ISDN Systems, 28(1–2):257– 265, December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1995&volume=28&
issue=1-2&aid=1520.

Rudin:1996:FMA

[795] J. P. Courtiat, P. Dembinski, G. J. Holzmann, L. Logrippo, Harry Rudin, and Pamela Zave. Formal methods after 15 years: status and trends: A paper based on contributions of the panelists at the FORmal TEchnique ’95 Conference, Montréal, October 1995. Computer Networks and ISDN Systems, 18(13):1845–1855, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Ong:1996:DRC

[796] Lyndon Y. Ong and Mischa Schwartz. Design of resource control protocols for Release 2/3 Broadband ISDN services. Computer Networks and ISDN Systems, 28(3):269– 282, January 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=3&aid=1418.

Sobrinho:1996:PPA

[797] João L. Sobrinho and José M. Brázio. Proposal and performance analysis of a multiple-access protocol for highspeed wireless LANs. Computer Networks and ISDN Systems, 28(3):283– 305, January 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=3&aid=1422.

Onvural:1996:SUS

[798] Raif O. Onvural, Hal Sandick, and Rao Cherukuri. Structure and use of signaling in B-ISDNs. Computer Networks and ISDN Systems, 28(3): 307–323, January 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=3&aid=1426.

Bruneel:1996:AAT

[799] Herwig Bruneel and Sabine Wittevrongel. An approximate analytical technique for the performance evaluation of ATM switching elements with burst routing. Computer Networks and ISDN Systems, 28(3):325– 343, January 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=3&aid=1427.

Hughes:1996:FRA

[800] C. J. Hughes. Feedback restricted access queues for controlling cell loss in ATM networks. Computer Networks and ISDN Systems, 28(3):345– 350, January 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=3&aid=1428.

Goransson:1996:BRC

[801] Paul Goransson. Bandwidth reservation on a commercial router. Computer Networks and ISDN Systems, 28 (3):351–370, January 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/

                                        browse.cgi?year=1996&volume=28&
                                        issue=3&aid=1429.

Akyildiz:1996:CFM

[802] Ian F. Akyildiz and David A. Levine. A collision-free MAC protocol for optical star LANs. Computer Networks and ISDN Systems, 28(3):371– 390, January 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=3&aid=1424.

Reiss:1996:PAA

[803] L. Kurt Reiss and Lazaros F. Merakos. Performance analysis of an adaptive bandwidth reservation scheme for ATM virtual path traffic. Computer Networks and ISDN Systems, 28(3): 391–400, January 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=3&aid=1425.

Economides:1996:AVC

[804] A. A. Economides, P. A. Ioannou, and J. A. Silvester. Adaptive virtual circuit routing. Computer Networks and ISDN Systems, 28(3):401– 409, January 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=3&aid=1430.

Ramamurthy:1996:AVB

[805] G. Ramamurthy and B. Sengupta. An analysis of a variable bit rate multiplexer using loss priorities. Computer Networks and ISDN Systems, 28 (3):411–423, January 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=3&aid=1433.

Hermanns:1996:PII

[806] Oliver Hermanns and Marko Schuba. Performance investigations of the IP multicast architecture. Computer Networks and ISDN Systems, 28(4):429– 439, February 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=4&aid=1498.

Siegl:1996:HNM

[807] Manfred R. Siegl and Georg Trausmuth. Hierarchical network management: a concept and its prototype in SNMPv2. Computer Networks and ISDN Systems, 28(4):441– 452, February 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=4&aid=1499.

Radhakrishnan:1996:FTS

[808] S. Radhakrishnan, S. V. Raghavan, and Ashok K. Agrawala. A flexible traffic shaper for high speed networks: design and comparative study with leaky bucket. Computer Networks and ISDN Systems, 28(4):453– 469, February 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=4&aid=1500.

Drury:1996:ATM

[809] David M. Drury. ATM traffic management and the impact of ATM switch design. Computer Networks and ISDN Systems, 28(4):471– 479, February 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=4&aid=1501.

Nentwig:1996:HCM

[810] Lutz Nentwig, Sonia Manhart, and Kurt Sandkuhl. Hotline and consulting in a metropolitan area network: the HotCon approach to integrated services. Computer Networks and ISDN Systems, 28(4):481– 490, February 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=4&aid=1502.

Wilbur:1996:RTV

[811] Sylvia Wilbur and Sarom Ing. Realtime video for informal workgroup communication: A survey of recent advances. Computer Networks and ISDN Systems, 28(4):491–497 (or 491– 498??), February 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=4&aid=1503.

Grimm:1996:SPO

[812] Rüdiger Grimm and Thomas Hetschold. Security policies in OSI-management experiences from the DeTeBerkom project BMSec. Computer Networks and ISDN Systems, 28(4):499– 511, February 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=4&aid=1504.

Sameshima:1996:SDI

[813] Yoshiki Sameshima and Peter T. Kirstein. Secure document interchange: A secure user agent. Computer Networks and ISDN Systems, 28(4):513– 523 (or 513–524??), February 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/cgi-bin/
cas/tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=4&aid=1505.

Fluckiger:1996:WWW

[814] François Fluckiger. From World-Wide Web to Information Superhighway. Computer Networks and ISDN Systems, 28(4):525–534, February 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

                                        http://www.elsevier.com/cgi-bin/
                                        cas/tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=4&aid=1536.

Rio:1996:FBM

[815] Miguel Rio, António Costa, Joaquim Macedo, and Vasco Freitas. A framework for broadcasting and management of URIs. Computer Networks and ISDN Systems, 28(4):535– 542, February 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=4&aid=1506.

Jose:1996:PME

[816] Rui J. P. José, António Costa, Joaquim Macedo, and Vasco Freitas. Providing multiple external views on directory user interfaces. Computer Networks and ISDN Systems, 28(4):543– 550, February 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=4&aid=1507.

Barker:1996:XID

[817] Paul Barker. X.500 Index DSAs and scaling issues for an indexed white pages directory service. Computer Networks and ISDN Systems, 28(4): 551–562, February 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/

browse.cgi?year=1996&volume=28&
issue=4&aid=1508.

Jonas:1996:IFS

[818] Karl Jonas, Helmut Jungblut, Joachim Kaeber, Manfred Kaul, Ina Müller, Horst Santo, Josef Schäfer, and Ralf Wegner. The Information Footprint: a satellite-based information on demand teleservice. Computer Networks and ISDN Systems, 28(4):563– 573, February 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=4&aid=1509.

Behringer:1996:TOE

[819] M. H. Behringer. Technical options for a European high-speed backbone. Computer Networks and ISDN Systems, 28(4):575–581, February 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/cgi-bin/
cas/tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=4&aid=1510.

Nussbacher:1996:LLM

[820] Hank Nussbacher. Lessons learned from a MAN. Computer Networks and ISDN Systems, 28(4):583– 588, February 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=4&aid=1511.

Mills:1996:GMH

[821] Peter Mills and Jim Strom. G-MING: A high performance multiservice telecommunications infrastructure for the Greater Manchester educational community. Computer Networks and ISDN Systems, 28(4):589– 597, February 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=4&aid=1512.

Lipper:1996:SSP

[822] E. H. Lipper. Switching system performance issues for universal personal communications. Computer Networks and ISDN Systems, 28(5): 603–611, March 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=5&aid=1490.

Wirth:1996:TID

[823] Patricia E. Wirth. Teletraffic implications of database architectures in mobile and personal communications. Computer Networks and ISDN Systems, 28(5):613–618, March 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

                                        http://www.elsevier.com/cgi-bin/
                                        cas/tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=5&aid=1491.

Williams:1996:RIA

[824] Philip M. Williams. Requirements and issues for automatic focused overload control. Computer Networks and ISDN Systems, 28(5): 619–625, March 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=5&aid=1492.

Yan:1996:DNR

[825] James Yan. Dimensioning network resources for IN services. Computer Networks and ISDN Systems, 28(5): 627–633, March 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=5&aid=1493.

Smith:1996:ABR

[826] Avril Smith, John Adams, and Geoff Tagg. Available Bit Rate — A new service for ATM. Computer Networks and ISDN Systems, 28(5): 635–640, March 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=5&aid=1494.

Saito:1996:RMC

[827] Hiroshi Saito. Resource management and charging in ATM networks. Computer Networks and ISDN Systems, 28(5):641–644, March 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/

                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=5&aid=1495.

Bolot:1996:CQT

[828] Jean-Chrysostome Bolot. Cost-quality tradeoffs in the Internet. Computer Networks and ISDN Systems, 28(5): 645–651, March 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=5&aid=1496.

Kim:1996:AFB

[829] Jaime Bae Kim, Rahul Simha, and Tatsuya Suda. Analysis of a finite buffer queue with heterogeneous Markov Modulated Arrival processes: a study of traffic burstiness and priority packet discarding. Computer Networks and ISDN Systems, 28(5): 653–673, March 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=5&aid=1432.

Floreani:1996:ASC

[830] D. J. Floreani and A. J. Dadej. Application of the stratification concept to radio networks and their gateways. Computer Networks and ISDN Systems, 28(5): 675–687, March 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/

browse.cgi?year=1996&volume=28&
issue=5&aid=1431.

Gaiti:1996:PII

[831] D. Gaïti. A proposal for integrating intelligent management in the intelligent network conceptual model. Computer Networks and ISDN Systems, 28(5):689–699, March 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=5&aid=1434.

Chrysochos:1996:UCL

[832] I. Chrysochos, M. Koukias, I. Papanikos, and G. Kokkinakis. Upgrading of cable line concentrators to accept ISDN subscribers. Computer Networks and ISDN Systems, 28(5): 701–707, March 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=5&aid=1473.

Jerman-Blazic:1996:TSK

[833] Borka Jerman-Blažič, Denis Trček, Tomaž Klobučar, and Franc Bračun. A tool for support of key distribution and validity certificate check in global Directory service. Computer Networks and ISDN Systems, 28(5): 709–717, March 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=5&aid=1474.

Dempsey:1996:RBE

[834] Bert J. Dempsey, Jörg Liebeherr, and Alfred C. Weaver. On retransmissionbased error control for continuous media traffic in packet-switching networks. Computer Networks and ISDN Systems, 28(5):719–736, March 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

                                        http://www.elsevier.com/cgi-bin/
                                        cas/tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=5&aid=1475.

Bernabei:1996:TAM

[835] Francesco Bernabei, Laura Gratta, and Marco Listanti. Throughput analysis of Multihop ShuffleNets in a hot spot traffic scenario: impact of routing strategies. Computer Networks and ISDN Systems, 28(6): 743–772, April 10, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=6&aid=1477.

Papadimitriou:1996:WSN

[836] Georgios I. Papadimitriou and Dimitris G. Maritsas. WDM star networks: hybrid random access and reservation protocols with high throughput and low delay. Computer Networks and ISDN Systems, 28(6): 773–787, April 10, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=6&aid=1478.

Wei:1996:FRD

[837] Lu Wei and Peter Hansson. Further research on dynamic time slice in ATM switch. Computer Networks and ISDN Systems, 28(6): 789–798, April 10, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=6&aid=1486.

MacGregor:1996:SEE

[838] James MacGregor, Eric Lee, and Frank Safayeni. Some effects of electronic mail use on the quality of relationships between different organizational functions. Computer Networks and ISDN Systems, 28(6): 799–809, April 10, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=6&aid=1489.

Ramamurthy:1996:PCC

[839] G. Ramamurthy and B. Sengupta. A predictive congestion control policy for broadband integrated wide area networks. Computer Networks and ISDN Systems, 28(6): 811–834, April 10, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=6&aid=1476.

Bianchi:1996:APN

[840] Giuseppe Bianchi and Achille Pattavina. Architecture and performance of non-blocking ATM switches with shared internal queueing. Computer Networks and ISDN Systems, 28(6): 835–853, April 10, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=6&aid=1479.

Akyildiz:1996:BRR

[841] Ian F. Akyildiz, Jörg Liebeherr, and Debapriya Sarkar. Bandwidth regulation of real-time traffic classes in internetworks. Computer Networks and ISDN Systems, 28(6): 855–872, April 10, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=6&aid=1487.

Mukherjee:1996:PQI

[842] Amarnath Mukherjee. A proof of quasiindependence of sliding window flow control and go-back- n error recovery under independent packet errors. Computer Networks and ISDN Systems, 28(6):873–887, April 10, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=6&aid=1488.

Markatos:1996:MMC

[843] Evangelos P. Markatos. Main memory caching of Web documents. Computer Networks and ISDN Systems, 28 (7–11):893–905, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1588.

Dingle:1996:WCC

[844] Adam Dingle and Tomáš Pártl. Web cache coherence. Computer Networks and ISDN Systems, 28(7–11): 907–920, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1573.

Baentsch:1996:IAL

[845] Michael Baentsch, Georg Molter, and Peter Sturm. Introducing application-level replication and naming into today’s Web. Computer Networks and ISDN Systems, 28(7– 11):921–930, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1583.

Dossick:1996:WAL

[846] Stephen E. Dossick and Gail E. Kaiser. WWW access to legacy client/ server applications. Computer Networks and ISDN Systems, 28(7–11):

931–940, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1576.

Ciancarini:1996:PAC

[847] Paolo Ciancarini, Andreas Knoche, Robert Tolksdorf, and Fabio Vitali. PageSpace: an architecture to coordinate distributed applications on the Web. Computer Networks and ISDN Systems, 28(7–11): 941–952, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1598.

Riva:1996:LSH

[848] Alberto Riva and Marco Ramoni. Lisp-Web: a specialized HTTP server for distributed AI applications. Computer Networks and ISDN Systems, 28 (7–11):953–961, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1615.

Woodruff:1996:IDW

[849] Allison Woodruff, Paul M. Aoki, Eric Brewer, Paul Gauthier, and Lawrence A. Rowe. An investigation of documents from the World Wide Web. Computer Networks and ISDN Systems, 28(7–11): 963–980, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1617.

Pitkow:1996:SWD

[850] James E. Pitkow and R. Kipp Jones. Supporting the Web: A distributed hyperlink database system. Computer Networks and ISDN Systems, 28 (7–11):981–991, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1627.

Bray:1996:MW

[851] Tim Bray. Measuring the Web. Computer Networks and ISDN Systems, 28 (7–11):993–1005, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1614.

Yan:1996:UAP

[852] Tak Woon Yan, Matthew Jacobsen, Hector García-Molina, and Umeshwar Dayal. From user access patterns to dynamic hypertext linking. Computer Networks and ISDN Systems, 28(7– 11):1007–1014, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1604.

Creech:1996:AOL

[853] Michael L. Creech. Author-oriented link management. Computer Networks and ISDN Systems, 28(7–11): 1015–1025, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1592.

Carr:1996:OIS

[854] Leslie Carr, Gary Hill, David De Roure, Wendy Hall, and Hugh Davis. Open information services. Computer Networks and ISDN Systems, 28(7–11): 1027–1036, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1603.

Brown:1996:DAO

[855] Marc H. Brown and Marc A. Najork. Distributed active objects. Computer Networks and ISDN Systems, 28(7– 11):1037–1052, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1586.

Trevor:1996:EDM

[856] Jonathan Trevor, Richard Bentley, and Gerrit Wildgruber. Exorcising daemons: a modular and lightweight approach to deploying applications on the Web. Computer Networks and ISDN Systems, 28(7–11):1053– 1062, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1590.

Schickler:1996:PBS

[857] Matthew A. Schickler, Murray S. Mazer, and Charles Brooks. Panbrowser support for annotations and other meta-information on the World Wide Web. Computer Networks and ISDN Systems, 28(7–11):1063– 1074, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1613.

Bonhomme:1996:IRH

[858] Stéphane Bonhomme and Cécile Roisin. Interactively restructuring HTML documents. Computer Networks and ISDN Systems, 28(7–11): 1075–1084, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1595.

Mea:1996:HGS

[859] Vincenzo Della Mea, Carlo Alberto Beltrami, Vito Roberto, and Davide Brunato. HTML generation and semantic markup for telepathology. Computer Networks and ISDN Systems, 28 (7–11):1085–1094, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1602.

Paoli:1996:EWT

[860] Jean Paoli. Extending the Web’s tag set using SGML: authoring new tags with Grif Symposia. Computer Networks and ISDN Systems, 28(7–11): 1095–1104, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1625.

Capps:1996:EGM

[861] Michael Capps, Brian Ladd, and David Stotts. Enhanced graph models in the Web: multi-client, multi-head, multi-tail browsing. Computer Networks and ISDN Systems, 28(7–11): 1105–1112, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1589.

Thau:1996:DCA

[862] Robert Thau. Design considerations for the Apache Server API. Computer Networks and ISDN Systems, 28 (7–11):1113–1122, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/

browse.cgi?year=1996&volume=28&
issue=7-11&aid=1601.

Meyer:1996:BIS

[863] Eric A. Meyer and Peter E. Murray. Borealis image server. Computer Networks and ISDN Systems, 28 (7–11):1123–1137, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1619.

Hadjiefthymiades:1996:GFD

[864] Stathes P. Hadjiefthymiades and Drakoulis I. Martakos. A generic framework for the deployment of structured databases on the World Wide Web. Computer Networks and ISDN Systems, 28(7–11):1139– 1148, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1587.

Duan:1996:DDA

[865] Nick N. Duan. Distributed database access in a corporate environment using Java. Computer Networks and ISDN Systems, 28(7–11):1149– 1156, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1600.

Kamiya:1996:GSP

[866] Kenichi Kamiya, Martin Röscheisen, and Terry Winograd. Grassroots: A system providing a uniform framework for communicating, structuring, sharing information, and organizing people. Computer Networks and ISDN Systems, 28(7–11): 1157–1174, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1579.

Crandall:1996:IEI

[867] Mike Crandall and Mark C. Swenson. Integrating electronic information through a corporate Web. Computer Networks and ISDN Systems, 28 (7–11):1175–1186, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1593.

Bogen:1996:DWA

[868] Manfred Bogen, Michael Lenz, and Susanne Zier. Deutsche Welle: On the air. Computer Networks and ISDN Systems, 28(7–11):1187– 1196, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1597.

Pays:1996:IPS

[869] Paul-André Pays and Fabrice de Comarmond. An intermediation and payment system technology. Computer Networks and ISDN Systems, 28(7– 11):1197–1206, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1629.

Yeh:1996:SNC

[870] Ping-Jer Yeh, Bih-Horng Chen, Ming-Chih Lai, and Shyan-Ming Yuan. Synchronous navigation control for distance learning on the Web. Computer Networks and ISDN Systems, 28 (7–11):1207–1218, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1618.

Goldberg:1996:WWW

[871] Murray W. Goldberg, Sasan Salari, and Paul Swoboda. World Wide Web - Course tool: An environment for building WWW-based courses. Computer Networks and ISDN Systems, 28 (7–11):1219–1231, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1574.

Hauck:1996:SHG

[872] Franz J. Hauck. Supporting hierarchical guided tours in the World Wide Web. Computer Networks and ISDN Systems, 28(7–11):1233– 1242, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1605.

Yang:1996:AIO

[873] Jack Jingshuang Yang and Gail E. Kaiser. An architecture for integrating OODBs with WWW. Computer Networks and ISDN Systems, 28(7– 11):1243–1254, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1599.

Ingham:1996:FBL

[874] David Ingham, Steve Caughey, and Mark Little. Fixing the “Broken-Link’’ problem: the W3Objects approach. Computer Networks and ISDN Systems, 28(7–11):1255– 1268, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1622.

Merle:1996:CGO

[875] Philippe Merle, Christophe Gransart, and Jean-Marc Geib. CorbaWeb: A generic object navigator. Computer Networks and ISDN Systems, 28(7– 11):1269–1281, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1624.

Bowers:1996:WQA

[876] Neil Bowers. Weblint: quality assurance for the World Wide Web. Computer Networks and ISDN Systems, 28 (7–11):1283–1290, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1591.

Crespo:1996:WBB

[877] Arturo Crespo and Eric A. Bier. Web-Writer: A browser-based editor for constructing Web applications. Computer Networks and ISDN Systems, 28 (7–11):1291–1306, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1616.

Perrone:1996:WSE

[878] Corrina Perrone, David Clark, and Alexander Repenning. WebQuest: Substantiating education in edutainment through interactive learning games. Computer Networks and ISDN Systems, 28(7–11):1307– 1319, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/

                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1606.

Maarek:1996:AOB

[879] Yoelle S. Maarek and Israel Z. Ben Shaul. Automatically organizing bookmarks per contents. Computer Networks and ISDN Systems, 28(7–11): 1321–1334, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1577.

Douglis:1996:WQN

[880] Fred Douglis, Thomas Ball, Yih-Farn Chen, and Eleftherios Koutsofios. WebGUIDE: querying and navigating changes in Web repositories. Computer Networks and ISDN Systems, 28 (7–11):1335–1344, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1612.

Jones:1996:NNH

[881] Kirsten L. Jones. nif-T-nav: A hierarchical navigator for WWW pages. Computer Networks and ISDN Systems, 28(7–11):1345–1354, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/comnet/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=28&issue=7-11&aid=1581.

Thistlewaite:1996:AFH

[882] Paul Thistlewaite and Steve Ball. Active FORMs [in HTML]. Computer Networks and ISDN Systems, 28(7– 11):1355–1364, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1596.

Rouaix:1996:WNA

[883] François Rouaix. A Web navigator with applets in Caml. Computer Networks and ISDN Systems, 28(7–11): 1365–1371, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1585.

Perret:1996:MAP

[884] Stéphane Perret and Andrzej Duda. Mobile assistant programming for efficient information access on the WWW. Computer Networks and ISDN Systems, 28(7–11):1373–1383, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/comnet/
cas_sub/browse/browse.cgi?year=
1996&volume=28&issue=7-11&aid=1609.

Jacobs:1996:FHF

[885] Stephan Jacobs, Michael Gebhardt, Stefanie Kethers, and Wojtek Rzasa. Filling HTML forms simultaneously: CoWeb-architecture and functionality. Computer Networks and ISDN Systems, 28(7–11):1385–1395, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/comnet/
cas_sub/browse/browse.cgi?year=
1996&volume=28&issue=7-11&aid=1607.

Bolot:1996:PEW

[886] Jean-Chrysostome Bolot and Philipp Hoschka. Performance engineering of the World Wide Web: Application to dimensioning and cache design. Computer Networks and ISDN Systems, 28 (7–11):1397–1405, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1626.

Smith:1996:UNW

[887] Neil G. Smith. The UK national Web cache — The state of the art. Computer Networks and ISDN Systems, 28 (7–11):1407–1414, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1582.

Neal:1996:HOC

[888] Donald Neal. The Harvest Object Cache in New Zealand. Computer Networks and ISDN Systems, 28(7– 11):1415–1430, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/

browse.cgi?year=1996&volume=28&
issue=7-11&aid=1630.

Schilit:1996:TLC

[889] Bill N. Schilit, Fred Douglis, David M. Kristol, Paul Krzyzanowski, James Sienicki, and John A. Trotter. TeleWeb: loosely connected access to the World Wide Web. Computer Networks and ISDN Systems, 28(7–11): 1431–1444, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&

issue=7-11&aid=1575. http://garuda.imag.fr/www5cd/www172/overview.htm.

Fox:1996:RWL

[890] Armando Fox and Eric A. Brewer. Reducing WWW latency and bandwidth requirements by real-time distillation. Computer Networks and ISDN Systems, 28(7–11):1445– 1456, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1580. cache storage;

client-server systems; computer communications software; data compression; Internet; network servers; realtime systems; network latency; bandwidth requirements; real-time distillation; Pythia proxy mechanism; World Wide Web; real-time refinement; statistical models; metered cellular phone service; transcoding; client-side rendering; data representation; client display constraints; content optimization; PPP; Point-to-Point Protocol; image loading; added value.

Lamm:1996:RTG

[891] Stephen E. Lamm, Daniel A. Reed, and Will H. Scullin. Real-time geographic visualization of World Wide Web traffic. Computer Networks and ISDN Systems, 28(7–11): 1457–1468, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1608.

Richmond:1996:EOS

[892] Alice Richmond. Enticing online shoppers to buy — a human behavior study. Computer Networks and ISDN Systems, 28(7–11):1469– 1480, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1611.

Pant:1996:BWS

[893] Somendra Pant and Cheng Hsu. Business on the Web: strategies and economics. Computer Networks and ISDN Systems, 28(7–11): 1481–1492, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1594.

Kohda:1996:UAW

[894] Youji Kohda and Susumu Endo. Ubiquitous advertising on the WWW:

Merging advertisement on the browser. Computer Networks and ISDN Systems, 28(7–11):1493–1499, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/comnet/
cas_sub/browse/browse.cgi?year=
1996&volume=28&issue=7-11&aid=1623.

Chang:1996:ABW

[895] Jin W. Chang and Colin T. Scott. Agent-based workflow: TRP Support Environment (TSE). Computer Networks and ISDN Systems, 28(7–11):1501–1511 (or 1501– 1512??), May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1621.

Sandewall:1996:TWW

[896] Erik Sandewall. Towards a worldwide data base. Computer Networks and ISDN Systems, 28(7–11): 1513–1522, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=7-11&aid=1628.

Mathews:1996:EMP

[897] G. Jason Mathews and Barry E. Jacobs. Electronic management of the peer review process. Computer Networks and ISDN Systems, 28(7–11): 1523–1538, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1610.

Gobbetti:1996:VSL

[898] Enrico Gobbetti and Andrea O. Leone. Virtual Sardinia: A largescale hypermedia regional information system. Computer Networks and ISDN Systems, 28(7–11):1539– 1546, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1584.

England:1996:RRT

[899] Paul England, Robert Allen, and Ron Underwood. RAVE: Real-time services for the Web. Computer Networks and ISDN Systems, 28(7–11): 1547–1558, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=28&
                                        issue=7-11&aid=1620.

Katkere:1996:IVW

[900] Arun Katkere, Jennifer Schlenzig, Amarnath Gupta, and Ramesh Jain. Interactive video on WWW: beyond VCR-like interfaces. Computer Networks and ISDN Systems, 28(7–11): 1559–1572, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/

browse.cgi?year=1996&volume=28&
issue=7-11&aid=1578.

Braek:1996:SB

[901] Rolv Bræk. SDL Basics. Computer Networks and ISDN Systems, 28(12): 1585–1602, June 10, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=12&aid=1543.

Sarma:1996:IS

[902] A. Sarma. Introduction to SDL-92. Computer Networks and ISDN Systems, 28(12):1603–1615, June 10, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/cgi-bin/
cas/tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=12&aid=1544.

Verhaard:1996:IZ

[903] L. Verhaard. An introduction to Z.105. Computer Networks and ISDN Systems, 28(12):1617– 1628, June 10, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=12&aid=1545.

Rudolph:1996:TMS

[904] Ekkart Rudolph, Peter Graubmann, and Jens Grabowski. Tutorial on Message Sequence Charts. Computer Networks and ISDN Systems, 28(12): 1629–1641, June 10, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=12&aid=1546.

Mauw:1996:FMS

[905] S. Mauw. The formalization of Message Sequence Charts. Computer Networks and ISDN Systems, 28(12): 1643–1657, June 10, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=12&aid=1547.

Hogrefe:1996:VSS

[906] Dieter Hogrefe. Validation of SDL systems. Computer Networks and ISDN Systems, 28(12):1659– 1667, June 10, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=12&aid=1548.

Cavalli:1996:TMS

[907] Ana R. Cavalli, Byoung-Moon Chin, and Kilnam Chon. Testing methods for SDL systems. Computer Networks and ISDN Systems, 28(12): 1669–1683, June 10, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=12&aid=1549.

Reed:1996:MRT

[908] Rick Reed. Methodology for real time systems. Computer Networks and ISDN Systems, 28(12):1685– 1701, June 10, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=12&aid=1550.

Cheng:1996:RDA

[909] K. E. Cheng. A requirements definition and assessment framework for SDL tools. Computer Networks and ISDN Systems, 28(12): 1703–1715, June 10, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=12&aid=1551.

Jain:1996:CCT

[910] Raj Jain. Congestion control and traffic management in ATM networks: Recent advances and A survey. Computer Networks and ISDN Systems, 28(13): 1723–1738, October 15, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://www.

cis.ohio-state.edu/~jain/papers/
cnis.htm;   http://www.elsevier.
com/cgi-bin/cas/tree/store/comnet/
cas_sub/browse/browse.cgi?year=
1996&volume=28&issue=13&aid=1565.

Hsu:1996:ACM

[911] Ivy Hsu and Jean Walrand. Admission control for multi-class ATM traffic with overflow constraints. Computer Networks and ISDN Systems, 28(13): 1739–1751, October 15, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1996/28/13/1564.pdf.

Jung:1996:TUQ

[912] Jae-Il Jung. Translation of user’s QoS requirements into ATM performance parameters in B-ISDN. Computer Networks and ISDN Systems, 28 (13):1753–1767, October 15, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1996/28/13/1563.pdf.

Tennenhouse:1996:VIP

[913] David Tennenhouse, Butler Lampson, Sharon Eisner Gillett, and Jennifer Steiner Klein. Virtual infrastructure: Putting information infrastructure on the technology curve. Computer Networks and ISDN Systems, 28 (13):1769–1790, October 15, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1996/28/13/1562.pdf.

Patel:1996:HSS

[914] B. Patel, F. Schaffa, and M. Willebeek-LeMair. The Helix switch: A single chip cell switch design. Computer Networks and ISDN Systems, 28(13): 1791–1807, October 15, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=13&aid=1534.

Zhou:1996:EAM

[915] Bin Zhou and M. Atiquzzaman. Efficient analysis of Multistage Interconnection Networks using finite outputbuffered switching elements. Computer Networks and ISDN Systems, 28(13): 1809–1829, October 15, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=13&aid=1556.

Peyravian:1996:DDU

[916] Mohammad Peyravian and Chin-Tau Lea. Deriving deadlock and unspecified reception free protocol converters from message mapping sets. Computer Networks and ISDN Systems, 28 (13):1831–1844, October 15, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1996&volume=28&
issue=13&aid=1561.

Courtiat:1996:FMA

[917] Jean-Pierre Courtiat, Piotr Dembinski, Gerard J. Holzmann, Luigi Logrippo, Harry Rudin, and Pamela Zave. Formal methods after 15 years: Status and trends. A paper based on contributions of the panelists at the FORmal TEchnique ’95 Conference, Montréal, October 1995. Computer Networks and ISDN Systems, 28(13): 1845–1855, October 15, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1996/28/13/1636.pdf.

Mathews:1996:WBD

[918] Gary Jason Mathews and Syed S. Towheed. WWW-based data systems for interactive manipulation of science data. Computer Networks and ISDN Systems, 28(13):1857– 1864, October 15, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1996/28/13/1557.pdf.

Apostolopoulos:1996:DNB

[919] Nicolas Apostolopoulos, Albert Geukes, and Stefan Zimmermann. DI-ALECT — Network-based digital interactive lectures. Computer Networks and ISDN Systems, 28(14):1873– 1886, November 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1996/28/14/1646.pdf.

McLoughlin:1996:WIB

[920] Henry McLoughlin. WEST: An Internet based education delivery and support environment. Computer Networks and ISDN Systems, 28(14):1887– 1890, November 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1996/28/14/1648.pdf.

Kolletzki:1996:SIB

[921] Stephan Kolletzki. Secure Internet banking with Privacy Enhanced Mail — A protocol for reliable exchange of secured order forms. Computer Networks and ISDN Systems, 28(14):1891–1899 (or 1891– 1900??), November 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1996/28/14/1642.pdf.

Gehrke:1996:MPK

[922] Michael Gehrke and Thomas Hetschold. Management of a public key certification infrastructure — Experiences from the DeTeBerkom project BM-Sec. Computer Networks and ISDN Systems, 28(14):1901–1914, November 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1996/
28/14/1647.pdf.

Schill:1996:IAE

[923] Alexander Schill, Sabine Kühn, and Frank Breiter. Internetworking over ATM: Experiences with IP/ IPng and RSVP. Computer Networks and ISDN Systems, 28(14):1915– 1927, November 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1996/28/14/1639.pdf.

Barnes:1996:IMR

[924] Jeremy Barnes, Dave Ginsburg, Dave Newson, and Dave Pratt. IP multicast of real-time MPEG over ATM. Computer Networks and ISDN Systems, 28(14):1929– 1937, November 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1996/28/14/1645.pdf.

Meulemans:1996:YPS

[925] Nils Meulemans. A Yellow Pages service based on X.500. Computer Networks and ISDN Systems, 28(14): 1939–1946, November 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1996/28/14/1644.pdf.

Ferreira:1996:CE

[926] Joao Nuno Ferreira, Alf Hansen, Tomaz Klobucar, Klaus-Peter Kossakowski, Manuel Medina, Damir Rajnovic, Olaf Schjelderup, and Don Stikvoort. CERTs in Europe. Computer Networks and ISDN Systems, 28(14):1947–1952, November 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/comnet/sub/1996/
                                        28/14/1649.pdf.

Hoffmann:1996:BWA

[927] Gertraud Hoffmann. B-WiN — The ATM-based high-speed network for the DFN community. Computer Networks and ISDN Systems, 28(14):1953– 1960, November 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1996/28/14/1638.pdf.

Bohm:1996:WMT

[928] Markus Böhm, Wolfgang Deiters, Manuel Friedrich, Frank Lindert, and Wolfgang Schulze. Workflow management as teleservice. Computer Networks and ISDN Systems, 28(14):1961–1969 (or 1961– 1970??), November 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1996/28/14/1640.pdf.

Horsch:1996:CWM

[929] Alexander Horsch, Karin Eberle, Axel Findling, Birgit Kraus, Vania Pentcheva-Spiridonov, and Artavazd Tarhanjan. Collaborative work with medical images in a university hospital environment: Three pilot projects. Computer Networks and ISDN Systems, 28(14):1971– 1978, November 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1996/28/14/1641.pdf.

Bogen:1996:WWA

[930] Manfred Bogen, Guido Hansen, and Michael Lenz. W3Gate — A Web access for outsiders. Computer Networks and ISDN Systems, 28(14):1979– 1990, November 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1996/28/14/1637.pdf.

Chalmers:1996:HES

[931] Alistair Chalmers and Christopher Duxbury. The hidden economic and societal issues of policies on advanced networking. Computer Networks and ISDN Systems, 28(14):1991– 1998, November 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1996/28/14/1643.pdf.

Cavalli:1996:SFM

[932] Ana R. Cavalli, Jean Philippe Favreau, and Marc Phalippou. Standardization of formal methods in conformance testing of communication protocols. Computer Networks and ISDN Systems, 29(1):3–14, December 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1996/29/1/1568.pdf.

Kristoffersen:1996:TTF

[933] Finn Kristoffersen and Thomas Walter. TTCN: Towards a formal semantics and validation of test suites. Computer Networks and ISDN Systems, 29 (1):15–47, December 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=29&
                                        issue=1&aid=1569.

Tretmans:1996:CTL

[934] Jan Tretmans. Conformance testing with labelled transition systems: Implementation relations and test generation. Computer Networks and ISDN Systems, 29(1):49– 79, December 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1996&volume=29&
                                        issue=1&aid=1570.

Petrenko:1996:FCT

[935] A. Petrenko, G. v. Bochmann, and M. Yao. On fault coverage of tests for finite state specifications. Computer Networks and ISDN Systems, 29 (1):81–106, December 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1996/29/1/1572.pdf.

Ates:1996:TSG

[936] Ahmet F. Ates and Behcet Sarikaya. Test sequence generation and timed testing. Computer Networks and ISDN Systems, 29(1):107–131, December 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/comnet/
cas_sub/browse/browse.cgi?year=
1996&volume=29&issue=1&aid=1571.

Doeringer:1997:AAM

[937] W. Doeringer, D. Dykeman, M. Peters, H. Sandick, K. Vu, and J. Derby. Access architecture for a multiprotocol broadband backbone. Computer Networks and ISDN Systems, 29(2): 137–155, January 27, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/2/1632.pdf.

Chowdhury:1997:DPT

[938] Shyamal Chowdhury and Bhaskar Sengupta. Design and performance of a threshold-based load balancing scheme in an ATM switch. Computer Networks and ISDN Systems, 29(2):157– 164, January 27, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/

                                        browse.cgi?year=1997&volume=29&
                                        issue=2&aid=1634.

Meempat:1997:TBR

[939] Gopalakrishnan Meempat. Throughput behavior of reliable stream protocols in admission controlled bearer networks. Computer Networks and ISDN Systems, 29(2):165– 179, January 27, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=2&aid=1635.

Simmons:1997:PCA

[940] Jane M. Simmons. Proof of correctness of ATM retransmission scheme. Computer Networks and ISDN Systems, 29 (2):181–194, January 27, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1997/29/2/1559.pdf.

Cheng:1997:MAS

[941] Tee-Hiang Cheng. A multichannel ATM switch with output buffering. Computer Networks and ISDN Systems, 29(2):195–208, January 27, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/comnet/sub/1997/
                                        29/2/1552.pdf.

Cheng:1997:DAM

[942] Tee-Hiang Cheng. Design and analysis of a multichannel transmission scheme. Computer Networks and ISDN Systems, 29(2):209– 220, January 27, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/2/1553.pdf.

Huang:1997:WFC

[943] Nen-Fu Huang and Shiann-Tsong Sheu. A waste-free congestion control scheme for dual bus highspeed networks. Computer Networks and ISDN Systems, 29(2):221– 235, January 27, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/2/1558.pdf.

Peyravian:1997:AEG

[944] Mohammad Peyravian and Raif Onvural. Algorithm for efficient generation of link-state updates in ATM networks. Computer Networks and ISDN Systems, 29(2):237– 247, January 27, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/2/1633.pdf.

Rubin:1997:AFN

[945] Izhak Rubin and James C.-H. Wu. Analysis of an FDDI network supporting stations with single-packet buffers. Computer Networks and ISDN Systems, 29(2):249–266, January 27, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/cgi-bin/
cas/tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=2&aid=1560.

Leonard:1997:ILD

[946] Luc Léonard and Guy Leduc. An introduction to ET-LOTOS for the description of time-sensitive systems. Computer Networks and ISDN Systems, 29 (3):271–292, February 1, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1997/29/3/cmn1631.pdf.

Landry:1997:MAT

[947] Randall Landry and Ioannis Stavrakakis. Multiplexing ATM traffic streams with time-scale-dependent arrival processes. Computer Networks and ISDN Systems, 29(3):293–313, February 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/comnet/sub/1997/
                                        29/3/cmn1650.pdf.

Saha:1997:DIN

[948] Abheek Saha, Maharaj Mukherjee, and P. Dhar. Design and implementation of a Network Service Access Point (NSAP) for OSI-compatibility. Computer Networks and ISDN Systems, 29 (3):315–324, February 1, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1997/29/3/cmn1651.pdf.

Habib:1997:NNC

[949] I. Habib, A. Tarraf, and T. Saadawi. A neural network controller for congestion control in ATM multiplexers. Computer Networks and ISDN Systems, 29(3):325–334, February 1,

1997.  CODEN CNISE9.  ISSN

0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1997/
29/3/cmn1652.pdf.

Szabo:1997:ECP

[950] Csaba Szabo. An Ethernet compatible protocol to support real time traffic and multimedia applications. Computer Networks and ISDN Systems, 29 (3):335–342, February 1, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/3/cmn1653.pdf.

Rubin:1997:VQC

[951] Izhak Rubin and Kirk K. Chang. Voice quality control under dynamic encoding algorithms in ATM networks. Computer Networks and ISDN Systems, 29 (3):343–355, February 1, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=3&aid=1654.

Ferguson:1997:FVR

[952] Michael J. Ferguson. Formalization and validation of the Radio Link Protocol (RLP1). Computer Networks and ISDN Systems, 29(3):357– 372, February 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=3&aid=1655.

Medhi:1997:MND

[953] D. Medhi. Models for network design, servicing and monitoring of ATM networks based on the virtual path concept. Computer Networks and ISDN Systems, 29(3):373– 386, February 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=3&aid=1656.

Turner:1997:SA

[954] K. J. Turner and G. Scollo. Specification architecture. Computer Networks and ISDN Systems, 29(4):393– 395, March 14, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1997/29/4/cmn1676.pdf.

Turner:1997:SAI

[955] Kenneth J. Turner. Specification architecture illustrated in a communications context. Computer Networks and ISDN Systems, 29(4):397–411 (or 397–412??), March 14, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1997/29/4/cmn1658.pdf.

Quartel:1997:RBD

[956] Dick A. C. Quartel, Luís Ferreira Pires, Marten J. van Sinderen, Henry M. Franken, and Chris A. Vissers. On the role of basic design concepts in behaviour structuring. Computer Networks and ISDN Systems, 29(4): 413–436, March 14, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/4/cmn1659.pdf.

Turner:1997:RAS

[957] Kenneth J. Turner. Relating architecture and specification. Computer Networks and ISDN Systems, 29(4): 437–456, March 14, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/4/cmn1660.pdf.

Sinnott:1997:AAS

[958] Richard O. Sinnott and Kenneth J. Turner. Applying the architectural semantics of ODP to develop a trader specification. Computer Networks and ISDN Systems, 29(4): 457–471, March 14, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/4/cmn1661.pdf.

Blair:1997:SAM

[959] G. S. Blair, L. Blair, and J. B. Stefani. A specification architecture for multimedia systems in Open Distributed Processing. Computer Networks and ISDN Systems, 29(4): 473–500, March 14, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/4/cmn1662.pdf.

Faci:1997:SMS

[960] Mohammed Faci, Luigi Logrippo, and Bernard Stepien. Structural models for specifying telephone systems. Computer Networks and ISDN Systems, 29 (4):501–528, March 14, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/4/cmn1663.pdf.

Albanese:1997:EI

[961] Celina S. Albanese, Joseph A. Bannister, George Polyzos, Malathi Veeraraghavan, and Martina Zitterbart. Editorial introduction. Computer Networks and ISDN Systems, 29(5): 535–537, April 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/5/cmn1670.pdf.

Steenhaut:1997:SAC

[962] Kris Steenhaut, Kurt Degieter, Wouter Brissinck, and Erik Dirkx. Scheduling and admission control policies: A case study for ATM. Computer Networks and ISDN Systems, 29(5): 539–554, April 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/5/cmn1671.pdf.

Cheng:1997:QSB

[963] Lily Cheng. Quality of services based on both call admission and cell scheduling. Computer Networks and ISDN Systems, 29(5): 555–567, April 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/5/cmn1672.pdf.

Srinidhi:1997:ASA

[964] Saragur M. Srinidhi, William H. Thesling, and Vijaya K. Konangi. An adaptive scheme for admission control in ATM networks. Computer Networks and ISDN Systems, 29(5): 569–582, April 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/5/1673.pdf.

Krishnan:1997:RBC

[965] Ram Krishnan. Rate-based control schemes for ABR traffic — Design principles and performance comparison. Computer Networks and ISDN Systems, 29(5):583–593, April 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1997/
29/5/1674.pdf.

Hogan:1997:AFS

[966] M. Hogan, R. Keaney, E. Wilson, M. Hudson, M. Bickerstaff, D. Skellern, and M. Krischer. An architectural framework for the support of integrated services by broadband customer premises equipment. Computer Networks and ISDN Systems, 29(5):595–610, April 1, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/5/1675.pdf.

Akyildiz:1997:PSE

[967] I. F. Akyildiz, I. Joe, R. M. Fujimoto, and I. Nikolaidis. Parallel simulation of end-to-end ATM models.

Computer Networks and ISDN Systems, 29(6):617–630, May 1, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=6&aid=1664.

Meempat:1997:CAC

[968] Gopalakrishnan Meempat and Malur K. Sundareshan. Combined access control and routing for connection oriented traffic in high-speed networks: A two-level approach. Computer Networks and ISDN Systems, 29(6): 631–652, May 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=6&aid=1665.

Marsan:1997:FIC

[969] M. Ajmone Marsan, C. Casetti, and F. Neri. The fairness issue in the CRMA-II MAC protocol. Computer Networks and ISDN Systems, 29(6): 653–673, May 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=6&aid=1669.

Marasli:1997:ASP

[970] Rahmi Marasli, Paul D. Amer, and Phillip T. Conrad. An analytic study of partially ordered transport services. Computer Networks and ISDN Systems, 29(6):675–699, May 1, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=6&aid=1666.

Hong:1997:ATD

[971] Sung Jo Hong and Hideaki Takagi. Analysis of transmission delay for a structured-priority packetswitching system. Computer Networks and ISDN Systems, 29(6): 701–715, May 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=6&aid=1668.

Hamdi:1997:DPP

[972] M. Hamdi and W. C. Ming. Dynamic p_{i}-persistent protocol with reduced station hardware. Computer Networks and ISDN Systems, 29(6): 717–722, May 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=6&aid=1667.

Bochmann:1997:F

[973] Gregor v. Bochmann, Rachida Dssouli, and Omar Rafiq. Forte ’95. Computer Networks and ISDN Systems, 29 (7):731–733, August 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=7&aid=1743.

Dabbous:1997:HPP

[974] Walid S. Dabbous. High-performance protocol architecture. Computer Networks and ISDN Systems, 29(7):735– 744, August 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=7&aid=1744.

Diaz:1997:MMM

[975] Michel Diaz and Philippe Owezarski. From multimedia models to multimedia transport protocols. Computer Networks and ISDN Systems, 29(7): 745–758, August 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=7&aid=1746.

Khoumsi:1997:TFM

[976] Ahmed Khoumsi and Kassem Saleh. Two formal methods for the synthesis of discrete event systems. Computer Networks and ISDN Systems, 29 (7):759–780, August 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=7&aid=1745.

Choi:1997:PAL

[977] Bong Dae Choi, Chul Geun Park, and Dan Keun Sung. Performance analysis of a Leaky Bucket scheme with a threshold in the data buffer. Computer Networks and ISDN Systems, 29 (7):781–795, August 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=7&aid=1657.

Tsai:1997:ICD

[978] I-Fei Tsai and Rong-Hong Jan. Internetting connectionless data networks with a wide area public ATM network. Computer Networks and ISDN Systems, 29(7):797– 810, August 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=7&aid=1677.

Lau:1997:TDI

[979] Wing cheong Lau and San qi Li. Traffic distortion and inter-source cross-correlation in high-speed integrated networks. Computer Networks and ISDN Systems, 29(7):811– 830, August 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=7&aid=1678.

Pointek:1997:NRR

[980] Julie Pointek, Forrest Shull, Roseanne Tesoriero, and Ashok Agrawala. Net-Dyn revisited: a replicated study of network dynamics. Computer Networks and ISDN Systems, 29(7):831– 840, August 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=7&aid=1679.

Clark:1997:PSW

[981] Russell J. Clark and Mostafa H. Ammar. Providing scalable Web services using multicast communication. Computer Networks and ISDN Systems, 29 (7):841–858, August 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=7&aid=1680.

Anonymous:1997:CPV

[982] Anonymous. Credits. Computer Networks and ISDN Systems, 29(8–13): v–??, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Mojsa:1997:WEC

[983] Tomasz Mojsa and Krzysztof Zieliński. Web enabled, CORBA driven, distributed VideoTalk environment on the Java Platform. Computer Networks and ISDN Systems, 29(8–13): 865–873, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1681.

Yu:1997:HPA

[984] Jin Yu and Yuanyuan Xiang. Hypermedia presentation and authoring system. Computer Networks and ISDN Systems, 29(8–13):875– 886, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1736.

Katkere:1997:CCI

[985] Arun Katkere, Jennifer Schlenzig, and Ramesh Jain. Content-centric interactive video on the World Wide Web. Computer Networks and ISDN Systems, 29(8–13):887– 895, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1714.

Eliens:1997:JW

[986] Anton Eliëns, Martijn van Welie, Jacco van Ossenbruggen, and Bastiaan Schönhage. Jamming (on) the Web. Computer Networks and ISDN Systems, 29(8–13):897– 903, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1722.

Day:1997:NST

[987] Mark Day, John F. Patterson, and David Mitchell. The notification service transfer protocol (NSTP): Infrastructure for synchronous groupware. Computer Networks and ISDN Systems, 29(8–13):905–915, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1733.

Gall:1997:PJB

[988] Ulrich Gall and Franz J. Hauck. Promondia: A Java-based framework for real-time group communication in the Web. Computer Networks and ISDN Systems, 29(8–13): 917–926, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1682.

Edwards:1997:HSW

[989] Nigel Edwards and Owen Rees. High security Web servers and gateways. Computer Networks and ISDN Systems, 29(8–13):927–938, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1729.

Herzberg:1997:MCP

[990] Amir Herzberg and Hilik Yochai. MiniPay: Charging per click on the Web. Computer Networks and ISDN Systems, 29(8–13):939– 951, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1741.

Chu:1997:RTM

[991] Yang-Hua Chu, Joan Feigenbaum, Brian LaMacchia, Paul Resnick, and Martin Strauss. REFEREE: Trust management for Web applications. Computer Networks and ISDN Systems, 29(8–13):953–964, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1684.

Yoshida:1997:MDW

[992] Akitoshi Yoshida. MOWS: Distributed Web and cache server in Java. Computer Networks and ISDN Systems, 29(8–13):965–975, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1700.

Wooster:1997:PCE

[993] Roland P. Wooster and Marc Abrams. Proxy caching that estimates page load delays. Computer Networks and ISDN Systems, 29(8–13):977– 986, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/

                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1716.

Nabeshima:1997:JCP

[994] Masaaki Nabeshima. The Japan Cache Project: an experiment on domain cache. Computer Networks and ISDN Systems, 29(8–13):987– 995, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1709.

Scheuermann:1997:CDC

[995] Peter Scheuermann, Junho Shim, and Radek Vingralek. A case for delay-conscious caching of Web documents. Computer Networks and ISDN Systems, 29(8–13):997–1005, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1707.

Caughey:1997:FOC

[996] S. J. Caughey, D. B. Ingham, and M. C. Little. Flexible open caching for the Web. Computer Networks and ISDN Systems, 29(8–13):1007– 1017, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1690.

Damani:1997:OIT

[997] Om P. Damani, P. Emerald Chung, Yennun Huang, Chandra Kintala, and Yi-Min Wang. ONE-IP: Techniques for hosting a service on a cluster of machines. Computer Networks and ISDN Systems, 29(8–13):1019– 1027, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1705.

Ben-Shaul:1997:WEB

[998] Israel Ben-Shaul and Shlomit Ifergan. WebRule: An event-based framework for active collaboration among Web servers. Computer Networks and ISDN Systems, 29(8–13):1029– 1040, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1727.

Szmurlo:1997:NAM

[999] Maurice Szmurlo and Jacques Madelaine. A network of asynchronous micro-servers as a framework for server development. Computer Networks and ISDN Systems, 29(8–13):1041– 1051, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1683.

Kong:1997:PSU

[1000] Keith Kong and Dipak Ghosal. Pseudoserving: A user-responsible paradigm for Internet access. Computer Networks and ISDN Systems, 29(8–13): 1053–1064, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1710.

Bathelt:1997:AES

[1001] Manfred Bathelt, Ulrich Gall, Bernd Hindel, and Christian Kurzke. Accessing embedded systems via WWW: The ProWeb toolset. Computer Networks and ISDN Systems, 29(8–13): 1065–1073, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1699.

Bickmore:1997:DDI

[1002] Timothy W. Bickmore and Bill N. Schilit. Digestor: Device-independent access to the World Wide Web. Computer Networks and ISDN Systems, 29 (8–13):1075–1082, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

                                        http://www.elsevier.com/cgi-bin/
                                        cas/tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1701.

Parnes:1997:MPF

[1003] Peter Parnes, Mattias Mattsson, Kåre Synnes, and Dick Schefström. The mWeb presentation framework. Computer Networks and ISDN Systems, 29 (8–13):1083–1090, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/cgi-bin/
cas/tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1712.

Liao:1997:WMW

[1004] Tie Liao. WebCanal: A multicast Web application. Computer Networks and ISDN Systems, 29(8–13):1091– 1102, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1687.

Keller:1997:BSO

[1005] Richard M. Keller, Shawn R. Wolfe, James R. Chen, Joshua L. Rabinowitz, and Nathalie Mathe. A bookmarking service for organizing and sharing URLs. Computer Networks and ISDN Systems, 29(8–13):1103– 1114, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1703.

Vitali:1997:EHP

[1006] Fabio Vitali, Chao-Min Chiu, and Michael Bieber. Extending HTML in a principled way with displets. Computer Networks and ISDN Systems, 29(8–13): 1115–1128, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1686.

Hartman:1997:IBH

[1007] John H. Hartman, Todd A. Proebsting, and Rajesh Sundaram. Indexbased hyperlinks. Computer Networks and ISDN Systems, 29(8–13): 1129–1135, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1717.

Sakata:1997:MMR

[1008] Tsuyoshi Sakata, Hiroyuki Tada, and Tomohisa Ohtake. Metadata mediation: Representation and protocol. Computer Networks and ISDN Systems, 29(8–13):1137–1146, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1735.

Wulfekuhler:1997:FSF

[1009] Marilyn R. Wulfekuhler and William F. Punch. Finding salient features for personal Web page categories. Computer Networks and ISDN Systems, 29(8–13): 1147–1156, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/

browse.cgi?year=1997&volume=29&
issue=08-13&aid=1685.

Broder:1997:SCW

[1010] Andrei Z. Broder, Steven C. Glassman, Mark S. Manasse, and Geoffrey Zweig. Syntactic clustering of the Web. Computer Networks and ISDN Systems, 29(8–13):1157– 1166, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1706.

Takahashi:1997:ADW

[1011] Kenji Takahashi and Eugene Liang.

Analysis and design of Web-based information systems. Computer Networks and ISDN Systems, 29(8–13): 1167–1180, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1715.

Mukherjea:1997:TMW

[1012] Sougata Mukherjea, Kyoji Hirata, and Yoshinori Hara. Towards a multimedia World-Wide Web information retrieval engine. Computer Networks and ISDN Systems, 29(8–13):1181– 1191, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1721.

Shakes:1997:DRS

[1013] Jonathan Shakes, Marc Langheinrich, and Oren Etzioni. Dynamic reference sifting: A case study in the homepage domain. Computer Networks and ISDN Systems, 29(8–13): 1193–1204, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1723.

Spertus:1997:PMS

[1014] Ellen Spertus. ParaSite: Mining structural information on the Web. Computer Networks and ISDN Systems, 29(8–13):1205–1215, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1708.

Chang:1997:CME

[1015] Chia-Hui Chang and Ching-Chi Hsu. Customizable multi-engine search tool with clustering. Computer Networks and ISDN Systems, 29(8–13):1217– 1224, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1728.

Marchiori:1997:QCI

[1016] Massimo Marchiori. The quest for correct information on the Web: Hyper search engines. Computer Networks and ISDN Systems, 29(8–13): 1225–1236, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1711.

LaMacchia:1997:IFC

[1017] Brian A. LaMacchia. The Internet Fish Construction Kit. Computer Networks and ISDN Systems, 29(8–13): 1237–1248, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1698.

Ngu:1997:SLA

[1018] Daniel Siaw Weng Ngu and Xindong Wu. SiteHelper: a localized agent that helps incremental exploration of the World Wide Web. Computer Networks and ISDN Systems, 29(8–13): 1249–1255, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1730.

Carriere:1997:WSV

[1019] S. Jeromy Carrière and Rick Kazman. WebQuery: Searching and visualizing the Web through connectivity. Computer Networks and ISDN Systems, 29(8–13):1257–1267, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1737.

Maarek:1997:WSD

[1020] Yoelle S. Maarek, Michal Jacovi, Menachem Shtalhaim, Sigalit Ur, Dror Zernik, and Israel Z. Ben-Shaul. Web-Cutter: A system for dynamic and tailorable site mapping. Computer Networks and ISDN Systems, 29(8–13): 1269–1279, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1725.

Little:1997:CRW

[1021] M. C. Little, S. K. Shrivastava, S. J. Caughey, and D. B. Ingham. Constructing reliable Web applications using atomic actions. Computer Networks and ISDN Systems, 29(8–13): 1281–1290, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1695.

Hadjiefthymiades:1997:IPC

[1022] Stathes P. Hadjiefthymiades and Drakoulis I. Martakos. Improving the performance of CGI compliant database gateways. Computer Networks and ISDN Systems, 29(8–13): 1291–1304, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1726.

Arocena:1997:AWQ

[1023] Gustavo O. Arocena, Alberto O.

Mendelzon, and George A. Mihaila. Applications of a Web query language. Computer Networks and ISDN Systems, 29(8–13):1305–1315, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1718.

Mills:1997:PWW

[1024] Tim Mills, Ken Moody, and Kerry Rodden. Providing world wide access to historical sources. Computer Networks and ISDN Systems, 29(8–13): 1317–1325, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1731.

Potmesil:1997:MAV

[1025] Michael Potmesil. Maps alive:

Viewing geospatial information on the WWW. Computer Networks and ISDN Systems, 29(8–13):1327– 1342, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1688.

Pitkow:1997:SRU

[1026] James Pitkow. In search of reliable usage data on the WWW. Computer Networks and ISDN Systems, 29(8–13): 1343–1355, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1696.

Graham-Cumming:1997:HMY

[1027] John Graham-Cumming. Hits and miss-es: A year watching the Web. Computer Networks and ISDN Systems, 29(8–13):1357–1365, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1697.

Beckett:1997:ASU

[1028] Dave J. Beckett. 30% accessible — a survey of the UK Wide Web. Computer Networks and ISDN Systems, 29(8–13): 1367–1375, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1734.

Raman:1997:CSS

[1029] T. V. Raman. Cascaded speech style sheets. Computer Networks and ISDN Systems, 29(8–13):1377– 1383, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1740.

Lau:1997:WBP

[1030] Raymond Lau, Giovanni Flammia, Christine Pao, and Victor Zue. WebGALAXY: Beyond point and click — a conversational interface to a browser. Computer Networks and ISDN Systems, 29(8–13):1385–1393, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1704.

James:1997:AAH

[1031] Frankie James. AHA: Audio HTML access. Computer Networks and ISDN Systems, 29(8–13):1395–1404, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1720.

Ingham:1997:SHM

[1032] D. B. Ingham, S. J. Caughey, and M. C. Little. Supporting highly manageable Web services. Computer Networks and ISDN Systems, 29(8–13): 1405–1416, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/

                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1719.

Reilly:1997:EMN

[1033] James Reilly, Petri Niska, Luca Deri, and Dieter Gantenbein. Enabling mobile network managers. Computer Networks and ISDN Systems, 29(8–13): 1417–1428, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1693.

Gellersen:1997:WOO

[1034] Hans-Werner Gellersen, Robert Wicke, and Martin Gaedke. WebComposition: An object-oriented support system for the Web engineering lifecycle. Computer Networks and ISDN Systems, 29(8–13):1429– 1437, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1713.

Vanderheiden:1997:AAA

[1035] Gregg C. Vanderheiden. Anywhere, anytime (+ anyone) access to the nextgeneration WWW. Computer Networks and ISDN Systems, 29(8–13): 1439–1446, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1742.

Sakagami:1997:LPP

[1036] Hidekazu Sakagami and Tomonari Kamba. Learning personal preferences on online newspaper articles from user behaviors. Computer Networks and ISDN Systems, 29(8–13): 1447–1455, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1691.

Banatre:1997:PQS

[1037] Michel Banâtre, Valérie Issarny, Frédéric Leleu, and Boris Charpiot. Providing quality of service over the Web: A newspaperbased approach. Computer Networks and ISDN Systems, 29(8–13):1457– 1465, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1692.

Shimada:1997:ISC

[1038] Takao Shimada, Naoko Iwami, Takewo Tomokane, Mitsuo Hayashi, and Yasuyuki Kuwahara. Interactive scaling control mechanism for World-Wide Web systems. Computer Networks and ISDN Systems, 29(8–13):1467– 1477, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1732.

Pieper:1997:BDT

[1039] Michael Pieper and Dirk Hermsdorf. BSCW for disabled teleworkers: usability evaluation and interface adaptation of an Internet-based cooperation environment. Computer Networks and ISDN Systems, 29(8–13): 1479–1487, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1689.

Bachiochi:1997:USD

[1040] D. Bachiochi, M. Berstene, E. Chouinard, N. Conlan, M. Danchak, T. Furey, C. Neligon, and D. Way. Usability studies and designing navigational aids for the World Wide Web. Computer Networks and ISDN Systems, 29(8–13): 1489–1496, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1702.

Phanouriou:1997:TCL

[1041] Constantinos Phanouriou and Marc Abrams. Transforming commandline driven systems to Web applications. Computer Networks and ISDN Systems, 29(8–13):1497–1505, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=08-13&aid=1724.

Crespo:1997:RIL

[1042] Arturo Crespo, Bay-Wei Chang, and Eric A. Bier. Responsive interaction for a large Web application: The meteor shower architecture in the Web-Writer II Editor. Computer Networks and ISDN Systems, 29(8–13): 1507–1517, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1739.

Seligmann:1997:MM

[1043] Dorée Duncan Seligmann, Cati Laporte, and Stephan Vladimir Bugaj. The message is the medium. Computer Networks and ISDN Systems, 29(8–13): 1519–1530, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1694.

Girgensohn:1997:SII

[1044] Andreas Girgensohn and Alison Lee.

Seamless integration of interactive forms into the Web. Computer Networks and ISDN Systems, 29(8–13): 1531–1542, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=08-13&aid=1738.

Santo:1997:VGW

[1045] Harold P. Santo. Visualization and graphics on the World Wide Web. Computer Networks and ISDN Systems, 29(14):1555–1557, October 15, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1997/
29/14/1764.pdf.

Imine:1997:NAP

[1046] Mohamed Imine and Hiroshi Nagahashi. A new approach in polynomial and Bézier parametric piecewise modeling. Computer Networks and ISDN Systems, 29(14):1559– 1570, October 15, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=14&aid=1747.

Lee:1997:PBM

[1047] Jintae Lee. Physically-based modeling of brush painting. Computer Networks and ISDN Systems, 29(14): 1571–1576, October 15, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=14&aid=1748.

Brown:1997:JCA

[1048] Marc H. Brown and Roope Raisamo. JCAT: Collaborative active textbooks using Java. Computer Networks and ISDN Systems, 29(14):1577– 1586, October 15, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/14/1765.pdf.

Brown:1997:FAS

[1049] P. J. C. Brown. A fast algorithm for selective refinement of terrain meshes. Computer Networks and ISDN Systems, 29(14):1587– 1599, October 15, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=14&aid=1749.

Thalmann:1997:SVW

[1050] Daniel Thalmann, Christian Babski, Tolga Capin, Nadia Magnenat Thalmann, and Igor Sunday Pandzic. Sharing VLNET worlds on the Web. Computer Networks and ISDN Systems, 29 (14):1601–1610, October 15, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/14/1750.pdf.

Rau-Chaplin:1997:GSW

[1051] Andrew Rau-Chaplin, Brian MacKay-Lyons, Timmy Doucette, Jedrzej Gajewski, Xiangqun Hu, and Peter Spierenburg. Graphics support for a World-Wide-Web based architectural design service. Computer Networks and ISDN Systems, 29(14):1611– 1623, October 15, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=14&aid=1751.

Venkataraman:1997:PWL

[1052] Kartik Venkataraman and Tim Poston. Piece-wise linear morphing and rendering with 3D textures. Computer Networks and ISDN Systems, 29(14): 1625–1633, October 15, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=14&aid=1752.

Lee:1997:WWW

[1053] Chungnan Lee, Tong yee Lee, Tain chi Lu, and Yao tsung Chen. A World-Wide Web based distributed animation environment. Computer Networks and ISDN Systems, 29(14): 1635–1644, October 15, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1997/29/14/1753.pdf.

Stachel:1997:CSH

[1054] Hellmuth Stachel. Coordinates — A survey on higher geometry. Computer Networks and ISDN Systems, 29(14): 1645–1654, October 15, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=14&aid=1754.

Marini:1997:VRW

[1055] D. Marini, M. Rossi, L. Moltedo, and O. Salvetti. Virtual reality and Web tools to convey the visual information of ancient monuments. Computer Networks and ISDN Systems, 29(14):

1655–1660, October 15, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=14&aid=1755.

Sojka:1997:TSE

[1056] E. Sojka. Two simple and efficient algorithms for Jordan sorting and polygon cutting and clipping. Computer Networks and ISDN Systems, 29(14): 1661–1673, October 15, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=14&aid=1756.

Moutsopoulos:1997:DML

[1057] Konstantinos Moutsopoulos and Duncan Gillies. Deformable models for laparoscopic surgery simulation. Computer Networks and ISDN Systems, 29 (14):1675–1683, October 15, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/14/1763.pdf.

Coninx:1997:HCI

[1058] K. Coninx, F. Van Reeth, and E. Flerackers. 2D human-computer interaction techniques in immersive virtual environments. Computer Networks and ISDN Systems, 29(14):1685– 1693, October 15, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/14/1758.pdf.

Mace:1997:TCL

[1059] Pierre Macé. Tensorial calculus of line and plane in homogeneous coordinates. Computer Networks and ISDN Systems, 29(14):1695– 1704, October 15, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=14&aid=1759.

Rokita:1997:AIP

[1060] Przemyslaw Rokita. Application of image processing techniques in computer graphics algorithms. Computer Networks and ISDN Systems, 29(14): 1705–1713, October 15, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1997/29/14/1760.pdf.

Bechmann:1997:CFF

[1061] D. Bechmann, Y. Bertrand, and S. Thery. Continuous free form deformation. Computer Networks and ISDN Systems, 29(14):1715– 1725, October 15, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=14&aid=1757.

Patow:1997:EMR

[1062] Gustavo A. Patow and Claudio A. Delrieux. Environment mapped refraction models for low cost scanline rendering. Computer Networks and ISDN Systems, 29(14):1727– 1736, October 15, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=14&aid=1762.

Lucas:1997:NUC

[1063] Laurent Lucas, Didier Gillard, and Yannick Rémion. A new unsupervised cube-based algorithm for isosurface generation. Computer Networks and ISDN Systems, 29(14):1737– 1744, October 15, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1997/29/14/1761.pdf.

Benford:1997:VPW

[1064] Steve Benford, Dave Snowdon, Chris Brown, and Gail Reynard. Visualising and populating the Web: Collaborative virtual environments for browsing, searching and inhabiting Webspace. Computer Networks and ISDN Systems, 29(15):1751– 1761, November 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=15&aid=1769.

Kohn:1997:PGB

[1065] Daniel M. Kohn. Providing global broadband Internet access using lowearth-orbit satellites. Computer Networks and ISDN Systems, 29(15):1763– 1768, November 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=15&aid=1783.

West:1997:UIB

[1066] E. J. W. West. Using the Internet for business — Web oriented routes to market and existing IT infrastructures. Computer Networks and ISDN Systems, 29(15):1769– 1776, November 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=15&aid=1771.

Cahoon:1997:EIN

[1067] Christine Cahoon, George Dunn, Majella McCarron, and George Munroe. Environment to inspire network users (ETINU). the development of a complete online support system. Computer Networks and ISDN Systems, 29(15): 1777–1785, November 1, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=15&aid=1770.

Grebner:1997:UIM

[1068] Robert Grebner. Use of instructional material in universal teleteaching environments. Computer Networks and ISDN Systems, 29(15):1787– 1797, November 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/

tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=15&aid=1772.

Dalton:1997:AMG

[1069] C. I. Dalton and J. F. Griffin. Applying military grade security to the Internet. Computer Networks and ISDN Systems, 29(15):1799– 1808, November 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=15&aid=1775.

Steves:1997:PST

[1070] Douglas H. Steves, Chris Edmondson-Yurkanan, and Mohamed Gouda. Properties of secure transaction protocols. Computer Networks and ISDN Systems, 29(15):1809– 1821, November 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=15&aid=1776.

Koops:1997:CRE

[1071] Bert-Jaap Koops. Crypto regulation in Europe. Some key trends and issues. Computer Networks and ISDN Systems, 29(15):1823– 1831, November 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=15&aid=1773.

Horn:1997:SVT

[1072] Uwe Horn and Bernd Girod. Scalable video transmission for the Internet. Computer Networks and ISDN Systems, 29(15):1833–1842, November 1, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/comnet/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=29&issue=15&aid=1768.

Schonwalder:1997:NMD

[1073] J. Schönwälder. Network management by delegation — From research prototypes towards standards. Computer Networks and ISDN Systems, 29(15): 1843–1852, November 1, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=15&aid=1774.

Noh:1997:RSS

[1074] Tai H. Noh, Dhadesugoor R. Vaman, and Xuedao Gu. Reconfiguration for service and self-healing in ATM networks based on virtual paths. Computer Networks and ISDN Systems, 29(16):1857– 1867, December 15, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=16&aid=1766.

Weinstock:1997:OOA

[1075] Jonathan Weinstock and Rajiv Tewari. An object-oriented approach to the management of distributed application systems. Computer Networks and ISDN Systems, 29(16):1869– 1879, December 15, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=16&aid=1767.

LaCorte:1997:APC

[1076] A. La Corte, A. Lombardo, and G. Schembra. An analytical paradigm to calculate multiplexer performance in an ATM multimedia environment. Computer Networks and ISDN Systems, 29(16):1881–1900, December 15, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=16&aid=1777.

Ahmadvand:1997:DHL

[1077] Nima Ahmadvand and Terence D.

Todd. Dual-hop LANs using a wavelength routing cross connect. Computer Networks and ISDN Systems, 29(16): 1901–1917, December 15, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=16&aid=1781.

Lee:1997:PSQ

[1078] Hoon Lee and Yoshiaki Nemoto. Providing the statistical QoS objectives in high-speed networks. Computer Networks and ISDN Systems, 29(16):

1919–1931, December 15, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=16&aid=1778.

Lee:1997:SRS

[1079] Shie-Jue Lee, Yu-Wen Lai, Jian-Jou Lai, and Chi-Jong Chang. A slot reuse scheme for CRMA networks. Computer Networks and ISDN Systems, 29(16): 1933–1951, December 15, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=16&aid=1782.

Giacomazzi:1997:DPE

[1080] Paolo Giacomazzi and Achille Pattavina. Design and performance evaluation of the ATM multicast shuffleout switch. Computer Networks and ISDN Systems, 29(16):1953– 1968, December 15, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/browse/
                                        browse.cgi?year=1997&volume=29&
                                        issue=16&aid=1780.

Chan:1997:RAA

[1081] Henry C. B. Chan and Victor C. M. Leung. Reservation arbitrated access — a bandwidth on demand protocol for multiplexing variable bit rate isochronous traffic over dual bus metropolitan area networks. Computer Networks and ISDN Systems, 29(16): 1969–1987, December 15, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/browse/
browse.cgi?year=1997&volume=29&
issue=16&aid=1779.

Matta:1998:FTD

[1082] Ibrahim Matta and A. Udaya Shankar.

Fast time-dependent evaluation of multi-service networks. Computer Networks and ISDN Systems, 29(17–18): 1999–2019, February 26, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/29/17-18/1792.pdf.

Mehaoua:1998:PAC

[1083] Ahmed Mehaoua and Raouf Boutaba.

Performance analysis of cell discarding techniques for best effort video communications over ATM networks. Computer Networks and ISDN Systems, 29(17–18):2021– 2037, February 26, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/29/17-18/1796.pdf.

Present:1998:OSA

[1084] D. Présent, C. Fayet, and G. Pujolle. An optimal solution for ATM switches. Computer Networks and ISDN Systems, 29(17–18):2039– 2052, February 26, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/29/17-18/1799.pdf.

Coelho:1998:GSA

[1085] Rosângela Coelho and Samir Tohmé. A generic smoothing algorithm for realtime variable bit rate video traffic. Computer Networks and ISDN Systems, 29(17–18):2053–2066, February 26, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/comnet/sub/1998/
                                        29/17-18/1798.pdf.

Srivastava:1998:ESA

[1086] Mani Srivastava, Partho Mishra, Prathima Agrawal, and Giao Nguyen. Ethersim: a simulator for applicationlevel performance modeling of wireless and mobile ATM networks. Computer Networks and ISDN Systems, 29(17– 18):2067–2090, February 26, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/29/17-18/1795.pdf.

An:1998:TIT

[1087] Liping An, Nirwan Ansari, and Ambalavanar Arulambalam. TCP/ IP traffic over ATM networks with FMMRA ABR flow and congestion control. Computer Networks and ISDN Systems, 29(17–18):2091– 2102, February 26, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/29/17-18/1789.pdf.

Dag:1998:EAT

[1088] Tamer Dag and Ioannis Stavrakakis. Evaluation of ABR traffic management under various system time scales.

Computer Networks and ISDN Systems, 29(17–18):2103–2117, February 26, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1998/
29/17-18/1794.pdf.

Wu:1998:PTS

[1089] Cheng-Shong Wu, Jin-Chyang Jiau, and Kim-Joan Chen. Providing traffic smoothness and service guarantees within ATM networks through a ratecontrolled service discipline. Computer Networks and ISDN Systems, 29(17– 18):2119–2138, February 26, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/29/17-18/1788.pdf.

Lai:1998:PAR

[1090] Yuan-Cheng Lai and Ying-Dar Lin.

Performance analysis of rate-based flow control under a variable number of sources. Computer Networks and ISDN Systems, 29(17–18):2139– 2152, February 26, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/29/17-18/1793.pdf.

Fonseca:1998:MOP

[1091] Nelson L. S. Fonseca and John A. Silvester. Modelling the output process of an ATM multiplexer with correlated priorities. Computer Networks and ISDN Systems, 29(17–18):2153– 2164, February 26, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/29/17-18/1786.pdf.

Marsan:1998:ISE

[1092] M. Ajmone Marsan, A. Bianco, C. Casetti, C. F. Chiasserini, A. Francini, R. Lo Cigno, and M. Munafò. An integrated simulation environment for the analysis of ATM networks at multiple time scales. Computer Networks and ISDN Systems, 29(17–18):2165– 2185, February 26, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/29/17-18/1791.pdf.

Veque:1998:MMR

[1093] Véronique Vèque and Jalel Ben-Othman. MRAP: A multiservices resource allocation policy for wireless ATM network. Computer Networks and ISDN Systems, 29(17–18):2187– 2200, February 26, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/29/17-18/1797.pdf.

Chong:1998:FOR

[1094] Song Chong, Ramesh Nagarajan, and Yung-Terng Wang. First-order ratebased flow control with dynamic queue threshold for high-speed wide-area ATM networks. Computer Networks and ISDN Systems, 29(17–18):2201– 2212, February 26, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/29/17-18/1785.pdf.

Marchiori:1998:LWM

[1095] Massimo Marchiori. The limits of Web metadata, and beyond. Computer Networks and ISDN Systems, 30 (1–7):1–9, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1896.pdf.

Liechti:1998:SGF

[1096] Olivier Liechti, Mark J. Sifer, and Tadao Ichikawa. Structured graph format: XML metadata for describing Web site structure. Computer Networks and ISDN Systems, 30(1– 7):11–21, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1850.pdf.

Waugh:1998:SMS

[1097] Andrew Waugh. Specifying metadata standards for metadata tool configuration. Computer Networks and ISDN Systems, 30(1–7):23– 32, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1913.pdf.

Underwood:1998:UCP

[1098] George M. Underwood, Paul P. Maglio, and Rob Barrett. User-centered push for timely information delivery. Computer Networks and ISDN Systems, 30(1–7):33–41, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1894.pdf.

Wray:1998:DVE

[1099] Mike Wray and Rycharde Hawkes. Distributed virtual environments and VRML: an event-based architecture. Computer Networks and ISDN Systems, 30(1–7):43–51, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/comnet/sub/1998/
                                        30/1-7/1836.pdf.

Sakagami:1998:EPP

[1100] Hidekazu Sakagami, Tomonari Kamba, Atsushi Sugiura, and Yoshiyuki Koseki. Effective personalization of pushtype systems — visualizing information freshness. Computer Networks and ISDN Systems, 30(1–7): 53–63, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1871.pdf.

Chakrabarti:1998:ARC

[1101] Soumen Chakrabarti, Byron Dom, Prabhakar Raghavan, Sridhar Rajagopalan, David Gibson, and Jon Kleinberg. Automatic resource compilation by analyzing hyperlink structure and associated text. Computer Networks and ISDN Systems, 30(1– 7):65–74, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1898.pdf.

Green:1998:ALG

[1102] Stephen J. Green. Automated link generation: Can we do better than term repetition? Computer Networks and ISDN Systems, 30(1–7): 75–84, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1834.pdf.

DeTroyer:1998:WUC

[1103] O. M. F. De Troyer and C. J. Leune. WSDM: a user centered design method for Web sites. Computer Networks and ISDN Systems, 30(1–7):85–94, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1853.pdf.

Lawrence:1998:INM

[1104] Steve Lawrence and C. Lee Giles. Inquirus, the NECI meta search engine. Computer Networks and ISDN Systems, 30(1–7):95–105, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1998/
30/1-7/1906.pdf.

Brin:1998:ALS

[1105] Sergey Brin and Lawrence Page. The anatomy of a large-scale hypertextual Web search engine. Computer Networks and ISDN Systems, 30(1– 7):107–117, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1921.pdf.

Miller:1998:SFC

[1106] Robert C. Miller and Krishna Bharat. SPHINX: a framework for creating personal, site-specific Web crawlers. Computer Networks and ISDN Systems, 30 (1–7):119–130, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1875.pdf.

Juhne:1998:AJB

[1107] Jesper Jühne, Anders T. Jensen, and Kaj Grønbæk. Ariadne: a Javabased guided tour system for the World Wide Web. Computer Networks and ISDN Systems, 30(1–7): 131–139, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1935.pdf.

Byers:1998:FTI

[1108] David Byers. Full-text indexing of non-textual resources. Computer Networks and ISDN Systems, 30(1–7): 141–148, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1870.pdf.

Li:1998:FCW

[1109] Wen-Syan Li and Junho Shim. Facilitating complex Web queries through visual user interfaces and query relaxation. Computer Networks and ISDN Systems, 30(1–7): 149–159, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1936.pdf.

Cho:1998:ECT

[1110] Junghoo Cho, Hector García-Molina, and Lawrence Page. Efficient crawling through URL ordering. Computer Networks and ISDN Systems, 30 (1–7):161–172, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1919.pdf.

Ardo:1998:RDW

[1111] Anders Ardö and Sigfrid Lundberg. A regional distributed WWW search and indexing service — the DESIRE way. Computer Networks and ISDN Systems, 30(1–7): 173–183, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1900.pdf.

Krishnamurthy:1998:PSI

[1112] Balachander Krishnamurthy and Craig E.

Wills. Piggyback server invalidation for proxy cache coherency. Computer Networks and ISDN Systems, 30(1– 7):185–193, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1844.pdf.

Krishnan:1998:UCO

[1113] P. Krishnan and Binay Sugla. Utility of co-operating Web proxy caches. Computer Networks and ISDN Systems, 30 (1–7):195–203, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1892.pdf.

Nishikawa:1998:MBA

[1114] Norifumi Nishikawa, Takafumi Hosokawa, Yasuhide Mori, Kenichi Yoshida, and Hiroshi Tsuji. Memory-based architecture for distributed WWW caching proxy. Computer Networks and ISDN Systems, 30(1–7): 205–214, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1928.pdf.

Yu:1998:PSC

[1115] Philip S. Yu and Edward A. Mac-Nair. Performance study of a collaborative method for hierarchical caching in proxy servers. Computer Networks and ISDN Systems, 30(1–7): 215–224, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1829.pdf.

Ciancarini:1998:ERE

[1116] Paolo Ciancarini, Alfredo Rizzi, and Fabio Vitali. An extensible rendering engine for XML and HTML. Computer Networks and ISDN Systems, 30(1–7):225–237 (or 225–238??), April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1926.pdf.

Kristensen:1998:TRX

[1117] Anders Kristensen. Template resolution in XML/HTML. Computer Networks and ISDN Systems, 30(1– 7):239–249, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1837.pdf.

Koppen:1998:PAT

[1118] Eckhart Köppen and Gustaf Neumann. A practical approach towards active hyperlinked documents. Computer Networks and ISDN Systems, 30 (1–7):251–258, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1881.pdf.

Kistler:1998:WPL

[1119] Thomas Kistler and Hannes Marais.

WebL — a programming language for the Web. Computer Networks and ISDN Systems, 30(1–7): 259–270, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1832.pdf.

Yong:1998:JIM

[1120] Leong Kok Yong, Liu Hai, and Oliver P. Wu. Java Input Method Engine. Computer Networks and ISDN Systems, 30(1–7):271–279, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1998/
30/1-7/1915.pdf.

Barta:1998:JOO

[1121] Robert A. Barta and Markus W. Schranz. JESSICA: an objectoriented hypermedia publishing processor. Computer Networks and ISDN Systems, 30(1–7):281–290, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/comnet/sub/1998/
                                        30/1-7/1882.pdf.

Brusilovsky:1998:WBE

[1122] Peter Brusilovsky, John Eklund, and Elmar Schwarz. Web-based education for all: a tool for development adaptive courseware. Computer Networks and ISDN Systems, 30(1–7): 291–300, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1893.pdf.

Robinson:1998:DCA

[1123] David A. Robinson, Callum R. Lester, and Neil M. Hamilton. Delivering computer assisted learning across the WWW. Computer Networks and ISDN Systems, 30(1–7): 301–307, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1876.pdf.

Hinton:1998:HPH

[1124] S. Hinton. From home page to home site: effective Web resource discovery at the ANU. Computer Networks and ISDN Systems, 30(1–7): 309–316, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1828.pdf.

Hersovici:1998:SSA

[1125] Michael Hersovici, Michal Jacovi, Yoelle S. Maarek, Dan Pelleg, Menachem Shtalhaim, and Sigalit Ur. The shark-search algorithm. An application: tailored Web site mapping. Computer Networks and ISDN Systems, 30 (1–7):317–326, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1849.pdf.

Rangarajan:1998:WTC

[1126] Sampath Rangarajan, Shalini Yajnik, and Pankaj Jalote. WCP — a tool for consistent on-line update of documents in a WWW server. Computer Networks and ISDN Systems, 30 (1–7):327–336, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1858.pdf.

Simonson:1998:VAU

[1127] J. Simonson, D. Berleant, X. Zhang, M. Xie, and H. Vo. Version augmented URIs for reference permanencevia an Apache module design. Computer Networks and ISDN Systems, 30 (1–7):337–345, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1879.pdf.

Hunt:1998:NDC

[1128] Guerney D. H. Hunt, Germán S. Goldszmidt, Richard P. King, and Rajat Mukherjee. Network Dispatcher: a connection router for scalable Internet services. Computer Networks and ISDN Systems, 30(1–7): 347–357, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1899.pdf.

Chalmers:1998:OTA

[1129] Matthew Chalmers, Kerry Rodden, and Dominique Brodbeck. The order of things: activity-centred information access. Computer Networks and ISDN Systems, 30(1–7): 359–367, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1880.pdf.

Pringle:1998:WTP

[1130] Glen Pringle, Lloyd Allison, and David L. Dowe. What is a tall poppy among Web pages? Computer Networks and ISDN Systems, 30 (1–7):369–377, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1872.pdf.

Bharat:1998:TMR

[1131] Krishna Bharat and Andrei Broder. A technique for measuring the relative size and overlap of public Web search engines. Computer Networks and ISDN Systems, 30(1–7):

379–388, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1937.pdf.

Khare:1998:ODS

[1132] Rohit Khare and Adam Rifkin. The origin of (document) species. Computer Networks and ISDN Systems, 30 (1–7):389–397, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1920.pdf.

Martinez:1998:HRC

[1133] Kirk Martinez, John Cupitt, and Steve Perry. High resolution colorimetric image browsing on the Web. Computer Networks and ISDN Systems, 30 (1–7):399–405, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1873.pdf.

Low:1998:DAR

[1134] Colin Low and Laurent Babarit. Distributed 3D audio rendering. Computer Networks and ISDN Systems, 30 (1–7):407–415, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1912.pdf.

Rousseau:1998:SMW

[1135] Franck Rousseau and Andrzej Duda.

Synchronized multimedia for the WWW. Computer Networks and ISDN Systems, 30(1–7):417–429, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/comnet/sub/1998/
                                        30/1-7/1833.pdf.

Almeroth:1998:IMJ

[1136] Kevin C. Almeroth and Mostafa H. Ammar. The Interactive Multimedia Jukebox (IMJ): a new paradigm for the on-demand delivery of audio/video. Computer Networks and ISDN Systems, 30(1–7): 431–441, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1930.pdf.

Lu:1998:EPM

[1137] Qi Lu, Matthias Eichstaedt, and Daniel Ford. Efficient profile matching for large scale Webcasting. Computer Networks and ISDN Systems, 30(1– 7):443–455, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1923.pdf.

Schechter:1998:UPP

[1138] Stuart Schechter, Murali Krishnan, and Michael D. Smith. Using path profiles to predict HTTP requests. Computer Networks and ISDN Systems, 30 (1–7):457–467, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1917.pdf.

Bharat:1998:CSF

[1139] Krishna Bharat, Andrei Broder, Monika Henzinger, Puneet Kumar, and Suresh Venkatasubramanian. The Connectivity Server: fast access to linkage information on the Web. Computer Networks and ISDN Systems, 30 (1–7):469–477, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1938.pdf.

Ionesco:1998:UIR

[1140] Vlad Ionesco. Using an intranet for real-time production management: experiences and effects. Computer Networks and ISDN Systems, 30(1–7): 479–488, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1918.pdf.

Chang:1998:PEW

[1141] Bay-Wei Chang. In-place editing of Web pages: Sparrow communityshared documents. Computer Networks and ISDN Systems, 30(1–7): 489–498, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1929.pdf.

Appelt:1998:EEN

[1142] Wolfgang Appelt, Elke Hinrichs, and Gerd Woetzel. Effectiveness and efficiency: the need for tailorable user interfaces on the Web. Computer Networks and ISDN Systems, 30(1– 7):499–508, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1830.pdf.

Barrett:1998:INP

[1143] Rob Barrett and Paul P. Maglio. Intermediaries: new places for producing and manipulating Web content. Computer Networks and ISDN Systems, 30 (1–7):509–518, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1895.pdf.

Bulterman:1998:GGI

[1144] Dick C. A. Bulterman, Lynda Hardman, Jack Jansen, K. Sjoerd Mullender, and Lloyd Rutledge. GRiNS: A GRaphical INterface for creating and playing SMIL documents. Computer Networks and ISDN Systems, 30 (1–7):519–529, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1939.pdf.

Abadi:1998:SWT

[1145] Martín Abadi, Andrew Birrell, Raymie Stata, and Edward Wobber. Secure Web tunneling. Computer Networks and ISDN Systems, 30(1–7): 531–539, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1859.pdf.

Naor:1998:SAA

[1146] Moni Naor and Benny Pinkas. Secure accounting and auditing on the Web. Computer Networks and ISDN Systems, 30(1–7):541– 550, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1927.pdf.

Pitkow:1998:SWC

[1147] James E. Pitkow. Summary of WWW characterizations. Computer Networks and ISDN Systems, 30(1–7): 551–558, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1877.pdf.

Stanfield:1998:CCU

[1148] Allison Stanfield. Cyber courts:

using the Internet to assist court processes. Computer Networks and ISDN Systems, 30(1–7):559– 566, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1878.pdf.

Enguix:1998:DQW

[1149] Carlos F. Enguix. Database querying on the World Wide Web: UniGuide, an object-relational search engine for Australian universities. Computer Networks and ISDN Systems, 30 (1–7):567–572, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1891.pdf.

Varela:1998:WAJ

[1150] Carlos A. Varela and Gul A. Agha.

What after Java? From objects to actors. Computer Networks and ISDN Systems, 30(1–7): 573–577, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1890.pdf.

Nguyen:1998:OOM

[1151] Thuy-Linh Nguyen, Xindong Wu, and Sayed Sajeev. Object-oriented modeling of multimedia documents. Computer Networks and ISDN Systems, 30 (1–7):578–582, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1852.pdf.

Mason:1998:CNE

[1152] Jon Mason. Communities, networks, and education. Computer Networks and ISDN Systems, 30(1–7): 583–586, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1922.pdf.

Quentin-Baxter:1998:HLE

[1153] Megan Quentin-Baxter. Hypermedia learning environments limit access to information. Computer Networks and ISDN Systems, 30(1–7): 587–590, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1941.pdf.

Alexandre:1998:JBP

[1154] Thomas J. Alexandre. A Javabased platform for intellectual property protection on the World Wide Web. Computer Networks and ISDN Systems, 30(1–7):591–593, April 1,

1998.  CODEN CNISE9.  ISSN

0169-7552 (print), 1879-2324 (elec-

tronic).  URL http://www.cl.
cam.ac.uk/~fapp2/steganography/
bibliography/072101.html;   http:
//www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1863.pdf.

Alexandre:1998:VMC

[1155] Thomas J. Alexandre. 3D visualization of multimedia content on the World Wide Web. Computer Networks and ISDN Systems, 30(1–7): 594–596, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1862.pdf.

Allen:1998:WYC

[1156] Matthew Allen. Are we yet cyborgs? University students and the practical consequences of humanmachine subjectivity. Computer Networks and ISDN Systems, 30(1–7): 597–600, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1924.pdf.

Almeida:1998:IGC

[1157] Vírgilio F. Almeida, Márcio G. Cesário, Rodrigo C. Fonseca, Wagner Meira, Jr., and Cristina D. Murta. The influence of geographical and cultural issues on the cache proxy server workload. Computer Networks and ISDN Systems, 30 (1–7):601–603, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1909.pdf.

Amato:1998:MDS

[1158] Giuseppe Amato, Fausto Rabitti, and Pasquale Savino. Multimedia document search on the Web. Computer Networks and ISDN Systems, 30 (1–7):604–606, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1907.pdf.

Bogen:1998:GGO

[1159] Manfred Bogen, Marion Borowski, Guido Hansen, and Michael Lenz. German Government online. Computer Networks and ISDN Systems, 30(1– 7):607–609, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1904.pdf.

Bristow:1998:PSS

[1160] Paul Bristow and Matthew White. A publishing system to support secondary school teaching and learning. Computer Networks and ISDN Systems, 30(1–7):610–611, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/comnet/sub/1998/
                                        30/1-7/1934.pdf.

Cao:1998:SHM

[1161] Ling Cao, Mun-Kew Leong, Ying Lu, and Hwee-Boon Low. Searching heterogeneous multilingual bibliographic sources. Computer Networks and ISDN Systems, 30(1–7): 612–615, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1874.pdf.

Carr:1998:AIL

[1162] L. Carr, H. C. Davis, D. DeRoure, and W. Hall. Applicationindependent link processing. Computer Networks and ISDN Systems, 30 (1–7):616–618, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1883.pdf.

Caughey:1998:MGB

[1163] Steve Caughey, David Ingham, and Paul Watson. Metabroker: a generic broker for electronic commerce. Computer Networks and ISDN Systems, 30 (1–7):619–620, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1905.pdf.

Chang:1998:IQE

[1164] Chia-Hui Chang and Ching-Chi Hsu.

Integrating query expansion and conceptual relevance feedback for personalized Web information retrieval. Computer Networks and ISDN Systems, 30 (1–7):621–623, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1887.pdf.

DeRoure:1998:OFC

[1165] D. DeRoure, W. Hall, S. Reich, A. Pikrakis, G. Hill, and M. Stairmand. An open framework for collaborative distributed information management. Computer Networks and ISDN Systems, 30(1–7): 624–625, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1867.pdf.

Disch:1998:VMV

[1166] Andreas Disch, Philip Jacob, and Michael Münchhofen. Viper — a 3Dmodeling and visualization toolkit for Web-based applications. Computer Networks and ISDN Systems, 30(1– 7):626–628, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1842.pdf.

Fenech:1998:UPE

[1167] Tino Fenech. Using perceived ease of use and perceived usefulness to predict acceptance of the World Wide Web. Computer Networks and ISDN Systems, 30(1–7): 629–630, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1839.pdf.

Fraternali:1998:WDT

[1168] Piero Fraternali. Web development tools: a survey. Computer Networks and ISDN Systems, 30(1–7): 631–633, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1835.pdf.

Goeller:1998:WBC

[1169] Karen E. Goeller. Web-based collaborative learning: a perspective on the future. Computer Networks and ISDN Systems, 30(1–7): 634–635, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1940.pdf.

Green:1998:NPL

[1170] David G. Green, Paul Bristow, Jeff Ash, Larry Benton, Pam Milliken, and David Newth. Network publishing languages. Computer Networks and ISDN Systems, 30(1–7): 636–637, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1925.pdf.

Huang:1998:WNV

[1171] Mao Lin Huang, Peter Eades, and Robert F. Cohen. WebOFDAV — navigating and visualizing the Web on-line with animated context swapping. Computer Networks and ISDN Systems, 30 (1–7):638–642, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1865.pdf.

Inder:1998:PAA

[1172] Robert Inder, Matthew Hurst, and Toshikazu Kato. A prototype agent to assist shoppers. Computer Networks and ISDN Systems, 30(1–7): 643–645, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1856.pdf.

Jenkins:1998:ACW

[1173] Charlotte Jenkins, Mike Jackson, Peter Burden, and Jon Wallis. Automatic classification of Web resources using Java and Dewey Decimal Classification. Computer Networks and ISDN Systems, 30(1–7): 646–648, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1846.pdf.

Joerding:1998:IMP

[1174] Tanja Joerding and Klaus Meissner. Intelligent multimedia presentations in the Web: fun without annoyance. Computer Networks and ISDN Systems, 30 (1–7):649–650, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1857.pdf.

Khare:1998:TMW

[1175] Rohit Khare and Adam Rifkin. Trust management on the World Wide Web. Computer Networks and ISDN Systems, 30(1–7):651– 653, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1902.pdf.

Kurzke:1998:WUP

[1176] Christian Kurzke, Michael Galle, and Manfred Bathelt. WebAssist: a user profile specific information retrieval assistant. Computer Networks and ISDN Systems, 30(1–7): 654–655, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1903.pdf.

Lee:1998:GAQ

[1177] Dongwon Lee, Divesh Srivastava, and Dimitra Vista. Generating advanced query interfaces. Computer Networks and ISDN Systems, 30(1–7): 656–657, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1901.pdf.

Lowder:1998:WAS

[1178] Jason Lowder and Xindong Wu. Wide area selection as a hyperdocument search interface. Computer Networks and ISDN Systems, 30(1–7): 658–659, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1854.pdf.

Maly:1998:UWT

[1179] K. Maly, C. M. Overstreet, A. González, M. Denbar, R. Cutaran, N. Karunaratne, and C. J. Srinivas. Use of Web technology for interactive remote instruction. Computer Networks and ISDN Systems, 30(1–7): 660–662, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1855.pdf.

Miller:1998:CLR

[1180] Michael Miller and L. Jay Wantz. COOL links: ride the wave. Computer Networks and ISDN Systems, 30 (1–7):663–665, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1910.pdf.

Milosavljevic:1998:DCW

[1181] Maria Milosavljevic and Jon Oberlander. Dynamic catalogues on the WWW. Computer Networks and ISDN Systems, 30(1–7):666– 668, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1868.pdf.

Ohsawa:1998:PTH

[1182] Yukio Ohsawa, Kenki Matsuda, and Masahiko Yachida. Personal and temporary hyper bridges: 2-D interface for undefined topics. Computer Networks and ISDN Systems, 30(1–7): 669–671, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1843.pdf.

Pavett:1998:VMC

[1183] Steve Pavett, Nihal Samarawera, Neil M. Hamilton, and Gorry Fairhurst. Video Medi-CAL: supporting MPEG-2 media based computer assisted learning on an intranet. Computer Networks and ISDN Systems, 30(1–7): 672–675, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1884.pdf.

Raggett:1998:ASB

[1184] Dave Raggett. Adding style and behaviour to Web pages with a dash of Spice. Computer Networks and ISDN Systems, 30(1–7): 676–678, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1932.pdf.

Raggett:1998:AMW

[1185] Dave Raggett and Davy Batsalle.

Adding math to Web pages with EzMath. Computer Networks and ISDN Systems, 30(1–7):679–681, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1998/
30/1-7/1931.pdf.

Sember:1998:WCF

[1186] P. P. Sember, T. R. Mueller, N. Baker, M. C. Flower, B. Raskutti, and W. Wen. Web cache focusing devices. Computer Networks and ISDN Systems, 30(1–7): 682–684, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1860.pdf.

Shin:1998:RAB

[1187] Myung-Ki Shin, Jae-Yong Lee, Jung-Sook Bae, and Jin-Ho Hahm. The RTMW application: bringing multicast audio/video to the Web. Computer Networks and ISDN Systems, 30 (1–7):685–687, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1861.pdf.

Sugiura:1998:ISA

[1188] Atsushi Sugiura and Yoshiyuki Koseki. Internet Scrapbook: automating Web browsing tasks by programmingby-demonstration. Computer Networks and ISDN Systems, 30(1–7): 688–690, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1886.pdf.

Sutton:1998:GEM

[1189] Stuart A. Sutton. Gateway to Educational Materials (GEM): metadata for networked information discovery and retrieval. Computer Networks and ISDN Systems, 30(1–7): 691–693, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1897.pdf.

vanOssenbruggen:1998:RMM

[1190] Jacco van Ossenbruggen, Anton Eliëns, Lloyd Rutledge, and Lynda Hardman. Requirements for multimedia markup and style sheets on the World Wide Web. Computer Networks and ISDN Systems, 30(1–7): 694–696, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1866.pdf.

Vantaggiato:1998:AEM

[1191] Antonio Vantaggiato. Automatic exams management with the Common Lisp HTTP Server. Computer Networks and ISDN Systems, 30(1–7): 697–698, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1847.pdf.

Wang:1998:DIE

[1192] Feng-Hsu Wang. On the design of an intelligent exploratory environment for geographic climates on WWW. Computer Networks and ISDN Systems, 30 (1–7):699–700, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1885.pdf.

Wantz:1998:EMU

[1193] L. Jay Wantz and Michael Miller. Educational metadata in use. Computer Networks and ISDN Systems, 30 (1–7):701–703, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1911.pdf.

Watters:1998:WHA

[1194] C. Watters and M. Shepherd. The Web and hypertext applications. Computer Networks and ISDN Systems, 30 (1–7):704–706, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1831.pdf.

Whittington:1998:TBL

[1195] Dave Whittington and Lorna Campbell. Task-based learning environments in a virtual university. Computer Networks and ISDN Systems, 30 (1–7):707–709, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1848.pdf.

Wolf:1998:USM

[1196] Heiner Wolf and Konrad Froitzheim. User space meets document space. Computer Networks and ISDN Systems, 30(1–7):710–712, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1998/
30/1-7/1864.pdf.

Ball:1998:XDT

[1197] Steve Ball. XML and the desperate Tcl hacker. Computer Networks and ISDN Systems, 30(1–7): 713–715, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1869.pdf.

Bannon:1998:TNY

[1198] David R. Bannon and Nicholas J. Hoogenraad. Targeting next year’s students. Computer Networks and ISDN Systems, 30(1–7):716–717, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1998/
30/1-7/1840.pdf.

Gibson:1998:LTT

[1199] Julie Gibson and Philip Rutherford.

Learners are teachers too in our virtual classroom. Computer Networks and ISDN Systems, 30(1–7): 718–720, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1888.pdf.

Lu:1998:AT

[1200] Qi Lu and Lev Stesin. Audio ticker. Computer Networks and ISDN Systems, 30(1–7):721– 722, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1841.pdf.

Man:1998:HSC

[1201] Shirley Ng Wai Man. High speed connection between vBNS and SingAREN in Singapore. Computer Networks and ISDN Systems, 30(1–7): 723–726, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1950.pdf.

OReilly:1998:ALT

[1202] Meg O’Reilly and Karey Patterson. Assessing learners through the WWW. Computer Networks and ISDN Systems, 30(1–7):727– 729, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1908.pdf.

Raggett:1998:CYW

[1203] Dave Raggett. Clean up your Web pages with HP’s HTML Tidy. Computer Networks and ISDN Systems, 30 (1–7):730–732, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1933.pdf.

Roadknight:1998:VCB

[1204] Chris Roadknight and Ian Marshall. Variations in cache behavior. Computer Networks and ISDN Systems, 30 (1–7):733–735, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1845.pdf.

Rodriguez-Mula:1998:CVF

[1205] Gerard Rodríguez-Mulà, Hector García-Molina, and Andreas Paepcke. Collaborative value filtering on the Web. Computer Networks and ISDN Systems, 30(1–7):736–738, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1998/
30/1-7/1851.pdf.

Seiler:1998:EIA

[1206] R. J. Seiler, A. M. Seiler, J. M. Ireland, A. M. Guy, and K. N. Woodward. Enhancing Internet access for people with disabilities. Computer Networks and ISDN Systems, 30(1– 7):739–741, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1914.pdf.

Serbedzija:1998:WSE

[1207] Nikola B. Serbedzija. The Web supercomputing environment. Computer Networks and ISDN Systems, 30(1– 7):742–744, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1838.pdf.

Shave:1998:DLC

[1208] Chuck Shave. Developing on-line courses — teacher support needs. Computer Networks and ISDN Systems, 30 (1–7):745–746, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1916.pdf.

Susaki:1998:ISS

[1209] Seiji Susaki and Tatsuya Muramoto.

Information sharing system on the WWW with interactive communication. Computer Networks and ISDN Systems, 30(1–7):747–749, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1998/
30/1-7/1889.pdf.

Cranor:1998:LSR

[1210] Lorrie Faith Cranor. Laws, selfregulation, and P3P: will W3C’s privacy platform help make the Web safe for privacy? Computer Networks and ISDN Systems, 30(1–7): 751–753, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1946.pdf.

Bieber:1998:WHB

[1211] Michael Bieber. Will hypertext become the Web’s missing link? Computer Networks and ISDN Systems, 30 (1–7):754–756, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1943.pdf.

Brewer:1998:WAT

[1212] Judy Brewer. Web accessibility: technical and business challenges, implementation successes. Computer Networks and ISDN Systems, 30(1–7): 757–758, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1945.pdf.

Brewer:1998:AUD

[1213] Judy Brewer. Accessibility and universal design: advantages and impact for business, government and developers. Computer Networks and ISDN Systems, 30(1–7): 759–760, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1944.pdf.

Ashman:1998:MLI

[1214] Helen Ashman and Hugh Davis. Missing the 404: link integrity on the World Wide Web. Computer Networks and ISDN Systems, 30(1–7): 761–762, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/1-7/1942.pdf.

Mason:1998:XRP

[1215] Jon Mason. XML and RDF: the promise and the reality of new Web architectures. Computer Networks and ISDN Systems, 30(1–7): 763–765, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1947.pdf.

Hafid:1998:QSN

[1216] Abdelhakim Hafid, Gregor von Bochmann, and Rachida Dssouli. A quality of service negotiation approach with future reservations (NAFUR): a detailed study. Computer Networks and ISDN Systems, 30(8): 777–794, May 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/8/1784.pdf.

Sun:1998:RLC

[1217] Quan Sun and Horst Langendörfer.

Routing for low call blocking in highspeed general topology networks. Computer Networks and ISDN Systems, 30(8):795–804, May 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/8/1787.pdf.

Bourdin:1998:PCB

[1218] H. Bourdin, A. Ferreira, and K. Marcus. A performance comparison between graph and hypergraph topologies for passive star WDM lightwave networks. Computer Networks and ISDN Systems, 30(8):

805–819, May 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/8/1800.pdf.

Houck:1998:TMA

[1219] David J. Houck and Wai Sum Lai. Traffic modeling and analysis of hybrid fiber-coax systems. Computer Networks and ISDN Systems, 30(8): 821–834, May 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/8/1801.pdf.

Xiong:1998:AMA

[1220] Yijun Xiong and Lorne Mason. Analysis of multicast ATM switching networks using CRWR scheme. Computer Networks and ISDN Systems, 30(8):835–854, May 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/8/1802.pdf.

Gotzhein:1998:TFD

[1221] Reinhard Gotzhein and Jan Bredereke. Trends in formal description techniques. Computer Networks and ISDN Systems, 30(9–10): 863–864, May 27, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/9-10/1805.pdf.

Fischer:1998:FMB

[1222] Stefan Fischer and Stefan Leue. Formal methods for broadband and multimedia systems. Computer Networks and ISDN Systems, 30(9–10): 865–899, May 27, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/9-10/1806.pdf.

Hermanns:1998:SPA

[1223] H. Hermanns, U. Herzog, and V. Mertsiotakis. Stochastic process algebras — between LOTOS and Markov chains. Computer Networks and ISDN Systems, 30(9–10): 901–924, May 27, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/9-10/1807.pdf.

Brinksma:1998:POM

[1224] Ed Brinksma, Joost-Pieter Katoen, Rom Langerak, and Diego Latella. Partial order models for quantitative extensions of LOTOS. Computer Networks and ISDN Systems, 30(9–10): 925–950, May 27, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/9-10/1808.pdf.

deVeciana:1998:RAM

[1225] G. de Veciana and R. Baldick. Resource allocation in multi-service networks via pricing: statistical multiplexing. Computer Networks and ISDN Systems, 30 (9–10):951–962, May 27, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/9-10/1803.pdf.

Atlasis:1998:LSR

[1226] Antonios F. Atlasis and Athanasios V. Vasilakos. LB-SELA: rate-based access control for ATM networks. Computer Networks and ISDN Systems, 30 (9–10):963–980, May 27, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/9-10/1809.pdf.

Labourdette:1998:TOR

[1227] Jean-François P. Labourdette. Traffic optimization and reconfiguration management of multiwavelength multihop broadcast lightwave networks. Computer Networks and ISDN Systems, 30 (9–10):981–998, May 27, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/9-10/1811.pdf.

Ryden:1998:OOC

[1228] Tobias Rydén. Optimal overload control of systems with access buffers — a difficult problem. Computer Networks and ISDN Systems, 30(9–10): 999–1005, May 27, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/9-10/1804.pdf.

Kabatepe:1998:DAF

[1229] Mete Kabatepe and Kenneth S. Vastola. A delay analysis of FDQ (Fair Distributed Queue) and DQDB networks. Computer Networks and ISDN Systems, 30(11): 1013–1027, June 22, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/11/1810.pdf.

Peyravian:1998:DNC

[1230] Mohammad Peyravian and Ajay D.

Kshemkalyani. Decentralized network connection preemption algorithms. Computer Networks and ISDN Systems, 30(11):1029– 1043, June 22, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/11/1812.pdf.

Tuok:1998:FSU

[1231] Randall Tuok and Luigi Logrippo. Formal specification and use case generation for a mobile telephony system. Computer Networks and ISDN Systems, 30(11):1045–1063, June 22, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1998/
30/11/1813.pdf.

Marsan:1998:DSA

[1232] M. Ajmone Marsan, A. Fumagalli, E. Leonardi, F. Neri, and P. Poggiolini. Daisy: a scalable alloptical packet network with multifiber ring topology. Computer Networks and ISDN Systems, 30(11): 1065–1082, June 22, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/11/1824.pdf.

Sharma:1998:EWP

[1233] Neeraj K. Sharma. Effect of windowing policies for input buffered ATM switch. Computer Networks and ISDN Systems, 30(11):1083– 1090, June 22, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/11/1826.pdf.

Bonjour:1998:IAN

[1234] Dominique Bonjour, Omar Elloumi, and Hossam Afifi. Internet applications over native ATM. Computer Networks and ISDN Systems, 30(12): 1097–1110, July 13, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/12/1815.pdf.

Tai:1998:STS

[1235] Kuo-Chung Tai and Yu-Chiou Young. Synchronizable test sequences of finite state machines. Computer Networks and ISDN Systems, 30(12): 1111–1134, July 13, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/12/1827.pdf.

Varvarigos:1998:ERC

[1236] Emmanouel A. Varvarigos and Vishal Sharma. An efficient reservation connection control protocol for gigabit networks. Computer Networks and ISDN Systems, 30(12): 1135–1156, July 13, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/12/1823.pdf.

Govindan:1998:RSI

[1237] Ramesh Govindan, Cengiz Alaettinoglu, Kannan Varadhan, and Deborah Estrin. Route servers for interdomain routing. Computer Networks and ISDN Systems, 30(12): 1157–1174, July 13, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/12/1822.pdf.

Michard:1998:ARD

[1238] Alain Michard, Vassilis Christophides, Michel Scholl, Mike Stapleton, Dale Sutcliffe, and Anne-Marie Vercoustre. The Aquarelle resource discovery system. Computer Networks and ISDN Systems, 30(13):1185– 1200, August 3, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/13/1821.pdf.

Steinmann:1998:RMM

[1239] Uwe Steinmann and David Shearer.

Reusing multi-media components: a catalogue implementation. Computer Networks and ISDN Systems, 30(13): 1201–1207, August 3, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/13/1819.pdf.

Selley:1998:DDI

[1240] Gordon Selley and Mark Gaved. Development of the Design Information Centre: Multimedia Assets for Industrial Design (MAID). Computer Networks and ISDN Systems, 30(13): 1209–1222, August 3, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/13/1816.pdf.

Hofstetter:1998:MAL

[1241] Isabella Hofstetter. Multimedia applications for local newspapers and local information. Computer Networks and ISDN Systems, 30(13): 1223–1232, August 3, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/13/1954.pdf.

Corona:1998:MBP

[1242] Alberto G. Corona. Multimedia Broker project: a tool for electronic information publishers. Computer Networks and ISDN Systems, 30(13): 1233–1236, August 3, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/13/1817.pdf.

terStal:1998:TOC

[1243] W. G. ter Stal, J.-H. Beijert, G. de Bruin, J. van Gent, F. M. G. de Jong, W. Kraaij, K. Netter, and G. Smart. Twenty-One: crosslanguage disclosure and retrieval of multimedia documents on sustainable development. Computer Networks and ISDN Systems, 30(13): 1237–1248, August 3, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/13/1818.pdf.

Faux:1998:IAP

[1244] Ivor Faux, Elke Radeke, Franz-Josef Stewing, Ger v.d. Broek, Pieter Kesteloot, and Andreas Sabin. Intelligent access, publishing and collaboration in global engineering networking. Computer Networks and ISDN Systems, 30(13):1249– 1262, August 3, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/13/1820.pdf.

Stephenson:1998:EPR

[1245] G. A. Stephenson. Electronic publishing resources on the Web. Computer Networks and ISDN Systems, 30 (13):1263–1271, August 3, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/13/1955.pdf.

Uyar:1998:DSA

[1246] M. Ümit Uyar. Dual-state augmentation for minimizing conformance test costs. Computer Networks and ISDN Systems, 30(14):1277– 1294, August 31, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/14/1825.pdf.

Chatterjee:1998:MEI

[1247] Samir Chatterjee, Cherian S. Thachenkary, and Joseph L. Katz. Modeling the economic impacts of broadband residential services. Computer Networks and ISDN Systems, 30(14):1295– 1310, August 31, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/14/1951.pdf.

deLima:1998:PMS

[1248] Heliomar Medeiros de Lima and Otto Carlos Muniz Bandeira Duarte. Pointto-multipoint SR ARQ scheme with accumulative acknowledgment for satellite communications. Computer Networks and ISDN Systems, 30(14):1311– 1325, August 31, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/14/1952.pdf.

Berger:1998:STM

[1249] Arthur W. Berger, Maurizio Naldi, Livia De Giovanni, and Manuel Villén-Altamirano. Standardization of traffic measurements and models for broadband networks: open issues. Computer Networks and ISDN Systems, 30 (14):1327–1340, August 31, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/14/1953.pdf.

Jacob:1998:CPS

[1250] Lillykutty Jacob and Anurag Kumar. Comparative performance of scheduling strategies for switching and multiplexing in a hub based ATM network: a simulation study. Computer Networks and ISDN Systems, 30(14): 1341–1354, August 31, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/14/1962.pdf.

Kimbler:1998:CAS

[1251] Kristofer Kimbler, Carla Capellmann, and Hugo Velthuijsen. Comprehensive approach to service interaction handling. Computer Networks and ISDN Systems, 30(15):1363– 1387, September 16, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/15/1958.pdf.

Turner:1998:ADI

[1252] Kenneth J. Turner. An architectural description of intelligent network features and their interactions. Computer Networks and ISDN Systems, 30(15): 1389–1419, September 16, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/15/1959.pdf.

Tsang:1998:NOP

[1253] Simon Tsang and Evan H. Magill.

The network operator’s perspective: detecting and resolving feature interaction problems. Computer Networks and ISDN Systems, 30(15):1421– 1441, September 16, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/15/1960.pdf.

Ciulli:1998:CEB

[1254] Nicola Ciulli, Stefano Giordano, and Daniele Sparano. A cooperative environment based on augmented reality: from telepresence to performance issues. Computer Networks and ISDN Systems, 30(16–18):1447– 1455, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/1973.pdf.

Fromme:1998:CCM

[1255] Michael Fromme, Bernd Böker, Lutz Grüneberg, and Helmut Pralle. A conference control management system for the Internet: Confman 2.0. Computer Networks and ISDN Systems, 30(16–18):1457–1465, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/1991.pdf.

Berzsenyi:1998:DIV

[1256] Miklós Berzsenyi, István Vajk, and Hui Zhang. Design and implementation of a video on-demand system. Computer Networks and ISDN Systems, 30(16–18):1467–1473, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/1974.pdf.

Dixon:1998:DPD

[1257] Anne Dixon. Distributed publishing: delivering publisher-neutral information to the research community. Computer Networks and ISDN Systems, 30(16–18):1475–1478, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/1992.pdf.

Worsfold:1998:SGF

[1258] Emma Worsfold. Subject gateways: fulfilling the DESIRE for knowledge. Computer Networks and ISDN Systems, 30(16–18):1479– 1489, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/1993.pdf.

Nick:1998:EIB

[1259] Achim Nick, Jürgen Koenemann, and Elmar Schalück. ELFI: information brokering for the domain of research funding. Computer Networks and ISDN Systems, 30(16–18):1491– 1500, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/1975.pdf.

Schunter:1998:SRS

[1260] Matthias Schunter, Michael Waidner, and Dale Whinnett. A status report on the SEMPER framework for secure electronic commerce. Computer Networks and ISDN Systems, 30(16–18): 1501–1510, September 30, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/1994.pdf.

Verschuren:1998:SAS

[1261] Ton Verschuren. Smart access: strong authentication on the web. Computer Networks and ISDN Systems, 30(16–18):1511–1519, September 30,

1998.  CODEN CNISE9.  ISSN

0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1998/
30/16-18/2009.pdf.

Berger:1998:PPP

[1262] Andreas Berger. Privacy protection for public directory services. Computer Networks and ISDN Systems, 30(16–18):1521–1529, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/1995.pdf.

Sousa:1998:FDT

[1263] Pedro Sousa and Vasco Freitas. A framework for the development of tolerant real-time applications. Computer Networks and ISDN Systems, 30(16–18):1531–1541, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1998/
30/16-18/1996.pdf.

RuizPinar:1998:MSB

[1264] Francisco Javier Ruiz Piñar, Ángel Fernández del Campo, Carlos Miguel Nieto, León Vidaller Sisó, Antonio Martínez Mas, and Juan Antonio Carral Pelayo. Multimedia systems based on satellite technology. Computer Networks and ISDN Systems, 30(16–18):1543– 1549, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/1997.pdf.

Zhang:1998:DNM

[1265] Dianlong Zhang and Werner Zorn. Developing network management applications in an application-oriented way using mobile agent. Computer Networks and ISDN Systems, 30(16–18): 1551–1557, September 30, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/1976.pdf.

Dixon:1998:BLS

[1266] Tim Dixon, Ingrid Melve, Rui Meneses, and Ton Verschuren. Building large-scale information services: tools and experiences from the DE-SIRE project. Computer Networks and ISDN Systems, 30(16–18):1559– 1569, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/1998.pdf.

Valkenburg:1998:SCP

[1267] Peter Valkenburg, Dave Beckett, Martin Hamilton, and Simon Wilkinson. Standards in the CHIC-Pilot distributed indexing architecture. Computer Networks and ISDN Systems, 30(16–18):1571–1578, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/2026.pdf.

Lee:1998:IBF

[1268] Matthew K. O. Lee. Internet-based financial EDI: towards a theory of its organizational adoption. Computer Networks and ISDN Systems, 30(16–18):

1579–1588, September 30, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/16-18/2016.pdf.

Watkins:1998:CES

[1269] Judy Watkins. COPEARMS and ERMS: safeguarding intellectual property rights in the digital age. Computer Networks and ISDN Systems, 30(16–18):1589–1595, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/comnet/sub/1998/
                                        30/16-18/1999.pdf.

Pordesch:1998:NSA

[1270] Ulrich Pordesch. Negotiating security among end users: concept and test in a simulation study for the public health service. Computer Networks and ISDN Systems, 30(16–18): 1597–1605, September 30, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/16-18/2000.pdf.

Horneffer:1998:MPA

[1271] Martin Horneffer. Methods for performance-analysis of Internet access points. Computer Networks and ISDN Systems, 30(16–18):1607– 1615, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/16-18/1977.pdf.

Mohacsi:1998:TII

[1272] János Mohácsi, Szabolcs Szigeti, and Tamás Máray. Testing IPv6 implementations. Computer Networks and ISDN Systems, 30(16–18):1617– 1625, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/1978.pdf.

Breiter:1998:UAR

[1273] F. Breiter, S. Kühn, E. Robles, and A. Schill. The usage of advance reservation mechanisms in distributed multimedia applications. Computer Networks and ISDN Systems, 30(16–18): 1627–1635, September 30, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/2001.pdf.

Bonkowski:1998:OIS

[1274] Christian Bonkowski, Marion Borowski, Guido Hansen, and Michael Lenz. Online information services for the German government’s audiences. Computer Networks and ISDN Systems, 30(16–18):1637–1647, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1998/
30/16-18/2002.pdf.

Strom:1998:RED

[1275] Jim Strom and George Neisser. Reengineering the desktop: taking the knowledge-base to the user to support education and city communities. Computer Networks and ISDN Systems, 30(16–18):1649–1656, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/2003.pdf.

Belov:1998:EIL

[1276] Serge D. Belov, Sergey V. Bredikhin, Serge P. Kovalyov, Serge A. Kulagin, Sam L. Musher, Natalie G. Scherbakova, and Igor V. Shabalnikov. The emerging Internet landscape in Siberia. Computer Networks and ISDN Systems, 30(16–18):1657– 1662, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/2004.pdf.

Gruntorad:1998:THS

[1277] Jan Gruntorád. TEN-34CZ: Highspeed academic network on microwave links. Computer Networks and ISDN Systems, 30(16–18):1663– 1667, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/2010.pdf.

Nakonieczny:1998:PSB

[1278] Mscislaw Nakonieczny, Stanislaw Starzak, Maciej Stroiński, and Jan Weglarz. Polish scientific broadband network: POL34. Computer Networks and ISDN Systems, 30(16–18): 1669–1676, September 30, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/2028.pdf.

Greenhalgh:1998:PNT

[1279] Chris Greenhalgh, Steve Benford, Adrian Bullock, Nico Kuijpers, and Kurt Donkers. Predicting network traffic for collaborative virtual environments. Computer Networks and ISDN Systems, 30(16–18):1677– 1685, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/2017.pdf.

Grimm:1998:LTB

[1280] C. Grimm, J.-S. Vöckler, and H. Pralle.

Load and traffic balancing in large scale cache meshes. Computer Networks and ISDN Systems, 30(16–18): 1687–1695, September 30, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/2005.pdf.

Cozzani:1998:TSM

[1281] Irene Cozzani and Stefano Giordano.

Traffic sampling methods for end-to-end QoS evaluation in large heterogeneous networks. Computer Networks and ISDN Systems, 30(16–18): 1697–1706, September 30, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/16-18/2006.pdf.

Klein:1998:TSH

[1282] Arno Klein. Teleteaching scenarios for high bandwidth networks. Computer Networks and ISDN Systems, 30(16–18):1707–1716, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/16-18/2007.pdf.

Allen:1998:WIM

[1283] Robert Allen. The Web: interactive and multimedia education. Computer Networks and ISDN Systems, 30(16–18):1717–1727, September 30, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/16-18/2008.pdf.

Lee:1998:FSA

[1284] David Lee, K. K. Ramakrishnan, and W. Melody Moh. A formal specification of the ATM ABR rate control scheme. Computer Networks and ISDN Systems, 30(19):1735– 1748, October 14, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/19/1967.pdf.

invited paper.

Golmie:1998:ASM

[1285] N. Golmie, Y. Saintillan, and D. Su. ABR switch mechanisms: design issues and performance evaluation. Computer Networks and ISDN Systems, 30 (19):1749–1761, October 14, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/19/1968.pdf.

Iliadis:1998:PTT

[1286] Ilias Iliadis and Daniel Orsatti. Performance of TCP traffic and ABR congestion control mechanisms. Computer Networks and ISDN Systems, 30(19): 1763–1774, October 14, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/19/1969.pdf.

Narvaez:1998:ABS

[1287] Paolo Narváez and Kai-Yeung Siu. Acknowledgment bucket scheme for regulating TCP flow over ATM. Computer Networks and ISDN Systems, 30 (19):1775–1791, October 14, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/19/1972.pdf.

Ren:1998:MMA

[1288] Wenge Ren, Kai-Yeung Siu, Hiroshi Suzuki, and Masayuki Shinohara. Multipoint-to-multipoint ABR service in ATM. Computer Networks and ISDN Systems, 30(19):1793– 1810, October 14, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/19/1971.pdf.

Kalyanaraman:1998:DCV

[1289] Shivkumar Kalyanaraman, Raj Jain, Jianping Jiang, Rohit Goyal, and Sonia Fahmy. Design considerations for the virtual source/virtual destination (VS/VD) feature in the ABR service of ATM networks. Computer Networks and ISDN Systems, 30(19): 1811–1824, October 14, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/19/1970.pdf.

Baerten:1998:UVJ

[1290] Herbert Baerten and Frank Van Reeth. Using VRML and JAVA to visualize 3D algorithms in computer graphics education. Computer Networks and ISDN Systems, 30(20–21):1833– 1839, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/20-21/2013.pdf.

Champciaux:1998:CBU

[1291] Laurent Champciaux. Classification: a basis for understanding tools in declarative modelling. Computer Networks and ISDN Systems, 30(20–21):1841– 1852, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/20-21/1986.pdf.

Chan:1998:SML

[1292] Stephen C. F. Chan, Vincent T. Y. Ng, and Albert S. F. Au. A solid modeling library for the World Wide Web. Computer Networks and ISDN Systems, 30(20–21):1853– 1863, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/20-21/2021.pdf.

Choi:1998:CBS

[1293] Jeongdan Choi, Jinseong Choi, Chanjong Park, and Donghyun Kim. A cell-based shared virtual world management mechanism in the cyber mall system. Computer Networks and ISDN Systems, 30(20–21):1865– 1874, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/20-21/2025.pdf.

Colin:1998:WMD

[1294] Christian Colin, Emmanuel Desmontils, Jean-Yves Martin, and Jean-Philippe Mounier. Working modes with a declarative modeler. Computer Networks and ISDN Systems, 30(20–21): 1875–1886, November 12, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/20-21/1984.pdf.

Conci:1998:FIA

[1295] Aura Conci and Claudia Belmiro Proença. A fractal image analysis system for fabric inspection based on a box-counting method. Computer Networks and ISDN Systems, 30(20–21): 1887–1895, November 12, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/20-21/2019.pdf.

Dai:1998:HLC

[1296] Ping Dai. A high level class library for dynamic virtual world modelling. Computer Networks and ISDN Systems, 30(20–21):1897– 1905, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/20-21/2020.pdf.

Felez:1998:VRA

[1297] Jesús Felez, Carlos Vera, and M. Luisa Martínez. Virtual reality applied to traffic accident analysis. Computer Networks and ISDN Systems, 30(20–21):1907–1914, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/comnet/sub/1998/
                                        30/20-21/1989.pdf.

Gradinscak:1998:SCB

[1298] Zlatko Gradinscak. A study on computer-based geometric modelling in engineering graphics. Computer Networks and ISDN Systems, 30(20–21): 1915–1922, November 12, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/20-21/2015.pdf.

Heleno:1998:AAV

[1299] Paulo Heleno and Manuel Próspero dos Santos. Artificial animals in virtual ecosystems. Computer Networks and ISDN Systems, 30(20–21):1923– 1932, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/20-21/1982.pdf.

Herder:1998:TWS

[1300] Jens Herder. Tools and widgets for spatial sound authoring. Computer Networks and ISDN Systems, 30(20–21): 1933–1940, November 12, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/20-21/1983.pdf.

Huerta:1998:MMU

[1301] Joaquín Huerta, Miguel Chover, José Ribelles, and Ricardo Quirós. Multiresolution modeling using binary space partitioning trees. Computer Networks and ISDN Systems, 30(20–21): 1941–1950, November 12, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/20-21/1979.pdf.

Jones:1998:TCM

[1302] Huw Jones. Trends in computermediated art and design: examples from the centre for electronic arts. Computer Networks and ISDN Systems, 30(20–21):1951– 1960, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/20-21/2029.pdf.

Kardassevitch:1998:HRP

[1303] Cyril Kardassevitch, Mathias Paulin, Jean Pierre Jessel, and René Caubet. A hierarchical radiosity platform using efficient data structures and VRML 97. Computer Networks and ISDN Systems, 30(20–21):1961– 1974, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/20-21/2011.pdf.

McAllister:1998:SIA

[1304] David F. McAllister, Robert D. Rodman, Donald L. Bitzer, and Andrew S. Freeman. Speaker independence in automated lip-sync for audio-video communication. Computer Networks and ISDN Systems, 30(20–21):1975– 1980, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/20-21/2024.pdf.

Mohler:1998:MCM

[1305] James L. Mohler. Migrating course materials to the World Wide Web: a case study of the Department of Technical Graphics at Purdue University. Computer Networks and ISDN Systems, 30(20–21):1981– 1990, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/20-21/1980.pdf.

Ng:1998:THK

[1306] Vincent Ng, Stephen Chan, Chak Man Ng, and Fergus Tang. Touring Hong Kong via the WWW. Computer Networks and ISDN Systems, 30(20–21): 1991–2001, November 12, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/20-21/2022.pdf.

Oikarinen:1998:MIP

[1307] Jarkko T. Oikarinen and Lasse J. Jyrkinen. Maximum intensity projection by 3-dimensional seed filling in view lattice. Computer Networks and ISDN Systems, 30(20–21):2003– 2014, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/20-21/1985.pdf.

Park:1998:ACG

[1308] Dong-Gyu Park, Sang-Hyun Park, Hoon-Hee Lee, and Hwan-Gue Cho. An automatic crowd generation system using image processing techniques. Computer Networks and ISDN Systems, 30(20–21):2015–2023, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/20-21/1981.pdf.

Reitan:1998:VCI

[1309] Paula J. Reitan. 3D visualization of color image histograms. Computer Networks and ISDN Systems, 30(20–21): 2025–2035, November 12, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/20-21/2023.pdf.

Rodrigues:1998:MST

[1310] M. A. F. Rodrigues, D. F. Gillies, and P. Charters. Modelling and simulation of the tongue during laryngoscopy. Computer Networks and ISDN Systems, 30(20–21):2037– 2045, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/20-21/1990.pdf.

Rokita:1998:CCG

[1311] Przemyslaw Rokita. Compositing computer graphics and real world video sequences. Computer Networks and ISDN Systems, 30(20–21):2047– 2057, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/20-21/2014.pdf.

Tong:1998:NPC

[1312] Franky H. F. Tong and David Zhang. A new progressive colour image transmission scheme for the World Wide Web. Computer Networks and ISDN Systems, 30(20–21):2059– 2064, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/20-21/2018.pdf.

VanReeth:1998:UNM

[1313] Frank Van Reeth and Karin Coninx. Using networked multimedia in computer graphics education. Computer Networks and ISDN Systems, 30(20–21):2065–2073, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/20-21/2012.pdf.

Danzig:1998:NAD

[1314] Peter Danzig. NetCache architecture and deployment. Computer Networks and ISDN Systems, 30(22–23): 2081–2091, November 25, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/22-23/2056.pdf.

Afonso:1998:QWC

[1315] Manuel Afonso, Alexandre Santos, and Vasco Freitas. QoS in Web caching. Computer Networks and ISDN Systems, 30(22–23):2093–2103, November 25, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/22-23/2061.pdf.

Melve:1998:WKY

[1316] Ingrid Melve. When to kill your siblings: cache mesh relation analysis. Computer Networks and ISDN Systems, 30(22–23):2105–2111, November 25, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/22-23/2058.pdf.

Kangasharju:1998:DIS

[1317] Jussi Kangasharju, Young Gap Kwon, and Antonio Ortega. Design and implementation of a soft caching proxy. Computer Networks and ISDN Systems, 30(22–23):2113–2121, November 25, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/22-23/2060.pdf.

Marshall:1998:LCP

[1318] Ian Marshall and Chris Roadknight.

Linking cache performance to user behaviour. Computer Networks and ISDN Systems, 30(22–23):2123– 2130, November 25, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/22-23/2050.pdf.

Huffaker:1998:VGT

[1319] Bradley Huffaker, Jaeyeon Jung, Evi Nemeth, Duane Wessels, and K. Claffy. Visualization of the growth and topology of the NLANR caching hierarchy. Computer Networks and ISDN Systems, 30(22–23): 2131–2139, November 25, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/22-23/2045.pdf.

Beck:1998:IDS

[1320] Micah Beck and Terry Moore. The Internet2 Distributed Storage Infrastructure project: an architecture for Internet content channels. Computer Networks and ISDN Systems, 30(22–23): 2141–2148, November 25, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/22-23/2051.pdf.

Reddy:1998:ECB

[1321] Mike Reddy and Graham P. Fletcher. Exp1: a comparison between a simple adaptive caching agent using document life histories and existing cache techniques. Computer Networks and ISDN Systems, 30(22–23):2149– 2153, November 25, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/22-23/2044.pdf.

Rousskov:1998:CD

[1322] Alex Rousskov and Duane Wessels. Cache digests. Computer Networks and ISDN Systems, 30(22–23):2155– 2168, November 25, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/22-23/2057.pdf.

Michel:1998:AWC

[1323] Scott Michel, Khoi Nguyen, Adam Rosenstein, Lixia Zhang, Sally Floyd, and Van Jacobson. Adaptive Web caching: towards a new global caching architecture. Computer Networks and ISDN Systems, 30(22–23):2169– 2177, November 25, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/22-23/2052.pdf.

Almeida:1998:MPP

[1324] Jussara Almeida and Pei Cao. Measuring proxy performance with the Wisconsin Proxy Benchmark. Computer Networks and ISDN Systems, 30(22–23):2179–2192, November 25, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/comnet/sub/1998/
30/22-23/2053.pdf.

Legedza:1998:UNL

[1325] Ulana Legedza and John Guttag. Using network-level support to improve cache routing. Computer Networks and ISDN Systems, 30(22–23):2193– 2201, November 25, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/22-23/2054.pdf.

Kurcewicz:1998:FAW

[1326] Michal Kurcewicz, Wojtek Sylwestrzak, and Adam Wierzbicki. A filtering algorithm for Web caches. Computer Networks and ISDN Systems, 30(22–23): 2203–2209, November 25, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/22-23/2047.pdf.

Russell:1998:CCF

[1327] Mark Russell and Tim Hopkins. CFTP: a caching FTP server. Computer Networks and ISDN Systems, 30(22–23): 2211–2222, November 25, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/22-23/2046.pdf.

Rodriguez:1998:IWC

[1328] Pablo Rodriguez, Keith W. Ross, and Ernst W. Biersack. Improving the WWW: caching or multicast? Computer Networks and ISDN Systems, 30(22–23):2223–2243, November 25, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/22-23/2043.pdf.

Touch:1998:LPC

[1329] Joe Touch and Amy S. Hughes. LSAM proxy cache: a multicast distributed virtual cache. Computer Networks and ISDN Systems, 30(22–23):2245– 2252, November 25, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/22-23/2059.pdf.

Rabinovich:1998:AHC

[1330] Michael Rabinovich, Jeff Chase, and Syam Gadde. Not all hits are created equal: cooperative proxy caching over a wide-area network. Computer Networks and ISDN Systems, 30(22–23): 2253–2259, November 25, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/22-23/2055.pdf.

Kurcewicz:1998:DWC

[1331] Michal Kurcewicz, Wojtek Sylwestrzak, and Adam Wierzbicki. A distributed WWW cache. Computer Networks and ISDN Systems, 30(22–23):2261– 2267, November 25, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/22-23/2049.pdf.

Grimm:1998:RRC

[1332] C. Grimm, J.-S. Vöckler, and H. Pralle. Request routing in cache meshes. Computer Networks and ISDN Systems, 30(22–23):2269– 2278, November 25, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/22-23/2048.pdf.

Wills:1998:ENS

[1333] Craig E. Wills, David C. Brown, Bertrum V. Dunskus, and Jonathan Kemble. Evaluating network serviceability. Computer Networks and ISDN Systems, 30(24):2283– 2291, December 14, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/24/1963.pdf.

Kalyanaraman:1998:UIL

[1334] Shivkumar Kalyanaraman, Raj Jain, Rohit Goyal, Sonia Fahmy, and Pradeep Samudra. Use-it or loseit policies for the available bit rate (ABR) service in ATM networks. Computer Networks and ISDN Systems, 30 (24):2293–2308, December 14, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

                                        http://www.cis.ohio-state.edu/~
                                        jain/papers/uili.htm;   http://
                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/24/1964.pdf.

McKeown:1998:QCI

[1335] Nick McKeown and Thomas E. Anderson. A quantitative comparison of iterative scheduling algorithms for input-queued switches. Computer Networks and ISDN Systems, 30(24):2309– 2326, December 14, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/24/1966.pdf.

Almesberger:1998:UQS

[1336] Werner Almesberger, Leena Chandran-Wadia, Silvia Giordano, Jean-Yves Le Boudec, and Rolf Schmid. Using quality of service can be simple: Arequipa with renegotiable ATM connections. Computer Networks and ISDN Systems, 30(24):2327– 2336, December 14, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        comnet/sub/1998/30/24/2030.pdf.

Iun:1998:PAN

[1337] Dennis P. C. Iun and Xi-Ren Cao.

Performance analysis of a nonblocking ATM switch with a bufferless internal speedup fabric. Computer Networks and ISDN Systems, 30(24):2337– 2347, December 14, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/24/2062.pdf.

Mohan:1998:PRW

[1338] G. Mohan, C. Siva Ram Murthy, and R. Vijayashankar. Permutation routing in wavelength-routed wrapped-around shuffle networks using fewer wavelengths. Computer Networks and ISDN Systems, 30(24):2349– 2358, December 14, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/24/2063.pdf.

Li:1998:PIC

[1339] Chung-Sheng Li, Yoram Ofek, Adrian Segall, and Khosrow Sohraby. Pseudoisochronous cell forwarding. Computer Networks and ISDN Systems, 30(24): 2359–2371, December 14, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/24/2064.pdf.

Kundaeli:1998:CSC

[1340] H. N. Kundaeli. Comparison of some CCITT and computed design parameters. Computer Networks and ISDN Systems, 30(24):2373– 2376, December 14, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/store/
comnet/sub/1998/30/24/1965.pdf.

Amer:1993:ECS

[1341] K. Amer, K. Christensen, and T. Toher. Experiments with client/server multimedia on token ring. In IEEE Computer Society [2041], pages 2–6. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Gibbon:1993:RDD

[1342] J. F. Gibbon and T. D. C. Little. Real-time data delivery for multimedia networks. In IEEE Computer Society [2041], pages 7–16. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Doganata:1993:PAV

[1343] Y. N. Doganata, A. Ruiz, and A. N. Tantawi. Performance analysis of voice and video services in isochronous LANs and systems. In IEEE Computer Society [2041], pages 17–24. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Bisdikian:1993:EPT

[1344] C. C. Bisdikian, B. Patel, F. Schaffa, and M. Willebeek-LeMair. On the effectiveness of priorities in token ring for multimedia traffic. In IEEE Computer Society [2041], pages 25–33. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Armitage:1993:UCL

[1345] G. J. Armitage and K. M. Adams.

Using the common LAN to introduce ATM connectivity. In IEEE Computer Society [2041], pages 34–43. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Tsai:1993:DAH

[1346] Z. Tsai, K. Yu, and F. Lai. Design and analysis of a hierarchical and modular local ATM switch. In IEEE Computer Society [2041], pages 44– 51. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Brown:1993:CAC

[1347] T. A. Brown and D. C. Kostick. “… and CNM for all:” customer network management services for broadband data services. In IEEE Computer Society [2041], pages 52–61. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Amley:1993:TTS

[1348] C. Amley, J. R. Bumblis, and L. Gale.

A technical trial of the SMDS Customer Network Management (CNM) by 3M and US West. In IEEE Computer Society [2041], pages 62– 69. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Adam:1993:EVN

[1349] J. F. Adam, H. H. Houh, and D. L. Tennenhouse. Experience with the VuNet: A network architecture for a distributed multimedia system. In IEEE Computer Society [2041], pages 70– 76. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Foldvik:1993:DLA

[1350] R. G. Foldvik and W. Walling. Distance learning and ATM: A trial at the University of Minnesota. In IEEE Computer Society [2041], pages 77– 80. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Chatterjee:1993:DCC

[1351] S. Chatterjee and M. A. Bassiouni. Dynamic congestion control schemes in high-speed multimedia networks. In IEEE Computer Society [2041], pages 81–90. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Dempsey:1993:NEC

[1352] B. J. Dempsey, J. Liebeherr, and A. C.

Weaver. A new error control scheme for packetized voice over high-speed local area networks. In IEEE Computer Society [2041], pages 91–101. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Martini:1993:MWI

[1353] P. Martini and M. Ruemekasten.

MAN/WAN integration — the ATMto-DQDB case. In IEEE Computer Society [2041], pages 102–109. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

OShea:1993:RD

[1354] S. O’Shea and J. Finucane. Reactive DQDB. In IEEE Computer Society [2041], pages 110–119. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Jing:1993:HBS

[1355] W. Jing and M. Paterakis. How the Bernoulli slot occupancy approximation in the node DQDB analytical model influences the throughput-delay performance. In IEEE Computer Society [2041], pages 120–129. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Pach:1993:OTP

[1356] A. R. Pach, S. Palazzo, and D. Panno. Overload throughput performance of different erasure node protocols for DQDB networks. In IEEE Computer Society [2041], pages 130–139. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Nguyen:1993:RAF

[1357] G. D. Nguyen. Reliability analysis for FDDI dual homing networks. In IEEE Computer Society [2041], pages 140– 146. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Hiles:1993:ECT

[1358] W. S. Hiles and D. T. Marlow. Experimentation on the concentrator tree with loopback. In IEEE Computer Society [2041], pages 147–156. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Hamdaoui:1993:INC

[1359] M. Hamdaoui and P. Ramanathan. Improved non-real-time communication in FDDI networks with real-time traffic. In IEEE Computer Society [2041], pages 157–166. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Dobosiewicz:1993:NCC

[1360] W. Dobosiewicz, J. Ottensmeyer, and R. Wittenberg. New contention control mechanisms for FDDI. In IEEE Computer Society [2041], pages 167– 175. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Barilovits:1993:FPS

[1361] S. Barilovits and J. Kadambi. FDDI-2 packet services and latency adjustment buffers. In IEEE Computer Society [2041], pages 176–185. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Malcolm:1993:GSM

[1362] N. Malcolm and W. Zhao. Guaranteeing synchronous messages with arbitrary deadline constraints in an FDDI network. In IEEE Computer Society [2041], pages 186–195. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Keshtgary:1993:BAF

[1363] M. Keshtgary and A. P. Jayasumana.

Bandwidth allocation in FDDI-II for isochronous, synchronous, and asynchronous traffic. In IEEE Computer Society [2041], pages 196–204. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Yuang:1993:AAD

[1364] M. C. Yuang and Y. R. Haung. Assessment and alternative design of FDDI-II MAC protocol. In IEEE Computer Society [2041], pages 205–213. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Lamb:1993:ELS

[1365] J. P. Lamb and R. E. Papritz. Enterprise LAN standards and guidelines: A case study. In IEEE Computer Society [2041], pages 214–221. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Padakanti:1993:MHD

[1366] V. M. Padakanti and S. Kakumanu. A multi-level hierarchical distributed failsafe architecture for metropolitan area networks. In IEEE Computer Society [2041], pages 222–229. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Soriano:1993:ISS

[1367] M. Soriano, J. Forne, J. L. Melus, and F. Recacha. Implementation of a security system in a local area network environment. In IEEE Computer Society [2041], pages 230–237. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Ocheltree:1993:CFC

[1368] K. B. Ocheltree, T. C. Tsai, R. Montalvo, and A. Leff. A comparison of Fibre Channel and 802 MAC services. In IEEE Computer Society [2041], pages 238–246. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Varma:1993:UCI

[1369] A. Varma, S. Murthy, and R. Bryant.

Using camp-on to improve the performance of a fiber channel switch. In IEEE Computer Society [2041], pages 247–255. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Chang:1993:OLB

[1370] E. S. Chang. Optical link bandwidth optimization of the Fibre Channel networks. In IEEE Computer Society [2041], pages 256–261. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Tsai:1993:SDV

[1371] T.-C. Tsai, A. Leff, K. B. Ocheltree, and R. Montalvo. A simulation design for verification and development of Fibre Channel. In IEEE Computer Society [2041], pages 262–273. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Williams:1993:EEH

[1372] K. A. Williams and D. H. C. Du. Efficient embedding of a hypercube in an irregular WDM network. In IEEE Computer Society [2041], pages 274– 283. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Cho:1993:SAP

[1373] W. Cho and S.-B. Lee. Slotted Aloha protocols for high-speed optical fiber local area networks using WDM passive star topology. In IEEE Computer Society [2041], pages 284–290. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Yener:1993:CPI

[1374] B. Yener, Y. Ofek, and M. Yung. Configuration and performance issues in the MetaNet design. In IEEE Computer Society [2041], pages 291–299. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Gagnaire:1993:AFM

[1375] M. Gagnaire. Analysis of fairness on the metaring gigabit/s network. In IEEE Computer Society [2041], pages 300–309. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Takizawa:1993:GCP

[1376] M. Takizawa, M. Takamura, and A. Nakamura. Group communication protocol for large group. In IEEE Computer Society [2041], pages 310– 319. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Street:1993:VMD

[1377] F. Street and A. C. Weaver. A video mail distribution system for networked personal computers. In IEEE Computer Society [2041], pages 320–326. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Ahn:1993:LBM

[1378] S. Ahn and D. H. C. Du. A load balancing multicast tree approach for groupbased multimedia applications. In IEEE Computer Society [2041], pages 327–336. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Huang:1993:DIM

[1379] J.-H. Huang, C.-C. Yang, W.-H. Tseng, and C.-W. Lee. Design and implementation of multimedia conference system on broadcast networks. In IEEE Computer Society [2041], pages 337– 343. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Seshan:1993:IHP

[1380] S. Seshan and R. H. Katz. Interfacing a high performance disk array file server to a gigabit LAN. In IEEE Computer Society [2041], pages 344– 353. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Khalil:1993:DNA

[1381] K. M. Khalil. A dual network architecture for high-performance distributed computing environments. In IEEE Computer Society [2041], pages 354– 360. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Nahrstedt:1993:AAN

[1382] K. Nahrstedt and J. M. Smith. An application-driven approach to networked multimedia systems. In IEEE Computer Society [2041], pages 361– 368. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Schmidt:1993:LSF

[1383] D. C. Schmidt, B. Stiller, T. Suda, and A. N. Tantawy. Language support for flexible, application-tailored protocol configuration. In IEEE Computer Society [2041], pages 369–379. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Jong:1993:ADP

[1384] S. G. Jong and Y. O. Chin. Algorithms on dynamic priority scheduling for heterogeneous traffic in ATM. In IEEE Computer Society [2041], pages 380– 387. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Khasnabish:1993:IBC

[1385] B. Khasnabish. Integrated bandwidth and congestion management in ATM networks using a simple learning algorithm and a bucket bank. In IEEE Computer Society [2041], pages 388– 396. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Sato:1993:CTM

[1386] Y. Sato and K.-I. Sato. Comparison of traffic management schemes for ATM networks. In IEEE Computer Society [2041], pages 397–402. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Buddhikot:1993:SAG

[1387] M. M. Buddhikot, S. Kapoor, and G. M. Parulkar. Simulation of an ATM-FDDI gateway. In IEEE Computer Society [2041], pages 403–413. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Michel:1993:PEO

[1388] J. R. Michel, A. S. Waterman, and A. C. Weaver. Performance evaluation of an off-host communications architecture. In IEEE Computer Society [2041], pages 414–422. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Atwood:1993:ECX

[1389] J. W. Atwood and G. C. K. Chung. Error control in the Xpress Transfer Protocol. In IEEE Computer Society [2041], pages 423–431. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Harju:1993:ALL

[1390] J. Harju, P. Raatikainen, T. Reinikka, and J. Zidbeck. Analysis of a link layer protocol for high speed data bus. In IEEE Computer Society [2041], pages 432–439. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Ott:1993:BBR

[1391] J. M. Ott and A. P. Jayasumana. Bandwidth balancing of the Register Insertion Bus (RIB) fiber optic local area network. In IEEE Computer Society [2041], pages 440–451. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Perez-Jimenez:1993:EIL

[1392] R. Perez-Jimenez, J. Martin-Bernardo, Melian, V. M., and J. Ruiz-Alzola. Equalizer for an IR-wireless LAN using RBF neural networks. In IEEE Computer Society [2041], pages 461– 466. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Izumoto:1993:SWL

[1393] T. Izumoto, H. Shigeno, T. Yokoyama, and Y. Matsushita. A self-organizing wireless LAN system. In IEEE Computer Society [2041], pages 467–473. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Song:1993:RCP

[1394] J. Song and O. W. W. Yang. Rotation Counter Protocol — A token-ring protocol for integrated services. In IEEE Computer Society [2041], pages 474– 483. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Li:1993:NAO

[1395] M. Li and Z. Tsai. A novel approach to output process characterization for token ring networks supporting priorities. In IEEE Computer Society [2041], pages 484–492. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Doganata:1993:PSH

[1396] Y. N. Doganata and C. C. Bisdikian. Performance study of a high performance information/control frame multiplexer. In IEEE Computer Society [2041], pages 493–502. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Werahera:1993:TTC

[1397] P. Werahera, A. P. Jayasumana, and D. C. Boes. Token-cycle time characteristics of FDDI networks with multiple classes of traffic. In IEEE Computer Society [2041], pages 503–512. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Ng:1993:PHN

[1398] J. K. Y. Ng and J. W. S. Liu. Performance of high-speed networks for multimedia applications. In IEEE Computer Society [2041], pages 513–522. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Sathaye:1994:FDH

[1399] S. Sathaye, K. K. Ramakrishnan, and H. Yang. FIFO design for a high-speed network interface. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 2– 11. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Frank:1994:FCF

[1400] M. Frank and R. Wittenberg. Flow control functions for high-speed transport protocols in local workstation environments. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 12–19. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Mazyad:1994:PEH

[1401] A. S. Mazyad and C. Y. R. Chen.

Performance evaluation of HIPPI interconnection system using a campon strategy. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 20–29. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Leff:1994:ARP

[1402] A. Leff, K. B. Ocheltree, and T. Tsai. An approach for rapid protocol development: Experiences with the Proposed Fibre Channel Standard. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 30–39. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Cronin:1994:CHS

[1403] W. J. Cronin, J. D. Hutchison, K. Ramakrishnan, K., and H. Yang. A comparison of high speed LAN technologies. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 40– 51. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Wittenberg:1994:MFW

[1404] R. Wittenberg and P. Martini. Measurements in a FDDI workstation cluster. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 52–58. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Cavalieri:1994:EFC

[1405] S. Cavalieri, A. Di Stefano, and O. Mirabella. Exploiting FDDI communication features to connect Field-Buses in process control environment. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 59–66. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Albert:1994:IFL

[1406] B. Albert and A. P. Jayasumana. Interconnection of FDDI LANs. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 67–72. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Popovich:1994:SSM

[1407] S. Popovich and M. Alam. A simulation study of the migration from a CSMA/ CD network to FDDI. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 73–81. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Lin:1994:PPD

[1408] H.-J. Lin and G. Campbell. PDQRAP — Prioritized Distributed Queueing Random Access Protocol. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 82–91. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Kumar:1994:DPA

[1409] L. N. Kumar and C. Douligeris. The dynamic 3-Tier protocol: An access remedial scheme to control unfairness in DQDB MANs. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 92–101. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Huang:1994:ACR

[1410] E. Y. Huang. Analysis of cyclic reservation multiple access protocol. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 102–107. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Gagnaire:1994:CUM

[1411] M. Gagnaire and A. Thelen. Comparison of UPC mechanisms in a DQDB/ B-ISDN gateway. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 108–119. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Barilovits:1994:NOR

[1412] S. Barilovits and J. Kadambi. 100VG-AnyLAN: Network operation and realtime capabilities. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 120–128. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Raha:1994:PEA

[1413] A. Raha, N. Malcolm, and W. Zhao.

Performance evaluation of admission policies in ATM based embedded realtime systems. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 129–138. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Barghi:1994:RSH

[1414] H. Barghi, W. T. Cronenwett, and J. G. Bredeson. Real-time study of a high performance protocol for local/ metropolitan ring networks. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 139–145. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Shin:1994:LAF

[1415] K. G. Shin, L.-H. Ng, and T. P. Monaghan. A layered approach to fault-tolerance and timeline issues in SAFENET. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 146–155. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Zheng:1994:RCA

[1416] Q. Zheng, K. G. Shin, and C. Shen. Real-time communication in ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 156– 165. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Khalil:1994:PCT

[1417] K. M. Khalil and Y. S. Sun. Performance considerations for TCP/IP in wide area networks. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 166–175. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Moldeklev:1994:TIB

[1418] K. Moldeklev, E. Klovning, and O. Kure. TCP/IP behavior in a highspeed local ATM network environment. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 176–185. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Jarman:1994:USI

[1419] D. Jarman and R. Y. Lee. A user solution to Internetworking. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 186–192. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Salwen:1994:WCR

[1420] H. Salwen. WAN characterization for the rest of the decade. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 193–195. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Barnett:1994:TCR

[1421] L. Barnett. A topology-aware collision resolution algorithm. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 196–205. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Whetten:1994:PSE

[1422] B. Whetten, S. Steinberg, and D. Ferrari. The packet starvation effect in CSMA/CD LANs and a solution. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 206–217. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Hughes:1994:IPR

[1423] L. Hughes and M. Thomson. Implosionavoidance protocols for reliable group communications. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 218–227. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Ramakrishnan:1994:ECE

[1424] K. K. Ramakrishnan and H. Yang. The Ethernet capture effect: Analysis and solution. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 228–241. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Christensen:1994:ETF

[1425] K. Christensen, F. Noel, and N. Strole.

Emerging trends — full-duplex and the switched LAN. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 242–248. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Lin:1994:ILL

[1426] T.-P. Lin, Y.-T. Wang, and N.-F.

Huang. Interconnection of largescale LANs via a two-stage switching HUB for multimedia applications. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 249–256. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Hong:1994:MCN

[1427] C. Hong, T. Yoneda, K. Tanaka, and S. Honda. Multimedia communication networking architecture model for high speed network. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 257–265. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Ma:1994:PCS

[1428] Q. Ma and P. Steenkiste. Performance of circuit switched LANs under different traffic conditions. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 266–277. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Rooholamini:1994:MAC

[1429] R. Rooholamini and V. Cherkassky. Moving ATM closer to multimedia applications. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 278–287. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Chang:1994:OPC

[1430] S.-P. Chang, P. Chang, and P. Landsberg. An overview of the Pipelined Common Buffer Architecture (PCBA) for memory based packet/cell switching systems. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 288–297. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Battou:1994:ATP

[1431] A. Battou and G. Nguyen. Analysis of a threshold priority queueing system with applications to ATM. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 298–306. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Moors:1994:ITS

[1432] T. Moors, N. Clarke, and G. Mercankosk. Implementing traffic shaping. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 307–315. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Patel:1994:PCE

[1433] B. V. Patel and C. C. Bisdikian. Performance characteristics of end-stations in an ATM network as viewed by applications and network. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 316–323. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Brustoloni:1994:EBS

[1434] J. C. Brustoloni. Exposed buffering and sub-datagram flow control for ATM LANs. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 324–344. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Camarda:1994:BMA

[1435] P. Camarda and V. A. Signore. Bandwidth management for ATM-Based LANs. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 345– 355. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Willebeek-LeMair:1994:MCU

[1436] M. H. Willebeek-LeMair, D. D. Kandlur, and Z.-Y. Shae. On multipoint control units for videoconferencing. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 356–364. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Harju:1994:QPD

[1437] J. Harju, V.-P. Kosonen, and C. Li. Quality and performance of a desktop video conferencing system in the network of interconnected LANs. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 365–371. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Elsaadany:1994:PCS

[1438] A. Elsaadany, M. Singhal, and M. T.

Liu. Priority communication schemes on local area networks for multimedia traffic. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 372–379. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Nussbaumer:1994:MDD

[1439] J.-P. Nussbaumer, B. V. Patel, and F. Schaffa. Multimedia delivery on demand: Capacity analysis and implications. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 380–386. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Amer:1994:PAM

[1440] K. Amer. Performance analysis of multiple priorities on token ring LANs to support multimedia. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 387–403. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Amley:1994:NMI

[1441] C. Amley. Network management’s impact on managed networks. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 404–410. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Heindel:1994:TSM

[1442] L. E. Heindel and K. J. Schlieber. Total systems management in open architecture provisioning systems. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 411–416. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Jeng:1994:NMP

[1443] S.-C. Jeng, T.-W. Chin, J.-L. Lin, and S.-S. Hung. Network management platform approach and its application — KView II. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 417–424. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Jha:1994:CPL

[1444] S. K. Jha and B. R. Howarth. Capacity planning of LAN using network management. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 425–431. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Hurley:1994:LEF

[1445] R. T. Hurley, J. P. Black, and J. W.

Wong. Limited effects of finite storage on a beneficial file migration policy. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 432–439. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Lamb:1994:LOE

[1446] J. Lamb and T. Cusato. LAN-based office for the enterprise, A case study. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 440–447. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Yuang:1994:BHT

[1447] M. C. Yuang, J. C. Liu, and C. L. Shay.

BATS: A high-performance transport system for broadband applications. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 448–455. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Dempsey:1994:ESP

[1448] B. J. Dempsey, M. T. Lucas, and A. C. Weaver. An empirical study of packet voice distribution over a campus-wide network. In IEEE Computer Society. Technical Committee on Computer Communications [2043], pages 456–

465. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-

02. IEEE catalog number 94TH8004.

Eichmann:1994:EWA

[1449] Dave Eichmann. Ethical Web agents. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

                                        314X. URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/Agents/
                                        eichmann.ethical/eichmann.html.

Graber:1994:WSA

[1450] Jeffrey Graber. Web search agents for effective data distillation. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Spetka:1994:TRB

[1451] Scott Spetka. The TkWWW robot: Beyond browsing. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-

                                        314X. URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/Agents/
                                        spetka/spetka.html.

Whitehead:1994:AFE

[1452] Steven D. Whitehead. Auto-FAQ:

an experiment in cyberspace leveraging. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/Agents/
whitehead/whitehead.html.

Gutfreund:1994:WOS

[1453] Yechezkal-Shimon Gutfreund, John Nicol, Russel Sasnett, and Vincent Phuah. WWWinda: An orchestration service for WWW browsers and accessories. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/Agents/
gutfreund/gutfreund.html.

Oostendorp:1994:TIW

[1454] W. F. Punch K. A. Oostendorp and R. W. Wiggins. A tool for individualizing the Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/Agents/
oostendorp/oostendorp.html.

Ford:1994:EBE

[1455] Andrew Ford. The electronic Beowulf:

From early Anglo-Saxon text to hypertext. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Arts/ford/
beowulf-paper.html.

McLaughlin:1994:AGE

[1456] Margaret L. McLaughlin. Art galleries and exhibits on the World Wide Web.

In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Arts/mclaughlin.
art/mclaughlin_Art.html.

Harris:1994:PCH

[1457] Marty Harris. Presentation: Cultural heritage and cyberspace. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Frett:1994:MPP

[1458] John J. Frett, Elizabeth O. Mackenzie, and Dewey Caron. A multimedia presentation of plants in the University of Delaware Botanic Gardens utilizing Mosaic. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Arts/mackenzie/
mackenzie.html.

Petruzzelli:1994:CTI

[1459] Jim Petruzzelli. Critical thinking with the interactive text: Decentering textual authority with Mosaic. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Arts/petruzzelli/
petruzzelli.html.

Greenberg:1994:OPM

[1460] Hope A. Greenberg. The Ovid Project: Metamorphosizing the metamorphoses. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Arts/greenberg/
greenberg.html.

Brenner:1994:FRI

[1461] Paul Brenner and Maria Roussos. The file room: An interactive computer project addressing cultural censorship. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Arts/brenner/
roussos.html.

Godwin-Jones:1994:LLW

[1462] Robert Godwin-Jones. Language learning and the World Wide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Arts/godwin-
jones/godwin.html.

Fischer:1994:GWC

[1463] Elizabeth Fischer. Graffiti on the Web:

A cultural interchange A lighthearted romp with an artist in Webland that stops being lighthearted at the end. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Arts/fischer/
fischer.html.

Humm:1994:CBA

[1464] Kathryn M. Humm. Chesley Bonestell Art Gallery. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Arts/humm/
bonestell.html.

Levine:1994:DCM

[1465] David A. Levine and Mark D. Pesce.

From database to cyberspace, Mosaic as middle ground. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Gollifer:1994:IFI

[1466] Sue Gollifer. ‘imaging the future’: Issues of identity and authenticity of a work of art. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Arts/gollifer/
gollifer.html.

Jennings:1994:HPL

[1467] Donald Jennings, Peter Damon, Maia Good, and Ryszard Pisarski. How to present lots of volatile information on the World Wide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Astronomy/
jennings/jennings.html.

Egret:1994:WAR

[1468] Daniel Egret and Andre’ Heck. The Web in astronomy and related space sciences. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Astronomy/
egret/egret.html.

Strom:1994:EPD

[1469] Karen M. Strom. Electronic publication and data distribution for the star formation group in the five college Astronomy Department. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Astronomy/
strom/strom.html.

Richmond:1994:IAI

[1470] Alan Richmond and Nick White. Implementation of an Astrophysics Information System on the World-Wide Web. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Astronomy/
richmond/richmond.html.

Jackson:1994:AT

[1471] Robert E. Jackson. AstroWeb tools.

In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Astronomy/
jackson/jackson.html.

Jordan:1994:TIP

[1472] James M. Jordan, Mark Cresitello-Dittmar, and Jesse Allen. Tailoring information presentation with words and pictures at the COSSC. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Astronomy/
jordan/jordan.html.

Antia:1994:EAW

[1473] Behram Antia, Brett Stroozas, William Boyd, Carol Christian, and Kuei Chen. EUV astrophysics and the World Wide Web. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Astronomy/
antia/antia.html.

Towheed:1994:NUW

[1474] Syed S. Towheed. NASA’s use of the World Wide Web to deliver Shoemaker-Levy 9 collision data in near-real time. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Astronomy/
                                        towheed/towheed.html.

Slavney:1994:PSD

[1475] Susan Slavney and Thomas C. Stein. Planetary science data distribution on the Web: The Magellan standard products catalog. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Astronomy/
                                        slavney/slavneystein.html.

Quin:1994:DCH

[1476] Liam Quin. Design considerations for an HTML editor. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Sullivan:1994:HME

[1477] Brian Sullivan, Jonathan Ryan Day, Jeff Spitulnik, and Elliot Soloway. HTML made easy: HTML Claris XTND translator. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Autools/sullivan/
                                        HTMLMadeEasy.html.

Punin:1994:SHE

[1478] John R. Punin and Mukkai S. Krishnamoorthy. A simple HTML editor. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Autools/krishnamoorthy/
user.html.

Jones:1994:TLA

[1479] Kennie Jones. TOPS on-line — automating the construction and maintenance of HTML pages. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Autools/jones/
kjones.html.

Kasper:1994:UWR

[1480] Theresa Kasper. Untangling the Web — the role of text retrieval in a hypertext environment. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Autools/kasper/
kasper.html.

Kent:1994:CWA

[1481] Robert E. Kent and Christian Neuss.

Creating a Web analysis and visualization environment. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Autools/kent/
kent.html.

Suttor:1994:SHB

[1482] Jeff Suttor. SGML, HTML and beyond. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Freese:1994:TSD

[1483] Eric Freese. The transformation of SGML documents for presentation on the WWW. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Autools/freese/
freese.html.

Sperberg-McQueen:1994:HMM

[1484] C. M. Sperberg-McQueen and Robert F. Goldstein. HTML to the max: A manifesto for adding SGML intelligence to the World-Wide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Autools/sperberg-
mcqueen/sperberg.html.

Myka:1994:PPD

[1485] Andreas Myka. Putting paper documents in the World-Wide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/myka/
myka.html.

Rung:1994:EIM

[1486] Steven F. Rung. Electronic imaging of a Macintosh page. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Autools/rung/
rung.html.

Pasewark:1994:SFH

[1487] Richard Pasewark. Standardized formatting for HTML data for print output. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Thralls:1994:BHA

[1488] Robert Keith Thralls, Sr. Building HTML application systems: Converting existing MS-Windows applications to HTML. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WebProd/thralls/
WhitePaper.html.

Plewe:1994:GPU

[1489] Brandon S. Plewe. The GeoWeb Project: Using WAIS and the World Wide Web to aid location of distributed data sets. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Autools/plewe/
plewe.html.

Baggott:1994:N

[1490] James Baggott and Sharon E. Dennis. NetBiochem. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/BioChem/baggott/
baggott.html.

Rzepa:1994:CWH

[1491] Henry S. Rzepa and Christopher Leach. Chemistry in the Web:

Hyperactive molecules. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/BioChem/rzepa/
rzepa.html.

Epstein:1994:WEH

[1492] Jonathan A. Epstein, Jonathan A. Kans, and Gregory D. Schuler. WWW Entrez: A hypertext retrieval tool for molecular biology. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/BioChem/epstein/
WWW_entrez.html.

FitzGerald:1994:WCE

[1493] Peter C. FitzGerald and Robert A. Pearlstein. The Web as a computational engine for chemistry and molecular biology. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/BioChem/pearlstein/
fitzgerald.html.

Robertson:1994:VFD

[1494] David W. Robertson, William E. Johnston, and Wing Nip. Virtual frog dissection: Interactive 3D graphics via the Web. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/BioChem/robertson/
robertson.html.

McLay:1994:BFB

[1495] Michael McLay. Birds of a feather on browser ancillary tools. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Bosak:1994:ICH

[1496] Jon Bosak and Terry Allen. Issues and concerns with HTML for real-world publishing. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Behlendorf:1994:PW

[1497] Brian Behlendorf. Pesce and Wired.

In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

DeSai:1994:NI

[1498] ?. DeSai. Navigation issues. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Ng:1994:GDG

[1499] Jason Ng. GSQL: A database gateway. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Zurko:1994:SW

[1500] Mary Ellen Zurko. Security on the Web. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Rees:1994:SCT

[1501] Michael Rees and Tak Woo. A synchronous collaboration tool for WWW. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CSCW/rees/
SynColTol.html.

Scharf:1994:UMR

[1502] Ronald Scharf, Stefan Hartmann, and Werner Wolz. Using Mosaic for remote test system control supports distributed engineering. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CSCW/scharf/
scharf.html.

Goldberg:1994:BWE

[1503] Ken Goldberg, Michael Mascha, Steven Gentner, Nick Rothenberg, Carl Sutter, and Jeff Wiegley. Beyond the Web: Excavating the real world via Mosaic. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CSCW/goldberg/
goldberg.html.

Weymouth:1994:UAR

[1504] Terry E. Weymouth, Dan Atkins, Gary Olson, Bob Clauer, Atul Prakash, Thomas Finholt, and Craig E. Rasmussen. The upper atmospheric research collaboratory: UARC. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CSCW/weymouth/
Weymouth.html.

Frivold:1994:EWS

[1505] Thane J. Frivold, Ruth E. Lang, and Martin W. Fong. Extending WWW for synchronous collaboration. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CSCW/frivold/
frivold.html.

Donath:1994:SW

[1506] Judith Donath and Niel Robertson. The sociable Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CSCW/donath/
SociableWeb.html.

Hurwitz:1994:WAP

[1507] Roger Hurwitz, Larry Koskinen, and John C. Mallery. Wide-area policy discussion: The town meeting on the Vice President’s National Performance Review. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

McLaughlin:1994:JCM

[1508] Margaret L. McLaughlin. The Journal of Computer-Mediated Communication. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CSCW/mclaughlin.
journal/mclaughlin_Jour.html.

Halama:1994:IEB

[1509] James R. Halama, Anne Kreft, and Robert E. Henkin. An interactive electronic bulletin board implementation for Mosaic and HTTP server. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CSCW/halama/
halama.html.

Frega:1994:MBB

[1510] N. Frega and A. Volpentesta. A multimedia bulletin board in WWW environment. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CSCW/volpentesta/
giuda.html.

Mallery:1994:WAC

[1511] Tim Berners-Lee John Mallery. Widearea collaboration and cooperative computing. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN

0309-314X. URL http://www.ncsa.
uiuc.edu/SDG/IT94/Proceedings/WWW2_
Proceedings.html.

Quinn:1994:GRC

[1512] Christine A. Quinn. From grass roots to corporate image — the maturation of the Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/Campus.
Infosys/quinn/quinn.html.

Sherwood:1994:TSA

[1513] Kaitlin Duck Sherwood. Technical and sociological aspects of developing campus-wide Webs: UIUC College of Engineering. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/Campus.
Infosys/sherwood/sherwood.html.

Lechner:1994:CWI

[1514] Pieter Lechner and Kent Wada.

Campus-wide information systems, based on WWW and beyond. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Roden:1994:MCW

[1515] Peter Roden and Suzana Lisanti.

MIT’s campus-wide information system: Scaling up for World Wide Web support. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/Campus.
Infosys/roden/roden.html.

Lavenant:1994:PMW

[1516] M. G. Lavenant and John A. Kruper.

Phoenicia: A model for World-Wide Web based campus information systems. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/Campus.
Infosys/kruper/lavenant.html.

Vekar:1994:CVP

[1517] John Vekar and Saul Schwartz. The Canadian virtual policy and administration community. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/Campus.
Infosys/vekar/vekar.html.

Taricani:1994:CAS

[1518] Ellen Taricani. College of Agricultural Sciences WWW server at Penn State. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Welz:1994:MBW

[1519] Gary Welz. The media business on the WWW. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

                                        314X. URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/ComEc/
                                        welz/welzwwwconf.html.

Kling:1994:ECW

[1520] Arnold Kling. The economic consequences of the World Wide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-

                                        314X. URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/ComEc/
                                        kling/webeconc.html.

Houh:1994:WCS

[1521] Henry Houh. Where are commercial services on the Web headed? In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Dratva:1994:WBH

[1522] Richard Dratva. WWW-based home banking services in Switzerland: A case study. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

                                        314X. URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/ComEc/
                                        dratva/dratva.html.

MacKie-Mason:1994:SFA

[1523] Jeffrey K. MacKie-Mason and Hal R.

Varian. Some FAQs about usage-based pricing. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/ComEc/
varian/useFAQs.html.

Newman:1994:CTW

[1524] Clifford Newman. Commercial transactions on the Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Neches:1994:FWY

[1525] Anna-Lena Neches, Paul Postel, and Craig Milo Rogers. Finding what you want to buy using the Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Ford:1994:OIL

[1526] Shelley G. Ford and Robert C.

Stern. OmniPort: Integrating legacy data into the WWW. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CorInfSys/
stern/stern-ford.html.

Pingleton:1994:UMW

[1527] Mike Pingleton and Tom Fischer.

Utilizing Mosaic and the WWW in an operations environment. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CorInfSys/
fischer/fischer.html.

Mayer-Kress:1994:HMI

[1528] Gottfried Mayer-Kress, William Bender, and John Bazik. Hyper-media on the Internet as a tool for approaching global problems: A teleconferencing experiment. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CorInfSys/
gmk/Mayer-Kress/HungerConf.html.

Smithson:1994:ICS

[1529] Brian Smithson and Barbara Singer. An information clearinghouse server for industry consortia. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CorInfSys/
smithson/smithson.html.

Wherry:1994:URI

[1530] Phillip S. Wherry, Mark A. Krause, and Richard K. Sciambi. Understanding return on investment for information discovery and retrieval technology. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Morey:1994:CCI

[1531] Kathy Morey. Corporate communication integration through Mosaic, WAIS, and Gopher. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Prah:1994:MCD

[1532] Chris Prah and Diane DiGiovanni. Mosaic as corporate data collector and dispenser. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CorInfSys/
prah/prah.html.

Gullapalli:1994:CMR

[1533] Sridhar Gullapalli and Peter Will. Coordinating MEMS research via the Web. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Brien:1994:UMW

[1534] John Brien, Peggy Havukainen, Matthew Krenmayr, Antonio Artese, and Mark Silva. Using Mosaic/ WWW in a business school: New opportunities for information sharing and teaching. issues in implementing organizational-wide networked multimedia. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CorInfSys/
brien/Chicago.html.

Grady:1994:WFA

[1535] Nancy Grady, Jim Kohl, Phil Papadopoulos, Forrest Hoffman, Mike Leuze, and Kimberly Barnes. WWW financial automated management environment at ORNL (Oak Ridge National Laboratory). In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Woo:1994:LCC

[1536] Sherman Woo, Ryan Martens, and Mike Gull. The loss of control in the creation of community: Leveraging the Web for business re-engineering. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Neches:1994:OPI

[1537] Anna-Lena Neches, Cesar Pina, Paul Postel, and Craig Milo Rogers. Online procurement at ISI: Transitioning to the Web. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Garver:1994:NSF

[1538] David Garver. The National Science Foundation (NSF) FastLane Project. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

McAlister:1994:T

[1539] Matt McAlister. A+ texture. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Kelly:1994:MWB

[1540] M. Kelly and M. Wilens. Mercury:

a WWW-based corporate information. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Bednarcyk:1994:LWI

[1541] Lauren A. Bednarcyk and Kevin D.

Bond. A local Web for information delivery. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CorInfSys/
bednarcyk/bednarcyk.html.

Filiatrault:1994:AWW

[1542] David W. Filiatrault, Craig A. Lightle, James E. Simmons, and Jerrold M. Grochow. AMS on the World-Wide-Web. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CorInfSys/
grochow/grochow.html.

Huynh:1994:CCM

[1543] Minh Huynh, Laird Popkin, and Matthew Stecker. Constructing a corporate memory infrastructure from Internet discovery technologies. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CorInfSys/
huynh/cmi.html.

Ivler:1994:DHA

[1544] J. M. Ivler. Developing in-house applications using Mosaic and HTTPD or internal WWW support of strike fighter development. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.

html. Original paper withdrawn at author’s request.

Petersen:1994:MSG

[1545] Neil A. Petersen. Mosaic streamlines government computer contract. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Holland:1994:AWW

[1546] William H. Holland. AT&T and the World Wide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CorInfSys/
holland/mospap.html.

Berners-Lee:1994:HSY

[1547] Tim Berners-Lee. HTML specifications: (Yuri’s workshop on HTML and SGML). In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Brown:1994:H

[1548] Jeremy Brown, Jake Harris, Lara Karbiner, Massimiliano Poletto, Andre’ DeHon, Ian Eslick, and Thomas F. Knight, Jr. HyperCode. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/brown/
hypercode/hypercode.html.

Hastings:1994:PCI

[1549] Edwin Hastings and Dilip H. Kumar. Providing customers information using the WEB and CORBA. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/hastings/
hastings.html.

Seidman:1994:HES

[1550] James L. Seidman. An HTML extension to support client-side image maps. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/seidman/
seidman.html.

Singh:1994:AM

[1551] Guy Singh and Roger Binns. An API to Mosaic. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/singh/
ixiwww94.html.

Slater:1994:EWC

[1552] Alan Falconer Slater. Extending W3 clients. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/slater/
ew3c.html.

Thrift:1994:PIE

[1553] Philip Thrift. Preprocessing instructions: Embedding external notations in HTML. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/thrift/
PPI.html.

Crocker:1994:WSI

[1554] Glenn A. Crocker. web2mush: Serving interactive resources to the Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/crocker/
tech.html.

Donszelmann:1994:GWW

[1555] M. Donszelmann and K. Rodden. Gateways for World-Wide Web in the ‘online’ data acquisition system of the DELPHI experiment at CERN. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/donszelmann/
paper.html.

Frystyk:1994:TUL

[1556] Henrik Frystyk. Towards a uniform library of common code: A presentation of the World Wide Web library. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/frystyk/
LibraryPaper.html.

Gessler:1994:PMW

[1557] Stefan Gessler and Andreas Kotulla. PDAs as mobile WWW browsers. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/gessler/
www_pda.html.

Meschkat:1994:UID

[1558] Steffen Meschkat. User interface documents. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Mogul:1994:IWL

[1559] Jeffrey C. Mogul and Venkata N. Padmanabhan. Improving WWW latency. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/mogul/
HTTPLatency.html.

Pinckney:1994:DDE

[1560] Tom Pinckney, M. Frans Kaashoek, and Joshua A. Tauber. Dynamic documents: Extensibility and adaptability in the WWW. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/pinckney/
dd.html.

Chhabra:1994:ICD

[1561] Bhavna Chhabra, Darren R. Hardy, Allan Hundhausen, David Merkel, John D. Noble, and Michael F. Schwartz. Integrating complex data access methods into the Mosaic/ WWW environment. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/schwartz/
schwartz.html.

Soo:1994:LMH

[1562] Jonathan Soo. Live multimedia over HTTP. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/DDay/soo/www94a.
                                        html.

Weber:1994:WSK

[1563] Jay Weber. The Webmaster’s starter kit. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/DDay/weber/
                                        WeberPaper.html.

Uhler:1994:IRT

[1564] Stephen Uhler. Incorporating realtime audio on the Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/DDay/uhler/
                                        uhler.html.

Varela:1994:PDW

[1565] Carlos A. Varela and Caroline C. Hayes. Providing data on the Web: From examples to programs. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/DDay/varela/
                                        paper.html.

Shigechika:1994:DIM

[1566] Noriyuki Shigechika, Jun Murai, Shin Yoshimura, and Shigeya Suzuki. The design and implementation of multilingual extension for WWW. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Mallery:1994:DFP

[1567] John C. Mallery and Benjamin Renaud. Dynamic form processing: A general framework for interactivity. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Claffy:1994:WTC

[1568] Kimberly C. Claffy and Hans-Werner Braun. Web traffic characterization: an assessment of the impact of caching documents from NCSA’s Web server. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/claffy/
main.html.

Sedayao:1994:MWK

[1569] Jeff Sedayao. “Mosaic will kill my network!” studying network traffic patterns of Mosaic use. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/sedayao/
mos_traf_paper.html.

Thompson:1994:NMC

[1570] Dave Thompson. NCSA Mosaic CCI.

In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Pitkow:1994:SYR

[1571] Jim Pitkow and Mimi Recker. A simple yet robust caching algorithm based on dynamic access patterns. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/DDay/pitkow/
                                        caching.html.

Boutell:1994:TSS

[1572] Thomas Boutell. Techniques for serverside dynamic document generation. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/DDay/boutell/
                                        servergen.html.

Fullton:1994:RDU

[1573] Jim Fullton and Kevin Gamiel. Recent developments in URI implementation. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

VanHeyningen:1994:UCS

[1574] Marc D. VanHeyningen. The Unified Computer Science Technical Report Index: Lessons in indexing diverse resources. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/DDay/vanheyningen/
                                        paper.html.

Clucas:1994:IVC

[1575] Jean Clucas and Velvin (Val) Watson. Interactive visualization of computational fluid dynamics using Mosaic. In Anonymous [2042], page ??

CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/clucas/
clucas.html.

Moline:1994:ISV

[1576] Jodi Moline. Inline support for vector files: Engineering drawings on the WWW. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/moline/
moline.html.

Spero:1994:WSP

[1577] Simon E. Spero. WWW-SPEED: A 4 person panel discussing issues covered by the charter of the www-speed mailing list. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Spero:1994:HB

[1578] Simon E. Spero. HTTP benchmarks.

In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Spero:1994:HPP

[1579] Simon E. Spero. HTTP performance problems. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Spero:1994:MMD

[1580] Simon E. Spero. MDMA — multithreaded daemon for multimedia access. In Anonymous [2042], page ??

CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Spero:1994:BBG

[1581] Simon E. Spero. BGI — Binary Gateway Interface. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

DeHon:1994:GCC

[1582] Andre’ DeHon, Jeremy Brown, Ian Eslick, Jake Harris, Lara Karbiner, and Thomas F. Knight, Jr. Global cooperative computing. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/dehon.
global/gcc_www94.html.

Radev:1994:RWS

[1583] Dragomir R. Radev. RENDEZVOUS: A WWW synchronization system. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/radev/
radev.html.

Spasojevic:1994:UWA

[1584] Mirjana Spasojevic, C. Mic Bowman, and Alfred Spector. Using a widearea file system within the World Wide Web. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/spasojevic/
paper.html.

Vanzyl:1994:VC

[1585] ?. Vanzyl. Value of compression. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Patrashin:1994:MFM

[1586] Michael Patrashin. Managing the fonts in multilingual WWW. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Viles:1994:SWS

[1587] Charles L. Viles and James C.

French. A summary of WWW server ability and latency measurements: March to May 1994. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

McLay:1994:EMS

[1588] Michael McLay. Extending Mosaic to support a research and engineering framework. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Fischer:1994:UEA

[1589] Michael Meyer Michael Fischer and Michael Witbrock. User extensibility in Amiga Mosaic. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/DDay/witbrock/
                                        amosaic/.

Boisvert:1994:WGV

[1590] Ronald F. Boisvert. A Web gateway to a virtual mathematical software repository. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/DDay/boisvert/
                                        boisvert.html.

Fabozzi:1994:UMD

[1591] Donald J. Fabozzi II. The use of Mosaic as a documentation tool for large, graphically-based simulation software. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/DDay/fabozzi/
                                        fabozzi.paper.html.

Werkman:1994:SRL

[1592] Keith Werkman and Jeff Poulin. Software reuse libraries with Mosaic. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/DDay/werkman/
                                        www94.html.

Bean:1994:BWS

[1593] Greg Bean. Building Web servers for microcomputer foundations. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/DDay/bean/
                                        bean.html.

Collins:1994:CRW

[1594] Julia A. Collins. Climate research and the Web: Integrating World Wide Web and Mosaic capabilities into the Climate Diagnostics Center. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/EarSci/
collins/collins.html.

Boston:1994:ISD

[1595] Tony Boston and David Stockwell.

Interactive species distribution reporting, mapping, and modelling using the World Wide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/EarSci/
boston/species_paper.html.

Meyer:1994:MHE

[1596] Theodore Meyer, Ramachandran Suresh, Douglas Ilg, and Bruce Moxon. Mosaic, HDF and EOS-DIS: Providing access to earth sciences data. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/EarSci/
meyer/ilg.html.

Nebert:1994:SDM

[1597] Douglas D. Nebert. Serving digital map information through the World Wide Web and wide-area information server technology. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Soreide:1994:MAR

[1598] N. N. Soreide, L. McCarty, and D. C. McClurg. Mosaic access to realtime data from the TOGA-TAO array of moored buoys. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-

                                        314X. URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/EarSci/
                                        soreide/soreide.html.

Baker:1994:UIH

[1599] C. Bruce Baker and Danny E. Brinegar. The use of the information highway to explore climate variability. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

                                        314X. URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/EarSci/
                                        baker/Baker94.html.

Hankin:1994:LVE

[1600] Steve Hankin and Jerry Davison. Live visualization and extraction of climate data with Mosaic and FER-RET. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

                                        314X. URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/EarSci/
                                        hankin/hankin.html.

Zimmerman:1994:ICW

[1601] Frederick Zimmerman and Peter Young. Information for a changing world: CIESIN’s use of WWW to facilitate access to data and information relevant to human interactions in the environment. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-

                                        314X. URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/EarSci/
                                        zimmerman/zimmerman.html.

Perron:1994:LWC

[1602] Daniel Perron. Learning on the WWW: A case study. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/perron/
perron.html.

Hong:1994:UWT

[1603] Jack Hong, George Toye, and Larry Lefier. Using the WWW for a teambased engineering design class. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/hong/
hong.html.

Blumberg:1994:MES

[1604] Roger B. Blumberg. MendelWeb:

An electronic science/math/history resource for the WWW. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/blumberg.
mendelweb/MendelWeb94.blumberg.
html.

Smeaton:1994:ISW

[1605] Calum Smeaton and Alan Slater. Integrating simulations and W3 courseware. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/smeaton/
ise_www/ise_www94.html.

Panoff:1994:BRB

[1606] R. M. Panoff, M. South, and B. Davenport. Building a research-based interactive discovery gallery for education: The NCSA Bridge Project. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Hilts:1994:BAW

[1607] Kathryn H. Hilts. Bell Atlantic– West Virginia world school. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Thurber:1994:DSK

[1608] Bonnie Thurber and Bob Davis. Designing a server with a K–8 school in mind. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Educ/thurber/
                                        thurber.html.

Wray:1994:OIM

[1609] Robert E. Wray, III, Ronald Chong, Joseph Phillips, Seth Rogers, William Walsh, and John Laird. Organizing information in Mosaic: A classroom experiment. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Educ/wray/
                                        wrayre-experiment-mosaic94.html.

Keeler:1994:ESS

[1610] Michael J. Keeler. The Earth System Science Community Curriculum Testbed. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Educ/keeler/
                                        ESSCCTestbed/ESSCCTOC.html.

Ridge:1994:UIS

[1611] Pat Ridge. Uniting international students with WWW and NCSA Mosaic. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Clement:1994:WST

[1612] John Clement and Janice Abrahams. A Web of students, teachers, projects, and resources: Current K–12 use in the United States. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/rebelsky/
rebelsky.html.

Gomez:1994:SCK

[1613] Louis Gomez. A sampling of collaborative K–12 projects on the Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Smith:1994:UIR

[1614] Jane Smith. Useful information resources for education. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Rebelsky:1994:WRI

[1615] Samuel A. Rebelsky. A Web of resources for introductory computer science. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Fox:1994:IET

[1616] E. Fox and N. Dwight Barnette. Improving education through a CS digital library with three types of WWW servers. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/barnette/
barnette.html.

Hurley:1994:CPC

[1617] S. Hurley, A. D. Marshall, S. N. McIntosh-Smith, and N. M. Stephens. Courseware for parallel computing using Mosaic and the World Wide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/hurley/
hurley.html.

Schmitz:1994:WES

[1618] John Schmitz. World explorer system. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Stagg:1994:VCD

[1619] Josef Stagg. The virtual community: Distance education and community master planning. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Dybvik:1994:CWM

[1620] Per Einar Dybvik and Hakon W.

Lie. Combining WWW/mosaic with realtime multimedia conferencing in distance education. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/lie/lie.
html.

Rice:1994:UMU

[1621] Judy Rice, Rosina Bignall, Dalinda Bond, and Phillip J. Windley. Uses of Mosaic in a university setting. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/rice.
university/article.html.

Kummerfeld:1994:ICC

[1622] Robert J. Kummerfeld and J. Kay.

An individualised course for the C programming language. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/kummerfeld/
kummerfeld.html.

Tewari:1994:ICD

[1623] Rajiv Tewari. Innovative courseware development with Mosaic and WWW. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/tewari/
tewari.html.

Ward:1994:ESI

[1624] Pat Ward and Kristopher Davis. Empowering students in the information age. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/ward/
ward.html.

Schmitz:1994:DS

[1625] John Schmitz, Marsha Woodbury, Mary Connors, Aaron Buckley, and Chip Aubry. Discovery system. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/schmitz.
discover/conference.html.

Nauta:1994:SDR

[1626] James M. Nauta. The School Development Resource System (SDRS) critical issues server. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Globus:1994:SUW

[1627] Al Globus and Chris Beaumont. Spinning a useful Weblet. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/globus/
RNR-94-017.html.

Begoc:1994:MTA

[1628] Nathalie Begoc, Bernard Bouchare, Alain Guittot, Yannick Lecomte, Jean-Francois Leprovost, and Jean Seguin. Mosaic and Tele-Amphi, hypertext hypermedia services with interactive work tools for interactive multimedia distance teaching. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/begoc/
awaiting.html.dueOct1.

Mueller:1994:UWI

[1629] Bernd Mueller. Using WWW as an information system to reduce the average period of study by better information providing and to relieve administration. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/mueller/
Paper.html.

Lieberman:1994:WWB

[1630] Jonathan A. Lieberman and Nadine O’Connor DiVito. WILT: A WWW based Interactive Language Teaching Tool. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/lieberman/
lieberman.html.

Rice:1994:HH

[1631] Judy Cossel Rice. Hypermedia for HyperKids. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/rice.
hyperkids/Hyper.html.

Spencer:1994:CIA

[1632] Stephan M. Spencer, Max L. Nibert, and Jean-Yves Sgro. Communicating information about virus structure and biology via the World Wide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Educ/spencer/
                                        spencer.html.

Sjolin:1994:WFE

[1633] Martin Sjolin. A WWW front end to an OODBMS. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Databases/
                                        sjolin/sjolin.html.

Maly:1994:WWA

[1634] Kurt Maly, J. French, A. Selman, and E. Fox. WATERS — the Wide Area Technical Report Server. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Databases/
                                        maly/maly.html.

Cohen:1994:RBI

[1635] Ellis S. Cohen. Review-based information services: Lessons learned from The Boston Restaurant List. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-

                                        314X. URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/InfOrg/
                                        cohen/info-serv.html.

VanHeyningen:1994:JLT

[1636] M. VanHeyningen. JW looking for title/link. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Vanderheiden:1994:DHI

[1637] Greg Vanderheiden. Design of human interfaces for bandwidth-limited users and people with disabilities. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

McKearney:1994:LWI

[1638] Stephen McKearney and Rowan Limb.

LITBASE — A WWW interface to documented databases. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Boucher:1994:LSD

[1639] Robert A. Boucher. Large-scale database publishing. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Eichmann:1994:MCS

[1640] Dave Eichmann. MORE: A case study in reengineering a database application for the Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Gelb:1994:RDM

[1641] Laurie Gelb. A representational data model for Web site administrators. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/InfOrg/
gelb/Gelb.html.

Likavec:1994:MLN

[1642] Jaromir Likavec. Mosaic launches new dimensions in the LAN management. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

                                        314X. URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/InfOrg/
                                        likavec/lanmanagement.html.

Joshi:1994:SPI

[1643] Rajeev Joshi and Hans Joseph. Selective presentation of information on the Web. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/EvolvingInter/
                                        joshi/Joseph-Joshi.html.

Proffer:1994:TTW

[1644] William J. Proffer. Taxing times: WWW access to federal and state tax forms, information and assistance. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Fin/proffer/
                                        proffer_paper.html.

Cooper:1994:IGB

[1645] Chris Cooper. The Internet grows up: Bringing commercial applications to end users. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Ginsburg:1994:EPC

[1646] Mark Ginsburg and Ajit Kambil. The EDGAR Project: A case study in disseminating financial data on the Internet. In Anonymous [2042], page ??

CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Fin/ginsburg/
edgar.html.

Frasca:1994:NMM

[1647] Robert Frasca. NETworth and Mosaic: The mutual fund distribution channel of the future. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Gee:1994:MDA

[1648] David A. Gee and Mike Boulter.

MosaicForms database access?: A palaeobotanic case study. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/GenSci/gee/
pfrwww.html.

Beck:1994:IFC

[1649] Bradley Carl Beck and Antonio M.

Baptista. An interactive forum for convection-diffusion problems. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/GenSci/
beck/beck-baptista.html.

Nielsen:1994:SUI

[1650] Jakob Nielsen and Darrell Sano. Sun-Web: User interface design for Sun Microsystem’s internal Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/HCI/nielsen/
sunweb.html.

Britton:1994:DUI

[1651] David R. Britton, Jr. and Arthur A. Reyes. Discovering usability improvements for Mosaic: Applications of the contextual inquiry technique with an expert user. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/HCI/britton/
                                        britton_reyes.html.

Randall:1994:CSI

[1652] Neil Randall. Countering the surfing impulse: Designing an attentionholding Web site. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Register:1994:IPI

[1653] Joel D. Register and John L. Gerone. Interfaces for public information and scientific research. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/HCI/register/
                                        register.html.

Hill:1994:CHU

[1654] Will Hill, Mark Rosenstein, and Larry Stead. Community and history-of-use navigation. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/HCI/hill/home-
                                        page.html.

Fontaine:1994:HII

[1655] Paul Fontaine. Human interface issues around accessibility by people with disabilities. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Glazer:1994:LLI

[1656] David Glazer. Lessons learned implementing a navigation server for the Web. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/HCI/glazer/
glazer.html.

Hsu:1994:AOH

[1657] Jui-Chieh Hsu, William Johnston, and John McCarthy. Active outlining for HTML documents: An X-Mosaic implementation. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/HCI/hsu/hsu.
html.

Seaman:1994:YWR

[1658] Gary Seaman. From the Yanomamo to Wudangshan: Representing ethnography on the World Wide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Bovingdon:1994:AND

[1659] Andrew J. Bovingdon. Adding a new dimension to the desktop. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Bayle:1994:JWU

[1660] Michael L. Bayle, Burton H. Lee, A. Goto, Y. Sakamoto, J. Hong, V. Lam, T. Takada, M. Washisaka, A. Yamazaki, and K. Kada. Japan Window: A US-Japan WWW collaboration for Japan information. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Int/lee/japan_
window.html.

Baker:1994:FGF

[1661] Thomas Baker, Inke Bruning, Lothar Klein, and Michael Lenz. Not free, but growing fast: The Web in Germany. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Int/baker/
web-in-germany.html.

Tsang:1994:IGA

[1662] Philip Tsang, James Henri, and Sandy Tse. Internet growth in Australia and Asia’s four dragons. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Int/tsang/
tsang.html.

Ridley:1994:IIA

[1663] Michael Ridley. Innovation and implementation: Adopting and managing World Wide Web services in academic libraries. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/LibApps/ridley/
ridley.html.

John:1994:IML

[1664] Nancy R. John and Edward J.

Valauskas. Introducing Mosaic to library end users to access Britannica Online. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Koopman:1994:SYO

[1665] Ann Koopman and Sharon Hay. Swim at your own risk — no librarian on duty: Large-scale application of Mosaic in an academic library. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/LibApps/hay/
WWWPap.html.

Davis:1994:NSL

[1666] Bob Davis, Jack Fritts, and Phyllis Geren. North Suburban Library System Internet training project. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/LibApps/davis/
nsls.html.

Smith:1994:SSB

[1667] Alastair G. Smith. South Seas BUBL:

A WWW information service for library and information studies. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/LibApps/smith.
                                        alastair/a_smith.html.

Jevec:1994:BIF

[1668] Tom Jevec. Bringing the Internet to faculty: Building partnerships to customize access to Internet resources. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Mockapetris:1994:COP

[1669] Paul Mockapetris, Ron Ohlander, and Craig Milo Rogers. Cities online: Public access to city services on the Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Kanfer:1994:RWC

[1670] Alaina G. Kanfer. The role of the Web in connecting a community: The development of Champaign County Network (CCNet). In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Cantrall:1994:OSA

[1671] Dan Cantrall. The Oregon State Archives on the World Wide Web: A case study in building a new user base for government services. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/

SDG/IT94/Proceedings/LocalCom/cantrall/
cantrall.html.

Cho:1994:IHI

[1672] Pagogh Cho, James Gutierrez, Anne LeMaistre, M. D., Cameron Payne, Russell Pruitt, and Charlotte Wilcox. IntelliTeaching — high impact learning. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/MedEd/
cho/ncsa-cho.html.

Davis:1994:ICB

[1673] Nancy B. Davis, Pagogh Cho, James Gutierrez, Cameron Payne, Russell Pruitt, Charlotte Wilcox, Anne LeMaistre, and M. D. Interactive computer based programs versus traditional methods for selfstudy. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/MedEd/
davis/nbdavis.html.

Espiritu:1994:LUM

[1674] Baltazar Espiritu, John A. Mc-Nulty, James R. Halama, Arcot J. Chandrasekhar, Ronald Price, Jr., and Brain Maggi. The Loyola University Medical Education Network. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/MedEd/
espiritu/LUMEN.html.

Frisby:1994:MEW

[1675] Anthony J. Frisby, James Baggott, Sharon Dennis, and Rodney B. Murray. Medical education and the World Wide Web. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

                                        314X. URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/MedEd/
                                        frisby/frisby.html.

Srinivas:1994:AMH

[1676] K. Srinivas, K. Gopinath, V. Jagannathan, R. Karinthi, Matthew Fuchs, Y. V. Reddy, George Almasi, and Tad Davis. Applications of Mosaic in health care delivery. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/MedTrack/srinivas/
                                        artemis.html.

Oliver:1994:DHN

[1677] Katie S. Oliver, Christine Remenyik, Alexander Crosby, and Jim Kazmer. Dynamic hypertext navigation and display management. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/MedTrack/oliver/
                                        OliverRem.html.

Willard:1994:WBM

[1678] K. E. Willard, J. H. Hallgren, and D. P. Connelly. W3 based medical information systems vs. custom client server applications. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/MedTrack/willard/
                                        UMHC_www/UMHC_www_paper.html.

Williamson:1994:DHI

[1679] Kenneth B. Williamson and Tommy Williams. Design of hypertext instructional modules in pathology: The transition from hypercard to the Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Sieburg:1994:CWS

[1680] Hans B. Sieburg, Stella Veretnik, Jay M. Otero, and Donald E. Mosier. Cybermouse WWW simulation server. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/MedTrack/sieburg/
sieburg.html.

Salomon:1994:EJB

[1681] Marc E. Salomon and David C. Martin. An electronic journal browser implemented in the World Wide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/MedTrack/salomon/
salomon.html.

Harbourt:1994:NWB

[1682] Anna Harbourt. NetCoach: A Webbased system for intelligent biomedical information retrieval. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Benton:1994:MSH

[1683] Vivian M. Benton, Dan Dwyer, Cordelia Baron Geiken, Greg McArthur, and Joshua Polterock. MetaCenter science highlights repository. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Meta/metainfo.
paper.html.

Cox:1994:RTI

[1684] Mark J. Cox and John E. F. Baruch. Robotic telescopes: An interactive exhibit on the World Wide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Museum/cox/
markcox.html.

Rasmussen:1994:GIM

[1685] Mary Rasmussen, Brad Blumenthal, Steve Borysewicz, and Paul Neumann. The global interactive museum using Mosaic. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/Museum/
rasmussen/global_museum.html.

Bohrer:1994:AFW

[1686] Rand Bohrer. Architecture of the future — A Web exhibition. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/Museum/
bohrer/bohrer.html.

Rakov:1994:ISM

[1687] Simon A. Rakov. The Internet sampler: A Mosaic-based museum kiosk about the Internet. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-

314X. URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/Museum/
rakov/rakov.html.

Macaulay:1994:WCG

[1688] Tyson Macaulay. WWW clearinghouse for government information. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/NationWideSites/
macaulay/macaulay.html.

Boyer:1994:IDD

[1689] Paul D. Boyer. Information discovery and distillation in government: An experience report. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/NationWideSites/
scott/index.html.

Fernandez:1994:UDH

[1690] Mark R. Fernandez, Jim Miller, Jeffrey L. Becklehimer, and Bill Maloy. The use of distributed hypermedia at The Naval Research Laboratory. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/NationWideSites/
fernandez/fernandez.html.

Slone:1994:WWW

[1691] Benjamin J. Slone III and Charles E.

Richardson. The World Wide Web and the nuclear industry. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/NationWideSites/
nelson/paper.html.

Nelson:1994:WWW

[1692] Michael L. Nelson and David J. Bianco.

The World Wide Web and technology transfer at NASA Langley. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Gottlich:1994:AWI

[1693] Gretchen Gottlich. Application of the WWW and Internet information technologies in a work environment. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Daddio:1994:NWE

[1694] Ernest Daddio and William Turnbull. NOAA on the Web — the experiences and future of one nation-wide user. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/NationWideSites/
turnbull/paper.html.

Arnett:1994:IAN

[1695] Nick Arnett. The Internet and the Anti-net: Two public internetworks are better than one. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Overviews/
arnett/Antinet.html.

Blumberg:1994:MPL

[1696] Roger B. Blumberg. Museums, public lands, and billboards: Toward a philosophy of the World Wide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Overviews/
blumberg.museums/WebPhil94.Blumberg.
html.

Duffy:1994:THR

[1697] Robert A. Duffy and Michael J.

Palmer. Taking the high road to institutional self-promotion: Substantive corporate communication on the Worldwide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Overviews/
duffy/main.html.

Munnecke:1994:WWW

[1698] Tom Munnecke. World Wide Web and the demise of the clockwork universe. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Overviews/
munnecke/abstract.html.

Fry:1994:PNM

[1699] Andrew L. Fry. Publishing in the new mass medium: Creating content on the Internet. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Overviews/
fry/fry.html.

Avalanche:1994:PIS

[1700] Avalanche. Product Information Sessions. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

EBT:1994:DSS

[1701] EBT. ‘DynaWeb’ SGML server for the Web with automated transformation into HTML for display. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Exoterica:1994:TTH

[1702] Exoterica. Transformation tools with HTML conversion runtime module. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Datalogics:1994:HP

[1703] Frame/Datalogics. HTML printing. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

HAL:1994:OHS

[1704] HAL. Olias HTML/SGML server for the Web. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

IDI:1994:FLS

[1705] IDI. With fall launch of SGML server for the Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

SoftQuad:1994:HHP

[1706] SoftQuad. HotMetal, HotMetal PRO commercial version, ongoing development and support of Raggett’s TIDY utility. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Dougherty:1994:PCW

[1707] Dale Dougherty, Robert Frasca, Paul Hoffman, Boris Putanec, and Kerry Schwartz. Panel: Commercial Web information providers. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Krauskopf:1994:NPC

[1708] Tim Krauskopf. Network publishing: The center of the Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Weibel:1994:ASP

[1709] Stuart Weibel, Eric Miller, Jean Godby, and Ralph LeVan. An architecture for scholarly publishing on the World Wide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Pub/weibel/
weibel_www_paper.html.

Davis:1994:DPW

[1710] Jim Davis and Carl Lagoze. “dropin” publishing with the World Wide Web. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Pub/davis/
                                        davis-lagoze.html.

Dougherty:1994:AP

[1711] Dale Dougherty. Adventures in publishing. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Weibel:1994:PDL

[1712] Stuart L. Weibel, Nick Arnett, Tony Barry, and Sue Davidsen. Panel: Delivery of library services on the Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Berrigan:1994:OWE

[1713] Ellen Berrigan. ONline WISCON-SIN: An electronic journal. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Kumar:1994:EJ

[1714] Kumar.  Electronic journal.  In

Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Martin:1994:OFU

[1715] Gerard Martin. Online fair use of copyrighted material: Issues and concerns. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Pub/martin/
WWW-Online-Essay.html.

Shklar:1994:III

[1716] Leon Shklar, Satish Thatte, Howard Marcus, and Amit Sheth. The Info-Harness information integration platform. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Searching/
shklar/shklar.html.

Pinkerton:1994:FWP

[1717] Brian Pinkerton. Finding what people want: Experiences with the WebCrawler. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN

0309-314X. URL http://www.ncsa.
uiuc.edu/SDG/IT94/Proceedings/Searching/
pinkerton/WebCrawler.html.

Doszkocs:1994:ACN

[1718] Tamas E. Doszkocs, Seth B. Widoff, and Bruno M. Vasta. Associative concept navigation in MEDLINE and other NLM databases via a Mosaic — forms — WWW interface combining natural language processing, expert systems and (un)conventional information retrieval techniques. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Searching/
doszkocs/doszkocs.html.

Lim:1994:UCI

[1719] Jong-Gyun Lim. Using coollists to index HTML documents in the Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Searching/
                                        lim/coollist.html.

Rodgers:1994:STA

[1720] R. P. C. Rodgers, Suresh Srinivasan, and Jim Fullton. Sourcerer: Thesaurus-assisted automated source identification for the World Wide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Searching/
                                        rodgers.sourcerer/paper.html.

Schwartz:1994:HID

[1721] Michael F. Schwartz, C. Mic Bowman, Peter B. Danzig, Darren R. Hardy, and Udi Manber. The Harvest information discovery and access system. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Searching/
                                        schwartz.harvest/schwartz.harvest.
                                        html.

Casey:1994:WTM

[1722] James Casey, Alberto Aimar, Nikos Drakos, Ian Hannell, Arash Khodabandeh, Paolo Palazzi, Bertrand Rousseau, and Mario Ruggier. WebLinker: A tool for managing cross references between WWW documents. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/

SDG/IT94/Proceedings/Searching/
casey/WLPaper_1.html.

Doemel:1994:WGH

[1723] Peter Doemel. WebMap — A graphical hypertext navigation tool. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Searching/
doemel/www-fall94.html.

Crossley:1994:WTW

[1724] David Crossley. WAIS through the Web — discovering environmental information. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Searching/
crossley/paper.html.

Beck:1994:WWP

[1725] Howard Beck, Amir Mobini, and Viswanath Kadambari. A word is worth 1000 pictures: Natural language access to digital libraries. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Searching/
beck/beckmain.html.

Clarkson:1994:IDH

[1726] Mike Clarkson. Intelligent distributed hypertext documentation. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Woelk:1994:IPI

[1727] Darrell Woelk and Christine Tomlinson. The InfoSleuth Project: Intelligent search management via semantic agents. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Searching/
                                        woelk/woelk.html.

Bra:1994:SAI

[1728] Paul De Bra and R. D. J. Post. Searching for arbitrary information in the World Wide Web: the fish-search for Mosaic. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Searching/
                                        debra/article.html.

Chen:1994:SRN

[1729] Hsinchun Chen and Bruce R. Schatz. Semantic retrieval for the NCSA Mosaic. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Searching/
                                        chen/chenschatz.html.

Schiffman:1994:WSB

[1730] Allan M. Schiffman. WWW security: The big picture. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Hallam-Baker:1994:SSE

[1731] Phillip M. Hallam-Baker, Henrik Frystyk Nielsen, and Ari Luotonen. SHEN security enhancements to HTTP. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Security/hallam-
baker/Shen/Documentation/paper.
html.

Dalva:1994:CSH

[1732] Dave Dalva. On client and server host security. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Altis:1994:PC

[1733] Kevin Altis. On proxies (not confirmed). In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Poler:1994:IWM

[1734] Ariel Poler. Improving WWW marketing through user information and non-intrusive communications. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Security/poler/
poler.html.

Lewontin:1994:DWP

[1735] Steve Lewontin and Mary Ellen Zurko.

The DCE Web Project: Providing authorization and other distributed services to the World Wide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Security/lewontin/
Web_DCE_Conf_94.html.

Barber:1994:UMD

[1736] David Barber. The University of Michigan Data System. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/SocialSci/
                                        barber/umdata.html.

Jacobsen:1994:PAD

[1737] Lynn Jacobsen, David Millman, and Walter Bourne. Providing access to a data library: SQL and full-text IR methods of automatically generating Web structure. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/SocialSci/
                                        jacobsen/jacobsen.html.

Rockwell:1994:T

[1738] Richard Rockwell.  title?  In

Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WWW2_Proceedings.
                                        html.

Kroeker:1994:IDC

[1739] Brian Kroeker, Hilde Colenbrander, Henry Fan, and George Phillips. Interactive database communication across the World Wide Web: Accessing statistics canada’s CANSIM database. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

                                        URL http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/SocialSci/
                                        kroeker/kroekeru.html.

Savage:1994:SSU

[1740] Dean Savage. Social science uses of Mosaic at the City University of New York. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/SocialSci/
savage/savage.html.

Goode:1994:INM

[1741] Susan Goode. Introduction to NCSA Mosaic. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Staff:1994:AFN

[1742] NCSA Development Staff. Advanced features of NCSA Mosaic. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Sheldrick:1994:H

[1743] Dennis Sheldrick, Michael Stefanini, and Michael Grobe. HTML. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Stevenson:1994:AH

[1744] Kim Stevenson. Advanced HTML.

In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Labovitz:1994:MCW

[1745] John Labovitz and Brian Golden. Management of content on WWW servers. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Hughes:1994:PMH

[1746] Kevin Hughes. Putting media into hypermedia. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN

0309-314X. URL http://www.ncsa.
uiuc.edu/SDG/IT94/Proceedings/WWW2_
Proceedings.html.

Schiffman:1994:SW

[1747] Allan Schiffman. Security on the Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Frystyk:1994:SAS

[1748] Henrik Frystyk and Carlos Varela. System administration/HTTPD servers for CERN httpd. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Jones:1994:SAW

[1749] Paul Jones. System administration of WWW servers. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Pesce:1994:VWW

[1750] Mark D. Pesce and Anthony Parisi. Visualization and the World Wide Web. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Meyer:1994:MBC

[1751] Tom Meyer, David Blair, and Suzanne Hader. A MOO-based collaborative hypermedia system for WWW. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/VR/meyer.waxweb/
meyer.html.

Ressler:1994:AUV

[1752] Sandy Ressler. Approaches using virtual environments with Mosaic. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/VR/ressler/
veMain.html.

Munzner:1994:VTW

[1753] Tamara Munzner, Paul Burchard, and Ed Chi. Visualization through the World Wide Web with Geomview, Cyberview, W3Kit, and WebOOGL. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/VR/munzner/
munzner-abstract.html.

Mason:1994:CMB

[1754] Joanna Mason, Marek Czernuszenko, Dana Plepys, and Thomas A. DeFanti.

CAVEview: Mosaic-based virtual reality. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/VR/mason/CAVEview.
paper.html.

Campbell:1994:IAC

[1755] Paul Campbell, Lawrence Fitzpatrick, David C. Macdonald, and Adrienne Griffith. Information access: A cornerstone of Web publishing. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WebProd/griffith/
witehtml.html.

Shapiro:1994:VPF

[1756] Ehud Shapiro. Virtual places — A foundation for human interaction. In Anonymous [2042], page ?? CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WebProd/shapiro/
demo1.html.

Penberthy:1994:IWA

[1757] Scott Penberthy. IBM WebExplorer: Advanced features for navigating the Web. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Dean:1994:WHA

[1758] Shaun Dean. What’s happening at Apple. In Anonymous [2042], page ?? CODEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

Pitkow:1995:RTW

[1759] J. E. Pitkow and C. M. Kehoe. Results from the third WWW user survey. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 1–14. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995.

US$39.95. URL http://www.ora.com/
gnn/bus/ora/item/wj1.html. The

World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Hurwitz:1995:OMW

[1760] R. Hurwitz and J. C. Mallery. The open meeting: A Web-based system for conferencing and collaboration. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 15– 36. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995.

US$39.95. URL http://www.ora.com/
gnn/bus/ora/item/wj1.html. The

World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Vitali:1995:UVP

[1761] F. Vitali and D. G. Durand. Using versioning to provide collaboration on the WWW. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 37–50. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Wittenburg:1995:GAB

[1762] K. Wittenburg, D. Das, W. Hill, and L. Stead. Group asynchronous browsing on the World Wide Web. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 51– 62. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995.

US$39.95. URL http://www.ora.com/
gnn/bus/ora/item/wj1.html. The

World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Bentley:1995:SCI

[1763] R. Bentley, T. Horstmann, K. Sikkel, and J. Trevor. Supporting collaborative information sharing with the WWW: The BSCW shared workspace system. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 63– 74. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995.

US$39.95. URL http://www.ora.com/
gnn/bus/ora/item/wj1.html. The

World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Rees:1995:WDO

[1764] O. Rees, N. Edwards, M. Madsen, and M. Beasley. A Web of distributed objects. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 75–88. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Ingham:1995:WBO

[1765] D. Ingham, M. Little, S. Caughey, and S. Shrivastava. W3Objects: Bringing object-oriented technology to the Web. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 89– 106. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995.

US$39.95. URL http://www.ora.com/
gnn/bus/ora/item/wj1.html. The

World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Malpani:1995:MWW

[1766] R. Malpani, J. Lorch, and D. Berger. Making World Wide Web caching servers cooperate. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 107–118. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

                                        URL http://www.ora.com/gnn/bus/
                                        ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Abrams:1995:CPL

[1767] M. Abrams, C. R. Standridge, G. Abdulla, and S. Williams. Caching proxies: Limitations and potentials. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 119– 134. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995.

                                        US$39.95. URL http://www.ora.com/
                                        gnn/bus/ora/item/wj1.html. The

World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Beckett:1995:ITU

[1768] D. Beckett. IAFA templates in use as Internet metadata. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 135–144. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Yuwono:1995:WWW

[1769] B. Yuwono, S. L. Y. Lam, J. H. Ying, and D. L. Lee. A World Wide Web resource discovery system. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 145–158. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Kamba:1995:KCI

[1770] T. Kamba, K. Bharat, and M. C. Albers. The Krakatoa Chronicle: An interactive personalized newspaper on the Web. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 159–170. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

                                        URL http://www.ora.com/gnn/bus/
                                        ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Gaines:1995:WMC

[1771] B. R. Gaines and M. L. G. Shaw. Web map: Concept mapping on the Web. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 171– 184. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995.

                                        US$39.95. URL http://www.ora.com/
                                        gnn/bus/ora/item/wj1.html. The

World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Hunter:1995:SAG

[1772] A. Hunter, I. Ferguson, and S. Hedges. Swoop: An application generator for ORACLE/WWW systems. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 185–194. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Selberg:1995:MES

[1773] E. Selberg and O. Etzioni. Multiengine search and comparison using the MetaCrawler. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 195–208. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Varela:1995:DBO

[1774] C. Varela, D. Nekhayev, P. Chandrasekharan, and C. Krishnan. DB: Browsing object-oriented databases over the Web. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 209–220. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

                                        URL http://www.ora.com/gnn/bus/
                                        ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Lai:1995:TNE

[1775] M.-C. Lai, B.-H. Chen, and S.-M. Yuan. Toward a new educational environment. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 221–230. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

                                        URL http://www.ora.com/gnn/bus/
                                        ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Hubler:1995:CIH

[1776] A. W. Hubler and A. M. Assad. Cyber-Prof: An intelligent human-computer interface for asynchronous wide area training and teaching. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 231–238. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

                                        URL http://www.ora.com/gnn/bus/
                                        ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Schroeder:1995:MTS

[1777] U. Schroeder, B. Tritsch, and A. Krjerriem-Jasnoch. A modular training system for education in the WWW environment. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 239–250. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Antchev:1995:WLE

[1778] K. Antchev, M. Luhtalahti, J. Multisilta, and S. Pohjolainen. A WWW learning environment for mathematics. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 251– 258. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995.

US$39.95. URL http://www.ora.com/
gnn/bus/ora/item/wj1.html. The

World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Schoenfeldinger:1995:WML

[1779] W. J. Schoenfeldinger. WWW meets Linda: Linda for global WWW-based transaction processing systems. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 259– 276. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995.

                                        US$39.95. URL http://www.ora.com/
                                        gnn/bus/ora/item/wj1.html. The

World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Barta:1995:IPG

[1780] R. A. Barta and M. Hauswirth. Interface-parasite gateways. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 277– 290. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995.

                                        US$39.95. URL http://www.ora.com/
                                        gnn/bus/ora/item/wj1.html. The

World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Lilley:1995:JDQ

[1781] C. Lilley. Not just decoration: Quality graphics for the Web. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 291–308. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

VanOssenbruggen:1995:BMW

[1782] J. Van Ossenbruggen and A. Eliens.

Bringing music to the Web. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 309–314. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Ang:1995:PVC

[1783] C. S. Ang, M. D. Doyle, and P. Brantley. Polymap: A versatile client-side image map for the Web. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 315–322. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Thompson:1995:TIM

[1784] R. Thompson and K. Shafer. Translating ISO 12083 mathematical markup for electronic documents. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 323–332. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Chen:1995:RTV

[1785] Z. Chen, S.-M. Tan, R. H. Campbell, and Y. Li. Real-time video and audio in the World Wide Web. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 333–348. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Barnard:1995:LWW

[1786] D. T. Barnard, L. Burnard, S. J.

DeRose, and D. G. Durand. Lessons for the World Wide Web from the text encoding initiative. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 349–358. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Lucco:1995:OUS

[1787] S. Lucco, O. Sharp, and R. Wahbe.

Omniware: A universal substrate for Web programming. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 359–368. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Yellin:1995:LLS

[1788] F. Yellin. Low level security in Java. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 369– 380. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995.

                                        US$39.95. URL http://www.ora.com/
                                        gnn/bus/ora/item/wj1.html. The

World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Weeks:1995:CBW

[1789] J. D. Weeks, A. Cain, and B. Sanderson. CCI-based Web security: A design using PGP. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 381–396. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

                                        URL http://www.ora.com/gnn/bus/
                                        ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Jones:1995:SWW

[1790] M. F. Jones and B. Schneier. Securing the World Wide Web: Smart tokens and their implementation. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 397–410. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Arisland:1995:ICA

[1791] K. O. Arisland, S. A. Johansen, and G. Roenning. Introducing Candleweb and (awe), bringing animation power to the World Wide Web. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 411–422. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Baker:1995:LCF

[1792] B. S. Baker and E. Grosse. Local control over filtered WWW access. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 423– 432. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995.

                                        US$39.95. URL http://www.ora.com/
                                        gnn/bus/ora/item/wj1.html. The

World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Ladd:1995:MHM

[1793] B. C. Ladd, M. V. Capps, P. D. Stotts, and R. Furuta. Multi-head multi-tail Mosaic. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 433–440. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

                                        URL http://www.ora.com/gnn/bus/
                                        ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Dardailler:1995:MGW

[1794] D. Dardailler. Mobile GUI on the Web. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 441– 450. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995.

                                        US$39.95. URL http://www.ora.com/
                                        gnn/bus/ora/item/wj1.html. The

World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Ayers:1995:UGH

[1795] E. Z. Ayers and J. T. Stasko. Using graphic history in browsing the World Wide Web. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 451–460. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Lingnau:1995:HBI

[1796] A. Lingnau, O. Drobnik, and P. Doemel.

An HTTP-based infrastructure for mobile agents. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 461–472. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95. URL http:

//www.ora.com/gnn/bus/ora/item/

wj1.html. The World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Davies:1995:JCI

[1797] J. Davies, R. Weeks, and M. Revett. Jasper: Communicating information agents for WWW. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 473–482. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

                                        URL http://www.ora.com/gnn/bus/
                                        ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Vidovic:1995:CWB

[1798] N. Vidovic and D. F. Vrsalovic. Constellation: A Web-based design framework for developing network applications. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 483–492. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

                                        URL http://www.ora.com/gnn/bus/
                                        ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Sollins:1995:LGI

[1799] K. R. Sollins and J. R. Van Dyke. Linking in a global information architecture. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 493–508. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Cuff:1995:CHP

[1800] L. D. Cuff. Commercial hypertext publishing: Electronic books using trails and the author-publisher-reader model. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 509– 518. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995.

US$39.95. URL http://www.ora.com/
gnn/bus/ora/item/wj1.html. The

World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Francis:1995:ISC

[1801] P. Francis, T. Kambayashi, S.-Y. Sato, and S. Shimizu. Ingrid: A selfconfiguring information navigation infrastructure. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 519–538. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

                                        URL http://www.ora.com/gnn/bus/
                                        ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Brooks:1995:ASP

[1802] C. Brooks, M. S. Mazer, S. Meeks, and J. Miller. Application-specific proxy servers as HTTP stream transducers. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 539– 548. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995.

                                        US$39.95. URL http://www.ora.com/
                                        gnn/bus/ora/item/wj1.html. The

World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Nicol:1995:DWI

[1803] G. T. Nicol. Dyna Web: Integrating large SGML repositories and the WWW. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 549–558. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

                                        URL http://www.ora.com/gnn/bus/
                                        ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Diaz:1995:RTD

[1804] A. Diaz, T. Isakowitz, V. Maiorana, and G. Gilabert. RMC: A tool to design WWW applications. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 559–566. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Ladd:1995:PWA

[1805] D. A. Ladd and C. Ramming. Programming the Web: An applicationoriented language for hypermedia service programming. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 567–586. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Peirce:1995:SSC

[1806] M. Peirce and D. O’Mahony. Scalable, secure, cash payment for WWW resources with the PayMe protocol set. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 587– 602. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995.

                                        US$39.95. URL http://www.ora.com/
                                        gnn/bus/ora/item/wj1.html. The

World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Glassman:1995:MPI

[1807] S. Glassman, M. Manasse, M. Abadi, and P. Gauthier. The Millicent protocol for inexpensive electronic commerce. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 603–618. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95. URL http:

                                        //www.ora.com/gnn/bus/ora/item/

wj1.html. The World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Kesseler:1995:SBA

[1808] M. Kesseler. A schema-based approach to HTML authoring. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 619–632. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Paoli:1995:REW

[1809] J. Paoli. Rules for extending a WWW client: The symposia API. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 633–646. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Carr:1995:DLS

[1810] L. Carr, D. De Roure, W. Hall, and G. Hill. The distributed link service: A tool for publishers, authors, and readers. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 647–656. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Decouchant:1995:SCA

[1811] D. Decouchant, V. Quint, and M. Romero Salcedo. Structured cooperative authoring on the World Wide Web. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 657–666. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95. URL http:

//www.ora.com/gnn/bus/ora/item/

wj1.html. The World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Dyreson:1995:BWP

[1812] C. E. Dyreson and A. M. Sloane. The boomerang white paper: A page as you like it. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 667–676. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Paulos:1995:WWW

[1813] E. Paulos and J. Canny. A World Wide Web telerobotic remote environment browser. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 677–690. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Gallagher:1995:DTW

[1814] J. Gallagher and G. Milkowski.

Data transport within the distributed oceanographic data system. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 691–702. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Port:1995:RTA

[1815] G. Port, C. Heath, T. Segall, and P. Merrick. Requirements for taking applications beyond the enterprise. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 703– 710. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995.

                                        US$39.95. URL http://www.ora.com/
                                        gnn/bus/ora/item/wj1.html. The

World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Neves:1995:CIO

[1816] F. L. Neves and J. N. Oliveira. Classifying Internet objects. In O’Reilly and Associates and Web Consortium (W3C) [2048], pages 711–722. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

                                        URL http://www.ora.com/gnn/bus/
                                        ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Foldvik:1995:MTA

[1817] R. G. Foldvik and D. Meyer. Moving towards ATM: LAN/WAN evolution and experimentation at the University of Oregon. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 3– 8. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Clark:1995:DAD

[1818] R. J. Clark, R. R. Hutchins, and S. Walker Register. Deploying ATM in a data network: An analysis of SVC requirements. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 9– 18. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Hou:1995:PHS

[1819] C.-J. Hou, C.-C. Han, and W.-C.

Chau. Priority-based high-speed switch scheduling for ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 19–28. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Lo:1995:FAA

[1820] C.-C. Lo and C.-Y. Chiu. A faulttolerant architecture for ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 29–36. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Yokotani:1995:CCM

[1821] T. Yokotani, T. Ichihashi, and M. Ishizaka. Congestion control of multi-cast connections in ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 38–47. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Sohn:1995:SSC

[1822] E. Sohn and T. Chung. The software structure of the call control for ATM systems. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 48–52. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Yuang:1995:ETR

[1823] M. C. Yuang and Y. G. Chen. An efficient terminal-pair reliability algorithm for network management. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 53–61. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Ascia:1995:DVF

[1824] G. Ascia, G. Ficili, and D. Panno.

Design of a VLSI fuzzy processor for ATM traffic sources management. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 62–72. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Jayasumana:1995:PPS

[1825] A. P. Jayasumana, S. A. Abd-Elmalak, and C. Vaishnav. Performance of a packet switched WDM network with dynamic wavelength selection. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 73–81. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Pavan:1995:EER

[1826] A. Pavan, J. Liu, A. Guha, and J. Pugaczewski. Experimental evaluation of real-time support on the Mercuri wide area ATM testbed. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 82–91. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Du:1995:OMR

[1827] D. H. C. Du, A. Pavan, S. Bhattacharya, and A. Guha. Optimal multihop routing: An iterative approach to TWDM embedding. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 92–101. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Kidambi:1995:IMA

[1828] K. R. Kidambi, R. Sankar, and J. Ottensmeyer. Internetworking MANs to ATM for broadband services support. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 102–113. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Pugaczewski:1995:MTC

[1829] J. Pugaczewski, W. Walling, and M. Wiklund. Multimedia traffic characterization from the high definition distance learning ATM trial. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 114–122. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Guizani:1995:DAA

[1830] M. Guizani. Design and analysis of all-optical ATM switching elements for high speed networks. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 123–131. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Ritter:1995:PCD

[1831] M. Ritter, S. Kornprobst, and F. Huebner. Performance comparison of design alternatives for source policing devices in ATM systems. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 132–139. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Mirfakhraei:1995:PAA

[1832] N. Mirfakhraei. Performance analysis of an ATM local computer network. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 140–149. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Jayasumana:1995:IFN

[1833] A. P. Jayasumana, Y. K. Malaiya, and R. V. Vogety. Interconnection of FDDI-II networks through an ATM backbone — an analysis. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 150–158. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Seo:1995:DCR

[1834] G. Seo, S.-M. Kang, D.-Y. Kim, and H.-B. Jeon. A distributed cycle reset protocol for the high-speed LAN/ MAN. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 159– 163. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Yasrebi:1995:EDO

[1835] M. Yasrebi. Experience with distributed objects in a portable and multithreaded library for a LAN/WAN gateway application. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 164–173. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Kumar:1995:RBC

[1836] S. Kumar and A. P. Jayasumana. Request based channel access protocol on folded bus topology. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 174–183. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Cullen:1995:IVO

[1837] C. Cullen. Internet vulnerabilities of one-time passcode mechanisms. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 184–186. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Hayashida:1995:MBR

[1838] Y. Hayashida, M. Ikegami, and N. Sugimachi. Multi-segment bandwidth reservation protocol for DQDB subnetwork. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 187–193. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Kim:1995:ITS

[1839] C. Kim and Jung Hoon Paik. An implementation of 100Mbps TAXI subscriber board. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 194– 201. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Elsaadany:1995:ACL

[1840] A. Elsaadany, M. Singhal, and M. T.

Liu. Alternative configurations for local area network design. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 202–208. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Esteve:1995:SSF

[1841] M. Esteve, C. Palau, and J. C. Guerri. A synchronization service framework for XTP. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 209– 218. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Zeino:1995:SAP

[1842] H. Zeino and M. Misson. A simulation architecture for a pico-cellular hybrid network. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 219– 226. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Barnett:1995:EFM

[1843] B. Barnett and A. Crapo. An expert fault manager using an object meta-model. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 227– 235. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

AlKasabi:1995:PAH

[1844] S. AlKasabi and S. Hariri. Performance analysis of a high-speed dynamically reconfigurable LAN. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 236–245. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Linder:1995:PXT

[1845] H. Linder, W. Stering, and U. Hofmann. Performance of XTP and TCP/IP in a satellite environment. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 246–253. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Hofmann:1995:QLB

[1846] U. Hofmann and M. Krajewski. Quasistatic load balancing in local area networks. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 254–?? ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Cooper:1995:CLL

[1847] S. H. Cooper and P. J. Teller. Configuring a large LAN for TCP/IP, AppleTalk, and IPX. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 264–270. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Bjoerkman:1995:RAT

[1848] N. Bjoerkman, A. Latour-Henner, O. Pers, and A. Miah. Real-time ATM traffic measurements and analysis. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 271–279. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Bjoerkman:1995:CLP

[1849] N. Bjoerkman, A. Latour-Henner, and U. Hansson. The cell loss process in ATM networks and its impact on quality of service. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 280– 288. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Saulnier:1995:DTB

[1850] E. T. Saulnier and B. J. Bortscheller. Data transfer bottlenecks over SPARCbased computer networks. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 289–295. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Feng:1995:GAD

[1851] F. Feng, S. Kamat, and W. Zhao.

Guaranteeing application-to-application deadlines in distributed real-time systems. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 296–306. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Pokam:1995:IMM

[1852] M. R. Pokam, J.-F. Guillaud, and G. Michel. Integrated multimedia in manufacturing networks using ATM. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 307–316. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Gupta:1995:PSA

[1853] S. Gupta and C. L. Williamson. A performance study of adaptive video coding algorithms for high speed networks. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 317–325. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Liu:1995:DDM

[1854] J. G. Liu, J. C. Liu, Y. G. Chen, and M. C. Yuang. DMTS: A distributed multimedia teleworking system. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 326–335. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Bolla:1995:QSM

[1855] R. Bolla, F. Davoli, and M. Marchese. Quality of service management and control of multimedia applications: A scenario and two simple strategies. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 336–?? ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Guizani:1995:NPO

[1856] M. Guizani. A new protocol for optical high-speed packet-switched networks. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 343–349. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Martini:1995:RCD

[1857] P. Martini and J. Ottensmeyer. Realtime communication in the demandpriority LAN the effects on normal priority traffic. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 350–?? ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Hiles:1995:AFM

[1858] W. S. Hiles and D. T. Marlow. Approximation of FDDI minimum reconfiguration time. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 360– 369. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Kosak:1995:BMF

[1859] C. Kosak, D. Eckhardt, T. Mummert, and P. Steenkiste. Buffer management and flow control in the credit net ATM host interface. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 370– 378. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Albert:1995:PAF

[1860] B. Albert and A. P. Jayasumana. Performance analysis of FDDI LANs using numerical methods. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 379–386. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Hayes:1995:SCS

[1861] W. Hayes and M. L. Molle. Solving capture in switched two-node Ethernets by changing only one node. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 387–396. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Rose:1995:SPM

[1862] O. Rose. Statistical properties of MPEG video traffic and their impact on traffic modeling in ATM systems. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 397–406. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Vasilakos:1995:ULA

[1863] A. V. Vasilakos, N. H. Loukas, and A. F. Atlasis. The use of learning algorithms in ATM networks call admission control problem: A methodology. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 407–412. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Chen:1995:OMR

[1864] Y. G. Chen, M. T. Yen, and M. C.

Yuang. Optimal multicast routing for ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 413–421. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Ramanujan:1995:MTS

[1865] R. S. Ramanujan, J. C. Bonney, J. L.

Mack, and C. R. Manson. Media transport service for video retrieval applications on ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 422–432. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Mirchandani:1995:AII

[1866] V. Mirchandani and D. Everitt. An architecture for internetworking integrated services LANs. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 433–442. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Pritty:1995:NCH

[1867] D. W. Pritty, D. N. Smeed, and N. L. Lawrie. A new class of high speed LAN access protocols based on the principle of timed packet release. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 443–452. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Bisdikian:1995:TBA

[1868] C. C. Bisdikian. Throughput behavior of the n-ary stack algorithm in mobile networks with capture. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 453–458. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Atwood:1995:DXS

[1869] J. W. Atwood and Y. Zhang. A definition of the XTP service and its formal specification. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 459– 468. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Barenco:1995:AVA

[1870] C. J. Barenco, E. Specialski, and L. Pacheco. Aspects of video and audio signals transmission in a FDDI network. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 469–??

ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Emerson:1995:PES

[1871] S. Emerson. Performance evaluation of switched fibre channel I/O system using FCP for SCSI. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 479–?? ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Emerson:1995:EDC

[1872] S. Emerson. Evaluation of a data communications model for switched fibre channel. In IEEE Computer Society. Technical Committee on Computer Communications [2046], pages 485–?? ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Varadharajan:1996:SHN

[1873] V. Varadharajan. Security in highspeed networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 2– 11. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Albert:1996:EDL

[1874] B. Albert and A. P. Jayasumana. The effects of data locality on FDDI-based schemes at gigabit speeds. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 12–18. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Lin:1996:CCH

[1875] Y.-D. Lin, T.-R. Huang, and Y.-C. Lai. Characterization and control of highly correlated traffic in high-speed networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 19–27. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Tan:1996:EBT

[1876] J. Tan, M. Sanchez, and C. Liang. Efficient buffering techniques for deflection routing in high-speed networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 28–39. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Barolli:1996:FPM

[1877] L. Barolli and K. Tanno. A fuzzy policing mechanism for still picture in ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 40–47. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Park:1996:DAN

[1878] B. S. Park and S. C. Kim. Design and analysis of a new fast packet switching fabric supporting multimedia traffic. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 48–53. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Bisdikian:1996:MGE

[1879] C. C. Bisdikian and F. A. Schaffa. The multimedia gateway: An evolution towards switched video services. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 54–58. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Lin:1996:EAV

[1880] M. Lin and D. Singer. An experimental ATM-based video-on-demand system on a Macintosh platform. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 59–68. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Yuang:1996:NIS

[1881] M. C. Yuang and P. L. Tien. A novel intra-media synchronization mechanism for multimedia communications. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 69–77. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Yan:1996:PAB

[1882] Z. Yan and Y.-C. Jenq. Performance analysis of a banyan-based ATM switching fabric with packet priority. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 78–87. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Cha:1996:MAN

[1883] Y.-W. Cha, J.-O. Kim, K.-H. Lee, and K.-J. Han. Multiaccess ATM network for residential and small business subscribers. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 88–93. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Tentov:1996:PAA

[1884] A. Tentov and A. Grnarov. Performance analysis of ATM switches with multistage packet switching interconnection networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 94–101. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Fan:1996:PFA

[1885] K. K. Fan and A. P. Jayasumana. Policing function in ATM network using multi-layer neural network. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 102–105. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Strayer:1996:CDT

[1886] W. T. Strayer. A class-chest for deriving transport protocols. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 106–115. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Lee:1996:DVA

[1887] C.-T. Lee and J. W. Harris. Designing a virtual access control configuration protocol for implementation over ISDN and shared-media networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 116– 125. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Singatwaria:1996:TBM

[1888] S. Singatwaria, R. Ahmad, and F. Halsall. Timer-based burst mode protocol for high-speed networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 126–135. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Huang:1996:DIM

[1889] N.-F. Huang, C.-A. Su, C.-W. Cheng, and C.-P. Wang. The design and implementation of a multicast real-time multimedia protocol. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 136–142. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Philp:1996:SBM

[1890] I. R. Philp, K. Nahrstedt, and J. W. S. Liu. Scheduling and buffer management for soft-real-time VBR traffic in packet-switched networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 143–152. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Lee:1996:PST

[1891] V. C. S. Lee, J. K. Y. Ng, K.-Y. Lam, and S.-L. Hung. Performance studies of transmitting real-time MPEG-I video in ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 153–160. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Duenkel:1996:IRT

[1892] C. Duenkel and D. Reschke. Intervalbased real-time transmission control. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 161–171. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Ahn:1996:MTA

[1893] S. Ahn and D. H. C. Du. A multicast tree algorithm considering maximum delay bound for real-time applications. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 172–181. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Yokotani:1996:PFC

[1894] T. Yokotani and T. Ichihashi. A proposal of flow control mechanism for multicast ABR and its performance. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 182– 190. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Harrison:1996:PSM

[1895] T. Harrison and C. Williamson. A performance study of multicast routing algorithms for ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 191–201. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Atwood:1996:RMX

[1896] J. W. Atwood, O. Catrina, J. Fenton, and W. T. Strayer. Reliable multicasting in the Xpress Transport Protocol. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 202–211. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Bauer:1996:ESM

[1897] D. Bauer, B. Stiller, and B. Plattner. An error-control scheme for a multicast protocol based on round-trip time calculations. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 212–221. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Gumbold:1996:SDR

[1898] M. Gumbold. Software distribution by reliable multicast. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 222–233. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Li:1996:AIL

[1899] G.-L. Li and F.-M. Li. An analysis of impacts of long-term traffic correlations on congestion control in ATM-based local computer networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 234–241. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Kim:1996:BBS

[1900] K.-B. Kim and H.-J. Kim. Backpressure buffering scheme to improve the cell loss property on the output buffered ATM switch. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 242–248. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Zheng:1996:ETT

[1901] Q. Zheng. An enhanced timed-round-robin traffic control scheme for ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 249–258. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Hah:1996:NCA

[1902] J. M. Hah, P. L. Tien, and M. C. Yuang. Neural-network-based call admission control in ATM networks with heterogeneous arrival. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 259–267. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Hussain:1996:AOB

[1903] S. S. Hussain and Y.-C. Jenq. Analysis and optimization of a banyan-based ATM switch by simulations. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 268–277. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Breker:1996:SSU

[1904] L. P. Breker and C. L. Williamson. A simulation study of usage-based pricing strategies for packet-switched networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 278–288. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Dowlatshahi:1996:CSN

[1905] M. Dowlatshahi and S. K. De. A comparative study of numerical and heuristic approaches to calculations of sustainable cell rate. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 289–294. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Plate:1996:PAF

[1906] C. Plate and J. Tan. Performance analysis of a fault-tolerant B-tree ATM switch. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 295– 304. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Valeroso:1996:ARS

[1907] E. S. Valeroso and M. Alam. Adaptive resource scheduling strategies and performance analysis of broadband networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 305– 315. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Chang:1996:SCM

[1908] S. G. Chang, D. Y. Suh, M. H. Kyung, and S. H. Park. A strategy for construction of a 40 mbps bundle of MPEG VBR bit streams. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 316–325. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Hong:1996:MSN

[1909] C. S. Hong, D. Kashiwa, Y. Koga, and Y. Matsushita. A multimedia service networking architecture and its applications in TINA-like model. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 326– 334. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Lee:1996:SAD

[1910] Y.-J. Lee, H.-J. Lee, W.-H. Ma, and D. H. C. Du. A systematic approach to design the network-based learning environment for home and office. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 335– 344. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Dempsey:1996:DBL

[1911] B. J. Dempsey and Y. Zhang. Destination buffering for low-bandwidth audio transmissions using redundancy-based error control. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 345–354. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Lin:1996:HNS

[1912] Y.-D. Lin, H.-Z. Lai, and Y.-C. Lai.

A hierarchical network storage architecture for video-on-demand services. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 355– 365. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Lee:1996:RAC

[1913] S. H. Lee. A reactive access control scheme at UNI for congestion control in ATM-based B-ISDN. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 366–373. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Talley:1996:GFC

[1914] T. Talley and K. Jeffay. A general framework for continuous media transmission control. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 374–383. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Lee:1996:PTE

[1915] C. T. Lee and J. W. Harris. Principle and technique for encapsulation of user control and data information in separate frames. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 384–395. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Christensen:1996:PEB

[1916] K. J. Christensen. Performance evaluation of the binary logarithmic arbitration method (BLAM). In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 396–403. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Wyss:1996:PLM

[1917] G. D. Wyss, H. K. Schriner, and T. R. Gaylor. Probabilistic logic modeling of network reliability for hybrid network architectures. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 404–413. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Tjhie:1996:DPE

[1918] D. D. Tjhie and H. Rzehak. Design and performance evaluation of network interconnection architectures. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 414–423. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Sato:1996:EPH

[1919] K. Sato and S. Takimoto. Evaluating performance of high-speed and long distance communications on HIPPI-ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 424–432. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Sun:1996:PAA

[1920] Y. S. Sun, C.-F. Ku, Y.-C. Pan, and C.-H. Wang. Performance analysis of application-level traffic shaping in a real-time multimedia conferencing system on Ethernets. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 433–442. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Park:1996:OIA

[1921] C. Park, S. M. Lee, and H. J. Kim.

The operation of IP and address resolution over the ATM LAN interworking unit. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 444–448. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Oh:1996:SPC

[1922] M.-K. Oh, D.-Y. Kim, S.-W. Ra, and C.-M. Han. A study on point-to-multipoint connection control in ATM switching system. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 449–457. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Feng:1996:MRH

[1923] F. Feng, C. Li, A. Raha, and S. Yu. Modeling and regulation of host traffic in ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 458–467. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Yang:1996:UPT

[1924] M.-R. Yang and G. K. Ma. Universal packet time slot — A new paradigm of designing an ATM switch. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 468–473. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Lee:1996:GCD

[1925] H. Lee. A gracious cell discard scheme in ATM multiplexer. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 474–485. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Wan:1996:TSL

[1926] P. J. Wan and A. Pavan. TWDM single-hop lightwave networks using multiple fixed transceivers at each station. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 486– 490. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Abbinante:1996:GNM

[1927] A. Abbinante and P. Camarda. GSN:

A new multihop architecture for optical networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 491–497. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Pavan:1996:NML

[1928] A. Pavan, P.-J. Wan, S.-R. Tong, and D. H. C. Du. A new multihop lightwave network based on the generalized 1De-Bruijn graph. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 498–507. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Abd-Elmalak:1996:PRW

[1929] S. A. Abd-Elmalak, C. Vaishnav, and A. P. Jayasumana. Performance of a robust WDM network with token-based reservations and a limited number of lasers at a station. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 508–517. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Camarda:1996:MMC

[1930] P. Camarda, G. Schiraldi, and F. Talucci. Mobility modeling in cellular communication networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 518–525. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Meggers:1996:PVC

[1931] J. Meggers, G. Bautz, and A. S.-B. Park. Providing video conferencing for the mobile user. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 526–534. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Bakshi:1996:PSC

[1932] B. S. Bakshi, P. Kirshna, D. K. Pradhan, and N. H. Vaidya. Providing seamless communication in mobile wireless networks. In IEEE Computer Society. Technical Committee on Computer Communications [2049], pages 535–543. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Hall:1997:CCC

[1933] J. Hall, R. Sabatino, S. Crosby, and I. Leslie. Counting the cycles: A comparative study of NFS performance over high speed networks. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 8–19. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Dempsey:1997:PAV

[1934] S. Dempsey, E. Fryland, B. Hazzard, and D. Cyganski. Performance of asynchronous, variable bit rate communications over ATM: A comparison with FDDI. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 20–27. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Tentov:1997:PAA

[1935] A. Tentov. Performance analysis of ATM switches with multistage multipath packet switching interconnection networks. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 28– 35. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Dowlatshahi:1997:NMP

[1936] M. Dowlatshahi and S. K. De. A numerical method for performance analysis of ATM multiplexers. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 36–43. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Pierre:1997:ICN

[1937] S. Pierre and A. Elgibaoui. Improving communication network topologies using tabu search. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 44–53. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Cam:1997:LIP

[1938] H. Cam. LAN-ATM Internetworking over a priority-based slotted-ring network. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 54– 63. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Kim:1997:DIB

[1939] D. J. Kim, H. Namgoong, and Y. C.

Lew. Design and implementation of bridge between CORBA and DCE. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 64–71. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Gayennet:1997:DSM

[1940] H. Gayennet, J. C. Lapayre, and M. Trehel. Distributed shared memory layer for cooperative work applications. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 72–79. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Mirchandani:1997:PIT

[1941] V. Mirchandani and D. Everitt. Performance of integrated telephony service over standalone LAN and ATM Internetworked LANs. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 80–88. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Aboelela:1997:FOM

[1942] E. Aboelela and C. Douligeris. Fuzzy optimization model for routing in B-ISDN. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 89– 98. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Beaubrun:1997:RAD

[1943] R. Beaubrun and S. Pierre. A routing algorithm for distributed communication networks. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 99–105. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Park:1997:NHR

[1944] C. Park, H. S. Choi, J. O. Kim, and J. H. Lee. Next hop resolution using classical IP over ATM. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 106–111. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Lin:1997:PCS

[1945] H. T. Lin and H. D. Hughes. Preserving cell sequence of handoff connections in wireless ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 112–122. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Bing:1997:MAT

[1946] B. Bing and R. Subramanian. Multiple access for third-generation personal communication systems. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 123–132. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Chintalapati:1997:ALM

[1947] R. V. J. Chintalapati, V. Kumar, and A. Datta. An adaptive location management algorithm for mobile computing. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 133– 141. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Wang:1997:EEC

[1948] J. Wang, K. B. Letaief, M. Hamdi, and X. R. Cao. Efficient estimation of cell loss and cell delay of nonblocking ATM switches. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 142–148. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Venkatesulu:1997:EFR

[1949] D. Venkatesulu and T. A. Gonsalves. Efficient fault-tolerant reliable broadcast in an extended LAN. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 149–158. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Qin:1997:OVP

[1950] Z. Qin, T. Lo, and F. F. Wu. The optimal virtual path design of ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 159– 169. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Harms:1997:VDH

[1951] J. J. Harms. Video distribution to the home using advanced reservation. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 170– 178. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Black:1997:PIV

[1952] R. Black, P. Barham, A. Donnelly, and N. Stratford. Protocol implementation in a vertically structured operating system. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 179–188. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Yeung:1997:OCR

[1953] K. H. Yeung and K. W. Ng. An optimal cache replacement algorithm for Internet systems. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 189–194. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Ma:1997:NER

[1954] K. T. Ma, R. Sankar, and K. J. Christensen. A new explicit rate-based congestion control scheme for ABR services. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 195– 203. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Karabeck:1997:NAP

[1955] R. Karabeck. A native ATM protocol architecture design and performance evaluation. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 204–210. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Molle:1997:SCC

[1956] M. Molle, M. Kalkunte, and J. Kadambi. Scaling CSMA/CD to 1Gb/s with frame bursting. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 211–219. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Sheu:1997:FOS

[1957] P. R. Sheu and C. M. Lin. A fast optimal slot reuse scheme for CRMA high speed networks. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 220–231. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Messier:1997:PMW

[1958] A. Messier, J. Robinson, and K. Pahlavan. Performance monitoring of a wireless campus area network. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 232–238. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Shigeno:1997:MPM

[1959] H. Shigeno, Y. Fujimaki, H. Omori, and K. Okada. MC02 platform for mobile and wireless computing. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 239–244. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Kwon:1997:SEA

[1960] T. Kwon and J. Song. Security and efficiency in authentication protocols resistant to password guessing attacks. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 245– 252. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Varadharajan:1997:SSP

[1961] V. Varadharajan, R. Shankaran, and M. Hitchens. Security services and public key infrastructure for ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 253– 263. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Guha:1997:ECF

[1962] A. Guha and T. S. Chang. The effectiveness of cell-level FEC for packet delivery in ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 264–273. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303.

LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Conrad:1997:DAP

[1963] C. Conrad and B. Stiller. The design of an application programming interface for QoS-based multimedia middleware. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 274– 283. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Chatterjee:1997:SDF

[1964] S. Chatterjee and W. Xiao. Selective discarding with FEC for ATMbased Internetworks. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 284–293. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Lindsay:1997:DTM

[1965] W. Lindsay and P. Ramanathan. DBP-M: A technique for meeting end-to-end (m, k)-firm guarantee requirements in point-to-point networks. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 294–305. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Cheon:1997:ESP

[1966] K. Cheon and S. S. Panwar. Early selective packet discard for alternating resource access of TCP over ATM-UBR. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 306– 316. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Molle:1997:REL

[1967] M. L. Molle and B. Yeger. Reducing the effects of large propagation delays on high speed IEEE 802.3 CSMA/CD networks using collision truncation. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 317–326. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Yamashita:1997:ECP

[1968] H. Yamashita, T. Suguri, and S. Kinoshita. An enhanced communication protocol for flexible interconnection. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 327– 336. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Frank:1997:PET

[1969] M. Frank and P. Martini. Practical experiences with a transport layer extension for end-to-end bandwidth regulation. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 337– 347. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Jeong:1997:ESM

[1970] T. I. Jeong, J. H. Jeong, and S. J.

Kim. An efficient scheduling mechanism using multiple thresholds for multimedia traffic in ATM switching nodes. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 348– 356. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

McKinley:1997:CGP

[1971] P. K. McKinley and Y. Huang. A centralized generic protocol for multipoint connections. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 357–368. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Miller:1997:RMP

[1972] C. K. Miller. Reliable multicast protocols: a practical view. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 369–378. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Johnstone:1997:RMU

[1973] G. S. Johnstone and G. D. Williams. Reliable multicast under error conditions. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 379– 387. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Fukuda:1997:RCR

[1974] K. Fukuda, N. Wakamiya, M. Murata, and H. Miyahara. Rate control of real-time MPEG-2 over ATM DBR service with bandwidth re-negotiation. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 388–397. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Ramanujan:1997:ASM

[1975] R. S. Ramanujan, J. A. Newhouse, M. Kaddoura, N., and A. Ahamad. Adaptive streaming of MPEG video over IP networks. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 398–409. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Balachandran:1997:NME

[1976] V. Balachandran, G. Bhat, S. Chakravarty, and A. Zhang. A network manager on Ethernets for distributed multimedia systems. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 410– 419. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Wong:1997:AIT

[1977] C. Wong and A. M. K. Cheng. An approach for imprecise transmission of TIFF image files through congested real-time ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 420–431. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Yokotani:1997:ULW

[1978] T. Yokotani, T. Ichihashi, and M. Yabe. An ultra light weight CAC mechanism (UL-CAC) for soft guarantee in ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 432–440. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Bonney:1997:DNS

[1979] J. Bonney, R. Ramanujan, and K. Thurber. Design of a network shared memory based workstation cluster. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 441–450. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Schelen:1997:AAA

[1980] O. Schelen and S. Pink. An agentbased architecture for advance reservations. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 451– 459. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Nour:1997:DRP

[1981] M. Nour, A. Hafid, and M. Gendreau. A dynamic routing procedure for connections with quality of service requirements. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 460–469. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Cao:1997:RAP

[1982] F. Cao. Reliability analysis of partitioned optical passive stars networks. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 470–475. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Wan:1997:NFC

[1983] P. J. Wan. On the number of fiber connections and star couplers in multi-star single-hop networks. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 476–481. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Abd-Elmalak:1997:PSR

[1984] S. A. Abd-Elmalak, C. Vaishnav, and A. P. Jayasumana. Performance of a symmetric robust WDM network when the channel access pattern at nodes is known. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 482–493. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Fragoulopoulos:1997:DEC

[1985] E. K. Fragoulopoulos, K. I. Mavrommatis, G. T. Kolyvas, and I. S. Venieris. Design and evaluation of a control plane architecture for the ATM access network. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 494–503. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Choi:1997:AED

[1986] Y. H. Choi. Algorithm-based error detection in ATM cell schedulers. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 504– 510. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Christensen:1997:RSA

[1987] K. J. Christensen and V. Ballingam.

Reduction of self-similarity by applicationlevel traffic shaping. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 511–518. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Mehaoua:1997:ASC

[1988] A. Mehaoua, R. Boutaba, and G. Pujolle. An adaptive and selective cell drop policy with dynamic data partitioning for best effort video over ATM. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 519– 529. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Fujii:1997:ESL

[1989] H. Fujii, S. Tanimoto, T. Hariu, and H. Miyashiro. An experimental study of logical office service over international ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 530–537. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Kunert:1997:IFT

[1990] O. Kunert and M. Zitterbart. Interconnecting FieldBuses through ATM. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 538–544. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Tzerefos:1997:CSS

[1991] P. Tzerefos, C. Smythe, I. Stergiou, and S. Cvetkovic. A comparative study of Simple Mail Transfer Protocol (SMTP), Post Office Protocol (POP) and X.400 electronic mail protocols. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 545– 554. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Malkin:1997:DVP

[1992] G. S. Malkin. Dial-in virtual private networks using layer 3 tunneling. In IEEE Computer Society. Technical Committee on Computer Communications [2052], pages 555– 561. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Tamura:1998:ERS

[1993] Y. Tamura, Y. Tobe, and H. Tokuda.

EFR: A retransmit scheme for TCP in wireless LANs. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 2– 11. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Khurana:1998:EHT

[1994] S. Khurana, A. Kahol, and A. P. Jayasumana. Effect of hidden terminals on the performance of IEEE 802.11 MAC protocol. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 12–20. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Al-Tawil:1998:NAP

[1995] K. Al-Tawil, A. Akrami, and H. Youssef.

A new authentication protocol for GSM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 21–30. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Lin:1998:QRA

[1996] C. R. Lin. QoS routing in ad hoc wireless networks. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 31–41. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Narvaez:1998:NTR

[1997] P. Narvaez and K.-Y. Siu. New techniques for regulating TCP flow over heterogeneous networks. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 42–51. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Nishida:1998:SSR

[1998] Y. Nishida. Smooth slow-start: Refining TCP slow-start for large-bandwidth with long-delay networks. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 52–60. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Petrovic:1998:ASA

[1999] S. S. Petrovic. ABR service in ATM networks: Performance comparison between adaptive stochastic and ERICA control schemes with guaranteed minimum cell rate. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 61–69. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Kalyanaraman:1998:PTA

[2000] S. Kalyanaraman, B. Vandalore, R. Jain, and R. Goyal. Performance of TCP over ABR with long-range dependent VBR background traffic over terrestrial and satellite ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 70–79. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Chen:1998:DQR

[2001] S. Chen and K. Nahrstedt. Distributed quality-of-service routing in high-speed networks based on selective probing. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 80–89. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Atlasis:1998:LTR

[2002] A. F. Atlasis, E. D. Baltatzis, Stassinopoulos, G. I., and I. Venieris. A linear-based trunk reservation routing algorithm for ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 90–98. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Oguchi:1998:RAR

[2003] N. Oguchi, Y.-M. Chen, J. Ogawa, and T. Tsuruoka. RISP: Address resolution protocol in network layer. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 99–108. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Ramanujan:1998:SAR

[2004] R. Ramanujan, S. Takkella, J. Bonney, and K. Thurber. Source-initiated adaptive routing algorithm (SARA) for autonomous wireless local area networks. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 109– 119. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Yokotani:1998:NAR

[2005] T. Yokotani, K. Murakami, T. Ichihashi, and S. Tonami. New approach of resource management to provide integrated services over ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 120– 129. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Boykin:1998:DST

[2006] M. S. Boykin and T. F. Znati. DQFS service: A tunable scheme to reduce scheduling overhead in integrated services networks. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 130–139. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Class:1998:MAS

[2007] C. Class and B. Stiller. A methodology to assess synchronization algorithms for distributed applications. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 140–149. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Banchs:1998:MMS

[2008] A. Banchs, W. Effelsberg, C. Tschudin, and V. Turau. Multicasting multimedia streams with active networks. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 150–161. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Koga:1998:SAU

[2009] Y. Koga, K. Shiomi, and Y. Matsushita. Service architecture for utilizing TINA-based public network services from the Internet environment. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 162– 171. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Schug:1998:HPI

[2010] K. Schug, A. Jayasumana, and P. Gopalakrishnan. High performance Integrated Network Communications Architecture (INCA). In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 172–181. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Dresler:1998:AMC

[2011] S. Dresler. ATM multipeer communication using a two-layer architecture. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 182– 189. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Harms:1998:PLD

[2012] J. J. Harms. Physical and logical design of ATM networks for multimedia traffic. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 190– 199. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Sisodia:1998:SMV

[2013] G. Sisodia, M. Hedley, S. De, and L. Guan. Source model for VBR coded video traffic in ATM networks. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 200– 207. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Nichols:1998:INS

[2014] K. M. Nichols. Improving network simulation with feedback. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 208–221. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Haenle:1998:PCR

[2015] C. Haenle and M. Hofmann. Performance comparison of reliable multicast protocols using the Network Simulator ns-2. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 222– 237. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Drobisz:1998:ASM

[2016] J. Drobisz and K. J. Christensen. Adaptive sampling methods to determine network traffic statistics including the Hurst parameter. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 238–249. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Sabbouh:1998:ROS

[2017] M. Sabbouh, K. Prasad, and W. Thompson. Replication in an object-oriented system. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 250– 259. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Weaver:1998:NPC

[2018] A. C. Weaver. A network protocol for cluster computing. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 260–269. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Hornfeldt:1998:CEA

[2019] J. Hornfeldt, M. Heumann, and W. P. Wilson. CommBridge — an enterprise application integration architecture for electronic government. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 270–277. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Mu:1998:NSC

[2020] Y. Mu and V. Varadharajan. A new scheme of credit based payment for electronic commerce. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 278–285. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Sobrinho:1998:EEQ

[2021] J. L. Sobrinho and A. S. Krishnakumar. EQuB — Ethernet quality of service using black bursts. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 286–296. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Kaminsky:1998:ODN

[2022] V. Kaminsky, L. Kacnelson, I. Gerlovin, and A. Zanis. On optimal design of network topology. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 297–304. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Ramanujan:1998:ARM

[2023] R. Ramanujan, A. Ahamad, and K. Thurber. An automated resource management tool for policy-driven QoS control within integrated services intranets. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 305– 314. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Frank:1998:FDL

[2024] M. Frank and P. Martini. Fairness and delay/loss study of an end-to-end bandwidth regulation scheme. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 315–325. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Papadimitriou:1998:OLC

[2025] G. I. Papadimitriou, A. N. Miliou, and A. S. Pomportsis. Optical logic circuits: A new approach to the control of fiber optic LANs. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 326– 335. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Christensen:1998:CGE

[2026] K. J. Christensen, M. Molle, and S. Li. Comparison of the gigabit Ethernet full-duplex repeater, CSMA/ CD, and 1000/100-Mbps switched Ethernet. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 336– 344. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Shin:1998:DIO

[2027] Y. Shin, J.-S. Ahn, H.-R. Lee, and D.-K. Jeong. Design and implementation of OTCA MAC protocol for high-speed point-to-point ring network. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 345– 352. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Varada:1998:DBM

[2028] S. Varada, Y. Yang, and D. Evans.

Data and buffer management in ATM systems. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 353– 363. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Norden:1998:NPH

[2029] S. Norden, G. Manimaran, and C. Siva Ram Murthy. New protocols for hard real-time communication in the switched LAN environment. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 364–373. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Abouaissa:1998:HAR

[2030] A. Abouaissa and A. Benslimane. Hierarchical architecture for real time causal delivery. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 374–383. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Jeong:1998:CLR

[2031] S.-H. Jeong and J. A. Copeland. Cell loss ratio and multiplexing gain of an ATM multiplexer for VBR voice sources. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 384– 389. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

Goyal:1998:PRG

[2032] R. Goyal, R. Jain, S. Fahmy, and B. Vandalore. Providing rate guarantees to TCP over the ATM GFR service. In IEEE Computer Society. Technical Committee on Computer Communications [2054], pages 390– 398. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.

IEEE:1985:CLC

[2033] IEEE, editor. 10th Conference on Local Computer Networks, Minneapolis Plaza Hotel, Minneapolis, Minnesota, October 7–9, 1985. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. ISBN 0-8186-0684-3 (paperback) 0-8186-4684-5 (microfiche) 0-8186-8684-7 (casebound). LCCN TK 5105.7 C66 1985. IEEE Catalog No. 85CH2222-8.

IEEE:1986:PCL

[2034] IEEE Computer Society. Technical Committee Computer Communications, editor. Proceedings / 11th Conference on Local Computer Networks, Minneapolis Plaza Hotel, Minneapolis, Minnesota, October 6–8, 1986. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. ISBN 0-8186-0730-0 (paperback), 0-8186-4730-2 (microfiche), 0-8186-8730-4 (casebound). LCCN TK 5105.7 C66 1986. IEEE Catalog no. 86CH2338-2. IEEE Computer Society order no. 730.

IEEE:1987:PCL

[2035] IEEE Computer Society. Technical Committee on Computer Communications, editor. Proceedings / 12th Conference on Local Computer Networks, Minneapolis Plaza Hotel, Minneapolis, Minnesota, October 5–7, 1987. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. ISBN 0-8186-0803-X, 0-8186-4803-1 (microfiche), 0-8186-8803-3 (casebound). LCCN TK 5105.7 C66 1987. IEEE Catalog no. 87CH2445-5. IEEE Computer Society order no. 803.

IEEE:1988:PCL

[2036] IEEE Computer Society. Technical Committee on Computer Communications, editor. Proceedings / 13th Conference on Local Computer Networks, Radisson Plaza Hotel, Minneapolis, Minnesota, October 10–12, 1988. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-4891-0 (microfiche). LCCN TK 5105.7 C66 1988a Microfiche.

IEEE:1989:CLC

[2037] IEEE Computer Society. Technical Committee Computer Communications, editor. 14th Conference on Local Computer Networks, Minneapolis, Minn., Oct. 10–12, 1989. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN 0-8186-1968-6. LCCN TK 5105.7 C66 1989. IEEE Computer Society order no. 1968. IEEE Catalog no. 89CH2746-6.

IEEE:1990:PCL

[2038] Institute of Electrical and Electronics Engineers, editor. Proceedings: 15th Conference on Local Computer Networks, Radisson Plaza Hotel, Minneapolis, Minnesota, September 30– October 3, 1990. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-2109-5. ISSN 0742-1303. LCCN TK5105.5.C66 1990. IEEE Computer Society order no. 2109. IEEE Catalog no. 90TH0335-0.

IEEE:1991:PCL

[2039] IEEE Computer Society. Technical Committee Computer Communications, editor. Proceedings / 16th Conference on Local Computer Networks, Minneapolis, Minnesota, October 14– 17, 1991. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-2370-5 (paperback), 0-8186-2371-3 (microfiche), 0-8186-2372-1 (casebound). LCCN TK 5105.7 C66 1991. IEEE Computer Society order no. 2370. IEEE Catalog no. 91TH0397-0.

IEEE:1992:PCL

[2040] IEEE Computer Society. Technical Committee on Computer Communications, editor. Proceedings / 17th Conference on Local Computer Networks, September 13–16, 1992, Minneapolis, Minnesota. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-3095-7 (paperback), 0-8186-3096-5 (microfiche), 0-8186-3097-3 (casebound). LCCN TK 5105.7 C66 1992. IEEE Computer Society Press order no. 3095. IEEE Catalog no. 92THO473-9.

IEEE:1993:PCL

[2041] IEEE Computer Society, editor. Proceedings: 18th Conference on Local Computer Networks, Minneapolis, Minnesota, September 19–22, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4510-5 (paperback), 0-8186-4511-3 (microfiche). LCCN TK 5105.7 C66 1993. IEEE Computer Society order no. 4510-02. IEEE Catalog no. 93TH0582-7.

Anonymous:1994:PIW

[2042] Anonymous, editor. Proceedings of the 2nd International World Wide Web conference, Mosaic and the Web, October 1994, Ramada-Congress Hotel, 520 South Michigan Avenue, Chicago, IL, volume 18(6) of Online & CDROM review: the international journal of online & optical information systems. Learned Information, Medford, NJ, USA, 1994. CO-DEN ONCDEW. ISSN 0309-314X.

URL http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/WWW2_Proceedings.
html.

IEEE:1994:PCL

[2043] IEEE Computer Society. Technical Committee on Computer Communications, editor. Proceedings / 19th Conference on Local Computer Networks, October 2–5, 1994, Minneapolis, Minnesota. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6680-3 (paperback), 0-8186-6681-1 (microfiche). LCCN TK5105.7 .C66 1994 Bar. IEEE Computer Society Press order number 6680-02. IEEE catalog number 94TH8004.

Plattner:1994:SPA

[2044] B. (Bernhard) Plattner, editor. Selected papers of the Annual Conference of the Internet Society/5th Joint European Networking Conference: June 3–17, 1994, Prague, Czech Republic, Computer networks and ISDN systems, v. 27, no. 3 (Dec. 1994) 0169-7552. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1994. ISBN ???? LCCN ????

Cailliau:1994:SPW

[2045] R. Cailliau, editor. Selected papers of the 1st World-Wide Web conference, 25–27 May 1994, Geneva, Switzerland, volume 27(2) of Computer Networks and ISDN Systems. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, November 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). LCCN ???? Published in Computer Networks and ISDN Systems, 27, No. 2.

IEEE:1995:PCL

[2046] IEEE Computer Society. Technical Committee on Computer Communications, editor. Proceedings: 20th Conference on Local Computer Networks, October 16–19, 1995, Minneapolis, Minnesota. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Kroemker:1995:PTI

[2047] D. Kroemker, editor. Proceedings of the Third International World-Wide Web Conference, 10–14 April, 1995, Darmstadt, Germany, volume 27(6) of Computer Networks and ISDN Systems. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). LCCN TK 5105.5 C647 v.27 no.6 1995.

OReilly:1995:WWW

[2048] O’Reilly and Associates and Web Consortium (W3C), editors. World Wide Web Journal: The Fourth International WWW Conference Proceedings. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1995. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95.

URL http://www.ora.com/gnn/bus/
ora/item/wj1.html. The World Wide

Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

IEEE:1996:ICL

[2049] IEEE Computer Society. Technical Committee on Computer Communications, editor. 21st IEEE Conference on Local Computer Networks: October 13–16, 1996, Minneapolis, Minnesota. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7617-5, 0-8186-7619-1 (microfiche). LCCN TK5105.5 .C66 1996. IEEE Computer Society Press order number PR07617. IEEE Order Plan number 96TB100073.

Anonymous:1996:FIW

[2050] Anonymous, editor. Fifth International World Wide Web Conference, May 6– 10, 1996, Paris, France, volume 28(7– 11) of Computer Networks and ISDN Systems. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, May 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). LCCN ???? Also published in Computer Networks and ISDN Systems, volume 28, number 7–11, May, 1996.

Anonymous:1997:PSI

[2051] Anonymous, editor. Papers from the Sixth International World Wide Web Conference, 7–11 April 1997, Santa Clara, volume 29(8–13) of Computer Networks and ISDN Systems. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). LCCN TK 5105.5 C647 v.29 no.8-13 1997.

IEEE:1997:PAC

[2052] IEEE Computer Society. Technical Committee on Computer Communications, editor. Proceedings, 22nd annual Conference on Local Computer Networks: LCN ’97: November 2–5, 1997, Minneapolis, Minnesota. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-8141-1, 0-8186-8142-X (casebound), 0-8186-8143-8 (microfiche). ISSN 0742-1303. LCCN TK5105.5 .C82 1997. IEEE Computer Society Press order number PR08141. IEEE Order Plan number 97TB100179.

Anonymous:1998:PIW

[2053] Anonymous, editor. Proceedings of the 7th International World Wide Web conference, April 1998, Brisbane, Australia, volume 30(1–7) of Computer Networks and ISDN Systems. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        decweb.ethz.ch/WWW7/00/.

IEEE:1998:LPA

[2054] IEEE Computer Society. Technical Committee on Computer Communications, editor. LCN’98: proceedings: 23rd Annual Conference on Local Computer Networks: October 11– 14, 1998, Lowell, Massachusetts. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8810-6, 0-8186-8818-1 (microfiche). LCCN TK5105.5 .C66 1998. IEEE Computer Society Press Order Number PR08810. IEEE Order Plan Catalog Number 98TB100260.